GNU TLS

Transport Layer Security Library for the GNU system for version 1.5.2, 21 September 2006

Nikos Mavroyanopoulos

Simon Josefsson (bug-gnutls@gnu.org)

This manual is last updated 21 September 2006 for version 1.5.2 of GNU TLS. Copyright (C) 2001, 2002, 2003, 2004, 2005, 2006 Free Software Foundation, Inc.

Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.2 or any later version published by the Free Software Foundation; with no Invariant Sections, no Front-Cover Texts, and no Back-Cover Texts. A copy of the license is included in the section entitled “GNU Free Documentation License”.

Table of Contents

1 Preface. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

2 The Library . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

2.1 General Idea . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2.2 Error handling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.3 Memory handling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.4 Callback functions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

3 Introduction to TLS . . . . . . . . . . . . . . . . . . . . . . . . 5

3.1 TLS layers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 3.2 The transport layer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.3 The TLS record protocol. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

3.3.1 Encryption algorithms used in the record layer . . . . . . . . . . . . 7 3.3.2 Compression algorithms used in the record layer . . . . . . . . . . 7 3.3.3 Weaknesses and countermeasures. . . . . . . . . . . . . . . . . . . . . . . . . 8

3.4 The TLS Alert Protocol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 3.5 The TLS Handshake Protocol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

3.5.1 TLS cipher suites . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 3.5.2 Client authentication. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 3.5.3 Resuming Sessions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 3.5.4 Resuming internals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

3.6 TLS Extensions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 3.6.1 Maximum fragment length negotiation . . . . . . . . . . . . . . . . . . 11 3.6.2 Server name indication . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

3.7 On SSL 2 and older protocols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

4 Authentication methods. . . . . . . . . . . . . . . . . . . 12

4.1 Certificate authentication . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 4.1.1 Authentication using X.509 certificates. . . . . . . . . . . . . . . . . . . 12 4.1.2 Authentication using OpenPGP keys . . . . . . . . . . . . . . . . . . . . . 12 4.1.3 Using certificate authentication . . . . . . . . . . . . . . . . . . . . . . . . . 12

4.2 Anonymous authentication . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 4.3 Authentication using SRP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 4.4 Authentication using PSK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 4.5 Authentication and credentials . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 4.6 Parameters stored in credentials . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

5 More on certificate authentication . . . . . . . . . 18

5.1 The X.509 trust model . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 5.1.1 X.509 certificates . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 5.1.2 Verifying X.509 certificate paths . . . . . . . . . . . . . . . . . . . . . . . . . 19 5.1.3 PKCS #10 certificate requests . . . . . . . . . . . . . . . . . . . . . . . . . . 21 5.1.4 PKCS #12 structures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

5.2 The OpenPGP trust model . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 5.2.1 OpenPGP keys . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 5.2.2 Verifying an OpenPGP key. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

5.3 Digital signatures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 5.3.1 Supported algorithms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 5.3.2 Trading security for interoperability . . . . . . . . . . . . . . . . . . . . . 24

6 How to use TLS in application protocols. … 26

6.1 Separate ports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 6.2 Upward negotiation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26

7 How to use GnuTLS in applications . . . . . . . . 28

7.1 Preparation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 7.1.1 Headers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 7.1.2 Version check. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 7.1.3 Building the source . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

7.2 Multi-threaded applications . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 7.3 Client examples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

7.3.1 Simple client example with anonymous authentication …. 30 7.3.2 Simple client example with X.509 certificate support . . . . . . 32 7.3.3 Obtaining session information . . . . . . . . . . . . . . . . . . . . . . . . . . 35 7.3.4 Verifying peer’s certificate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 7.3.5 Using a callback to select the certificate to use . . . . . . . . . . . 44 7.3.6 Client with Resume capability example . . . . . . . . . . . . . . . . . . 49 7.3.7 Simple client example with SRP authentication. . . . . . . . . . . 53

7.4 Server examples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 7.4.1 Echo Server with X.509 authentication . . . . . . . . . . . . . . . . . . . 56 7.4.2 Echo Server with X.509 authentication II . . . . . . . . . . . . . . . . 60 7.4.3 Echo Server with OpenPGP authentication . . . . . . . . . . . . . . . 67 7.4.4 Echo Server with SRP authentication . . . . . . . . . . . . . . . . . . . . 72 7.4.5 Echo Server with anonymous authentication . . . . . . . . . . . . . 75

7.5 Miscellaneous examples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 7.5.1 Checking for an alert. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 7.5.2 X.509 certificate parsing example . . . . . . . . . . . . . . . . . . . . . . . . 80 7.5.3 Certificate request generation . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 7.5.4 PKCS #12 structure generation . . . . . . . . . . . . . . . . . . . . . . . . . 84

7.6 Compatibility with the OpenSSL library . . . . . . . . . . . . . . . . . . . . . 87

8 Included programs . . . . . . . . . . . . . . . . . . . . . . . . 88

8.1 Invoking srptool. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 8.2 Invoking gnutls-cli. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 8.3 Invoking gnutls-cli-debug . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 8.4 Invoking gnutls-serv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 8.5 Invoking certtool . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91

9 Function reference . . . . . . . . . . . . . . . . . . . . . . . . 95

9.1 Core functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 9.2 X.509 certificate functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 9.3 GnuTLS-extra functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 9.4 OpenPGP functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 9.5 TLS Inner Application (TLS/IA) functions . . . . . . . . . . . . . . . . . . . 193 9.6 Error codes and descriptions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200

10 Certificate to XML convertion functions . . 206

10.1 An X.509 certificate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 10.2 An OpenPGP key. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209

11 All the supported ciphersuites in GnuTLS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211

12 Internal architecture of GnuTLS. . . . . . . . . 213

12.1 The TLS protocol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 12.2 TLS Handshake protocol. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 12.3 TLS authentication methods . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215 12.4 TLS Extension handling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 216 12.5 Certificate handling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 216

Appendix A Copying Information . . . . . . . . . . 218

A.1 GNU Free Documentation License . . . . . . . . . . . . . . . . . . . . . . . . . 218 A.2 GNU Lesser General Public License . . . . . . . . . . . . . . . . . . . . . . . . 224 A.3 GNU General Public License . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232

Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239

Function and Data Index . . . . . . . . . . . . . . . . . . . . 240

Bibliography. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244

1 Preface

This document tries to demonstrate and explain the GnuTLS library API. A brief introduction to the protocols and the technology involved, is also included so that an application programmer can better understand the GnuTLS purpose and actual offerings. Even if GnuTLS is a typical library software, it operates over several security and cryptographic protocols, which require the programmer to make careful and correct usage of them, otherwise he risks to offer just a false sense of security. Security and the network security terms are very general terms even for computer software thus cannot be easily restricted to a single cryptographic library. For that reason, do not consider a program secure just because it uses GnuTLS; there are several ways to compromise a program or a communication line and GnuTLS only helps with some of them. Although this document tries to be self contained, basic network programming and PKI knowlegde is assumed in most of it. A good introduction to networking can be found in [STEVENS] (See [Bibliography], page 244.) and for Public Key Infrastructure in [GUTPKI] (See [Bibliography], page 244.) . Updated versions of the GnuTLS software and this document will be available from

http://www.gnutls.org/ and http://www.gnu.org/software/gnutls/.

2 The Library

In brief GnuTLS can be described as a library which offers an API to access secure communication protocols. These protocols provide privacy over insecure lines, and were designed to prevent eavesdropping, tampering, or message forgery.

Technically GnuTLS is a portable ANSI C based library which implements the TLS 1.1 and SSL 3.0 protocols (See Chapter 3 [Introduction to TLS], page 5, for a more detailed description of the protocols), accompanied with the required framework for authentication and public key infrastructure. The library is available under the GNU Lesser GPL license^{1}. Important features of the GnuTLS library include:

• Support for TLS 1.0, TLS 1.1, and SSL 3.0 protocols.

• Support for both X.509 and OpenPGP certificates.

• Support for handling and verification of certificates.

• Support for SRP for TLS authentication.

• Support for PSK for TLS authentication.

• Support for TLS Extension mechanism.

• Support for TLS Compression Methods.

Additionally GnuTLS provides a limited emulation API for the widely used OpenSSL^{2} library, to ease integration with existing applications.

GnuTLS consists of three independent parts, namely the “TLS protocol part”, the “Certificate part”, and the “Crypto backend” part. The ‘TLS protocol part’ is the actual protocol implementation, and is entirely implemented within the GnuTLS library. The ‘Certificate part’ consists of the certificate parsing, and verification functions which is partially implemented in the GnuTLS library. The Libtasn1^{3}, a library which offers ASN.1 parsing capabilities, is used for the X.509 certificate parsing functions, and Opencdk^{4} is used for the OpenPGP key support in GnuTLS. The “Crypto backend” is provided by the Libgcrypt^{5} library.

In order to ease integration in embedded systems, parts of the GnuTLS library can be disabled at compile time. That way a small library, with the required features, can be generated.

1 A copy of the license is included in the distribution 2 http://www.openssl.org/ 3 ftp://ftp.gnupg.org/gcrypt/alpha/gnutls/libtasn1/ 4 ftp://ftp.gnupg.org/gcrypt/alpha/gnutls/opencdk/ 5 ftp://ftp.gnupg.org/gcrypt/alpha/libgcrypt/

2.1 General Idea

A brief description of how GnuTLS works internally is shown at the figure below. This section may be easier to understand after having seen the examples (see [examples], page 28).

As shown in the figure, there is a read-only global state that is initialized once by the global initialization function. This global structure, among others, contains the memory allocation functions used, and some structures needed for the ASN.1 parser. This structure is never modified by any GnuTLS function, except for the deinitialization function which frees all memory allocated in the global structure and is called after the program has permanently finished using GnuTLS. The credentials structure is used by some authentication methods, such as certificate authentication (see [Certificate Authentication], page 18). A credentials structure may contain certificates, private keys, temporary parameters for diffie hellman or RSA key exchange, and other stuff that may be shared between several TLS sessions. This structure should be initialized using the appropriate initialization functions. For example an application which uses certificate authentication would probably initialize the credentials, using the appropriate functions, and put its trusted certificates in this structure. The next step is to associate the credentials structure with each TLS session. A GnuTLS session contains all the required stuff for a session to handle one secure connection. This session calls directly to the transport layer functions, in order to communicate with the peer. Every session has a unique session ID shared with the peer. Since TLS sessions can be resumed, servers would probably need a database backend to hold the session’s parameters. Every GnuTLS session after a successful handshake calls the appropriate backend function (See [resume], page 10, for information on initialization) to store the newly negotiated session. The session database is examined by the server just after having received the client hello^{6}, and if the session ID sent by the client, matches a stored session, the stored session will be retrieved, and the new session will be a resumed one, and will share the same session ID with the previous one.

6 The first message in a TLS handshake

2.2 Error handling

In GnuTLS most functions return an integer type as a result. In almost all cases a zero or a positive number means success, and a negative number indicates failure, or a situation that some action has to be taken. Thus negative error codes may be fatal or not. Fatal errors terminate the connection immediately and further sends and receives will be disallowed. An example of a fatal error code is GNUTLS_E_DECRYPTION_FAILED. Non-fatal errors may warn about something, i.e., a warning alert was received, or indicate the some action has to be taken. This is the case with the error code GNUTLS_E_REHANDSHAKE returned by [gnutls record recv], page 128. This error code indicates that the server requests a rehandshake. The client may ignore this request, or may reply with an alert. You can test if an error code is a fatal one by using the [gnutls error is fatal], page 116. If any non fatal errors, that require an action, are to be returned by a function, these error codes will be documented in the function’s reference. See [Error Codes], page 200, for all the error codes.

2.3 Memory handling

GnuTLS internally handles heap allocated objects differently, depending on the sensitivity of the data they contain. However for performance reasons, the default memory functions do not overwrite sensitive data from memory, nor protect such objects from being written to the swap. In order to change the default behavior the [gnutls global set mem functions], page 118 function is available which can be used to set other memory handlers than the defaults. The Libgcrypt library on which GnuTLS depends, has such secure memory allocation functions available. These should be used in cases where even the system’s swap memory is not considered secure. See the documentation of Libgcrypt for more information.

2.4 Callback functions

There are several cases where GnuTLS may need some out of band input from your program. This is now implemented using some callback functions, which your program is expected to register. An example of this type of functions are the push and pull callbacks which are used to specify the functions that will retrieve and send data to the transport layer.

• [gnutls transport set push function], page 142

• [gnutls transport set pull function], page 141

Other callback functions such as the one set by [gnutls srp set server credentials function], page 139, may require more complicated input, including data to be allocated. These callbacks should allocate and free memory using the functions shown below.

• [gnutls malloc], page 121

• [gnutls free], page 116

3 Introduction to TLS

TLS stands for “Transport Layer Security” and is the successor of SSL, the Secure Sockets Layer protocol [SSL3] (See [Bibliography], page 244.) designed by Netscape. TLS is an Internet protocol, defined by IETF^{1}, described in RFC 2246 and also in [RESCOLA] (See [Bibliography], page 244.) . The protocol provides confidentiality, and authentication layers over any reliable transport layer. The description, below, refers to TLS 1.0 but also applies to TLS 1.1 [RFC4346] (See [Bibliography], page 244.) and SSL 3.0, since the differences of these protocols are minor. Older protocols such as SSL 2.0 are not discussed nor implemented in GnuTLS since they are not considered secure today.

3.1 TLS layers

TLS is a layered protocol, and consists of the Record Protocol, the Handshake Protocol and the Alert Protocol. The Record Protocol is to serve all other protocols and is above the transport layer. The Record protocol offers symmetric encryption, data authenticity, and optionally compression.

The Alert protocol offers some signaling to the other protocols. It can help informing the peer for the cause of failures and other error conditions. See [The Alert Protocol], page 8, for more information. The alert protocol is above the record protocol.

The Handshake protocol is responsible for the security parameters’ negotiation, the initial key exchange and authentication. See [The Handshake Protocol], page 8, for more information about the handshake protocol. The protocol layering in TLS is shown in the figure below.

1 IETF, or Internet Engineering Task Force, is a large open international community of network designers,

operators, vendors, and researchers concerned with the evolution of the Internet architecture and the smooth operation of the Internet. It is open to any interested individual.

3.2 The transport layer

TLS is not limited to one transport layer, it can be used above any transport layer, as long as it is a reliable one. A set of functions is provided and their purpose is to load to GnuTLS the required callbacks to access the transport layer.

• [gnutls transport set push function], page 142

• [gnutls transport set pull function], page 141

• [gnutls transport set ptr], page 141

• [gnutls transport set lowat], page 141

These functions accept a callback function as a parameter. The callback functions should return the number of bytes written, or -1 on error and should set errno appropriately.

GnuTLS currently only interprets the EINTR and EAGAIN errno values and returns the corresponding GnuTLS error codes GNUTLS_E_INTERRUPTED and GNUTLS_E_AGAIN. These values are usually returned by interrupted system calls, or when non blocking IO is used. All GnuTLS functions can be resumed (called again), if any of these error codes is returned. The error codes above refer to the system call, not the GnuTLS function, since signals do not interrupt GnuTLS’ functions.

For non blocking sockets or other custom made pull/push functions the [gnutls transport set lowat], page 141 must be called, with a zero low water mark value.

By default, if the transport functions are not set, GnuTLS will use the Berkeley Sockets functions. In this case GnuTLS will use some hacks in order for select to work, thus making it easy to add TLS support to existing TCP/IP servers.

3.3 The TLS record protocol

The Record protocol is the secure communications provider. Its purpose is to encrypt, authenticate and —optionally— compress packets. The following functions are available:

[gnutls record send], page 129:

To send a record packet (with application data).

[gnutls record recv], page 128:

To receive a record packet (with application data).

[gnutls record get direction], page 128:

To get the direction of the last interrupted function call.

As you may have already noticed, the functions which access the Record protocol, are quite limited, given the importance of this protocol in TLS. This is because the Record protocol’s parameters are all set by the Handshake protocol.

The Record protocol initially starts with NULL parameters, which means no encryption, and no MAC is used. Encryption and authentication begin just after the handshake protocol has finished.

3.3.1 Encryption algorithms used in the record layer

Confidentiality in the record layer is achieved by using symmetric block encryption algorithms like 3DES, AES^{2}, or stream algorithms like ARCFOUR_128^{3}. Ciphers are encryption algorithms that use a single, secret, key to encrypt and decrypt data. Block algorithms in TLS also provide protection against statistical analysis of the data. Thus, if you’re using the TLS protocol, a random number of blocks will be appended to data, to prevent eavesdroppers from guessing the actual data size. Supported cipher algorithms:

3DES_CBC 3DES_CBC is the DES block cipher algorithm used with triple encryption (EDE). Has 64 bits block size and is used in CBC mode.

ARCFOUR_128

ARCFOUR is a fast stream cipher.

ARCFOUR_40

This is the ARCFOUR cipher that is fed with a 40 bit key, which is considered weak.

AES_CBC AES or RIJNDAEL is the block cipher algorithm that replaces the old DES algorithm. Has 128 bits block size and is used in CBC mode. This is not officially supported in TLS.

Supported MAC algorithms:

MAC_MD5 MD5 is a cryptographic hash algorithm designed by Ron Rivest. Outputs 128 bits of data.

MAC_SHA SHA is a cryptographic hash algorithm designed by NSA. Outputs 160 bits of data.

3.3.2 Compression algorithms used in the record layer

The TLS record layer also supports compression. The algorithms implemented in GnuTLS can be found in the table below. All the algorithms except for DEFLATE which is referenced in [RFC3749] (See [Bibliography], page 244.) , should be considered as GnuTLS’ extensions^{4}, and should be advertised only when the peer is known to have a compliant client, to avoid interoperability problems. The included algorithms perform really good when text, or other compressible data are to be transfered, but offer nothing on already compressed data, such as compressed images, zipped archives etc. These compression algorithms, may be useful in high bandwidth TLS tunnels, and in cases where network usage has to be minimized. As a drawback, compression increases latency. The record layer compression in GnuTLS is implemented based on the proposal [RFC3749] (See [Bibliography], page 244.) . The supported compression algorithms are:

DEFLATE Zlib compression, using the deflate algorithm.

2 AES, or Advanced Encryption Standard, is actually the RIJNDAEL algorithm. This is the algorithm that

replaced DES.

3 ARCFOUR_128 is a compatible algorithm with RSA’s RC4 algorithm, which is considered to be a trade secret.

4 You should use [gnutls handshake set private extensions], page 119 to enable private extensions.

LZO LZO is a very fast compression algorithm. This algorithm is only available if the GnuTLS-extra library has been initialized and the private extensions are enabled.

3.3.3 Weaknesses and countermeasures

Some weaknesses that may affect the security of the Record layer have been found in TLS 1.0 protocol. These weaknesses can be exploited by active attackers, and exploit the facts that

1. TLS has separate alerts for “decryption failed” and “bad record mac”

2. The decryption failure reason can be detected by timing the response time.

3. The IV for CBC encrypted packets is the last block of the previous encrypted packet.

Those weaknesses were solved in TLS 1.1 [RFC4346] (See [Bibliography], page 244.) which is implemented in GnuTLS. For a detailed discussion see the archives of the TLS Working Group mailing list and the paper [CBCATT] (See [Bibliography], page 244.) .

3.4 The TLS Alert Protocol

The Alert protocol is there to allow signals to be sent between peers. These signals are mostly used to inform the peer about the cause of a protocol failure. Some of these signals are used internally by the protocol and the application protocol does not have to cope with them (see GNUTLS_A_CLOSE_NOTIFY), and others refer to the application protocol solely (see GNUTLS_A_USER_CANCELLED). An alert signal includes a level indication which may be either fatal or warning. Fatal alerts always terminate the current connection, and prevent future renegotiations using the current session ID. The alert messages are protected by the record protocol, thus the information that is included does not leak. You must take extreme care for the alert information not to leak to a possible attacker, via public log files etc.

[gnutls alert send], page 95:

To send an alert signal.

[gnutls error to alert], page 116:

To map a gnutls error number to an alert signal.

[gnutls alert get], page 95:

Returns the last received alert.

[gnutls alert get name], page 95:

Returns the name, in a character array, of the given alert.

3.5 The TLS Handshake Protocol

The Handshake protocol is responsible for the ciphersuite negotiation, the initial key exchange, and the authentication of the two peers. This is fully controlled by the application layer, thus your program has to set up the required parameters. Available functions to control the handshake protocol include:

[gnutls cipher set priority], page 109:

To set the priority of bulk cipher algorithms.

[gnutls mac set priority], page 121:

To set the priority of MAC algorithms.

[gnutls kx set priority], page 120:

To set the priority of key exchange algorithms.

[gnutls compression set priority], page 109:

To set the priority of compression methods.

[gnutls certificate type set priority], page 107:

To set the priority of certificate types (e.g., OpenPGP, X.509).

[gnutls protocol set priority], page 125:

To set the priority of protocol versions (e.g., SSL 3.0, TLS 1.0).

[gnutls set default priority], page 136:

To set some defaults in the current session. That way you don’t have to call each priority function, independently, but you have to live with the defaults.

[gnutls credentials set], page 110:

To set the appropriate credentials structures.

[gnutls certificate server set request], page 101:

To set whether client certificate is required or not.

[gnutls handshake], page 119:

To initiate the handshake.

3.5.1 TLS cipher suites

The Handshake Protocol of TLS negotiates cipher suites of the form TLS_DHE_RSA_WITH_ 3DES_CBC_SHA. The usual cipher suites contain these parameters:

• The key exchange algorithm. DHE_RSA in the example.

• The Symmetric encryption algorithm and mode 3DES_CBC in this example.

• The MAC^{5} algorithm used for authentication. MAC_SHA is used in the above example.

The cipher suite negotiated in the handshake protocol will affect the Record Protocol, by enabling encryption and data authentication. Note that you should not over rely on TLS to negotiate the strongest available cipher suite. Do not enable ciphers and algorithms that you consider weak. The priority functions, dicussed above, allow the application layer to enable and set priorities on the individual ciphers. It may imply that all combinations of ciphersuites are allowed, but this is not true. For several reasons, not discussed here, some combinations were not defined in the TLS protocol. The supported ciphersuites are shown in [ciphersuites], page 211.

3.5.2 Client authentication

In the case of ciphersuites that use certificate authentication, the authentication of the client is optional in TLS. A server may request a certificate from the client — using the [gnutls certificate server set request], page 101 function. If a certificate is to be requested from the client during the handshake, the server will send a certificate request message that contains a list of acceptable certificate signers. In GnuTLS the certificate signers list is constructed using the trusted Certificate Authorities by the server. That is the ones set using

5 MAC stands for Message Authentication Code. It can be described as a keyed hash algorithm. See RFC2104.

• [gnutls certificate set x509 trust file], page 106

• [gnutls certificate set x509 trust mem], page 106

Sending of the names of the CAs can be controlled using [gnutls certificate send x509 rdn sequence], page 101. The client, then, may send a certificate, signed by one of the server’s acceptable signers.

3.5.3 Resuming Sessions

The [gnutls handshake], page 119 function, is expensive since a lot of calculations are performed. In order to support many fast connections to the same server a client may use session resuming. Session resuming is a feature of the TLS protocol which allows a client to connect to a server, after a successful handshake, without the expensive calculations. This is achieved by using the previously established keys. GnuTLS supports this feature, and the example (see [ex:resume-client], page 50) illustrates a typical use of it. Keep in mind that sessions are expired after some time, for security reasons, thus it may be normal for a server not to resume a session even if you requested that. Also note that you must enable, using the priority functions, at least the algorithms used in the last session.

3.5.4 Resuming internals

The resuming capability, mostly in the server side, is one of the problems of a thread-safe TLS implementations. The problem is that all threads must share information in order to be able to resume sessions. The gnutls approach is, in case of a client, to leave all the burden of resuming to the client. I.e., copy and keep the necessary parameters. See the functions:

• [gnutls session get data], page 134

• [gnutls session get id], page 134

• [gnutls session set data], page 135

The server side is different. A server has to specify some callback functions which store, retrieve and delete session data. These can be registered with:

• [gnutls db set remove function], page 111

• [gnutls db set store function], page 112

• [gnutls db set retrieve function], page 112

• [gnutls db set ptr], page 111

It might also be useful to be able to check for expired sessions in order to remove them, and save space. The function [gnutls db check entry], page 110 is provided for that reason.

3.6 TLS Extensions

A number of extensions to the TLS protocol have been proposed mainly in [TLSEXT] (See [Bibliography], page 244.) . The extensions supported in GnuTLS are:

• Maximum fragment length negotiation

• Server name indication

and they will be discussed in the subsections that follow.

3.6.1 Maximum fragment length negotiation

This extension allows a TLS implementation to negotiate a smaller value for record packet maximum length. This extension may be useful to clients with constrained capabilities. See the [gnutls record set max size], page 129 and the [gnutls record get max size], page 128 functions.

3.6.2 Server name indication

A common problem in HTTPS servers is the fact that the TLS protocol is not aware of the hostname that a client connects to, when the handshake procedure begins. For that reason the TLS server has no way to know which certificate to send. This extension solves that problem within the TLS protocol, and allows a client to send the HTTP hostname before the handshake begins within the first handshake packet. The functions [gnutls server name set], page 133 and [gnutls server name get], page 132 can be used to enable this extension, or to retrieve the name sent by a client.

3.7 On SSL 2 and older protocols

One of the initial decisions in the GnuTLS development was to implement the known security protocols for the transport layer. Initially TLS 1.0 was implemented since it was the latest at that time, and was considered to be the most advanced in security properties. Later the SSL 3.0 protocol was implemented since it is still the only protocol supported by several servers and there are no serious security vulnerabilities known. One question that may arise is why we didn’t implement SSL 2.0 in the library. There are several reasons, most important being that it has serious security flaws, unacceptable for a modern security library. Other than that, this protocol is barely used by anyone these days since it has been deprecated since 1996. Other protocols such as Microsoft’s PCT 1 and PCT 2 were not implemented because they were also abandoned and deprecated by SSL 3.0 and later TLS 1.0.

4 Authentication methods

The TLS protocol provides confidentiality and encryption, but also offers authentication, which is a prerequisite for a secure connection. The available authentication methods in

GnuTLS are:

• Certificate authentication

• Anonymous authentication

• SRP authentication

• PSK authentication

4.1 Certificate authentication

4.1.1 Authentication using X.509 certificates

X.509 certificates contain the public parameters, of a public key algorithm, and an authority’s signature, which proves the authenticity of the parameters. See Section 5.1 [The X.509 trust model], page 18, for more information on X.509 protocols.

4.1.2 Authentication using OpenPGP keys

OpenPGP keys also contain public parameters of a public key algorithm, and signatures from several other parties. Depending on whether a signer is trusted the key is considered trusted or not. GnuTLS’s OpenPGP authentication implementation is based on the [TLSPGP] (See [Bibliography], page 244.) proposal. See Section 5.2 [The OpenPGP trust model], page 21, for more information about the OpenPGP trust model. For a more detailed introduction to OpenPGP and GnuPG see [GPGH] (See [Bibliography], page 244.) .

4.1.3 Using certificate authentication

In GnuTLS both the OpenPGP and X.509 certificates are part of the certificate authentication and thus are handled using a common API. When using certificates the server is required to have at least one certificate and private key pair. A client may or may not have such a pair. The certificate and key pair should be loaded, before any TLS session is initialized, in a certificate credentials structure. This should be done by using [gnutls certificate set x509 key file], page 104 or [gnutls certificate set openpgp key file], page 184 depending on the certificate type. In the X.509 case, the functions will also accept and use a certificate list that leads to a trusted authority. The certificate list must be ordered in such way that every certificate certifies the one before it. The trusted authority’s certificate need not to be included, since the peer should possess it already. As an alternative, a callback may be used so the server or the client specify the certificate and the key at the handshake time. That callback can be set using the functions:

• [gnutls certificate server set retrieve function], page 102

• [gnutls certificate client set retrieve function], page 99

Certificate verification is possible by loading the trusted authorities into the credentials structure by using [gnutls certificate set x509 trust file], page 106 or [gnutls certificate set openpgp keyring file], page 185 for openpgp keys. Note however that the peer’s certificate is not automatically verified, you should call [gnutls certificate verify peers2], page 107, after a successful handshake, to verify the signatures of the certificate. An alternative way, which reports a more detailed verification output, is to use [gnutls certificate get peers], page 101 to obtain the raw certificate of the peer and verify it using the functions discussed in Section 5.1 [The X.509 trust model], page 18. In a handshake, the negotiated cipher suite depends on the certificate’s parameters, so not all key exchange methods will be available with some certificates. GnuTLS will disable ciphersuites that are not compatible with the key, or the enabled authentication methods. For example keys marked as sign-only, will not be able to access the plain RSA ciphersuites, but only the DHE_RSA ones. It is recommended not to use RSA keys for both signing and encryption. If possible use the same key for the DHE_RSA and RSA_EXPORT ciphersuites, which use signing, and a different key for the plain RSA ciphersuites, which use encryption. All the key exchange methods shown below are available in certificate authentication. Note that the DHE key exchange methods are generally slower^{1} than plain RSA and require Diffie Hellman parameters to be generated and associated with a credentials structure, by the server. The RSA-EXPORT method also requires 512 bit RSA parameters, that should also be generated and associated with the credentials structure. See the functions:

• [gnutls dh params generate2], page 114

• [gnutls certificate set dh params], page 102

• [gnutls rsa params generate2], page 131

• [gnutls certificate set rsa export params], page 102

Sometimes in order to avoid bottlenecks in programs it is usefull to store and read parameters from formats that can be generated by external programs such as certtool. This is possible with GnuTLS by using the following functions:

• [gnutls dh params import pkcs3], page 115

• [gnutls rsa params import pkcs1], page 132

• [gnutls dh params export pkcs3], page 114

• [gnutls rsa params export pkcs1], page 131

Key exchange algorithms for OpenPGP and X.509 certificates:

RSA: The RSA algorithm is used to encrypt a key and send it to the peer. The certificate must allow the key to be used for encryption.

RSA_EXPORT:

The RSA algorithm is used to encrypt a key and send it to the peer. In the EXPORT algorithm, the server signs temporary RSA parameters of 512 bits — which are considered weak — and sends them to the client.

DHE_RSA: The RSA algorithm is used to sign Ephemeral Diffie Hellman parameters which are sent to the peer. The key in the certificate must allow the key to be used for signing. Note that key exchange algorithms which use Ephemeral Diffie Hellman parameters, offer perfect forward secrecy. That means that even if the private key used for signing is compromised, it cannot be used to reveal past session data.

1 It really depends on the group used. Primes with lesser bits are always faster, but also easier to break.

Values less than 768 should not be used today

DHE_DSS: The DSS algorithm is used to sign Ephemeral Diffie Hellman parameters which are sent to the peer. The certificate must contain DSA parameters to use this key exchange algorithm. DSS stands for Digital Signature Standard.

4.2 Anonymous authentication

The anonymous key exchange performs encryption but there is no indication of the identity of the peer. This kind of authentication is vulnerable to a man in the middle attack, but this protocol can be used even if there is no prior communication and trusted parties with the peer, or when full anonymity is required. Unless really required, do not use anonymous authentication. Available key exchange methods are shown below.

Note that the key exchange methods for anonymous authentication require Diffie Hellman parameters to be generated by the server and associated with an anonymous credentials structure.

Supported anonymous key exchange algorithms:

ANON_DH: This algorithm exchanges Diffie Hellman parameters.

4.3 Authentication using SRP

Authentication via the Secure Remote Password protocol, SRP^{2}, is supported. The SRP key exchange is an extension to the TLS protocol, and it is a password based authentication (unlike X.509 or OpenPGP that use certificates). The two peers can be identified using a single password, or there can be combinations where the client is authenticated using SRP and the server using a certificate.

The advantage of SRP authentication, over other proposed secure password authentication schemes, is that SRP does not require the server to hold the user’s password. This kind of protection is similar to the one used traditionally in the UNIX ‘/etc/passwd’ file, where the contents of this file did not cause harm to the system security if they were revealed. The SRP needs instead of the plain password something called a verifier, which is calculated using the user’s password, and if stolen cannot be used to impersonate the user. Check [TOMSRP] (See [Bibliography], page 244.) for a detailed description of the SRP protocol and the Stanford SRP libraries, which includes a PAM module that synchronizes the system’s users passwords with the SRP password files. That way SRP authentication could be used for all the system’s users.

The implementation in GnuTLS is based on paper [TLSSRP] (See [Bibliography], page 244.) . The supported SRP key exchange methods are:

SRP: Authentication using the SRP protocol.

SRP_DSS: Client authentication using the SRP protocol. Server is authenticated using a certificate with DSA parameters.

2 SRP is described in [RFC2945] (See [Bibliography], page 244.)

SRP_RSA: Client authentication using the SRP protocol. Server is authenticated using a certificate with RSA parameters.

If clients supporting SRP know the username and password before the connection, should initialize the client credentials and call the function [gnutls srp set client credentials], page 139. Alternatively they could specify a callback function by using the function [gnutls srp set client credentials function], page 138. This has the advantage that allows probing the server for SRP support. In that case the callback function will be called twice per handshake. The first time is before the ciphersuite is negotiated, and if the callback returns a negative error code, the callback will be called again if SRP has been negotiated. This uses a special TLS-SRP handshake idiom in order to avoid, in interactive applications, to ask the user for SRP password and username if the server does not negotiate an SRP ciphersuite. In server side the default behaviour of GnuTLS is to read the usernames and SRP verifiers from password files. These password files are the ones used by the Stanford srp libraries and can be specified using the [gnutls srp set server credentials file], page 139. If a different password file format is to be used, then the function [gnutls srp set server credentials function], page 139, should be called, in order to set an appropriate callback. Some helper functions such as

• [gnutls srp verifier], page 140

• [gnutls srp base64 encode], page 137

• [gnutls srp base64 decode], page 137

are included in GnuTLS, and can be used to generate and maintain SRP verifiers and password files. A program to manipulate the required parameters for SRP authentication is also included. See [srptool], page 88, for more information.

4.4 Authentication using PSK

Authentication using Pre-shared keys is a method to authenticate using usernames and binary keys. This protocol avoids making use of public key infrastructure and expensive calculations, thus it is suitable for constraint clients. The implementation in GnuTLS is based on paper [TLSPSK] (See [Bibliography], page 244.) . The supported PSK key exchange methods are:

PSK: Authentication using the PSK protocol.

DHE-PSK: Authentication using the PSK protocol and Diffie Hellman key exchange. This method offers perfect forward secrecy.

Clients supporting PSK should supply the username and key before the connection to the client credentials by calling the function [gnutls psk set client credentials], page 126. Alternatively they could specify a callback function by using the function [gnutls psk set client credentials function], page 126. This has the advantage that the callback will be called only if PSK has been negotiated. In server side the default behaviour of GnuTLS is to read the usernames and PSK keys from a password file. The password file should contain usernames and keys in hexadecimal format. The name of the password file can be stored to the credentials structure by calling [gnutls psk set server credentials file], page 127. If a different password file format is to be used, then the function [gnutls psk set server credentials function], page 127, should be used instead. Some helper functions such as:

• [gnutls hex encode], page 120

• [gnutls hex decode], page 119

are included in GnuTLS, and may be used to generate and maintain PSK keys.

4.5 Authentication and credentials

In GnuTLS every key exchange method is associated with a credentials type. So in order to enable to enable a specific method, the corresponding credentials type should be initialized and set using [gnutls credentials set], page 110. A mapping is shown below. Key exchange algorithms and the corresponding credential types:

Key exchange  Client credentials  Server credentials

KX_RSA
KX_DHE_RSA
KX_DHE_DSS

KX_RSA_EXPORT            CRD_CERTIFICATE        CRD_CERTIFICATE
KX_SRP_RSA               CRD_SRP                CRD_SRP
KX_SRP_DSS                                     CRD_CERTIFICATE
KX_SRP                   CRD_SRP                CRD_SRP
KX_ANON_DH               CRD_ANON               CRD_ANON
KX_PSK                   CRD_PSK                CRD_PSK

4.6 Parameters stored in credentials

Several parameters such as the ones used for Diffie-Hellman authentication are stored within the credentials structures, so all sessions can access them. Those parameters are stored in structures such as gnutls_dh_params_t and gnutls_rsa_params_t, and functions like [gnutls certificate set dh params], page 102 and [gnutls certificate set rsa export params], page 102 can be used to associate those parameters with the given credentials structure. Since those parameters need to be renewed from time to time and a global structure such as the credentials, may not be easy to modify since it is accessible by all sessions, an alternative interface is available using a callback function. This can be set using the [gnutls certificate set params function], page 102. An example is shown below.

     #include <gnutls.h>

gnutls_rsa_params_t rsa_params;
gnutls_dh_params_t dh_params;
/* This function will be called once a session requests DH
* or RSA parameters. The parameters returned (if any) will
* be used for the first handshake only.
*/
static int get_params( gnutls_session_t session,
     gnutls_params_type_t type,
     gnutls_params_st *st)
{
if (type == GNUTLS_PARAMS_RSA_EXPORT)
   st->params.rsa_export = rsa_params;
else if (type == GNUTLS_PARAMS_DH)
   st->params.dh = dh_params;
else return -1;
st->type = type;
/* do not deinitialize those parameters.
 */
st->deinit = 0;
return 0;
}
int main()
{
gnutls_certificate_credentials_t cert_cred;
initialize_params();
/* ...
 */
gnutls_certificate_set_params_function( cert_cred, get_params);
}

5 More on certificate authentication

5.1 The X.509 trust model

The X.509 protocols rely on a hierarchical trust model. In this trust model Certification Authorities (CAs) are used to certify entities. Usually more than one certification authorities exist, and certification authorities may certify other authorities to issue certificates as well, following a hierarchical model.

One needs to trust one or more CAs for his secure communications. In that case only the certificates issued by the trusted authorities are acceptable. See the figure above for a typical example. The API for handling X.509 certificates is described at section [sec:x509api], page 142. Some examples are listed below.

5.1.1 X.509 certificates

An X.509 certificate usually contains information about the certificate holder, the signer, a unique serial number, expiration dates and some other fields [RFC3280] (See [Bibliography], page 244.) as shown in the table below.

version: The field that indicates the version of the certificate.

serialNumber:

This field holds a unique serial number per certificate.

issuer: Holds the issuer’s distinguished name.

validity:

The activation and expiration dates.

subject: The subject’s distinguished name of the certificate.

extensions:

The extensions are fields only present in version 3 certificates.

The certificate’s subject or issuer name is not just a single string. It is a Distinguished name and in the ASN.1 notation is a sequence of several object IDs with their corresponding values. Some of available OIDs to be used in an X.509 distinguished name are defined in

‘gnutls/x509.h’.

The Version field in a certificate has values either 1 or 3 for version 3 certificates. Version 1 certificates do not support the extensions field so it is not possible to distinguish a CA from a person, thus their usage should be avoided. The validity dates are there to indicate the date that the specific certificate was activated and the date the certificate’s key would be considered invalid. Certificate extensions are there to include information about the certificate’s subject that did not fit in the typical certificate fields. Those may be e-mail addresses, flags that indicate whether the belongs to a CA etc. All the supported X.509 version 3 extensions are shown in the table below.

subject key id (2.5.29.14):

An identifier of the key of the subject.

authority key id (2.5.29.35):

An identifier of the authority’s key used to sign the certificate.

subject alternative name (2.5.29.17):

Alternative names to subject’s distinguished name.

key usage (2.5.29.15):

Constraints the key’s usage of the certificate.

extended key usage (2.5.29.37):

Constraints the purpose of the certificate.

basic constraints (2.5.29.19):

Indicates whether this is a CA certificate or not.

CRL distribution points (2.5.29.31):

This extension is set by the CA, in order to inform about the issued CRLs.

In GnuTLS the X.509 certificate structures are handled using the gnutls_x509_crt_t type and the corresponding private keys with the gnutls_x509_privkey_t type. All the available functions for X.509 certificate handling have their prototypes in ‘gnutls/x509.h’. An example program to demonstrate the X.509 parsing capabilities can be found at section [ex:x509-info], page 80.

5.1.2 Verifying X.509 certificate paths

Verifying certificate paths is important in X.509 authentication. For this purpose the function [gnutls x509 crt verify], page 177 is provided. The output of this function is the bitwise OR of the elements of the gnutls_certificate_status_t enumeration. A detailed description of these elements can be found in figure below. The function [gnutls certificate verify peers2], page 107 is equivalent to the previous one, and will verify the peer’s certificate in a TLS session.

CERT_INVALID:

The certificate is not signed by one of the known authorities, or the signature is invalid.

CERT_REVOKED:

The certificate has been revoked by its CA.

CERT_SIGNER_NOT_FOUND:

The certificate’s issuer is not known. This is the case when the issuer is not in the trusted certificates list.

GNUTLS_CERT_SIGNER_NOT_CA:

The certificate’s signer was not a CA. This may happen if this was a version 1 certificate, which is common with some CAs, or a version 3 certificate without the basic constrains extension.

GNUTLS_CERT_INSECURE_ALGORITHM:

The certificate was signed using an insecure algorithm such as MD2 or MD5. These algorithms have been broken and should not be trusted.

There is also to possibility to pass some input to the verification functions in the form of flags. For [gnutls x509 crt verify], page 177 the flags are passed straightforward, but [gnutls certificate verify peers2], page 107 depends on the flags set by calling [gnutls certificate set verify flags], page 103. All the available flags are part of the enumeration [gnutls certificate verify flags], page 20 and are explained in the table below.

GNUTLS_VERIFY_DISABLE_CA_SIGN:

If set a signer does not have to be a certificate authority. This flag should normaly be disabled, unless you know what this means.

GNUTLS_VERIFY_ALLOW_X509_V1_CA_CRT:

Allow only trusted CA certificates that have version 1. This is safer than GNUTLS VERIFY ALLOW ANY X509 V1 CA CRT, and should be used instead. That way only signers in your trusted list will be allowed to have certificates of version 1.

GNUTLS_VERIFY_ALLOW_ANY_X509_V1_CA_CRT:

Allow CA certificates that have version 1 (both root and intermediate). This is dangerous since those haven’t the basicConstraints extension. Must be used in combination with GNUTLS VERIFY ALLOW X509 V1 CA CRT.

GNUTLS_VERIFY_DO_NOT_ALLOW_SAME:

If a certificate is not signed by anyone trusted but exists in the trusted CA list do not treat it as trusted.

GNUTLS_VERIFY_ALLOW_SIGN_RSA_MD2:

Allow certificates to be signed using the old MD2 algorithm.

GNUTLS_VERIFY_ALLOW_SIGN_RSA_MD5:

Allow certificates to be signed using the broken MD5 algorithm.

Although the verification of a certificate path indicates that the certificate is signed by trusted authority, does not reveal anything about the peer’s identity. It is required to verify if the certificate’s owner is the one you expect. For more information consult [RFC2818] (See [Bibliography], page 244.) and section [ex:verify], page 37 for an example.

5.1.3 PKCS #10 certificate requests

A certificate request is a structure, which contain information about an applicant of a certificate service. It usually contains a private key, a distinguished name and secondary data such as a challenge password. GnuTLS supports the requests defined in PKCS #10 [RFC2986] (See [Bibliography], page 244.) . Other certificate request’s format such as PKIX’s [RFC4211] (See [Bibliography], page 244.) are not currently supported.

In GnuTLS the PKCS #10 structures are handled using the gnutls_x509_crq_t type. An example of a certificate request generation can be found at section [ex:crq], page 83.

5.1.4 PKCS #12 structures

A PKCS #12 structure [PKCS12] (See [Bibliography], page 244.) usually contains a user’s private keys and certificates. It is commonly used in browsers to export and import the user’s identities.

In GnuTLS the PKCS #12 structures are handled using the gnutls_pkcs12_t type. This is an abstract type that may hold several gnutls_pkcs12_bag_t types. The Bag types are the holders of the actual data, which may be certificates, private keys or encrypted data. An Bag of type encrypted should be decrypted in order for its data to be accessed.

An example of a PKCS #12 structure generation can be found at section [ex:pkcs12], page 84.

5.2 The OpenPGP trust model

The OpenPGP key authentication relies on a distributed trust model, called the “web of trust”. The “web of trust” uses a decentralized system of trusted introducers, which are the same as a CA. OpenPGP allows anyone to sign anyone’s else public key. When Alice signs Bob’s key, she is introducing Bob’s key to anyone who trusts Alice. If someone trusts Alice to introduce keys, then Alice is a trusted introducer in the mind of that observer.

For example: If David trusts Alice to be an introducer, and Alice signed Bob’s key, Dave also trusts Bob’s key to be the real one. There are some key points that are important in that model. In the example Alice has to sign Bob’s key, only if she is sure that the key belongs to Bob. Otherwise she may also make Dave falsely believe that this is Bob’s key. Dave has also the responsibility to know who to trust. This model is similar to real life relations. Just see how Charlie behaves in the previous example. Although he has signed Bob’s key - because he knows, somehow, that it belongs to Bob - he does not trust Bob to be an introducer. Charlie decided to trust only Kevin, for some reason. A reason could be that Bob is lazy enough, and signs other people’s keys without being sure that they belong to the actual owner.

5.2.1 OpenPGP keys

In GnuTLS the OpenPGP key structures [RFC2440] (See [Bibliography], page 244.) are handled using the gnutls_openpgp_key_t type and the corresponding private keys with the gnutls_openpgp_privkey_t type. All the prototypes for the key handling functions

can be found at ‘gnutls/openpgp.h’.

5.2.2 Verifying an OpenPGP key

The verification functions of OpenPGP keys, included in GnuTLS, are simple ones, and do not use the features of the “web of trust”. For that reason, if the verification needs are complex, the assistance of external tools like GnuPG and GPGME

(http://www.gnupg.org/related_software/gpgme/) is recommended.

There are two verification functions in GnuTLS, The [gnutls openpgp key verify ring], page 189 and the [gnutls openpgp key verify trustdb], page 190. The first one checks an OpenPGP key against a given set of public keys (keyring) and returns the key status. The key verification status is the same as in X.509 certificates, although the meaning and interpretation are different. For example an OpenPGP key may be valid, if the self signature is ok, even if no signers were found. The meaning of verification status is shown in the figure below. The latter function checks a GnuPG trust database for the given key. This function does not check the key signatures, only checks for disabled and revoked keys.

CERT_INVALID:

A signature on the key is invalid. That means that the key was modified by somebody, or corrupted during transport.

CERT_REVOKED:

The key has been revoked by its owner.

CERT_SIGNER_NOT_FOUND:

The key was not signed by a known signer.

GNUTLS_CERT_INSECURE_ALGORITHM:

The certificate was signed using an insecure algorithm such as MD2 or MD5. These algorithms have been broken and should not be trusted.

5.3 Digital signatures

In this section we will provide some information about digital signatures, how they work, and give the rationale for disabling some of the algorithms used. Digital signatures work by using somebody’s secret key to sign some arbitrary data. Then anybody else could use the public key of that person to verify the signature. Since the data may be arbitrary it is not suitable input to a cryptographic digital signature algorithm. For this reason and also for performance cryptographic hash algorithms are used to preprocess the input to the signature algorithm. This works as long as it is difficult enough to generate two different messages with the same hash algorithm output. In that case the same signature could be used as a proof for both messages. Nobody wants to sign an innocent message of donating 1 to Greenpeace and find out that he donated 1.000.000 to Bad Inc. For a hash algorithm to be called cryptographic the following three requirements must hold 1. Preimage resistance. That means the algorithm must be one way and given the output

of the hash function H(x), it is impossible to calculate x.

2. 2nd preimage resistance. That means that given a pair x, y with y = H(x) it is impossible to calculate an x′ such that y = H(x′).

3. Collision resistance. That means that it is impossible to calculate random x and x′ such H(x′) = H(x).

The last two requirements in the list are the most important in digital signatures. These protect against somebody who would like to generate two messages with the same hash output. When an algorithm is considered broken usually it means that the Collision resistance of the algorithm is less than brute force. Using the birthday paradox the brute force attack takes 2^{(hash size)/2} operations. Today colliding certificates using the MD5 hash algorithm have been generated as shown in [WEGER] (See [Bibliography], page 244.) .

There has been cryptographic results for the SHA-1 hash algorithms as well, although they are not yet critical. Before 2004, MD5 had a presumed collision strength of 2^{6}4, but it has been showed to have a collision strength well under 2^{5}0. As of November 2005, it is believed that SHA-1’s collision strength is around 2^{6}3. We consider this sufficiently hard so that we still support SHA-1. We anticipate that SHA-256/386/512 will be used in publiclydistributed certificates in the future. When 2^{6}3 can be considered too weak compared to the computer power available sometime in the future, SHA-1 will be disabled as well. The collision attacks on SHA-1 may also get better, given the new interest in tools for creating them.

5.3.1 Supported algorithms

The available digital signature algorithms in GnuTLS are listed below:

RSA RSA is public key cryptosystem designed by Ronald Rivest, Adi Shamir and Leonard Adleman. It can be used with any hash functions.

DSA DSA is the USA’s Digital Signature Standard. It uses only the SHA-1 hash algorithm.

The supported cryptographic hash algorithms are:

MD2 MD2 is a cryptographic hash algorithm designed by Ron Rivest. It is optimized for 8-bit processors. Outputs 128 bits of data. There are no known weaknesses of this algorithm but since this algorithm is rarely used and not really studied it should not be used today.

MD5 MD5 is a cryptographic hash algorithm designed by Ron Rivest. Outputs 128 bits of data. It is considered to be broken.

SHA-1 SHA is a cryptographic hash algorithm designed by NSA. Outputs 160 bits of data. It is also considered to be broken, though no practical attacks have been found.

RMD160 RIPEMD is a cryptographic hash algorithm developed in the framework of the EU project RIPE. Outputs 160 bits of data.

5.3.2 Trading security for interoperability

If you connect to a server and use GnuTLS’ functions to verify the certificate chain, and get a [GNUTLS CERT INSECURE ALGORITHM], page 20 validation error (see Section 5.1.2 [Verifying X.509 certificate paths], page 19), it means that somewhere in the certificate chain there is a certificate signed using RSA-MD2 or RSA-MD5. These two digital signature algorithms are considered broken, so GnuTLS fail when attempting to verify the certificate. In some situations, it may be useful to be able to verify the certificate chain anyway, assuming an attacker did not utilize the fact that these signatures algorithms are broken. This section will give help on how to achieve that. First, it is important to know that you do not have to enable any of the flags discussed here to be able to use trusted root CA certificates signed using RSA-MD2 or RSA-MD5. The only attack today is that it is possible to generate certificates with colliding signatures (collision resistance); you cannot generate a certificate that has the same signature as an already existing signature (2nd preimage resistance).

If you are using [gnutls certificate verify peers2], page 107 to verify the certificate chain, you can call [gnutls certificate set verify flags], page 103 with the GNUTLS_VERIFY_ALLOW_

SIGN_RSA_MD2 or GNUTLS_VERIFY_ALLOW_SIGN_RSA_MD5 flag, as in:
       gnutls_certificate_set_verify_flags (x509cred,
                                            GNUTLS_VERIFY_ALLOW_SIGN_RSA_MD5);

This will tell the verifier algorithm to enable RSA-MD5 when verifying the certificates. If you are using [gnutls x509 crt verify], page 177 or [gnutls x509 crt list verify], page 171,

you can pass the GNUTLS_VERIFY_ALLOW_SIGN_RSA_MD5 parameter directly in the flags

parameter. If you are using these flags, it may also be a good idea to warn the user when verification failure occur for this reason. The simplest is to not use the flags by default, and only fall back to using them after warning the user. If you wish to inspect the certificate chain yourself, you can use [gnutls certificate get peers], page 101 to extract the raw server’s certificate chain, then use [gnutls x509 crt import], page 170 to parse each of the certificates, and then use [gnutls x509 crt get signature algorithm], page 169 to find out the signing algorithm used for each certificate. If any of the intermediary certificates are using GNUTLS_SIGN_RSA_MD2 or GNUTLS_SIGN_RSA_MD5, you could present a warning.

6 How to use TLS in application protocols

This chapter is intended to provide some hints on how to use the TLS over simple custom made application protocols. The discussion below mainly refers to the TCP/IP transport layer but may be extended to other ones too.

6.1 Separate ports

Traditionally SSL was used in application protocols by assigning a new port number for the secure services. That way two separate ports were assigned, one for the non secure sessions, and one for the secured ones. This has the benefit that if a user requests a secure session then the client will try to connect to the secure port and fail otherwise. The only possible attack with this method is a denial of service one. The most famous example of this method is the famous “HTTP over TLS” or HTTPS protocol [RFC2818] (See [Bibliography], page 244.) . Despite its wide use, this method is not as good as it seems. This approach starts the TLS Handshake procedure just after the client connects on the —so called— secure port. That way the TLS protocol does not know anything about the client, and popular methods like the host advertising in HTTP do not work^{1}. There is no way for the client to say “I connected to YYY server” before the Handshake starts, so the server cannot possibly know which certificate to use. Other than that it requires two separate ports to run a single service, which is unnecessary complication. Due to the fact that there is a limitation on the available privileged ports, this approach was soon obsoleted.

6.2 Upward negotiation

Other application protocols^{2} use a different approach to enable the secure layer. They use something called the “TLS upgrade” method. This method is quite tricky but it is more flexible. The idea is to extend the application protocol to have a “STARTTLS” request, whose purpose it to start the TLS protocols just after the client requests it. This is a really neat idea and does not require an extra port. This method is used by almost all modern protocols and there is even the [RFC2817] (See [Bibliography], page 244.) paper which proposes extensions to HTTP to support it. The tricky part, in this method, is that the “STARTTLS” request is sent in the clear, thus is vulnerable to modifications. A typical attack is to modify the messages in a way that the client is fooled and thinks that the server does not have the “STARTTLS” capability. See a typical conversation of a hypothetical protocol:

(client connects to the server) CLIENT: HELLO I’M MR. XXX SERVER: NICE TO MEET YOU XXX CLIENT: PLEASE START TLS SERVER: OK *** TLS STARTS CLIENT: HERE ARE SOME CONFIDENTIAL DATA

1 See also the Server Name Indication extension on [serverind], page 11.

2 See LDAP, IMAP etc.

And see an example of a conversation where someone is acting in between:

(client connects to the server) CLIENT: HELLO I’M MR. XXX SERVER: NICE TO MEET YOU XXX CLIENT: PLEASE START TLS (here someone inserts this message) SERVER: SORRY I DON’T HAVE THIS CAPABILITY CLIENT: HERE ARE SOME CONFIDENTIAL DATA

As you can see above the client was fooled, and was dummy enough to send the confidential data in the clear. How to avoid the above attack? As you may have already thought this one is easy to avoid. The client has to ask the user before it connects whether the user requests TLS or not. If the user answered that he certainly wants the secure layer the last conversation should be:

(client connects to the server) CLIENT: HELLO I’M MR. XXX SERVER: NICE TO MEET YOU XXX CLIENT: PLEASE START TLS (here someone inserts this message) SERVER: SORRY I DON’T HAVE THIS CAPABILITY CLIENT: BYE (the client notifies the user that the secure connection was not possible)

This method, if implemented properly, is far better than the traditional method, and the security properties remain the same, since only denial of service is possible. The benefit is that the server may request additional data before the TLS Handshake protocol starts, in order to send the correct certificate, use the correct password file^{3}, or anything else!

3 in SRP authentication

7 How to use GnuTLS in applications

7.1 Preparation

To use GnuTLS, you have to perform some changes to your sources and your build system. The necessary changes are explained in the following subsections.

7.1.1 Headers

All the data types and functions of the GnuTLS library are defined in the header file ‘gnutls/gnutls.h’. This must be included in all programs that make use of the GnuTLS library. The extra functionality of the GnuTLS-extra library is available by including the header file ‘gnutls/extra.h’ in your programs.

7.1.2 Version check

It is often desirable to check that the version of ‘gnutls’ used is indeed one which fits all requirements. Even with binary compatibility new features may have been introduced but due to problem with the dynamic linker an old version is actually used. So you may want to check that the version is okay right after program startup. See the function [gnutls check version], page 108.

7.1.3 Building the source

If you want to compile a source file including the ‘gnutls/gnutls.h’ header file, you must make sure that the compiler can find it in the directory hierarchy. This is accomplished by adding the path to the directory in which the header file is located to the compilers include file search path (via the -I option). However, the path to the include file is determined at the time the source is configured. To solve this problem, GnuTLS ships with two small helper programs libgnutls-config and libgnutls-extra-config that knows about the path to the include file and other configuration options. The options that need to be added to the compiler invocation at compile time are output by the --cflags option to libgnutls-config. The following example shows how it can be used at the command line:

     gcc -c foo.c `libgnutls-config --cflags`

Adding the output of libgnutls-config --cflags to the compilers command line will ensure that the compiler can find the GnuTLS header file. A similar problem occurs when linking the program with the library. Again, the compiler has to find the library files. For this to work, the path to the library files has to be added to the library search path (via the -L option). For this, the option --libs to libgnutlsconfig can be used. For convenience, this option also outputs all other options that are required to link the program with the GnuTLS libararies. The example shows how to link ‘foo.o’ with the GnuTLS libraries to a program foo.

     gcc -o foo foo.o `libgnutls-config --libs`

Of course you can also combine both examples to a single command by specifying both options to ‘libgnutls-config’:

     gcc -o foo foo.c `libgnutls-config --cflags --libs`

7.2 Multi-threaded applications

Although the GnuTLS library is thread safe by design, some parts of the crypto backend, such as the random generator, are not. Since libgcrypt 1.1.92 there was an automatic detection of the thread library used by the application, so most applications wouldn’t need to do any changes to ensure thread-safety. Due to the unportability of the automatic thread detection, this was removed from later releases of libgcrypt, so applications have now to register callback functions to ensure proper locking in sensitive parts of libgcrypt. There are helper macros to help you properly initialize the libraries. Examples are shown below.

• POSIX threads

#include <gnutls.h>
#include <gcrypt.h>
#include <errno.h>
#include <pthread.h>
GCRY_THREAD_OPTION_PTHREAD_IMPL;
int main()
{
   /* The order matters.
    */
   gcry_control (GCRYCTL_SET_THREAD_CBS, &gcry_threads_pthread);
   gnutls_global_init();
}

• GNU PTH threads

#include <gnutls.h>
#include <gcrypt.h>
#include <errno.h>
#include <pth.h>
GCRY_THREAD_OPTION_PTH_IMPL;
int main()
{
   gcry_control (GCRYCTL_SET_THREAD_CBS, &gcry_threads_pth);
   gnutls_global_init();
}

• Other thread packages

/* The gcry_thread_cbs structure must have been
 * initialized.
 */
static struct gcry_thread_cbs gcry_threads_other = { ... };
int main()
{
   gcry_control (GCRYCTL_SET_THREAD_CBS, &gcry_threads_other);
}

7.3 Client examples

This section contains examples of TLS and SSL clients, using GnuTLS. Note that these examples contain little or no error checking.

7.3.1 Simple client example with anonymous authentication

The simplest client using TLS is the one that doesn’t do any authentication. This means no external certificates or passwords are needed to set up the connection. As could be expected, the connection is vulnerable to man-in-the-middle (active or redirection) attacks. However, the data is integrity and privacy protected.

#if HAVE_CONFIG_H
# include <config.h>
#endif
#include <stdio.h>
#include <stdlib.h>
#include <string.h>
#include <sys/types.h>
#include <sys/socket.h>
#include <arpa/inet.h>
#include <unistd.h>
#include <gnutls/gnutls.h>
/* A very basic TLS client, with anonymous authentication.
 */
#define MAX_BUF 1024
#define SA struct sockaddr
#define MSG "GET / HTTP/1.0\r\n\r\n"
extern int tcp_connect (void);
extern void tcp_close (int sd);
int
main (void)
{
  int ret, sd, ii;
  gnutls_session_t session;
  char buffer[MAX_BUF + 1];
  gnutls_anon_client_credentials_t anoncred;
  /* Need to enable anonymous KX specifically. */
  const int kx_prio[] = { GNUTLS_KX_ANON_DH, 0 };
  gnutls_global_init ();
  gnutls_anon_allocate_client_credentials (&anoncred);

/* Initialize TLS session
 */
gnutls_init (&session, GNUTLS_CLIENT);
/* Use default priorities */
gnutls_set_default_priority (session);
gnutls_kx_set_priority (session, kx_prio);
/* put the anonymous credentials to the current session
 */
gnutls_credentials_set (session, GNUTLS_CRD_ANON, anoncred);
/* connect to the peer
 */
sd = tcp_connect ();
gnutls_transport_set_ptr (session, (gnutls_transport_ptr_t) sd);
/* Perform the TLS handshake
 */
ret = gnutls_handshake (session);
if (ret < 0)
  {
    fprintf (stderr, "*** Handshake failed\n");
    gnutls_perror (ret);
    goto end;
  }
else
  {
    printf ("- Handshake was completed\n");
  }
gnutls_record_send (session, MSG, strlen (MSG));
ret = gnutls_record_recv (session, buffer, MAX_BUF);
if (ret == 0)
  {
    printf ("- Peer has closed the TLS connection\n");
    goto end;
  }
else if (ret < 0)
  {
    fprintf (stderr, "*** Error: %s\n", gnutls_strerror (ret));
    goto end;
  }

  printf ("- Received %d bytes: ", ret);
  for (ii = 0; ii < ret; ii++)
    {
      fputc (buffer[ii], stdout);
    }
  fputs ("\n", stdout);
  gnutls_bye (session, GNUTLS_SHUT_RDWR);
end:
  tcp_close (sd);
  gnutls_deinit (session);
  gnutls_anon_free_client_credentials (anoncred);
  gnutls_global_deinit ();
  return 0;
}

7.3.2 Simple client example with X.509 certificate support

Let’s assume now that we want to create a TCP client which communicates with servers that use X.509 or OpenPGP certificate authentication. The following client is a very simple TLS client, it does not support session resuming, not even certificate verification. The TCP functions defined in this example are used in most of the other examples below, without redefining them.

#if HAVE_CONFIG_H
# include <config.h>
#endif
#include <stdio.h>
#include <stdlib.h>
#include <string.h>
#include <sys/types.h>
#include <sys/socket.h>
#include <arpa/inet.h>
#include <unistd.h>
#include <gnutls/gnutls.h>
/* A very basic TLS client, with X.509 authentication.
 */
#define MAX_BUF 1024
#define CAFILE "ca.pem"

#define MSG "GET / HTTP/1.0\r\n\r\n"
extern int tcp_connect (void);
extern void tcp_close (int sd);
int
main (void)
{
int ret, sd, ii;
gnutls_session_t session;
char buffer[MAX_BUF + 1];
gnutls_certificate_credentials_t xcred;
/* Allow connections to servers that have OpenPGP keys as well.
 */
const int cert_type_priority[3] = { GNUTLS_CRT_X509,
  GNUTLS_CRT_OPENPGP, 0
};
gnutls_global_init ();
/* X509 stuff */
gnutls_certificate_allocate_credentials (&xcred);
/* sets the trusted cas file
 */
gnutls_certificate_set_x509_trust_file (xcred, CAFILE, GNUTLS_X509_FMT_PEM);
/* Initialize TLS session
 */
gnutls_init (&session, GNUTLS_CLIENT);
/* Use default priorities */
gnutls_set_default_priority (session);
gnutls_certificate_type_set_priority (session, cert_type_priority);
/* put the x509 credentials to the current session
 */
gnutls_credentials_set (session, GNUTLS_CRD_CERTIFICATE, xcred);
/* connect to the peer
 */
sd = tcp_connect ();
gnutls_transport_set_ptr (session, (gnutls_transport_ptr_t) sd);
/* Perform the TLS handshake
 */

ret = gnutls_handshake (session);
if (ret < 0)
  {
    fprintf (stderr, "*** Handshake failed\n");
    gnutls_perror (ret);
    goto end;
  }
else
  {
    printf ("- Handshake was completed\n");
  }
gnutls_record_send (session, MSG, strlen (MSG));
ret = gnutls_record_recv (session, buffer, MAX_BUF);
if (ret == 0)
  {
    printf ("- Peer has closed the TLS connection\n");
    goto end;
  }
else if (ret < 0)
  {
    fprintf (stderr, "*** Error: %s\n", gnutls_strerror (ret));
    goto end;
  }
printf ("- Received %d bytes: ", ret);
for (ii = 0; ii < ret; ii++)
  {
    fputc (buffer[ii], stdout);
  }
fputs ("\n", stdout);
gnutls_bye (session, GNUTLS_SHUT_RDWR);
end:
tcp_close (sd);
gnutls_deinit (session);
gnutls_certificate_free_credentials (xcred);
gnutls_global_deinit ();

}

7.3.3 Obtaining session information

Most of the times it is desirable to know the security properties of the current established session. This includes the underlying ciphers and the protocols involved. That is the purpose of the following function. Note that this function will print meaningful values only if called after a successful [gnutls handshake], page 119.

#if HAVE_CONFIG_H
# include <config.h>
#endif
#include <stdio.h>
#include <stdlib.h>
#include <gnutls/gnutls.h>
#include <gnutls/x509.h>
extern void print_x509_certificate_info (gnutls_session_t);
/* This function will print some details of the
 * given session.
 */
int
print_info (gnutls_session_t session)
{
  const char *tmp;
  gnutls_credentials_type_t cred;
  gnutls_kx_algorithm_t kx;
  /* print the key exchange's algorithm name
   */
  kx = gnutls_kx_get (session);
  tmp = gnutls_kx_get_name (kx);
  printf ("- Key Exchange: %s\n", tmp);
  /* Check the authentication type used and switch
   * to the appropriate.
   */
  cred = gnutls_auth_get_type (session);
  switch (cred)
    {
    case GNUTLS_CRD_SRP:
      printf ("- SRP session with username %s\n",
              gnutls_srp_server_get_username (session));
      break;

    case GNUTLS_CRD_ANON:       /* anonymous authentication */

    printf ("- Anonymous DH using prime of %d bits\n",
            gnutls_dh_get_prime_bits (session));
    break;

  case GNUTLS_CRD_CERTIFICATE:        /* certificate authentication */

    /* Check if we have been using ephemeral Diffie Hellman.
     */
    if (kx == GNUTLS_KX_DHE_RSA || kx == GNUTLS_KX_DHE_DSS)
      {
        printf ("\n- Ephemeral DH using prime of %d bits\n",
                gnutls_dh_get_prime_bits (session));
      }
    /* if the certificate list is available, then
     * print some information about it.
     */
    print_x509_certificate_info (session);

  }                           /* switch */

/* print the protocol's name (ie TLS 1.0)
 */
tmp = gnutls_protocol_get_name (gnutls_protocol_get_version (session));
printf ("- Protocol: %s\n", tmp);
/* print the certificate type of the peer.
 * ie X.509
 */
tmp =
  gnutls_certificate_type_get_name (gnutls_certificate_type_get (session));
printf ("- Certificate Type: %s\n", tmp);
/* print the compression algorithm (if any)
 */
tmp = gnutls_compression_get_name (gnutls_compression_get (session));
printf ("- Compression: %s\n", tmp);
/* print the name of the cipher used.
 * ie 3DES.
 */
tmp = gnutls_cipher_get_name (gnutls_cipher_get (session));
printf ("- Cipher: %s\n", tmp);
/* Print the MAC algorithms name.

   * ie SHA1
   */
  tmp = gnutls_mac_get_name (gnutls_mac_get (session));
  printf ("- MAC: %s\n", tmp);
  return 0;
}

7.3.4 Verifying peer’s certificate

A TLS session is not secure just after the handshake procedure has finished. It must be considered secure, only after the peer’s certificate and identity have been verified. That is, you have to verify the signature in peer’s certificate, the hostname in the certificate, and expiration dates. Just after this step you should treat the connection as being a secure one.

#if HAVE_CONFIG_H
# include <config.h>
#endif
#include <stdio.h>
#include <gnutls/gnutls.h>
#include <gnutls/x509.h>
/* This function will try to verify the peer's certificate, and
 * also check if the hostname matches, and the activation, expiration dates.
 */
void
verify_certificate (gnutls_session_t session, const char *hostname)
{
  unsigned int status;
  const gnutls_datum_t *cert_list;
  int cert_list_size, ret;
  gnutls_x509_crt_t cert;

  /* This verification function uses the trusted CAs in the credentials
   * structure. So you must have installed one or more CA certificates.
   */
  ret = gnutls_certificate_verify_peers2 (session, &status);
  if (ret < 0)
    {
      printf ("Error\n");
      return;
    }
  if (status & GNUTLS_CERT_INVALID)
    printf ("The certificate is not trusted.\n");

if (status & GNUTLS_CERT_SIGNER_NOT_FOUND)
  printf ("The certificate hasn't got a known issuer.\n");
if (status & GNUTLS_CERT_REVOKED)
  printf ("The certificate has been revoked.\n");

/* Up to here the process is the same for X.509 certificates and
 * OpenPGP keys. From now on X.509 certificates are assumed. This can
 * be easily extended to work with openpgp keys as well.
 */
if (gnutls_certificate_type_get (session) != GNUTLS_CRT_X509)
  return;
if (gnutls_x509_crt_init (&cert) < 0)
  {
    printf ("error in initialization\n");
    return;
  }
cert_list = gnutls_certificate_get_peers (session, &cert_list_size);
if (cert_list == NULL)
  {
    printf ("No certificate was found!\n");
    return;
  }
/* This is not a real world example, since we only check the first
 * certificate in the given chain.
 */
if (gnutls_x509_crt_import (cert, &cert_list[0], GNUTLS_X509_FMT_DER) < 0)
  {
    printf ("error parsing certificate\n");
    return;
  }
/* Beware here we do not check for errors.
 */
if (gnutls_x509_crt_get_expiration_time (cert) < time (0))
  {
    printf ("The certificate has expired\n");
    return;
  }
if (gnutls_x509_crt_get_activation_time (cert) > time (0))
  {

      printf ("The certificate is not yet activated\n");
      return;
    }
  if (!gnutls_x509_crt_check_hostname (cert, hostname))
    {
      printf ("The certificate's owner does not match hostname '%s'\n",
              hostname);
      return;
    }
  gnutls_x509_crt_deinit (cert);
  return;
}

An other example is listed below which provides a more detailed verification output.

#if HAVE_CONFIG_H
# include <config.h>
#endif
#include <stdio.h>
#include <stdlib.h>
#include <gnutls/gnutls.h>
#include <gnutls/x509.h>
/* All the available CRLs
 */
gnutls_x509_crl_t *crl_list;
int crl_list_size;
/* All the available trusted CAs
 */
gnutls_x509_crt_t *ca_list;
int ca_list_size;
static void verify_cert2 (gnutls_x509_crt_t crt,
                          gnutls_x509_crt_t issuer,
                          gnutls_x509_crl_t * crl_list, int crl_list_size);
static void verify_last_cert (gnutls_x509_crt_t crt,
                              gnutls_x509_crt_t * ca_list, int ca_list_size,
                              gnutls_x509_crl_t * crl_list,
                              int crl_list_size);

/* This function will try to verify the peer's certificate chain, and
 * also check if the hostname matches, and the activation, expiration dates.

*/
void
verify_certificate_chain (gnutls_session_t session,
                        const char *hostname,
                        const gnutls_datum_t * cert_chain,
                        int cert_chain_length)
{
int i;
gnutls_x509_crt_t *cert;
cert = malloc (sizeof (*cert) * cert_chain_length);
/* Import all the certificates in the chain to
 * native certificate format.
 */
for (i = 0; i < cert_chain_length; i++)
  {
    gnutls_x509_crt_init (&cert[i]);
    gnutls_x509_crt_import (cert[i], &cert_chain[i], GNUTLS_X509_FMT_DER);
  }
/* If the last certificate in the chain is self signed ignore it.
 * That is because we want to check against our trusted certificate
 * list.
 */
if (gnutls_x509_crt_check_issuer (cert[cert_chain_length - 1],
                                  cert[cert_chain_length - 1]) > 0
    && cert_chain_length > 0)
  {
    cert_chain_length--;
  }
/* Now verify the certificates against their issuers
 * in the chain.
 */
for (i = 1; i < cert_chain_length; i++)
  {
    verify_cert2 (cert[i - 1], cert[i], crl_list, crl_list_size);
  }
/* Here we must verify the last certificate in the chain against
 * our trusted CA list.
 */
verify_last_cert (cert[cert_chain_length - 1],
                  ca_list, ca_list_size, crl_list, crl_list_size);
/* Check if the name in the first certificate matches our destination!

 */
if (!gnutls_x509_crt_check_hostname (cert[0], hostname))
  {
    printf ("The certificate's owner does not match hostname '%s'\n",
            hostname);
  }
for (i = 0; i < cert_chain_length; i++)
  gnutls_x509_crt_deinit (cert[i]);
return;
}

/* Verifies a certificate against an other certificate
* which is supposed to be it's issuer. Also checks the
* crl_list if the certificate is revoked.
*/
static void
verify_cert2 (gnutls_x509_crt_t crt, gnutls_x509_crt_t issuer,
            gnutls_x509_crl_t * crl_list, int crl_list_size)
{
unsigned int output;
int ret;
time_t now = time (0);
size_t name_size;
char name[64];
/* Print information about the certificates to
 * be checked.
 */
name_size = sizeof (name);
gnutls_x509_crt_get_dn (crt, name, &name_size);
fprintf (stderr, "\nCertificate: %s\n", name);
name_size = sizeof (name);
gnutls_x509_crt_get_issuer_dn (crt, name, &name_size);
fprintf (stderr, "Issued by: %s\n", name);
/* Get the DN of the issuer cert.
 */
name_size = sizeof (name);
gnutls_x509_crt_get_dn (issuer, name, &name_size);
fprintf (stderr, "Checking against: %s\n", name);

/* Do the actual verification.
 */
gnutls_x509_crt_verify (crt, &issuer, 1, 0, &output);
if (output & GNUTLS_CERT_INVALID)
  {
    fprintf (stderr, "Not trusted");
    if (output & GNUTLS_CERT_SIGNER_NOT_FOUND)
      fprintf (stderr, ": no issuer was found");
    if (output & GNUTLS_CERT_SIGNER_NOT_CA)
      fprintf (stderr, ": issuer is not a CA");
    fprintf (stderr, "\n");
  }
else
  fprintf (stderr, "Trusted\n");

/* Now check the expiration dates.
 */
if (gnutls_x509_crt_get_activation_time (crt) > now)
  fprintf (stderr, "Not yet activated\n");
if (gnutls_x509_crt_get_expiration_time (crt) < now)
  fprintf (stderr, "Expired\n");
/* Check if the certificate is revoked.
 */
ret = gnutls_x509_crt_check_revocation (crt, crl_list, crl_list_size);
if (ret == 1)

  {                           /* revoked */

    fprintf (stderr, "Revoked\n");
  }
}

/* Verifies a certificate against our trusted CA list.
* Also checks the crl_list if the certificate is revoked.
*/
static void
verify_last_cert (gnutls_x509_crt_t crt,
                gnutls_x509_crt_t * ca_list, int ca_list_size,
                gnutls_x509_crl_t * crl_list, int crl_list_size)
{
unsigned int output;

int ret;
time_t now = time (0);
size_t name_size;
char name[64];
/* Print information about the certificates to
 * be checked.
 */
name_size = sizeof (name);
gnutls_x509_crt_get_dn (crt, name, &name_size);
fprintf (stderr, "\nCertificate: %s\n", name);
name_size = sizeof (name);
gnutls_x509_crt_get_issuer_dn (crt, name, &name_size);
fprintf (stderr, "Issued by: %s\n", name);
/* Do the actual verification.
 */
gnutls_x509_crt_verify (crt, ca_list, ca_list_size,
                        GNUTLS_VERIFY_ALLOW_X509_V1_CA_CRT, &output);
if (output & GNUTLS_CERT_INVALID)
  {
    fprintf (stderr, "Not trusted");
    if (output & GNUTLS_CERT_SIGNER_NOT_CA)
      fprintf (stderr, ": Issuer is not a CA\n");
    else
      fprintf (stderr, "\n");
  }
else
  fprintf (stderr, "Trusted\n");

/* Now check the expiration dates.
 */
if (gnutls_x509_crt_get_activation_time (crt) > now)
  fprintf (stderr, "Not yet activated\n");
if (gnutls_x509_crt_get_expiration_time (crt) < now)
  fprintf (stderr, "Expired\n");
/* Check if the certificate is revoked.
 */
ret = gnutls_x509_crt_check_revocation (crt, crl_list, crl_list_size);

  if (ret == 1)

    {                           /* revoked */

      fprintf (stderr, "Revoked\n");
    }
}

7.3.5 Using a callback to select the certificate to use

There are cases where a client holds several certificate and key pairs, and may not want to load all of them in the credentials structure. The following example demonstrates the use of the certificate selection callback.

#if HAVE_CONFIG_H
# include <config.h>
#endif
#include <stdio.h>
#include <stdlib.h>
#include <string.h>
#include <sys/types.h>
#include <sys/socket.h>
#include <arpa/inet.h>
#include <unistd.h>
#include <sys/stat.h>
#include <gnutls/gnutls.h>
#include <gnutls/x509.h>
#include <sys/types.h>
#include <sys/stat.h>
#include <fcntl.h>
/* A TLS client that loads the certificate and key.
 */
#define MAX_BUF 1024
#define MSG "GET / HTTP/1.0\r\n\r\n"
#define CERT_FILE "cert.pem"
#define KEY_FILE "key.pem"
#define CAFILE "ca.pem"
extern int tcp_connect (void);
extern void tcp_close (int sd);
static int cert_callback (gnutls_session_t session,
                          const gnutls_datum_t * req_ca_rdn, int nreqs,
                          const gnutls_pk_algorithm_t * sign_algos,
                          int sign_algos_length, gnutls_retr_st * st);

gnutls_x509_crt_t crt;
gnutls_x509_privkey_t key;
/* Helper functions to load a certificate and key
 * files into memory.
 */
static gnutls_datum
load_file (const char *file)
{
  FILE *f;
  gnutls_datum loaded_file = { NULL, 0 };
  long filelen;
  void *ptr;
  if (!(f = fopen(file, "r"))
      || fseek(f, 0, SEEK_END) != 0
      || (filelen = ftell(f)) < 0
      || fseek(f, 0, SEEK_SET) != 0
      || !(ptr = malloc((size_t)filelen))
      || fread(ptr, 1, (size_t)filelen, f) < (size_t)filelen)
    {
      return loaded_file;
    }
  loaded_file.data = ptr;
  loaded_file.size = (unsigned int)filelen;
  return loaded_file;
}
static void unload_file(gnutls_datum data)
{
  free(data.data);
}
/* Load the certificate and the private key.
 */
static void
load_keys (void)
{
  int ret;
  gnutls_datum_t data;
  data = load_file (CERT_FILE);
  if (data.data == NULL)
    {
      fprintf (stderr, "*** Error loading cert file.\n");
      exit (1);

  }
gnutls_x509_crt_init (&crt);
ret = gnutls_x509_crt_import (crt, &data, GNUTLS_X509_FMT_PEM);
if (ret < 0)
  {
    fprintf (stderr, "*** Error loading key file: %s\n",
             gnutls_strerror (ret));
    exit (1);
  }
unload_file (data);
data = load_file (KEY_FILE);
if (data.data == NULL)
  {
    fprintf (stderr, "*** Error loading key file.\n");
    exit (1);
  }
gnutls_x509_privkey_init (&key);
ret = gnutls_x509_privkey_import (key, &data, GNUTLS_X509_FMT_PEM);
if (ret < 0)
  {
    fprintf (stderr, "*** Error loading key file: %s\n",
             gnutls_strerror (ret));
    exit (1);
  }
unload_file (data);
}
int
main (void)
{
int ret, sd, ii;
gnutls_session_t session;
char buffer[MAX_BUF + 1];
gnutls_certificate_credentials_t xcred;
/* Allow connections to servers that have OpenPGP keys as well.
 */
gnutls_global_init ();
load_keys ();

/* X509 stuff */
gnutls_certificate_allocate_credentials (&xcred);
/* sets the trusted cas file
 */
gnutls_certificate_set_x509_trust_file (xcred, CAFILE, GNUTLS_X509_FMT_PEM);
gnutls_certificate_client_set_retrieve_function (xcred, cert_callback);
/* Initialize TLS session
 */
gnutls_init (&session, GNUTLS_CLIENT);
/* Use default priorities */
gnutls_set_default_priority (session);
/* put the x509 credentials to the current session
 */
gnutls_credentials_set (session, GNUTLS_CRD_CERTIFICATE, xcred);
/* connect to the peer
 */
sd = tcp_connect ();
gnutls_transport_set_ptr (session, (gnutls_transport_ptr_t) sd);
/* Perform the TLS handshake
 */
ret = gnutls_handshake (session);
if (ret < 0)
  {
    fprintf (stderr, "*** Handshake failed\n");
    gnutls_perror (ret);
    goto end;
  }
else
  {
    printf ("- Handshake was completed\n");
  }
gnutls_record_send (session, MSG, strlen (MSG));
ret = gnutls_record_recv (session, buffer, MAX_BUF);
if (ret == 0)
  {

    printf ("- Peer has closed the TLS connection\n");
    goto end;
  }
else if (ret < 0)
  {
    fprintf (stderr, "*** Error: %s\n", gnutls_strerror (ret));
    goto end;
  }
printf ("- Received %d bytes: ", ret);
for (ii = 0; ii < ret; ii++)
  {
    fputc (buffer[ii], stdout);
  }
fputs ("\n", stdout);
gnutls_bye (session, GNUTLS_SHUT_RDWR);
end:
tcp_close (sd);
gnutls_deinit (session);
gnutls_certificate_free_credentials (xcred);
gnutls_global_deinit ();
return 0;
}

/* This callback should be associated with a session by calling
* gnutls_certificate_client_set_retrieve_function( session, cert_callback),
* before a handshake.
*/
static int
cert_callback (gnutls_session_t session,
             const gnutls_datum_t * req_ca_rdn, int nreqs,
             const gnutls_pk_algorithm_t * sign_algos,
             int sign_algos_length, gnutls_retr_st * st)
{
char issuer_dn[256];
int i, ret;
size_t len;

gnutls_certificate_type_t type;
/* Print the server's trusted CAs
 */
if (nreqs > 0)
  printf ("- Server's trusted authorities:\n");
else
  printf ("- Server did not send us any trusted authorities names.\n");
/* print the names (if any) */
for (i = 0; i < nreqs; i++)
  {
    len = sizeof (issuer_dn);
    ret = gnutls_x509_rdn_get (&req_ca_rdn[i], issuer_dn, &len);
    if (ret >= 0)
      {
        printf ("   [%d]: ", i);
        printf ("%s\n", issuer_dn);
      }
  }
/* Select a certificate and return it.
 * The certificate must be of any of the "sign algorithms"
 * supported by the server.
 */
type = gnutls_certificate_type_get (session);
if (type == GNUTLS_CRT_X509)
  {
    st->type = type;
    st->ncerts = 1;
    st->cert.x509 = &crt;
    st->key.x509 = key;
    st->deinit_all = 0;
  }
else
  {
    return -1;
  }
return 0;
}

7.3.6 Client with Resume capability example

This is a modification of the simple client example. Here we demonstrate the use of session resumption. The client tries to connect once using TLS, close the connection and then try to establish a new connection using the previously negotiated data.

#if HAVE_CONFIG_H
# include <config.h>
#endif
#include <string.h>
#include <stdio.h>
#include <stdlib.h>
#include <gnutls/gnutls.h>
/* Those functions are defined in other examples.
 */
extern void check_alert (gnutls_session_t session, int ret);
extern int tcp_connect (void);
extern void tcp_close (int sd);
#define MAX_BUF 1024
#define CRLFILE "crl.pem"
#define CAFILE "ca.pem"
#define MSG "GET / HTTP/1.0\r\n\r\n"
int
main (void)
{
  int ret;
  int sd, ii;
  gnutls_session_t session;
  char buffer[MAX_BUF + 1];
  gnutls_certificate_credentials_t xcred;
  /* variables used in session resuming
   */
  int t;
  char *session_data;
  size_t session_data_size;
  gnutls_global_init ();
  /* X509 stuff */
  gnutls_certificate_allocate_credentials (&xcred);
  gnutls_certificate_set_x509_trust_file (xcred, CAFILE, GNUTLS_X509_FMT_PEM);

for (t = 0; t < 2; t++)

{                           /* connect 2 times to the server */

sd = tcp_connect ();
gnutls_init (&session, GNUTLS_CLIENT);
gnutls_set_default_priority (session);
gnutls_credentials_set (session, GNUTLS_CRD_CERTIFICATE, xcred);
if (t > 0)
  {
    /* if this is not the first time we connect */
    gnutls_session_set_data (session, session_data, session_data_size);
    free (session_data);
  }
gnutls_transport_set_ptr (session, (gnutls_transport_ptr_t) sd);
/* Perform the TLS handshake
 */
ret = gnutls_handshake (session);
if (ret < 0)
  {
    fprintf (stderr, "*** Handshake failed\n");
    gnutls_perror (ret);
    goto end;
  }
else
  {
    printf ("- Handshake was completed\n");
  }
if (t == 0)

  {                       /* the first time we connect */

    /* get the session data size */
    gnutls_session_get_data (session, NULL, &session_data_size);
    session_data = malloc (session_data_size);
    /* put session data to the session variable */
    gnutls_session_get_data (session, session_data, &session_data_size);
  }
else

  {                       /* the second time we connect */

    /* check if we actually resumed the previous session */
    if (gnutls_session_is_resumed (session) != 0)
      {
        printf ("- Previous session was resumed\n");
      }
    else
      {
        fprintf (stderr, "*** Previous session was NOT resumed\n");
      }
  }
/* This function was defined in a previous example
 */
/* print_info(session); */
gnutls_record_send (session, MSG, strlen (MSG));
ret = gnutls_record_recv (session, buffer, MAX_BUF);
if (ret == 0)
  {
    printf ("- Peer has closed the TLS connection\n");
    goto end;
  }
else if (ret < 0)
  {
    fprintf (stderr, "*** Error: %s\n", gnutls_strerror (ret));
    goto end;
  }
printf ("- Received %d bytes: ", ret);
for (ii = 0; ii < ret; ii++)
  {
    fputc (buffer[ii], stdout);
  }
fputs ("\n", stdout);
gnutls_bye (session, GNUTLS_SHUT_RDWR);
end:
tcp_close (sd);
gnutls_deinit (session);

}                           /* for() */

  gnutls_certificate_free_credentials (xcred);
  gnutls_global_deinit ();
  return 0;
}

7.3.7 Simple client example with SRP authentication

The following client is a very simple SRP TLS client which connects to a server and authenticates using a username and a password. The server may authenticate itself using a certificate, and in that case it has to be verified.

#if HAVE_CONFIG_H
# include <config.h>
#endif
#include <stdio.h>
#include <stdlib.h>
#include <string.h>
#include <gnutls/gnutls.h>
#include <gnutls/extra.h>
/* Those functions are defined in other examples.
 */
extern void check_alert (gnutls_session_t session, int ret);
extern int tcp_connect (void);
extern void tcp_close (int sd);
#define MAX_BUF 1024
#define USERNAME "user"
#define PASSWORD "pass"
#define CAFILE "ca.pem"
#define SA struct sockaddr
#define MSG "GET / HTTP/1.0\r\n\r\n"
const int kx_priority[] = { GNUTLS_KX_SRP, GNUTLS_KX_SRP_DSS,
  GNUTLS_KX_SRP_RSA, 0
};
int
main (void)
{
  int ret;
  int sd, ii;
  gnutls_session_t session;
  char buffer[MAX_BUF + 1];
  gnutls_srp_client_credentials_t srp_cred;

gnutls_certificate_credentials_t cert_cred;
gnutls_global_init ();
/* now enable the gnutls-extra library which contains the
 * SRP stuff.
 */
gnutls_global_init_extra ();
gnutls_srp_allocate_client_credentials (&srp_cred);
gnutls_certificate_allocate_credentials (&cert_cred);
gnutls_certificate_set_x509_trust_file (cert_cred, CAFILE,
                                        GNUTLS_X509_FMT_PEM);
gnutls_srp_set_client_credentials (srp_cred, USERNAME, PASSWORD);
/* connects to server
 */
sd = tcp_connect ();
/* Initialize TLS session
 */
gnutls_init (&session, GNUTLS_CLIENT);

/* Set the priorities.
 */
gnutls_set_default_priority (session);
gnutls_kx_set_priority (session, kx_priority);

/* put the SRP credentials to the current session
 */
gnutls_credentials_set (session, GNUTLS_CRD_SRP, srp_cred);
gnutls_credentials_set (session, GNUTLS_CRD_CERTIFICATE, cert_cred);
gnutls_transport_set_ptr (session, (gnutls_transport_ptr_t) sd);
/* Perform the TLS handshake
 */
ret = gnutls_handshake (session);
if (ret < 0)
  {
    fprintf (stderr, "*** Handshake failed\n");
    gnutls_perror (ret);
    goto end;

  }
else
  {
    printf ("- Handshake was completed\n");
  }
gnutls_record_send (session, MSG, strlen (MSG));
ret = gnutls_record_recv (session, buffer, MAX_BUF);
if (gnutls_error_is_fatal (ret) == 1 || ret == 0)
  {
    if (ret == 0)
      {
        printf ("- Peer has closed the GNUTLS connection\n");
        goto end;
      }
    else
      {
        fprintf (stderr, "*** Error: %s\n", gnutls_strerror (ret));
        goto end;
      }
  }
else
  check_alert (session, ret);
if (ret > 0)
  {
    printf ("- Received %d bytes: ", ret);
    for (ii = 0; ii < ret; ii++)
      {
        fputc (buffer[ii], stdout);
      }
    fputs ("\n", stdout);
  }
gnutls_bye (session, GNUTLS_SHUT_RDWR);
end:
tcp_close (sd);
gnutls_deinit (session);
gnutls_srp_free_client_credentials (srp_cred);
gnutls_certificate_free_credentials (cert_cred);
gnutls_global_deinit ();

  return 0;
}

7.4 Server examples

This section contains examples of TLS and SSL servers, using GnuTLS.

7.4.1 Echo Server with X.509 authentication

This example is a very simple echo server which supports X.509 authentication, using the RSA ciphersuites.

#if HAVE_CONFIG_H
# include <config.h>
#endif
#include <stdio.h>
#include <stdlib.h>
#include <errno.h>
#include <sys/types.h>
#include <sys/socket.h>
#include <arpa/inet.h>
#include <netinet/in.h>
#include <string.h>
#include <unistd.h>
#include <gnutls/gnutls.h>
#define KEYFILE "key.pem"
#define CERTFILE "cert.pem"
#define CAFILE "ca.pem"
#define CRLFILE "crl.pem"
/* This is a sample TLS 1.0 echo server, using X.509 authentication.
 */

#define SA struct sockaddr
#define SOCKET_ERR(err,s) if(err==-1) {perror(s);return(1);}
#define MAX_BUF 1024

#define PORT 5556               /* listen to 5556 port */

#define DH_BITS 1024
/* These are global */
gnutls_certificate_credentials_t x509_cred;
gnutls_session_t
initialize_tls_session (void)
{
  gnutls_session_t session;

gnutls_init (&session, GNUTLS_SERVER);
/* avoid calling all the priority functions, since the defaults
 * are adequate.
 */
gnutls_set_default_priority (session);
gnutls_credentials_set (session, GNUTLS_CRD_CERTIFICATE, x509_cred);
/* request client certificate if any.
 */
gnutls_certificate_server_set_request (session, GNUTLS_CERT_REQUEST);
gnutls_dh_set_prime_bits (session, DH_BITS);
return session;
}
static gnutls_dh_params_t dh_params;
static int
generate_dh_params (void)
{
/* Generate Diffie Hellman parameters - for use with DHE
 * kx algorithms. These should be discarded and regenerated
 * once a day, once a week or once a month. Depending on the
 * security requirements.
 */
gnutls_dh_params_init (&dh_params);
gnutls_dh_params_generate2 (dh_params, DH_BITS);
return 0;
}
int
main (void)
{
int err, listen_sd, i;
int sd, ret;
struct sockaddr_in sa_serv;
struct sockaddr_in sa_cli;
int client_len;
char topbuf[512];
gnutls_session_t session;
char buffer[MAX_BUF + 1];

int optval = 1;
/* this must be called once in the program
 */
gnutls_global_init ();
gnutls_certificate_allocate_credentials (&x509_cred);
gnutls_certificate_set_x509_trust_file (x509_cred, CAFILE,
                                        GNUTLS_X509_FMT_PEM);
gnutls_certificate_set_x509_crl_file (x509_cred, CRLFILE,
                                      GNUTLS_X509_FMT_PEM);
gnutls_certificate_set_x509_key_file (x509_cred, CERTFILE, KEYFILE,
                                      GNUTLS_X509_FMT_PEM);
generate_dh_params ();
gnutls_certificate_set_dh_params (x509_cred, dh_params);
/* Socket operations
 */
listen_sd = socket (AF_INET, SOCK_STREAM, 0);
SOCKET_ERR (listen_sd, "socket");
memset (&sa_serv, '\0', sizeof (sa_serv));
sa_serv.sin_family = AF_INET;
sa_serv.sin_addr.s_addr = INADDR_ANY;

sa_serv.sin_port = htons (PORT);      /* Server Port number */

setsockopt (listen_sd, SOL_SOCKET, SO_REUSEADDR, &optval, sizeof (int));
err = bind (listen_sd, (SA *) & sa_serv, sizeof (sa_serv));
SOCKET_ERR (err, "bind");
err = listen (listen_sd, 1024);
SOCKET_ERR (err, "listen");
printf ("Server ready. Listening to port '%d'.\n\n", PORT);
client_len = sizeof (sa_cli);
for (;;)
  {
    session = initialize_tls_session ();
    sd = accept (listen_sd, (SA *) & sa_cli, &client_len);
    printf ("- connection from %s, port %d\n",

        inet_ntop (AF_INET, &sa_cli.sin_addr, topbuf,
                   sizeof (topbuf)), ntohs (sa_cli.sin_port));
gnutls_transport_set_ptr (session, (gnutls_transport_ptr_t) sd);
ret = gnutls_handshake (session);
if (ret < 0)
  {
    close (sd);
    gnutls_deinit (session);
    fprintf (stderr, "*** Handshake has failed (%s)\n\n",
             gnutls_strerror (ret));
    continue;
  }
printf ("- Handshake was completed\n");
/* see the Getting peer's information example */
/* print_info(session); */
i = 0;
for (;;)
  {
    memset (buffer, 0, MAX_BUF + 1);
    ret = gnutls_record_recv (session, buffer, MAX_BUF);
    if (ret == 0)
      {
        printf ("\n- Peer has closed the GNUTLS connection\n");
        break;
      }
    else if (ret < 0)
      {
        fprintf (stderr, "\n*** Received corrupted "
                 "data(%d). Closing the connection.\n\n", ret);
        break;
      }
    else if (ret > 0)
      {
        /* echo data back to the client
         */
        gnutls_record_send (session, buffer, strlen (buffer));
      }
  }
printf ("\n");
/* do not wait for the peer to close the connection.
 */
gnutls_bye (session, GNUTLS_SHUT_WR);

      close (sd);
      gnutls_deinit (session);
    }
  close (listen_sd);
  gnutls_certificate_free_credentials (x509_cred);
  gnutls_global_deinit ();
  return 0;
}

7.4.2 Echo Server with X.509 authentication II

The following example is a server which supports X.509 authentication. This server supports the export-grade cipher suites, the DHE ciphersuites and session resuming.

#if HAVE_CONFIG_H
# include <config.h>
#endif
#include <stdio.h>
#include <stdlib.h>
#include <errno.h>
#include <sys/types.h>
#include <sys/socket.h>
#include <arpa/inet.h>
#include <netinet/in.h>
#include <string.h>
#include <unistd.h>
#include <gnutls/gnutls.h>
#define KEYFILE "key.pem"
#define CERTFILE "cert.pem"
#define CAFILE "ca.pem"
#define CRLFILE "crl.pem"
/* This is a sample TLS 1.0 echo server.
 * Export-grade ciphersuites and session resuming are supported.
 */
#define SA struct sockaddr
#define SOCKET_ERR(err,s) if(err==-1) {perror(s);return(1);}
#define MAX_BUF 1024

#define PORT 5556               /* listen to 5556 port */

#define DH_BITS 1024

/* These are global */
gnutls_certificate_credentials_t cert_cred;
static void wrap_db_init (void);
static void wrap_db_deinit (void);
static int wrap_db_store (void *dbf, gnutls_datum_t key, gnutls_datum_t data);
static gnutls_datum_t wrap_db_fetch (void *dbf, gnutls_datum_t key);
static int wrap_db_delete (void *dbf, gnutls_datum_t key);
#define TLS_SESSION_CACHE 50
gnutls_session_t
initialize_tls_session (void)
{
  gnutls_session_t session;
  gnutls_init (&session, GNUTLS_SERVER);
  /* Use the default priorities, plus, export cipher suites.
   */
  gnutls_set_default_export_priority (session);
  gnutls_credentials_set (session, GNUTLS_CRD_CERTIFICATE, cert_cred);
  /* request client certificate if any.
   */
  gnutls_certificate_server_set_request (session, GNUTLS_CERT_REQUEST);
  gnutls_dh_set_prime_bits (session, DH_BITS);
  if (TLS_SESSION_CACHE != 0)
    {
      gnutls_db_set_retrieve_function (session, wrap_db_fetch);
      gnutls_db_set_remove_function (session, wrap_db_delete);
      gnutls_db_set_store_function (session, wrap_db_store);
      gnutls_db_set_ptr (session, NULL);
    }
  return session;
}
gnutls_dh_params_t dh_params;
/* Export-grade cipher suites require temporary RSA
 * keys.
 */
gnutls_rsa_params_t rsa_params;

int
generate_dh_params (void)
{
/* Generate Diffie Hellman parameters - for use with DHE
 * kx algorithms. These should be discarded and regenerated
 * once a day, once a week or once a month. Depends on the
 * security requirements.
 */
gnutls_dh_params_init (&dh_params);
gnutls_dh_params_generate2 (dh_params, DH_BITS);
return 0;
}
static int
generate_rsa_params (void)
{
gnutls_rsa_params_init (&rsa_params);
/* Generate RSA parameters - for use with RSA-export
 * cipher suites. These should be discarded and regenerated
 * once a day, once every 500 transactions etc. Depends on the
 * security requirements.
 */
gnutls_rsa_params_generate2 (rsa_params, 512);
return 0;
}
int
main (void)
{
int err, listen_sd, i;
int sd, ret;
struct sockaddr_in sa_serv;
struct sockaddr_in sa_cli;
int client_len;
char topbuf[512];
gnutls_session_t session;
char buffer[MAX_BUF + 1];
int optval = 1;
char name[256];
strcpy (name, "Echo Server");

/* this must be called once in the program
 */
gnutls_global_init ();
gnutls_certificate_allocate_credentials (&cert_cred);
gnutls_certificate_set_x509_trust_file (cert_cred, CAFILE,
                                        GNUTLS_X509_FMT_PEM);
gnutls_certificate_set_x509_crl_file (cert_cred, CRLFILE,
                                      GNUTLS_X509_FMT_PEM);
gnutls_certificate_set_x509_key_file (cert_cred, CERTFILE, KEYFILE,
                                      GNUTLS_X509_FMT_PEM);
generate_dh_params ();
generate_rsa_params ();
if (TLS_SESSION_CACHE != 0)
  {
    wrap_db_init ();
  }
gnutls_certificate_set_dh_params (cert_cred, dh_params);
gnutls_certificate_set_rsa_export_params (cert_cred, rsa_params);
/* Socket operations
 */
listen_sd = socket (AF_INET, SOCK_STREAM, 0);
SOCKET_ERR (listen_sd, "socket");
memset (&sa_serv, '\0', sizeof (sa_serv));
sa_serv.sin_family = AF_INET;
sa_serv.sin_addr.s_addr = INADDR_ANY;

sa_serv.sin_port = htons (PORT);      /* Server Port number */

setsockopt (listen_sd, SOL_SOCKET, SO_REUSEADDR, &optval, sizeof (int));
err = bind (listen_sd, (SA *) & sa_serv, sizeof (sa_serv));
SOCKET_ERR (err, "bind");
err = listen (listen_sd, 1024);
SOCKET_ERR (err, "listen");
printf ("%s ready. Listening to port '%d'.\n\n", name, PORT);
client_len = sizeof (sa_cli);
for (;;)

{
session = initialize_tls_session ();
sd = accept (listen_sd, (SA *) & sa_cli, &client_len);
printf ("- connection from %s, port %d\n",
        inet_ntop (AF_INET, &sa_cli.sin_addr, topbuf,
                   sizeof (topbuf)), ntohs (sa_cli.sin_port));
gnutls_transport_set_ptr (session, (gnutls_transport_ptr_t) sd);
ret = gnutls_handshake (session);
if (ret < 0)
  {
    close (sd);
    gnutls_deinit (session);
    fprintf (stderr, "*** Handshake has failed (%s)\n\n",
             gnutls_strerror (ret));
    continue;
  }
printf ("- Handshake was completed\n");
/* print_info(session); */
i = 0;
for (;;)
  {
    memset (buffer, 0, MAX_BUF + 1);
    ret = gnutls_record_recv (session, buffer, MAX_BUF);
    if (ret == 0)
      {
        printf ("\n- Peer has closed the TLS connection\n");
        break;
      }
    else if (ret < 0)
      {
        fprintf (stderr, "\n*** Received corrupted "
                 "data(%d). Closing the connection.\n\n", ret);
        break;
      }
    else if (ret > 0)
      {
        /* echo data back to the client
         */
        gnutls_record_send (session, buffer, strlen (buffer));
      }
  }

      printf ("\n");
      /* do not wait for the peer to close the connection.
       */
      gnutls_bye (session, GNUTLS_SHUT_WR);
      close (sd);
      gnutls_deinit (session);
    }
  close (listen_sd);
  gnutls_certificate_free_credentials (cert_cred);
  gnutls_global_deinit ();
  return 0;
}

/* Functions and other stuff needed for session resuming.
 * This is done using a very simple list which holds session ids
 * and session data.
 */
#define MAX_SESSION_ID_SIZE 32
#define MAX_SESSION_DATA_SIZE 512
typedef struct
{
  char session_id[MAX_SESSION_ID_SIZE];
  int session_id_size;
  char session_data[MAX_SESSION_DATA_SIZE];
  int session_data_size;
} CACHE;
static CACHE *cache_db;
static int cache_db_ptr = 0;
static void
wrap_db_init (void)
{
  /* allocate cache_db */
  cache_db = calloc (1, TLS_SESSION_CACHE * sizeof (CACHE));
}

static void
wrap_db_deinit (void)
{
return;
}
static int
wrap_db_store (void *dbf, gnutls_datum_t key, gnutls_datum_t data)
{
if (cache_db == NULL)
  return -1;
if (key.size > MAX_SESSION_ID_SIZE)
  return -1;
if (data.size > MAX_SESSION_DATA_SIZE)
  return -1;
memcpy (cache_db[cache_db_ptr].session_id, key.data, key.size);
cache_db[cache_db_ptr].session_id_size = key.size;
memcpy (cache_db[cache_db_ptr].session_data, data.data, data.size);
cache_db[cache_db_ptr].session_data_size = data.size;
cache_db_ptr++;
cache_db_ptr %= TLS_SESSION_CACHE;
return 0;
}
static gnutls_datum_t
wrap_db_fetch (void *dbf, gnutls_datum_t key)
{
gnutls_datum_t res = { NULL, 0 };
int i;
if (cache_db == NULL)
  return res;
for (i = 0; i < TLS_SESSION_CACHE; i++)
  {
    if (key.size == cache_db[i].session_id_size &&
        memcmp (key.data, cache_db[i].session_id, key.size) == 0)
      {

          res.size = cache_db[i].session_data_size;
          res.data = gnutls_malloc (res.size);
          if (res.data == NULL)
            return res;
          memcpy (res.data, cache_db[i].session_data, res.size);
          return res;
        }
    }
  return res;
}
static int
wrap_db_delete (void *dbf, gnutls_datum_t key)
{
  int i;
  if (cache_db == NULL)
    return -1;
  for (i = 0; i < TLS_SESSION_CACHE; i++)
    {
      if (key.size == cache_db[i].session_id_size &&
          memcmp (key.data, cache_db[i].session_id, key.size) == 0)
        {
          cache_db[i].session_id_size = 0;
          cache_db[i].session_data_size = 0;
          return 0;
        }
    }
  return -1;
}

7.4.3 Echo Server with OpenPGP authentication

The following example is an echo server which supports OpenPGP key authentication. You can easily combine this functionality —that is have a server that supports both X.509 and OpenPGP certificates— but we separated them to keep these examples as simple as possible.

#if HAVE_CONFIG_H
# include <config.h>
#endif

#include <stdio.h>
#include <stdlib.h>
#include <errno.h>
#include <sys/types.h>
#include <sys/socket.h>
#include <arpa/inet.h>
#include <netinet/in.h>
#include <string.h>
#include <unistd.h>
#include <gnutls/gnutls.h>
/* Must be linked against gnutls-extra.
 */
#include <gnutls/extra.h>
#define KEYFILE "secret.asc"
#define CERTFILE "public.asc"
#define RINGFILE "ring.gpg"
/* This is a sample TLS 1.0-OpenPGP echo server.
 */

#define SA struct sockaddr
#define SOCKET_ERR(err,s) if(err==-1) {perror(s);return(1);}
#define MAX_BUF 1024

#define PORT 5556               /* listen to 5556 port */

#define DH_BITS 1024
/* These are global */
gnutls_certificate_credentials_t cred;
const int cert_type_priority[2] = { GNUTLS_CRT_OPENPGP, 0 };
gnutls_dh_params_t dh_params;
static int
generate_dh_params (void)
{
  /* Generate Diffie Hellman parameters - for use with DHE
   * kx algorithms. These should be discarded and regenerated
   * once a day, once a week or once a month. Depending on the
   * security requirements.
   */
  gnutls_dh_params_init (&dh_params);
  gnutls_dh_params_generate2 (dh_params, DH_BITS);

}
gnutls_session_t
initialize_tls_session (void)
{
gnutls_session_t session;
gnutls_init (&session, GNUTLS_SERVER);
/* avoid calling all the priority functions, since the defaults
 * are adequate.
 */
gnutls_set_default_priority (session);
gnutls_credentials_set (session, GNUTLS_CRD_CERTIFICATE, cred);
/* request client certificate if any.
 */
gnutls_certificate_server_set_request (session, GNUTLS_CERT_REQUEST);
gnutls_dh_set_prime_bits (session, DH_BITS);
return session;
}
int
main (void)
{
int err, listen_sd, i;
int sd, ret;
struct sockaddr_in sa_serv;
struct sockaddr_in sa_cli;
int client_len;
char topbuf[512];
gnutls_session_t session;
char buffer[MAX_BUF + 1];
int optval = 1;
char name[256];
strcpy (name, "Echo Server");
/* this must be called once in the program
 */
gnutls_global_init ();
gnutls_certificate_allocate_credentials (&cred);
gnutls_certificate_set_openpgp_keyring_file (cred, RINGFILE);

gnutls_certificate_set_openpgp_key_file (cred, CERTFILE, KEYFILE);
generate_dh_params ();
gnutls_certificate_set_dh_params (cred, dh_params);
/* Socket operations
 */
listen_sd = socket (AF_INET, SOCK_STREAM, 0);
SOCKET_ERR (listen_sd, "socket");
memset (&sa_serv, '\0', sizeof (sa_serv));
sa_serv.sin_family = AF_INET;
sa_serv.sin_addr.s_addr = INADDR_ANY;

sa_serv.sin_port = htons (PORT);      /* Server Port number */

setsockopt (listen_sd, SOL_SOCKET, SO_REUSEADDR, &optval, sizeof (int));
err = bind (listen_sd, (SA *) & sa_serv, sizeof (sa_serv));
SOCKET_ERR (err, "bind");
err = listen (listen_sd, 1024);
SOCKET_ERR (err, "listen");
printf ("%s ready. Listening to port '%d'.\n\n", name, PORT);
client_len = sizeof (sa_cli);
for (;;)
  {
    session = initialize_tls_session ();
    gnutls_certificate_type_set_priority (session, cert_type_priority);
    sd = accept (listen_sd, (SA *) & sa_cli, &client_len);
    printf ("- connection from %s, port %d\n",
            inet_ntop (AF_INET, &sa_cli.sin_addr, topbuf,
                       sizeof (topbuf)), ntohs (sa_cli.sin_port));
    gnutls_transport_set_ptr (session, (gnutls_transport_ptr_t) sd);
    ret = gnutls_handshake (session);
    if (ret < 0)
      {
        close (sd);
        gnutls_deinit (session);
        fprintf (stderr, "*** Handshake has failed (%s)\n\n",
                 gnutls_strerror (ret));
        continue;

  }
printf ("- Handshake was completed\n");
/* see the Getting peer's information example */
/* print_info(session); */
i = 0;
for (;;)
  {
    memset (buffer, 0, MAX_BUF + 1);
    ret = gnutls_record_recv (session, buffer, MAX_BUF);
    if (ret == 0)
      {
        printf ("\n- Peer has closed the GNUTLS connection\n");
        break;
      }
    else if (ret < 0)
      {
        fprintf (stderr, "\n*** Received corrupted "
                 "data(%d). Closing the connection.\n\n", ret);
        break;
      }
    else if (ret > 0)
      {
        /* echo data back to the client
         */
        gnutls_record_send (session, buffer, strlen (buffer));
      }
  }
printf ("\n");
/* do not wait for the peer to close the connection.
 */
gnutls_bye (session, GNUTLS_SHUT_WR);
close (sd);
gnutls_deinit (session);
}
close (listen_sd);
gnutls_certificate_free_credentials (cred);
gnutls_global_deinit ();

}

7.4.4 Echo Server with SRP authentication

This is a server which supports SRP authentication. It is also possible to combine this functionality with a certificate server. Here it is separate for simplicity.

#if HAVE_CONFIG_H
# include <config.h>
#endif
#include <stdio.h>
#include <stdlib.h>
#include <errno.h>
#include <sys/types.h>
#include <sys/socket.h>
#include <arpa/inet.h>
#include <netinet/in.h>
#include <string.h>
#include <unistd.h>
#include <gnutls/gnutls.h>
#include <gnutls/extra.h>
#define SRP_PASSWD "tpasswd"
#define SRP_PASSWD_CONF "tpasswd.conf"
#define KEYFILE "key.pem"
#define CERTFILE "cert.pem"
#define CAFILE "ca.pem"
/* This is a sample TLS-SRP echo server.
 */
#define SA struct sockaddr
#define SOCKET_ERR(err,s) if(err==-1) {perror(s);return(1);}
#define MAX_BUF 1024

#define PORT 5556               /* listen to 5556 port */

/* These are global */
gnutls_srp_server_credentials_t srp_cred;
gnutls_certificate_credentials_t cert_cred;
gnutls_session_t
initialize_tls_session (void)
{
  gnutls_session_t session;
  const int kx_priority[] = { GNUTLS_KX_SRP, GNUTLS_KX_SRP_DSS,
    GNUTLS_KX_SRP_RSA, 0

};
gnutls_init (&session, GNUTLS_SERVER);
gnutls_set_default_priority (session);
gnutls_kx_set_priority (session, kx_priority);
gnutls_credentials_set (session, GNUTLS_CRD_SRP, srp_cred);
/* for the certificate authenticated ciphersuites.
 */
gnutls_credentials_set (session, GNUTLS_CRD_CERTIFICATE, cert_cred);
/* request client certificate if any.
 */
gnutls_certificate_server_set_request (session, GNUTLS_CERT_IGNORE);
return session;
}
int
main (void)
{
int err, listen_sd, i;
int sd, ret;
struct sockaddr_in sa_serv;
struct sockaddr_in sa_cli;
int client_len;
char topbuf[512];
gnutls_session_t session;
char buffer[MAX_BUF + 1];
int optval = 1;
char name[256];
strcpy (name, "Echo Server");
/* these must be called once in the program
 */
gnutls_global_init ();

gnutls_global_init_extra ();  /* for SRP */

/* SRP_PASSWD a password file (created with the included srptool utility)
 */
gnutls_srp_allocate_server_credentials (&srp_cred);
gnutls_srp_set_server_credentials_file (srp_cred, SRP_PASSWD,
                                        SRP_PASSWD_CONF);
gnutls_certificate_allocate_credentials (&cert_cred);

gnutls_certificate_set_x509_trust_file (cert_cred, CAFILE,
                                        GNUTLS_X509_FMT_PEM);
gnutls_certificate_set_x509_key_file (cert_cred, CERTFILE, KEYFILE,
                                      GNUTLS_X509_FMT_PEM);
/* TCP socket operations
 */
listen_sd = socket (AF_INET, SOCK_STREAM, 0);
SOCKET_ERR (listen_sd, "socket");
memset (&sa_serv, '\0', sizeof (sa_serv));
sa_serv.sin_family = AF_INET;
sa_serv.sin_addr.s_addr = INADDR_ANY;

sa_serv.sin_port = htons (PORT);      /* Server Port number */

setsockopt (listen_sd, SOL_SOCKET, SO_REUSEADDR, &optval, sizeof (int));
err = bind (listen_sd, (SA *) & sa_serv, sizeof (sa_serv));
SOCKET_ERR (err, "bind");
err = listen (listen_sd, 1024);
SOCKET_ERR (err, "listen");
printf ("%s ready. Listening to port '%d'.\n\n", name, PORT);
client_len = sizeof (sa_cli);
for (;;)
  {
    session = initialize_tls_session ();
    sd = accept (listen_sd, (SA *) & sa_cli, &client_len);
    printf ("- connection from %s, port %d\n",
            inet_ntop (AF_INET, &sa_cli.sin_addr, topbuf,
                       sizeof (topbuf)), ntohs (sa_cli.sin_port));
    gnutls_transport_set_ptr (session, (gnutls_transport_ptr_t) sd);
    ret = gnutls_handshake (session);
    if (ret < 0)
      {
        close (sd);
        gnutls_deinit (session);
        fprintf (stderr, "*** Handshake has failed (%s)\n\n",
                 gnutls_strerror (ret));
        continue;
      }
    printf ("- Handshake was completed\n");

/* print_info(session); */
i = 0;
for (;;)
  {
    memset (buffer, 0, MAX_BUF + 1);
    ret = gnutls_record_recv (session, buffer, MAX_BUF);
    if (ret == 0)
      {
        printf ("\n- Peer has closed the GNUTLS connection\n");
        break;
      }
    else if (ret < 0)
      {
        fprintf (stderr, "\n*** Received corrupted "
                 "data(%d). Closing the connection.\n\n", ret);
        break;
      }
    else if (ret > 0)
      {
        /* echo data back to the client
         */
        gnutls_record_send (session, buffer, strlen (buffer));
      }
  }
printf ("\n");
/* do not wait for the peer to close the connection. */
gnutls_bye (session, GNUTLS_SHUT_WR);
close (sd);
gnutls_deinit (session);
}
close (listen_sd);
gnutls_srp_free_server_credentials (srp_cred);
gnutls_certificate_free_credentials (cert_cred);
gnutls_global_deinit ();
return 0;
}

7.4.5 Echo Server with anonymous authentication

This example server support anonymous authentication, and could be used to serve the example client for anonymous authentication.

#if HAVE_CONFIG_H
# include <config.h>
#endif
#include <stdio.h>
#include <stdlib.h>
#include <errno.h>
#include <sys/types.h>
#include <sys/socket.h>
#include <arpa/inet.h>
#include <netinet/in.h>
#include <string.h>
#include <unistd.h>
#include <gnutls/gnutls.h>
/* This is a sample TLS 1.0 echo server, for anonymous authentication only.
 */

#define SA struct sockaddr
#define SOCKET_ERR(err,s) if(err==-1) {perror(s);return(1);}
#define MAX_BUF 1024

#define PORT 5556               /* listen to 5556 port */

#define DH_BITS 1024
/* These are global */
gnutls_anon_server_credentials_t anoncred;
gnutls_session_t
initialize_tls_session (void)
{
  gnutls_session_t session;
  const int kx_prio[] = { GNUTLS_KX_ANON_DH, 0 };
  gnutls_init (&session, GNUTLS_SERVER);
  /* avoid calling all the priority functions, since the defaults
   * are adequate.
   */
  gnutls_set_default_priority (session);
  gnutls_kx_set_priority (session, kx_prio);
  gnutls_credentials_set (session, GNUTLS_CRD_ANON, anoncred);

gnutls_dh_set_prime_bits (session, DH_BITS);
return session;
}
static gnutls_dh_params_t dh_params;
static int
generate_dh_params (void)
{
/* Generate Diffie Hellman parameters - for use with DHE
 * kx algorithms. These should be discarded and regenerated
 * once a day, once a week or once a month. Depending on the
 * security requirements.
 */
gnutls_dh_params_init (&dh_params);
gnutls_dh_params_generate2 (dh_params, DH_BITS);
return 0;
}
int
main (void)
{
int err, listen_sd, i;
int sd, ret;
struct sockaddr_in sa_serv;
struct sockaddr_in sa_cli;
int client_len;
char topbuf[512];
gnutls_session_t session;
char buffer[MAX_BUF + 1];
int optval = 1;
/* this must be called once in the program
 */
gnutls_global_init ();
gnutls_anon_allocate_server_credentials (&anoncred);
generate_dh_params ();
gnutls_anon_set_server_dh_params (anoncred, dh_params);
/* Socket operations

 */
listen_sd = socket (AF_INET, SOCK_STREAM, 0);
SOCKET_ERR (listen_sd, "socket");
memset (&sa_serv, '\0', sizeof (sa_serv));
sa_serv.sin_family = AF_INET;
sa_serv.sin_addr.s_addr = INADDR_ANY;

sa_serv.sin_port = htons (PORT);      /* Server Port number */

setsockopt (listen_sd, SOL_SOCKET, SO_REUSEADDR, &optval, sizeof (int));
err = bind (listen_sd, (SA *) & sa_serv, sizeof (sa_serv));
SOCKET_ERR (err, "bind");
err = listen (listen_sd, 1024);
SOCKET_ERR (err, "listen");
printf ("Server ready. Listening to port '%d'.\n\n", PORT);
client_len = sizeof (sa_cli);
for (;;)
  {
    session = initialize_tls_session ();
    sd = accept (listen_sd, (SA *) & sa_cli, &client_len);
    printf ("- connection from %s, port %d\n",
            inet_ntop (AF_INET, &sa_cli.sin_addr, topbuf,
                       sizeof (topbuf)), ntohs (sa_cli.sin_port));
    gnutls_transport_set_ptr (session, (gnutls_transport_ptr_t) sd);
    ret = gnutls_handshake (session);
    if (ret < 0)
      {
        close (sd);
        gnutls_deinit (session);
        fprintf (stderr, "*** Handshake has failed (%s)\n\n",
                 gnutls_strerror (ret));
        continue;
      }
    printf ("- Handshake was completed\n");
    /* see the Getting peer's information example */
    /* print_info(session); */
    i = 0;
    for (;;)
      {

memset (buffer, 0, MAX_BUF + 1);
ret = gnutls_record_recv (session, buffer, MAX_BUF);
if (ret == 0)
{
  printf ("\n- Peer has closed the GNUTLS connection\n");
  break;
}
else if (ret < 0)
{
  fprintf (stderr, "\n*** Received corrupted "
           "data(%d). Closing the connection.\n\n", ret);
  break;
}
else if (ret > 0)
{
  /* echo data back to the client
   */
  gnutls_record_send (session, buffer, strlen (buffer));
}
}
printf ("\n");
/* do not wait for the peer to close the connection.
*/
gnutls_bye (session, GNUTLS_SHUT_WR);
close (sd);
gnutls_deinit (session);
}
close (listen_sd);
gnutls_anon_free_server_credentials (anoncred);
gnutls_global_deinit ();
return 0;
}

7.5 Miscellaneous examples

7.5.1 Checking for an alert

This is a function that checks if an alert has been received in the current session.

#if HAVE_CONFIG_H
# include <config.h>

#endif
#include <stdio.h>
#include <stdlib.h>
#include <gnutls/gnutls.h>
/* This function will check whether the given return code from
 * a gnutls function (recv/send), is an alert, and will print
 * that alert.
 */
void
check_alert (gnutls_session_t session, int ret)
{
  int last_alert;
  if (ret == GNUTLS_E_WARNING_ALERT_RECEIVED
      || ret == GNUTLS_E_FATAL_ALERT_RECEIVED)
    {
      last_alert = gnutls_alert_get (session);
      /* The check for renegotiation is only useful if we are
       * a server, and we had requested a rehandshake.
       */
      if (last_alert == GNUTLS_A_NO_RENEGOTIATION &&
          ret == GNUTLS_E_WARNING_ALERT_RECEIVED)
        printf ("* Received NO_RENEGOTIATION alert. "
                "Client Does not support renegotiation.\n");
      else
        printf ("* Received alert '%d': %s.\n", last_alert,
                gnutls_alert_get_name (last_alert));
    }
}

7.5.2 X.509 certificate parsing example

To demonstrate the X.509 parsing capabilities an example program is listed below. That program reads the peer’s certificate, and prints information about it.

#if HAVE_CONFIG_H
# include <config.h>
#endif
#include <stdio.h>
#include <stdlib.h>
#include <gnutls/gnutls.h>
#include <gnutls/x509.h>
static const char *

bin2hex (const void *bin, size_t bin_size)
{
static char printable[110];
const unsigned char *_bin = bin;
char *print;
size_t i;
if (bin_size > 50)
  bin_size = 50;
print = printable;
for (i = 0; i < bin_size; i++)
  {
    sprintf (print, "%.2x ", _bin[i]);
    print += 2;
  }
return printable;
}
/* This function will print information about this session's peer
* certificate.
*/
void
print_x509_certificate_info (gnutls_session_t session)
{
char serial[40];
char dn[128];
size_t size;
unsigned int algo, bits;
time_t expiration_time, activation_time;
const gnutls_datum_t *cert_list;
int cert_list_size = 0;
gnutls_x509_crt_t cert;
/* This function only works for X.509 certificates.
 */
if (gnutls_certificate_type_get (session) != GNUTLS_CRT_X509)
  return;
cert_list = gnutls_certificate_get_peers (session, &cert_list_size);
printf ("Peer provided %d certificates.\n", cert_list_size);
if (cert_list_size > 0)
  {

/* we only print information about the first certificate.
 */
gnutls_x509_crt_init (&cert);
gnutls_x509_crt_import (cert, &cert_list[0], GNUTLS_X509_FMT_DER);
printf ("Certificate info:\n");
expiration_time = gnutls_x509_crt_get_expiration_time (cert);
activation_time = gnutls_x509_crt_get_activation_time (cert);
printf ("\tCertificate is valid since: %s", ctime (&activation_time));
printf ("\tCertificate expires: %s", ctime (&expiration_time));
/* Print the serial number of the certificate.
 */
size = sizeof (serial);
gnutls_x509_crt_get_serial (cert, serial, &size);
size = sizeof (serial);
printf ("\tCertificate serial number: %s\n", bin2hex (serial, size));
/* Extract some of the public key algorithm's parameters
 */
algo = gnutls_x509_crt_get_pk_algorithm (cert, &bits);
printf ("Certificate public key: %s",
        gnutls_pk_algorithm_get_name (algo));
/* Print the version of the X.509
 * certificate.
 */
printf ("\tCertificate version: #%d\n",
        gnutls_x509_crt_get_version (cert));
size = sizeof (dn);
gnutls_x509_crt_get_dn (cert, dn, &size);
printf ("\tDN: %s\n", dn);
size = sizeof (dn);
gnutls_x509_crt_get_issuer_dn (cert, dn, &size);
printf ("\tIssuer's DN: %s\n", dn);
gnutls_x509_crt_deinit (cert);
}
}

7.5.3 Certificate request generation

The following example is about generating a certificate request, and a private key. A certificate request can be later be processed by a CA, which should return a signed certificate.

#if HAVE_CONFIG_H
# include <config.h>
#endif
#include <stdio.h>
#include <stdlib.h>
#include <string.h>
#include <gnutls/gnutls.h>
#include <gnutls/x509.h>
#include <time.h>
/* This example will generate a private key and a certificate
 * request.
 */
int
main (void)
{
  gnutls_x509_crq_t crq;
  gnutls_x509_privkey_t key;
  unsigned char buffer[10 * 1024];
  int buffer_size = sizeof (buffer);
  gnutls_global_init ();
  /* Initialize an empty certificate request, and
   * an empty private key.
   */
  gnutls_x509_crq_init (&crq);
  gnutls_x509_privkey_init (&key);
  /* Generate a 1024 bit RSA private key.
   */
  gnutls_x509_privkey_generate (key, GNUTLS_PK_RSA, 1024, 0);
  /* Add stuff to the distinguished name
   */
  gnutls_x509_crq_set_dn_by_oid (crq, GNUTLS_OID_X520_COUNTRY_NAME,
                                 0, "GR", 2);
  gnutls_x509_crq_set_dn_by_oid (crq, GNUTLS_OID_X520_COMMON_NAME,
                                 0, "Nikos", strlen ("Nikos"));

/* Set the request version.
 */
gnutls_x509_crq_set_version (crq, 1);
/* Set a challenge password.
 */
gnutls_x509_crq_set_challenge_password (crq, "something to remember here");
/* Associate the request with the private key
 */
gnutls_x509_crq_set_key (crq, key);
/* Self sign the certificate request.
 */
gnutls_x509_crq_sign (crq, key);
/* Export the PEM encoded certificate request, and
 * display it.
 */
gnutls_x509_crq_export (crq, GNUTLS_X509_FMT_PEM, buffer, &buffer_size);
printf ("Certificate Request: \n%s", buffer);

/* Export the PEM encoded private key, and
 * display it.
 */
buffer_size = sizeof (buffer);
gnutls_x509_privkey_export (key, GNUTLS_X509_FMT_PEM, buffer, &buffer_size);
printf ("\n\nPrivate key: \n%s", buffer);
gnutls_x509_crq_deinit (crq);
gnutls_x509_privkey_deinit (key);
return 0;
}

7.5.4 PKCS #12 structure generation

The following example is about generating a PKCS #12 structure.

#if HAVE_CONFIG_H
# include <config.h>
#endif

#include <stdio.h>
#include <stdlib.h>
#include <gnutls/gnutls.h>
#include <gnutls/pkcs12.h>
#define OUTFILE "out.p12"
/* This function will write a pkcs12 structure into a file.
* cert: is a DER encoded certificate
* pkcs8_key: is a PKCS #8 encrypted key (note that this must be
*  encrypted using a PKCS #12 cipher, or some browsers will crash)
* password: is the password used to encrypt the PKCS #12 packet.
*/
int
write_pkcs12 (const gnutls_datum_t * cert,
            const gnutls_datum_t * pkcs8_key, const char *password)
{
gnutls_pkcs12_t pkcs12;
int ret, bag_index;
gnutls_pkcs12_bag_t bag, key_bag;
char pkcs12_struct[10 * 1024];
int pkcs12_struct_size;
FILE *fd;
/* A good idea might be to use gnutls_x509_privkey_get_key_id()
 * to obtain a unique ID.
 */
gnutls_datum_t key_id = { "\x00\x00\x07", 3 };
gnutls_global_init ();
/* Firstly we create two helper bags, which hold the certificate,
 * and the (encrypted) key.
 */
gnutls_pkcs12_bag_init (&bag);
gnutls_pkcs12_bag_init (&key_bag);
ret = gnutls_pkcs12_bag_set_data (bag, GNUTLS_BAG_CERTIFICATE, cert);
if (ret < 0)
  {
    fprintf (stderr, "ret: %s\n", gnutls_strerror (ret));
    return 1;
  }
/* ret now holds the bag's index.
 */

bag_index = ret;
/* Associate a friendly name with the given certificate. Used
 * by browsers.
 */
gnutls_pkcs12_bag_set_friendly_name (bag, bag_index, "My name");
/* Associate the certificate with the key using a unique key
 * ID.
 */
gnutls_pkcs12_bag_set_key_id (bag, bag_index, &key_id);
/* use weak encryption for the certificate.
 */
gnutls_pkcs12_bag_encrypt (bag, password, GNUTLS_PKCS_USE_PKCS12_RC2_40);
/* Now the key.
 */
ret = gnutls_pkcs12_bag_set_data (key_bag,
                                  GNUTLS_BAG_PKCS8_ENCRYPTED_KEY,
                                  pkcs8_key);
if (ret < 0)
  {
    fprintf (stderr, "ret: %s\n", gnutls_strerror (ret));
    return 1;
  }
/* Note that since the PKCS #8 key is already encrypted we don't
 * bother encrypting that bag.
 */
bag_index = ret;
gnutls_pkcs12_bag_set_friendly_name (key_bag, bag_index, "My name");
gnutls_pkcs12_bag_set_key_id (key_bag, bag_index, &key_id);

/* The bags were filled. Now create the PKCS #12 structure.
 */
gnutls_pkcs12_init (&pkcs12);
/* Insert the two bags in the PKCS #12 structure.
 */
gnutls_pkcs12_set_bag (pkcs12, bag);
gnutls_pkcs12_set_bag (pkcs12, key_bag);

/* Generate a message authentication code for the PKCS #12
 * structure.
 */
gnutls_pkcs12_generate_mac (pkcs12, password);
pkcs12_struct_size = sizeof (pkcs12_struct);
ret =
  gnutls_pkcs12_export (pkcs12, GNUTLS_X509_FMT_DER, pkcs12_struct,
                        &pkcs12_struct_size);
if (ret < 0)
  {
    fprintf (stderr, "ret: %s\n", gnutls_strerror (ret));
    return 1;
  }
fd = fopen (OUTFILE, "w");
if (fd == NULL)
  {
    fprintf (stderr, "cannot open file\n");
    return 1;
  }
fwrite (pkcs12_struct, 1, pkcs12_struct_size, fd);
fclose (fd);
gnutls_pkcs12_bag_deinit (bag);
gnutls_pkcs12_bag_deinit (key_bag);
gnutls_pkcs12_deinit (pkcs12);
return 0;
}

7.6 Compatibility with the OpenSSL library

To ease GnuTLS’ integration with existing applications, a compatibility layer with the widely used OpenSSL library is included in the gnutls-openssl library. This compatibility layer is not complete and it is not intended to completely reimplement the OpenSSL API with GnuTLS. It only provides source-level compatibility. There is currently no attempt to make it binary-compatible with OpenSSL. The prototypes for the compatibility functions are in the ‘gnutls/openssl.h’ header file. Current limitations imposed by the compatibility layer include:

• Error handling is not thread safe.

8 Included programs

Included with GnuTLS are also a few command line tools that let you use the library for common tasks without writing an application. The applications are discussed in this chapter.

8.1 Invoking srptool

The ‘srptool’ is a very simple program that emulates the programs in the Stanford SRP libraries. It is intended for use in places where you don’t expect SRP authentication to be the used for system users. Traditionally libsrp used two files. One called ’tpasswd’ which holds usernames and verifiers, and ’tpasswd.conf’ which holds generators and primes. How to use srptool:

• To create tpasswd.conf which holds the g and n values for SRP protocol (generator and

a large prime), run:

         $ srptool --create-conf /etc/tpasswd.conf

• This command will create /etc/tpasswd and will add user ’test’ (you will also be

prompted for a password). Verifiers are stored by default in the way libsrp expects.

         $ srptool --passwd /etc/tpasswd \
             --passwd-conf /etc/tpasswd.conf -u test

• This command will check against a password. If the password matches the one in

/etc/tpasswd you will get an ok.

         $ srptool --passwd /etc/tpasswd \
             --passwd-conf /etc/tpasswd.conf --verify -u test

8.2 Invoking gnutls-cli

Simple client program to set up a TLS connection to some other computer. It sets up a TLS connection and forwards data from the standard input to the secured socket and vice versa.

GNU TLS test client
Usage:  gnutls-cli [options] hostname
     -d, --debug integer      Enable debugging
     -r, --resume             Connect, establish a session. Connect
                              again and resume this session.
     -s, --starttls           Connect, establish a plain session and
                              start TLS when EOF or a SIGALRM is
                              received.
     --crlf                   Send CR LF instead of LF.
     --x509fmtder             Use DER format for certificates to read
                              from.
     -f, --fingerprint        Send the openpgp fingerprint, instead
                              of the key.
     --disable-extensions     Disable all the TLS extensions.
     --xml                    Print the certificate information in

                         XML format.
--print-cert             Print the certificate in PEM format.
-p, --port integer       The port to connect to.
--recordsize integer     The maximum record size to advertize.
-V, --verbose            More verbose output.
--ciphers cipher1 cipher2...
                         Ciphers to enable.
--protocols protocol1 protocol2...
                         Protocols to enable.
--comp comp1 comp2...    Compression methods to enable.
--macs mac1 mac2...      MACs to enable.
--kx kx1 kx2...          Key exchange methods to enable.
--ctypes certType1 certType2...
                         Certificate types to enable.
--x509cafile FILE        Certificate file to use.
--x509crlfile FILE       CRL file to use.
--pgpkeyfile FILE        PGP Key file to use.
--pgpkeyring FILE        PGP Key ring file to use.
--pgptrustdb FILE        PGP trustdb file to use.
--pgpcertfile FILE       PGP Public Key (certificate) file to
                         use.
--x509keyfile FILE       X.509 key file to use.
--x509certfile FILE      X.509 Certificate file to use.
--srpusername NAME       SRP username to use.
--srppasswd PASSWD       SRP password to use.
--insecure               Don't abort program if server
                         certificate can't be validated.
-l, --list               Print a list of the supported
                         algorithms  and modes.
-h, --help               prints this help
-v, --version            prints the program's version number
--copyright              prints the program's license

8.3 Invoking gnutls-cli-debug

This program was created to assist in debugging GnuTLS, but it might be useful to extract a TLS server’s capabilities. It’s purpose is to connect onto a TLS server, perform some tests and print the server’s capabilities. If called with the ‘-v’ parameter a more checks will be performed. An example output is:

crystal:/cvs/gnutls/src$ ./gnutls-cli-debug localhost -p 5556
Resolving 'localhost'...
Connecting to '127.0.0.1:5556'...
Checking for TLS 1.1 support... yes
Checking fallback from TLS 1.1 to... N/A
Checking for TLS 1.0 support... yes
Checking for SSL 3.0 support... yes
Checking for version rollback bug in RSA PMS... no
Checking for version rollback bug in Client Hello... no
Checking whether we need to disable TLS 1.0... N/A

Chapter 8: Included programs 90

Checking whether the server ignores the RSA PMS version... no
Checking whether the server can accept Hello Extensions... yes
Checking whether the server can accept cipher suites not in SSL 3.0 spec... yes
Checking whether the server can accept a bogus TLS record version in the client hello... yes
Checking for certificate information... N/A
Checking for trusted CAs... N/A
Checking whether the server understands TLS closure alerts... yes
Checking whether the server supports session resumption... yes
Checking for export-grade ciphersuite support... no
Checking RSA-export ciphersuite info... N/A
Checking for anonymous authentication support... no
Checking anonymous Diffie Hellman group info... N/A
Checking for ephemeral Diffie Hellman support... no
Checking ephemeral Diffie Hellman group info... N/A
Checking for AES cipher support (TLS extension)... yes
Checking for 3DES cipher support... yes
Checking for ARCFOUR 128 cipher support... yes
Checking for ARCFOUR 40 cipher support... no
Checking for MD5 MAC support... yes
Checking for SHA1 MAC support... yes
Checking for ZLIB compression support (TLS extension)... yes
Checking for LZO compression support (GnuTLS extension)... yes
Checking for max record size (TLS extension)... yes
Checking for SRP authentication support (TLS extension)... yes
Checking for OpenPGP authentication support (TLS extension)... no

8.4 Invoking gnutls-serv

Simple server program that listens to incoming TLS connections.

GNU TLS test server
Usage: gnutls-serv [options]

-d, --debug integer      Enable debugging
-g, --generate           Generate Diffie Hellman Parameters.
-p, --port integer       The port to connect to.
-q, --quiet              Suppress some messages.
--nodb                   Does not use the resume database.
--http                   Act as an HTTP Server.
--echo                   Act as an Echo Server.
--dhparams FILE          DH params file to use.
--x509fmtder             Use DER format for certificates
--x509cafile FILE        Certificate file to use.
--x509crlfile FILE       CRL file to use.
--pgpkeyring FILE        PGP Key ring file to use.
--pgptrustdb FILE        PGP trustdb file to use.
--pgpkeyfile FILE        PGP Key file to use.
--pgpcertfile FILE       PGP Public Key (certificate) file to
                         use.
--x509keyfile FILE       X.509 key file to use.
--x509certfile FILE      X.509 Certificate file to use.
--x509dsakeyfile FILE    Alternative X.509 key file to use.
--x509dsacertfile FILE   Alternative X.509 certificate file to

                         use.
--srppasswd FILE         SRP password file to use.
--srppasswdconf FILE     SRP password conf file to use.
--ciphers cipher1 cipher2...
                         Ciphers to enable.
--protocols protocol1 protocol2...
                         Protocols to enable.
--comp comp1 comp2...    Compression methods to enable.
--macs mac1 mac2...      MACs to enable.
--kx kx1 kx2...          Key exchange methods to enable.
--ctypes certType1 certType2...
                         Certificate types to enable.
-l, --list               Print a list of the supported
                         algorithms  and modes.
-h, --help               prints this help
-v, --version            prints the program's version number
--copyright              prints the program's license

8.5 Invoking certtool

This is a program to generate X.509 certificates, certificate requests, CRLs and private keys. The program can be used interactively or non interactively by specifying the --template command line option. See below for an example of a template file. How to use certtool interactively:

• To generate parameters for Diffie Hellman key exchange, use the command:

    $ certtool --generate-dh-params --outfile dh.pem

• To generate parameters for the RSA-EXPORT key exchange, use the command:

    $ certtool --generate-privkey --bits 512 --outfile rsa.pem

• To create a self signed certificate, use the command:

    $ certtool --generate-privkey --outfile ca-key.pem
    $ certtool --generate-self-signed --load-privkey ca-key.pem \
       --outfile ca-cert.pem

Note that a self-signed certificate usually belongs to a certificate authority, that signs other certificates.

• To create a private key, run:

    $ certtool --generate-privkey --outfile key.pem

• To create a certificate request, run:

    $ certtool --generate-request --load-privkey key.pem \
      --outfile request.pem

• To generate a certificate using the previous request, use the command:

    $ certtool --generate-certificate --load-request request.pem \
       --outfile cert.pem \
       --load-ca-certificate ca-cert.pem --load-ca-privkey ca-key.pem

• To view the certificate information, use:

    $ certtool --certificate-info --infile cert.pem

• To generate a PKCS #12 structure using the previous key and certificate, use the

command:

    $ certtool --load-certificate cert.pem --load-privkey key.pem \
      --to-p12 --outder --outfile key.p12

Certtool’s template file format:

• Firstly create a file named ’cert.cfg’ that contains the information about the certificate. An example file is listed below.

• Then execute:

    $ certtool --generate-certificate cert.pem --load-privkey key.pem  \
       --template cert.cfg \
       --load-ca-certificate ca-cert.pem --load-ca-privkey ca-key.pem

An example certtool template file:

# X.509 Certificate options
#
# DN options
# The organization of the subject.
organization = "Koko inc."
# The organizational unit of the subject.
unit = "sleeping dept."
# The locality of the subject.
# locality =
# The state of the certificate owner.
state = "Attiki"
# The country of the subject. Two letter code.
country = GR
# The common name of the certificate owner.
cn = "Cindy Lauper"
# A user id of the certificate owner.
#uid = "clauper"
# If the supported DN OIDs are not adequate you can set
# any OID here.
# For example set the X.520 Title and the X.520 Pseudonym
# by using OID and string pairs.
#dn_oid = "2.5.4.12" "Dr." "2.5.4.65" "jackal"
# This is deprecated and should not be used in new

# certificates.
# pkcs9_email = "none@none.org"
# The serial number of the certificate
serial = 007
# In how many days, counting from today, this certificate will expire.
expiration_days = 700
# X.509 v3 extensions
# A dnsname in case of a WWW server.
#dns_name = "www.none.org"
# An IP address in case of a server.
#ip_address = "192.168.1.1"
# An email in case of a person
email = "none@none.org"
# An URL that has CRLs (certificate revocation lists)
# available. Needed in CA certificates.
#crl_dist_points = "http://www.getcrl.crl/getcrl/"
# Whether this is a CA certificate or not
#ca
# Whether this certificate will be used for a TLS client
#tls_www_client
# Whether this certificate will be used for a TLS server
#tls_www_server
# Whether this certificate will be used to sign data (needed
# in TLS DHE ciphersuites).
signing_key
# Whether this certificate will be used to encrypt data (needed
# in TLS RSA ciphersuites). Note that it is prefered to use different
# keys for encryption and signing.
#encryption_key
# Whether this key will be used to sign other certificates.
#cert_signing_key
# Whether this key will be used to sign CRLs.
#crl_signing_key

# Whether this key will be used to sign code.
#code_signing_key
# Whether this key will be used to sign OCSP data.
#ocsp_signing_key
# Whether this key will be used for time stamping.
#time_stamping_key

9 Function reference

9.1 Core functions

The prototypes for the following functions lie in ‘gnutls/gnutls.h’.

gnutls alert get name

[Function] const char * gnutls_alert_get_name (gnutls alert description t

   alert)

alert: is an alert number gnutls_session_t structure. This function will return a string that describes the given alert number or NULL. See

gnutls_alert_get().

gnutls alert get

[Function] gnutls_alert_description_t gnutls_alert_get (gnutls session t
   session)

session: is a gnutls_session_t structure. This function will return the last alert number received. This function should be called if GNUTLS E WARNING ALERT RECEIVED or GNUTLS E FATAL ALERT RECEIVED has been returned by a gnutls function. The peer may send alerts if he thinks some things were not right. Check gnutls.h for the available alert descriptions. If no alert has been received the returned value is undefined.

gnutls alert send appropriate

[Function] int gnutls_alert_send_appropriate (gnutls session t session, int
   err)

session: is a gnutls_session_t structure. err: is an integer Sends an alert to the peer depending on the error code returned by a gnutls function. This function will call gnutls_error_to_alert() to determine the appropriate alert to send. This function may also return GNUTLS E AGAIN, or GNUTLS E INTERRUPTED. If the return value is GNUTLS E INVALID REQUEST, then no alert has been sent to the peer. Returns zero on success.

gnutls alert send

[Function] int gnutls_alert_send (gnutls session t session, gnutls alert level t

level, gnutls alert description t desc) session: is a gnutls_session_t structure. level: is the level of the alert desc: is the alert description

This function will send an alert to the peer in order to inform him of something important (eg. his Certificate could not be verified). If the alert level is Fatal then the peer is expected to close the connection, otherwise he may ignore the alert and continue.

The error code of the underlying record send function will be returned, so you may also receive GNUTLS E INTERRUPTED or GNUTLS E AGAIN as well.

Returns 0 on success.

gnutls anon allocate client credentials

[Function] int gnutls_anon_allocate_client_credentials

(gnutls anon client credentials t * sc)

sc: is a pointer to an gnutls_anon_client_credentials_t structure.

This structure is complex enough to manipulate directly thus this helper function is provided in order to allocate it.

Returns a negative value in case of an error.

gnutls anon allocate server credentials

[Function] int gnutls_anon_allocate_server_credentials

(gnutls anon server credentials t * sc)

sc: is a pointer to an gnutls_anon_server_credentials_t structure.

This structure is complex enough to manipulate directly thus this helper function is provided in order to allocate it.

Returns a negative value in case of an error.

gnutls anon free client credentials

[Function] void gnutls_anon_free_client_credentials

(gnutls anon client credentials t sc)

sc: is an gnutls_anon_client_credentials_t structure.

This structure is complex enough to manipulate directly thus this helper function is provided in order to free (deallocate) it.

gnutls anon free server credentials

[Function] void gnutls_anon_free_server_credentials

(gnutls anon server credentials t sc)

sc: is an gnutls_anon_server_credentials_t structure.

This structure is complex enough to manipulate directly thus this helper function is provided in order to free (deallocate) it.

gnutls anon set params function

[Function] void gnutls_anon_set_params_function

(gnutls anon server credentials t res, gnutls params function * func) res: is a gnutls anon server credentials t structure func: is the function to be called This function will set a callback in order for the server to get the diffie hellman or RSA parameters for anonymous authentication. The callback should return zero on success.

gnutls anon set server dh params

[Function] void gnutls_anon_set_server_dh_params

(gnutls anon server credentials t res, gnutls dh params t dh_params) res: is a gnutls anon server credentials t structure dh params: is a structure that holds diffie hellman parameters. This function will set the diffie hellman parameters for an anonymous server to use. These parameters will be used in Anonymous Diffie Hellman cipher suites.

gnutls anon set server params function

[Function] void gnutls_anon_set_server_params_function

(gnutls anon server credentials t res, gnutls params function * func) res: is a gnutls certificate credentials t structure func: is the function to be called This function will set a callback in order for the server to get the diffie hellman parameters for anonymous authentication. The callback should return zero on success.

gnutls auth client get type

[Function] gnutls_credentials_type_t gnutls_auth_client_get_type

(gnutls session t session) session: is a gnutls_session_t structure. Returns the type of credentials that were used for client authentication. The returned information is to be used to distinguish the function used to access authentication data.

gnutls auth get type

[Function] gnutls_credentials_type_t gnutls_auth_get_type

(gnutls session t session) session: is a gnutls_session_t structure. Returns type of credentials for the current authentication schema. The returned information is to be used to distinguish the function used to access authentication data. Eg. for CERTIFICATE ciphersuites (key exchange algorithms: KX RSA, KX DHE RSA), the same function are to be used to access the authentication data.

gnutls auth server get type

[Function] gnutls_credentials_type_t gnutls_auth_server_get_type

(gnutls session t session) session: is a gnutls_session_t structure. Returns the type of credentials that were used for server authentication. The returned information is to be used to distinguish the function used to access authentication data.

gnutls bye

[Function] int gnutls_bye (gnutls session t session, gnutls close request t how)

session: is a gnutls_session_t structure. how: is an integer Terminates the current TLS/SSL connection. The connection should have been initiated using gnutls_handshake(). how should be one of GNUTLS SHUT RDWR, GNUTLS SHUT WR. In case of GNUTLS SHUT RDWR then the TLS connection gets terminated and further receives and sends will be disallowed. If the return value is zero you may continue using the connection. GNUTLS SHUT RDWR actually sends an alert containing a close request and waits for the peer to reply with the same message. In case of GNUTLS SHUT WR then the TLS connection gets terminated and further sends will be disallowed. In order to reuse the connection you should wait for an EOF from the peer. GNUTLS SHUT WR sends an alert containing a close request. Note that not all implementations will properly terminate a TLS connection. Some of them, usually for performance reasons, will terminate only the underlying transport layer, thus causing a transmission error to the peer. This error cannot be distinguished from a malicious party prematurely terminating the session, thus this behavior is not recommended. This function may also return GNUTLS E AGAIN or GNUTLS E INTERRUPTED;

cf. gnutls_record_get_direction().

gnutls certificate activation time peers

[Function] time_t gnutls_certificate_activation_time_peers

(gnutls session t session) session: is a gnutls session This function will return the peer’s certificate activation time. This is the creation time for openpgp keys. Returns (time t) -1 on error.

gnutls certificate allocate credentials

[Function] int gnutls_certificate_allocate_credentials

(gnutls certificate credentials t * res)

res: is a pointer to an gnutls_certificate_credentials_t structure.

This structure is complex enough to manipulate directly thus this helper function is provided in order to allocate it.

Returns 0 on success.

gnutls certificate client get request status

[Function] int gnutls_certificate_client_get_request_status

(gnutls session t session) session: is a gnutls session

This function will return 0 if the peer (server) did not request client authentication or 1 otherwise. Returns a negative value in case of an error.

gnutls certificate client set retrieve function

[Function] void gnutls_certificate_client_set_retrieve_function

(gnutls certificate credentials t cred, gnutls certificate client retrieve function

   * func)
cred: is a gnutls_certificate_credentials_t structure.

func: is the callback function

This function sets a callback to be called in order to retrieve the certificate to be used in the handshake. The callback’s function prototype is: int (*callback)(gnutls session t, const gnutls datum t* req ca dn, int nreqs, gnutls pk algorithm t* pk algos, int pk algos length, gnutls retr st* st);

st should contain the certificates and private keys.

req_ca_cert, is only used in X.509 certificates. Contains a list with the CA names that the server considers trusted. Normally we should send a certificate that is signed by one of these CAs. These names are DER encoded. To get a more meaningful value

use the function gnutls_x509_rdn_get().

pk_algos, contains a list with server’s acceptable signature algorithms. The certificate returned should support the server’s given algorithms.

If the callback function is provided then gnutls will call it, in the handshake, after the certificate request message has been received.

The callback function should set the certificate list to be sent, and return 0 on success. If no certificate was selected then the number of certificates should be set to zero. The value (-1) indicates error and the handshake will be terminated.

gnutls certificate expiration time peers

[Function] time_t gnutls_certificate_expiration_time_peers

(gnutls session t session) session: is a gnutls session

This function will return the peer’s certificate expiration time.

Returns (time t) -1 on error.

gnutls certificate free ca names

[Function] void gnutls_certificate_free_ca_names

(gnutls certificate credentials t sc)

sc: is an gnutls_certificate_credentials_t structure.

This function will delete all the CA name in the given credentials. Clients may call this to save some memory since in client side the CA names are not used.

CA names are used by servers to advertize the CAs they support to clients.

gnutls certificate free cas

[Function] void gnutls_certificate_free_cas (gnutls certificate credentials t

   sc)
sc: is an gnutls_certificate_credentials_t structure.

This function will delete all the CAs associated with the given credentials. Servers

that do not use gnutls_certificate_verify_peers2() may call this to save some

memory.

gnutls certificate free credentials

[Function] void gnutls_certificate_free_credentials

(gnutls certificate credentials t sc)

sc: is an gnutls_certificate_credentials_t structure.

This structure is complex enough to manipulate directly thus this helper function is provided in order to free (deallocate) it.

This function does not free any temporary parameters associated with this structure (ie RSA and DH parameters are not freed by this function).

gnutls certificate free crls

[Function] void gnutls_certificate_free_crls (gnutls certificate credentials t

   sc)
sc: is an gnutls_certificate_credentials_t structure.

This function will delete all the CRLs associated with the given credentials.

gnutls certificate free keys

[Function] void gnutls_certificate_free_keys (gnutls certificate credentials t

   sc)
sc: is an gnutls_certificate_credentials_t structure.

This function will delete all the keys and the certificates associated with the given credentials. This function must not be called when a TLS negotiation that uses the credentials is in progress.

gnutls certificate get ours

[Function] const gnutls_datum_t * gnutls_certificate_get_ours

(gnutls session t session) session: is a gnutls session This function will return the certificate as sent to the peer, in the last handshake. These certificates are in raw format. In X.509 this is a certificate list. In OpenPGP this is a single certificate. Returns NULL in case of an error, or if no certificate was used.

gnutls certificate get peers

[Function] const gnutls_datum_t * gnutls_certificate_get_peers

(gnutls session t session, unsigned int * list_size) session: is a gnutls session list size: is the length of the certificate list This function will return the peer’s raw certificate (chain) as sent by the peer. These certificates are in raw format (DER encoded for X.509). In case of a X.509 then a certificate list may be present. The first certificate in the list is the peer’s certificate, following the issuer’s certificate, then the issuer’s issuer etc. In case of OpenPGP keys a single key will be returned in raw format. Returns NULL in case of an error, or if no certificate was sent.

gnutls certificate send x509 rdn sequence

[Function] void gnutls_certificate_send_x509_rdn_sequence

(gnutls session t session, int status) session: is a pointer to a gnutls_session_t structure. status: is 0 or 1 If status is non zero, this function will order gnutls not to send the rdnSequence in the certificate request message. That is the server will not advertize it’s trusted CAs to the peer. If status is zero then the default behaviour will take effect, which is to advertize the server’s trusted CAs. This function has no effect in clients, and in authentication methods other than certificate with X.509 certificates.

gnutls certificate server set request

[Function] void gnutls_certificate_server_set_request (gnutls session t

session, gnutls certificate request t req) session: is an gnutls_session_t structure. req: is one of GNUTLS CERT REQUEST, GNUTLS CERT REQUIRE This function specifies if we (in case of a server) are going to send a certificate request message to the client. If req is GNUTLS CERT REQUIRE then the server will return an error if the peer does not provide a certificate. If you do not call this function then the client will not be asked to send a certificate.

gnutls certificate server set retrieve function

[Function] void gnutls_certificate_server_set_retrieve_function

(gnutls certificate credentials t cred, gnutls certificate server retrieve function

   * func)
cred: is a gnutls_certificate_credentials_t structure.

func: is the callback function This function sets a callback to be called in order to retrieve the certificate to be used in the handshake. The callback’s function prototype is: int (*callback)(gnutls session t, gnutls retr st* st); st should contain the certificates and private keys. If the callback function is provided then gnutls will call it, in the handshake, after the certificate request message has been received. The callback function should set the certificate list to be sent, and return 0 on success. The value (-1) indicates error and the handshake will be terminated.

gnutls certificate set dh params

[Function] void gnutls_certificate_set_dh_params

(gnutls certificate credentials t res, gnutls dh params t dh_params) res: is a gnutls certificate credentials t structure dh params: is a structure that holds diffie hellman parameters. This function will set the diffie hellman parameters for a certificate server to use. These parameters will be used in Ephemeral Diffie Hellman cipher suites. Note that only a pointer to the parameters are stored in the certificate handle, so if you deallocate the parameters before the certificate is deallocated, you must change the parameters stored in the certificate first.

gnutls certificate set params function

[Function] void gnutls_certificate_set_params_function

(gnutls certificate credentials t res, gnutls params function * func) res: is a gnutls certificate credentials t structure func: is the function to be called This function will set a callback in order for the server to get the diffie hellman or RSA parameters for certificate authentication. The callback should return zero on success.

gnutls certificate set rsa export params

[Function] void gnutls_certificate_set_rsa_export_params

(gnutls certificate credentials t res, gnutls rsa params t rsa_params) res: is a gnutls certificate credentials t structure rsa params: is a structure that holds temporary RSA parameters. This function will set the temporary RSA parameters for a certificate server to use. These parameters will be used in RSA-EXPORT cipher suites.

gnutls certificate set verify flags

[Function] void gnutls_certificate_set_verify_flags

(gnutls certificate credentials t res, unsigned int flags) res: is a gnutls certificate credentials t structure flags: are the flags This function will set the flags to be used at verification of the certificates. Flags

must be OR of the gnutls_certificate_verify_flags enumerations.

gnutls certificate set verify limits

[Function] void gnutls_certificate_set_verify_limits

(gnutls certificate credentials t res, unsigned int max_bits, unsigned int

   max_depth)

res: is a gnutls certificate credentials structure max bits: is the number of bits of an acceptable certificate (default 8200) max depth: is maximum depth of the verification of a certificate chain (default 5) This function will set some upper limits for the default verification function, gnutls_ certificate_verify_peers2(), to avoid denial of service attacks.

gnutls certificate set x509 crl file

[Function] int gnutls_certificate_set_x509_crl_file

(gnutls certificate credentials t res, const char * crlfile,

gnutls x509 crt fmt t type)

res: is an gnutls_certificate_credentials_t structure.

crlfile: is a file containing the list of verified CRLs (DER or PEM list) type: is PEM or DER This function adds the trusted CRLs in order to verify client or server certificates. In case of a client this is not required to be called if the certificates are not verified using gnutls_certificate_verify_peers2(). This function may be called multiple times. Returns the number of CRLs processed or a negative value on error.

gnutls certificate set x509 crl mem

[Function] int gnutls_certificate_set_x509_crl_mem

(gnutls certificate credentials t res, const gnutls datum t * CRL, gnutls x509 crt fmt t type)

res: is an gnutls_certificate_credentials_t structure.

CRL: is a list of trusted CRLs. They should have been verified before. type: is DER or PEM This function adds the trusted CRLs in order to verify client or server certificates. In case of a client this is not required to be called if the certificates are not verified using gnutls_certificate_verify_peers2(). This function may be called multiple times. Returns the number of CRLs processed or a negative value on error.

gnutls certificate set x509 crl

[Function] int gnutls_certificate_set_x509_crl

(gnutls certificate credentials t res, gnutls x509 crl t * crl_list, int

   crl_list_size)
res: is an gnutls_certificate_credentials_t structure.

crl list: is a list of trusted CRLs. They should have been verified before. crl list size: holds the size of the crl list This function adds the trusted CRLs in order to verify client or server certificates. In case of a client this is not required to be called if the certificates are not verified using gnutls_certificate_verify_peers2(). This function may be called multiple times. Returns 0 on success.

gnutls certificate set x509 key file

[Function] int gnutls_certificate_set_x509_key_file

(gnutls certificate credentials t res, const char * CERTFILE, const char * KEYFILE, gnutls x509 crt fmt t type)

res: is an gnutls_certificate_credentials_t structure.

CERTFILE: is a file that containing the certificate list (path) for the specified private key, in PKCS7 format, or a list of certificates KEYFILE: is a file that contains the private key type: is PEM or DER This function sets a certificate/private key pair in the gnutls certificate credentials t structure. This function may be called more than once (in case multiple keys/certificates exist for the server). Currently only PKCS-1 encoded RSA and DSA private keys are accepted by this function.

gnutls certificate set x509 key mem

[Function] int gnutls_certificate_set_x509_key_mem

(gnutls certificate credentials t res, const gnutls datum t * cert, const gnutls datum t * key, gnutls x509 crt fmt t type)

res: is an gnutls_certificate_credentials_t structure.

cert: contains a certificate list (path) for the specified private key key: is the private key type: is PEM or DER This function sets a certificate/private key pair in the gnutls certificate credentials t structure. This function may be called more than once (in case multiple keys/certificates exist for the server). Currently are supported: RSA PKCS-1 encoded private keys, DSA private keys. DSA private keys are encoded the OpenSSL way, which is an ASN.1 DER sequence of 6 INTEGERs - version, p, q, g, pub, priv.

Note that the keyUsage (2.5.29.15) PKIX extension in X.509 certificates is supported. This means that certificates intended for signing cannot be used for ciphersuites that require encryption.

If the certificate and the private key are given in PEM encoding then the strings that hold their values must be null terminated.

gnutls certificate set x509 key

[Function] int gnutls_certificate_set_x509_key

(gnutls certificate credentials t res, gnutls x509 crt t * cert_list, int cert_list_size, gnutls x509 privkey t key)

res: is an gnutls_certificate_credentials_t structure.

cert list: contains a certificate list (path) for the specified private key

cert list size: holds the size of the certificate list

key: is a gnutls x509 privkey t key

This function sets a certificate/private key pair in the gnutls certificate credentials t structure. This function may be called more than once (in case multiple keys/certificates exist for the server).

gnutls certificate set x509 simple pkcs12 file

[Function] int gnutls_certificate_set_x509_simple_pkcs12_file

(gnutls certificate credentials t res, const char * pkcs12file, gnutls x509 crt fmt t type, const char * password)

res: is an gnutls_certificate_credentials_t structure.

pkcs12file: filename of file containing PKCS12 blob.

type: is PEM or DER of the pkcs12file.

password: optional password used to decrypt PKCS12 file, bags and keys.

This function sets a certificate/private key pair and/or a CRL in the gnutls certificate credentials t structure. This function may be called more than once (in case multiple keys/certificates exist for the server).

MAC: ed PKCS12 files are supported. Encrypted PKCS12 bags are supported. Encrypted PKCS8 private keys are supported. However, only password based security, and the same password for all operations, are supported.

The private keys may be RSA PKCS1 or DSA private keys encoded in the OpenSSL way.

PKCS12 file may contain many keys and/or certificates, and there is no way to identify which key/certificate pair you want. You should make sure the PKCS12 file only contain one key/certificate pair and/or one CRL.

It is believed that the limitations of this function is acceptable for most usage, and that any more flexibility would introduce complexity that would make it harder to use this functionality at all.

Return value: Returns 0 on success, or an error code.

gnutls certificate set x509 trust file

[Function] int gnutls_certificate_set_x509_trust_file

(gnutls certificate credentials t res, const char * cafile,

gnutls x509 crt fmt t type)

res: is an gnutls_certificate_credentials_t structure.

cafile: is a file containing the list of trusted CAs (DER or PEM list) type: is PEM or DER This function adds the trusted CAs in order to verify client or server certificates. In case of a client this is not required to be called if the certificates are not verified using gnutls_certificate_verify_peers2(). This function may be called multiple times. In case of a server the names of the CAs set here will be sent to the client if a certificate request is sent. This can be disabled using gnutls_certificate_send_x509_rdn_

sequence().

Returns the number of certificates processed or a negative value on error.

gnutls certificate set x509 trust mem

[Function] int gnutls_certificate_set_x509_trust_mem

(gnutls certificate credentials t res, const gnutls datum t * ca, gnutls x509 crt fmt t type)

res: is an gnutls_certificate_credentials_t structure.

ca: is a list of trusted CAs or a DER certificate type: is DER or PEM This function adds the trusted CAs in order to verify client or server certificates. In case of a client this is not required to be called if the certificates are not verified using gnutls_certificate_verify_peers2(). This function may be called multiple times. In case of a server the CAs set here will be sent to the client if a certificate request is

sent. This can be disabled using gnutls_certificate_send_x509_rdn_sequence().

Returns the number of certificates processed or a negative value on error.

gnutls certificate set x509 trust

[Function] int gnutls_certificate_set_x509_trust

(gnutls certificate credentials t res, gnutls x509 crt t * ca_list, int

   ca_list_size)
res: is an gnutls_certificate_credentials_t structure.

ca list: is a list of trusted CAs ca list size: holds the size of the CA list This function adds the trusted CAs in order to verify client or server certificates. In case of a client this is not required to be called if the certificates are not verified using gnutls_certificate_verify_peers2(). This function may be called multiple times.

In case of a server the CAs set here will be sent to the client if a certificate request is

sent. This can be disabled using gnutls_certificate_send_x509_rdn_sequence().

Returns 0 on success.

gnutls certificate type get name

[Function] const char * gnutls_certificate_type_get_name

(gnutls certificate type t type) type: is a certificate type Returns a string (or NULL) that contains the name of the specified certificate type.

gnutls certificate type get

[Function] gnutls_certificate_type_t gnutls_certificate_type_get

(gnutls session t session) session: is a gnutls_session_t structure. Returns the currently used certificate type. The certificate type is by default X.509, unless it is negotiated as a TLS extension.

gnutls certificate type set priority

[Function] int gnutls_certificate_type_set_priority (gnutls session t
   session, const int * list)

session: is a gnutls_session_t structure. list: is a 0 terminated list of gnutls certificate type t elements. Sets the priority on the certificate types supported by gnutls. Priority is higher for types specified before others. After specifying the types you want, you must append a 0. Note that the certificate type priority is set on the client. The server does not use the cert type priority except for disabling types that were not specified. Returns 0 on success.

gnutls certificate verify peers2

[Function] int gnutls_certificate_verify_peers2 (gnutls session t session,

unsigned int * status) session: is a gnutls session status: is the output of the verification This function will try to verify the peer’s certificate and return its status (trusted, invalid etc.). The value of status should be one or more of the gnutls certificate status t enumerated elements bitwise or’d. To avoid denial of service attacks some default upper limits regarding the certificate key size and chain

size are set. To override them use gnutls_certificate_set_verify_limits().

Note that you must also check the peer’s name in order to check if the verified certificate belongs to the actual peer. Returns a negative error code on error and zero on success. This is the same as gnutls_x509_verify_certificate() and uses the loaded CAs in the credentials as trusted CAs.

Note that some commonly used X.509 Certificate Authorities are still using Version 1 certificates. If you want to accept them, you need to call gnutls_certificate_set_

verify_flags() with, e.g., GNUTLS_VERIFY_ALLOW_X509_V1_CA_CRT parameter.

gnutls certificate verify peers

[Function] int gnutls_certificate_verify_peers (gnutls session t session)

session: is a gnutls session

This function will try to verify the peer’s certificate and return its status (trusted, invalid etc.). However you must also check the peer’s name in order to check if the verified certificate belongs to the actual peer.

The return value should be one or more of the gnutls certificate status t enumerated elements bitwise or’d, or a negative value on error.

This is the same as gnutls_x509_verify_certificate().
Deprecated: Use gnutls_certificate_verify_peers2() instead.

gnutls check version

[Function] const char * gnutls_check_version (const char * req_version)

req version: the version to check

Check that the version of the library is at minimum the requested one and return the version string; return NULL if the condition is not satisfied. If a NULL is passed to this function, no check is done, but the version string is simply returned.

gnutls cipher get key size

[Function] size_t gnutls_cipher_get_key_size (gnutls cipher algorithm t

   algorithm)

algorithm: is an encryption algorithm

Returns the length (in bytes) of the given cipher’s key size. Returns 0 if the given cipher is invalid.

gnutls cipher get name

[Function] const char * gnutls_cipher_get_name (gnutls cipher algorithm t

   algorithm)

algorithm: is an encryption algorithm

Returns a pointer to a string that contains the name of the specified cipher or NULL.

gnutls cipher get

[Function] gnutls_cipher_algorithm_t gnutls_cipher_get (gnutls session t
   session)

session: is a gnutls_session_t structure.

Returns the currently used cipher.

gnutls cipher set priority

[Function] int gnutls_cipher_set_priority (gnutls session t session, const

int * list) session: is a gnutls_session_t structure. list: is a 0 terminated list of gnutls cipher algorithm t elements. Sets the priority on the ciphers supported by gnutls. Priority is higher for ciphers specified before others. After specifying the ciphers you want, you must append a 0. Note that the priority is set on the client. The server does not use the algorithm’s priority except for disabling algorithms that were not specified. Returns 0 on success.

gnutls cipher suite get name

[Function] const char * gnutls_cipher_suite_get_name

(gnutls kx algorithm t kx_algorithm, gnutls cipher algorithm t

   cipher_algorithm, gnutls mac algorithm t mac_algorithm)

kx algorithm: is a Key exchange algorithm cipher algorithm: is a cipher algorithm mac algorithm: is a MAC algorithm Returns a string that contains the name of a TLS cipher suite, specified by the given algorithms, or NULL. Note that the full cipher suite name must be prepended by TLS or SSL depending of the protocol in use.

gnutls compression get name

[Function] const char * gnutls_compression_get_name

(gnutls compression method t algorithm) algorithm: is a Compression algorithm Returns a pointer to a string that contains the name of the specified compression algorithm or NULL.

gnutls compression get

[Function] gnutls_compression_method_t gnutls_compression_get

(gnutls session t session) session: is a gnutls_session_t structure. Returns the currently used compression method.

gnutls compression set priority

[Function] int gnutls_compression_set_priority (gnutls session t session,

const int * list) session: is a gnutls_session_t structure. list: is a 0 terminated list of gnutls compression method t elements.

Sets the priority on the compression algorithms supported by gnutls. Priority is higher for algorithms specified before others. After specifying the algorithms you want, you must append a 0. Note that the priority is set on the client. The server does not use the algorithm’s priority except for disabling algorithms that were not specified.

TLS 1.0 does not define any compression algorithms except NULL. Other compression algorithms are to be considered as gnutls extensions.

Returns 0 on success.

gnutls credentials clear

[Function] void gnutls_credentials_clear (gnutls session t session)

session: is a gnutls_session_t structure.

Clears all the credentials previously set in this session.

gnutls credentials set

[Function] int gnutls_credentials_set (gnutls session t session,

gnutls credentials type t type, void * cred) session: is a gnutls_session_t structure.

type: is the type of the credentials

cred: is a pointer to a structure.

Sets the needed credentials for the specified type. Eg username, password - or public and private keys etc. The (void* cred) parameter is a structure that depends on the specified type and on the current session (client or server). [ In order to minimize memory usage, and share credentials between several threads gnutls keeps a pointer to cred, and not the whole cred structure. Thus you will have to keep the structure allocated until you call gnutls_deinit(). ]

For GNUTLS CRD ANON cred should be gnutls anon client credentials t in case of a client. In case of a server it should be gnutls anon server credentials t.

For GNUTLS CRD SRP cred should be gnutls srp client credentials t in case of a client, and gnutls srp server credentials t, in case of a server.

For GNUTLS CRD CERTIFICATE cred should be gnutls certificate credentials t.

gnutls db check entry

[Function] int gnutls_db_check_entry (gnutls session t session,
   gnutls datum t session_entry)

session: is a gnutls_session_t structure.

session entry: is the session data (not key)

This function returns GNUTLS E EXPIRED, if the database entry has expired or 0 otherwise. This function is to be used when you want to clear unnesessary session which occupy space in your backend.

gnutls db get ptr

[Function] void * gnutls_db_get_ptr (gnutls session t session)

session: is a gnutls_session_t structure.

Returns the pointer that will be sent to db store, retrieve and delete functions, as the first argument.

gnutls db remove session

[Function] void gnutls_db_remove_session (gnutls session t session)

session: is a gnutls_session_t structure.

This function will remove the current session data from the session database. This will prevent future handshakes reusing these session data. This function should be called if a session was terminated abnormally, and before gnutls_deinit() is called.

Normally gnutls_deinit() will remove abnormally terminated sessions.

gnutls db set cache expiration

[Function] void gnutls_db_set_cache_expiration (gnutls session t session,
   int seconds)

session: is a gnutls_session_t structure.

seconds: is the number of seconds.

Sets the expiration time for resumed sessions. The default is 3600 (one hour) at the time writing this.

gnutls db set ptr

[Function] void gnutls_db_set_ptr (gnutls session t session, void * ptr)

session: is a gnutls_session_t structure.

ptr: is the pointer

Sets the pointer that will be provided to db store, retrieve and delete functions, as the first argument.

gnutls db set remove function

[Function] void gnutls_db_set_remove_function (gnutls session t session,

gnutls db remove func rem_func) session: is a gnutls_session_t structure.

rem func: is the function.

Sets the function that will be used to remove data from the resumed sessions database. This function must return 0 on success.

The first argument to rem_func() will be null unless gnutls_db_set_ptr() has been called.

gnutls db set retrieve function

[Function] void gnutls_db_set_retrieve_function (gnutls session t session,

gnutls db retr func retr_func) session: is a gnutls_session_t structure. retr func: is the function. Sets the function that will be used to retrieve data from the resumed sessions database. This function must return a gnutls datum t containing the data on success, or a gnutls datum t containing null and 0 on failure. The datum’s data must be allocated using the function gnutls_malloc(). The first argument to retr_func() will be null unless gnutls_db_set_ptr() has been called.

gnutls db set store function

[Function] void gnutls_db_set_store_function (gnutls session t session,

gnutls db store func store_func) session: is a gnutls_session_t structure. store func: is the function Sets the function that will be used to store data from the resumed sessions database. This function must remove 0 on success. The first argument to store_func() will be null unless gnutls_db_set_ptr() has been called.

gnutls deinit

[Function] void gnutls_deinit (gnutls session t session)

session: is a gnutls_session_t structure. This function clears all buffers associated with the session. This function will also remove session data from the session database if the session was terminated abnormally.

gnutls dh get group

[Function] int gnutls_dh_get_group (gnutls session t session, gnutls datum t *

   raw_gen, gnutls datum t * raw_prime)

session: is a gnutls session raw gen: will hold the generator. raw prime: will hold the prime. This function will return the group parameters used in the last Diffie Hellman authentication with the peer. These are the prime and the generator used. This function should be used for both anonymous and ephemeral diffie Hellman. The output parameters must be freed with gnutls_free(). Returns a negative value in case of an error.

gnutls dh get peers public bits

[Function] int gnutls_dh_get_peers_public_bits (gnutls session t session)

session: is a gnutls session This function will return the bits used in the last Diffie Hellman authentication with the peer. Should be used for both anonymous and ephemeral diffie Hellman. Returns a negative value in case of an error.

gnutls dh get prime bits

[Function] int gnutls_dh_get_prime_bits (gnutls session t session)

session: is a gnutls session This function will return the bits of the prime used in the last Diffie Hellman authentication with the peer. Should be used for both anonymous and ephemeral diffie Hellman. Returns a negative value in case of an error.

gnutls dh get pubkey

[Function] int gnutls_dh_get_pubkey (gnutls session t session, gnutls datum t

   * raw_key)

session: is a gnutls session raw key: will hold the public key. This function will return the peer’s public key used in the last Diffie Hellman authentication. This function should be used for both anonymous and ephemeral diffie Hellman. The output parameters must be freed with gnutls_free(). Returns a negative value in case of an error.

gnutls dh get secret bits

[Function] int gnutls_dh_get_secret_bits (gnutls session t session)

session: is a gnutls session This function will return the bits used in the last Diffie Hellman authentication with the peer. Should be used for both anonymous and ephemeral diffie Hellman. Returns a negative value in case of an error.

gnutls dh params cpy

[Function] int gnutls_dh_params_cpy (gnutls dh params t dst,

gnutls dh params t src) dst: Is the destination structure, which should be initialized. src: Is the source structure This function will copy the DH parameters structure from source to destination.

gnutls dh params deinit

[Function] void gnutls_dh_params_deinit (gnutls dh params t dh_params)

dh params: Is a structure that holds the prime numbers This function will deinitialize the DH parameters structure.

gnutls dh params export pkcs3

[Function] int gnutls_dh_params_export_pkcs3 (gnutls dh params t params,

gnutls x509 crt fmt t format, unsigned char * params_data, size t *

   params_data_size)

params: Holds the DH parameters

format: the format of output params. One of PEM or DER.

params data: will contain a PKCS3 DHParams structure PEM or DER encoded

params data size: holds the size of params data (and will be replaced by the actual size of parameters)

This function will export the given dh parameters to a PKCS3 DHParams structure. This is the format generated by "openssl dhparam" tool. If the buffer provided is not long enough to hold the output, then GNUTLS E SHORT MEMORY BUFFER will be returned.

If the structure is PEM encoded, it will have a header of "BEGIN DH PARAME-TERS".

In case of failure a negative value will be returned, and 0 on success.

gnutls dh params export raw

[Function] int gnutls_dh_params_export_raw (gnutls dh params t params,

gnutls datum t * prime, gnutls datum t * generator, unsigned int * bits) params: Holds the DH parameters

prime: will hold the new prime

generator: will hold the new generator

bits: if non null will hold is the prime’s number of bits

This function will export the pair of prime and generator for use in the Diffie-Hellman key exchange. The new parameters will be allocated using gnutls_malloc() and will be stored in the appropriate datum.

gnutls dh params generate2

[Function] int gnutls_dh_params_generate2 (gnutls dh params t params,

unsigned int bits) params: Is the structure that the DH parameters will be stored

bits: is the prime’s number of bits

This function will generate a new pair of prime and generator for use in the Diffie-Hellman key exchange. The new parameters will be allocated using gnutls_malloc() and will be stored in the appropriate datum. This function is normally slow.

Note that the bits value should be one of 768, 1024, 2048, 3072 or 4096. Also note that the DH parameters are only useful to servers. Since clients use the parameters sent by the server, it’s of no use to call this in client side.

gnutls dh params import pkcs3

[Function] int gnutls_dh_params_import_pkcs3 (gnutls dh params t params,

const gnutls datum t * pkcs3_params, gnutls x509 crt fmt t format) params: A structure where the parameters will be copied to

pkcs3 params: should contain a PKCS3 DHParams structure PEM or DER encoded

format: the format of params. PEM or DER.

This function will extract the DHParams found in a PKCS3 formatted structure. This is the format generated by "openssl dhparam" tool.

If the structure is PEM encoded, it should have a header of "BEGIN DH PARAME-TERS".

In case of failure a negative value will be returned, and 0 on success.

gnutls dh params import raw

[Function] int gnutls_dh_params_import_raw (gnutls dh params t dh_params,

const gnutls datum t * prime, const gnutls datum t * generator) dh params: Is a structure that will hold the prime numbers

prime: holds the new prime

generator: holds the new generator

This function will replace the pair of prime and generator for use in the Diffie-Hellman key exchange. The new parameters should be stored in the appropriate gnutls datum.

gnutls dh params init

[Function] int gnutls_dh_params_init (gnutls dh params t * dh_params)

dh params: Is a structure that will hold the prime numbers

This function will initialize the DH parameters structure.

gnutls dh set prime bits

[Function] void gnutls_dh_set_prime_bits (gnutls session t session, unsigned
   int bits)

session: is a gnutls_session_t structure.

bits: is the number of bits

This function sets the number of bits, for use in an Diffie Hellman key exchange. This is used both in DH ephemeral and DH anonymous cipher suites. This will set the minimum size of the prime that will be used for the handshake.

In the client side it sets the minimum accepted number of bits. If a server sends a prime with less bits than that GNUTLS E DH PRIME UNACCEPTABLE will be returned by the handshake.

gnutls error is fatal

[Function] int gnutls_error_is_fatal (int error)

error: is an error returned by a gnutls function. Error should be a negative value. If a function returns a negative value you may feed that value to this function to see if it is fatal. Returns 1 for a fatal error 0 otherwise. However you may want to check the error code manually, since some non-fatal errors to the protocol may be fatal for you (your program). This is only useful if you are dealing with errors from the record layer or the handshake layer.

gnutls error to alert

[Function] int gnutls_error_to_alert (int err, int * level)

err: is a negative integer level: the alert level will be stored there Returns an alert depending on the error code returned by a gnutls function. All alerts sent by this function should be considered fatal. The only exception is when err == GNUTLS E REHANDSHAKE, where a warning alert should be sent to the peer indicating that no renegotiation will be performed. If the return value is GNUTLS E INVALID REQUEST, then there was no mapping to an alert.

gnutls fingerprint

[Function] int gnutls_fingerprint (gnutls digest algorithm t algo, const

gnutls datum t * data, void * result, size t * result_size) algo: is a digest algorithm data: is the data result: is the place where the result will be copied (may be null). result size: should hold the size of the result. The actual size of the returned result will also be copied there. This function will calculate a fingerprint (actually a hash), of the given data. The result is not printable data. You should convert it to hex, or to something else printable. This is the usual way to calculate a fingerprint of an X.509 DER encoded certificate. Note however that the fingerprint of an OpenPGP is not just a hash and cannot be calculated with this function. Returns a negative value in case of an error.

gnutls free

[Function] void gnutls_free (void * ptr)

This function will free data pointed by ptr. The deallocation function used is the one set by gnutls_global_set_mem_

functions().

gnutls global deinit

[Function] void gnutls_global_deinit ( void)

This function deinitializes the global data, that were initialized using gnutls_global_

init().

Note! This function is not thread safe. See the discussion for gnutls_global_init() for more information.

gnutls global init

[Function] int gnutls_global_init ( void)

This function initializes the global data to defaults. Every gnutls application has a global data which holds common parameters shared by gnutls session structures. You must call gnutls_global_deinit() when gnutls usage is no longer needed Returns zero on success. Note that this function will also initialize libgcrypt, if it has not been initialized before. Thus if you want to manually initialize libgcrypt you must do it before calling this function. This is useful in cases you want to disable libgcrypt’s internal lockings etc. This function increment a global counter, so that gnutls_global_deinit() only releases resources when it has been called as many times as gnutls_global_init(). This is useful when GnuTLS is used by more than one library in an application. This function can be called many times, but will only do something the first time. Note! This function is not thread safe. If two threads call this function simultaneously, they can cause a race between checking the global counter and incrementing it, causing both threads to execute the library initialization code. That would lead to a memory leak. To handle this, your application could invoke this function after aquiring a thread mutex. To ignore the potential memory leak is also an option.

gnutls global set log function

[Function] void gnutls_global_set_log_function (gnutls log func log_func)

log func: it’s a log function This is the function where you set the logging function gnutls is going to use. This function only accepts a character array. Normally you may not use this function since it is only used for debugging purposes. gnutls log func is of the form, void (*gnutls log func)( int level, const char*);

gnutls global set log level

[Function] void gnutls_global_set_log_level (int level)

level: it’s an integer from 0 to 9. This is the function that allows you to set the log level. The level is an integer between 0 and 9. Higher values mean more verbosity. The default value is 0. Larger values should only be used with care, since they may reveal sensitive information. Use a log level over 10 to enable all debugging options.

gnutls global set mem functions

[Function] void gnutls_global_set_mem_functions (gnutls alloc function
   alloc_func, gnutls alloc function secure_alloc_func,

gnutls is secure function is_secure_func, gnutls realloc function

   realloc_func, gnutls free function free_func)

alloc func: it’s the default memory allocation function. Like malloc(). secure alloc func: This is the memory allocation function that will be used for sensitive data. is secure func: a function that returns 0 if the memory given is not secure. May be NULL. realloc func: A realloc function free func: The function that frees allocated data. Must accept a NULL pointer. This is the function were you set the memory allocation functions gnutls is going to use. By default the libc’s allocation functions (malloc(), free()), are used by gnutls, to allocate both sensitive and not sensitive data. This function is provided to set the memory allocation functions to something other than the defaults (ie the gcrypt allocation functions). This function must be called before gnutls_global_init() is called.

gnutls handshake get last in

[Function] gnutls_handshake_description_t
   gnutls_handshake_get_last_in (gnutls session t session)

session: is a gnutls_session_t structure. Returns the last handshake message received. This function is only useful to check where the last performed handshake failed. If the previous handshake succeed or was not performed at all then no meaningful value will be returned. Check gnutls.h for the available handshake descriptions.

gnutls handshake get last out

[Function] gnutls_handshake_description_t
   gnutls_handshake_get_last_out (gnutls session t session)

session: is a gnutls_session_t structure. Returns the last handshake message sent. This function is only useful to check where the last performed handshake failed. If the previous handshake succeed or was not performed at all then no meaningful value will be returned. Check gnutls.h for the available handshake descriptions.

gnutls handshake set max packet length

[Function] void gnutls_handshake_set_max_packet_length (gnutls session t
   session, size t max)

session: is a gnutls_session_t structure. max: is the maximum number.

This function will set the maximum size of a handshake message. Handshake messages over this size are rejected. The default value is 16kb which is large enough. Set this to 0 if you do not want to set an upper limit.

gnutls handshake set private extensions

[Function] void gnutls_handshake_set_private_extensions (gnutls session t
   session, int allow)

session: is a gnutls_session_t structure. allow: is an integer (0 or 1) This function will enable or disable the use of private cipher suites (the ones that start with 0xFF). By default or if allow is 0 then these cipher suites will not be advertized nor used. Unless this function is called with the option to allow (1), then no compression algorithms, like LZO. That is because these algorithms are not yet defined in any RFC or even internet draft. Enabling the private ciphersuites when talking to other than gnutls servers and clients may cause interoperability problems.

gnutls handshake

[Function] int gnutls_handshake (gnutls session t session)

session: is a gnutls_session_t structure. This function does the handshake of the TLS/SSL protocol, and initializes the TLS connection. This function will fail if any problem is encountered, and will return a negative error code. In case of a client, if the client has asked to resume a session, but the server couldn’t, then a full handshake will be performed. The non-fatal errors such as GNUTLS E AGAIN and GNUTLS E INTERRUPTED interrupt the handshake procedure, which should be later be resumed. Call this

function again, until it returns 0; cf. gnutls_record_get_direction() and gnutls_
error_is_fatal().

If this function is called by a server after a rehandshake request then GNUTLS E GOT APPLICATION DATA or GNUTLS E WARNING ALERT RECEIVED may be returned. Note that these are non fatal errors, only in the specific case of a rehandshake. Their meaning is that the client rejected the rehandshake request.

gnutls hex decode

[Function] int gnutls_hex_decode (const gnutls datum t * hex_data, char *

   result, size t * result_size)

hex data: contain the encoded data result: the place where decoded data will be copied result size: holds the size of the result This function will decode the given encoded data, using the hex encoding used by PSK password files.

Note that hex data should be null terminated. Returns GNUTLS E SHORT MEMORY BUFFER if the buffer given is not long enough, or 0 on success.

gnutls hex encode

[Function] int gnutls_hex_encode (const gnutls datum t * data, char * result,

   size t * result_size)

data: contain the raw data result: the place where hex data will be copied result size: holds the size of the result This function will convert the given data to printable data, using the hex encoding, as used in the PSK password files. Returns GNUTLS E SHORT MEMORY BUFFER if the buffer given is not long enough, or 0 on success.

gnutls init

[Function] int gnutls_init (gnutls session t * session, gnutls connection end t

   con_end)

session: is a pointer to a gnutls_session_t structure. con end: is used to indicate if this session is to be used for server or client. Can be one of GNUTLS CLIENT and GNUTLS SERVER. This function initializes the current session to null. Every session must be initialized before use, so internal structures can be allocated. This function allocates structures which can only be free’d by calling gnutls_deinit(). Returns zero on success.

gnutls kx get name

[Function] const char * gnutls_kx_get_name (gnutls kx algorithm t

   algorithm)

algorithm: is a key exchange algorithm Returns a pointer to a string that contains the name of the specified key exchange algorithm or NULL.

gnutls kx get

[Function] gnutls_kx_algorithm_t gnutls_kx_get (gnutls session t session)

session: is a gnutls_session_t structure. Returns the key exchange algorithm used in the last handshake.

gnutls kx set priority

[Function] int gnutls_kx_set_priority (gnutls session t session, const int *

   list)

session: is a gnutls_session_t structure. list: is a 0 terminated list of gnutls kx algorithm t elements.

Sets the priority on the key exchange algorithms supported by gnutls. Priority is higher for algorithms specified before others. After specifying the algorithms you want, you must append a 0. Note that the priority is set on the client. The server does not use the algorithm’s priority except for disabling algorithms that were not specified. Returns 0 on success.

gnutls mac get name

[Function] const char * gnutls_mac_get_name (gnutls mac algorithm t

   algorithm)

algorithm: is a MAC algorithm Returns a string that contains the name of the specified MAC algorithm or NULL.

gnutls mac get

[Function] gnutls_mac_algorithm_t gnutls_mac_get (gnutls session t
   session)

session: is a gnutls_session_t structure. Returns the currently used mac algorithm.

gnutls mac set priority

[Function] int gnutls_mac_set_priority (gnutls session t session, const int *

   list)

session: is a gnutls_session_t structure. list: is a 0 terminated list of gnutls mac algorithm t elements. Sets the priority on the mac algorithms supported by gnutls. Priority is higher for algorithms specified before others. After specifying the algorithms you want, you must append a 0. Note that the priority is set on the client. The server does not use the algorithm’s priority except for disabling algorithms that were not specified. Returns 0 on success.

gnutls malloc

[Function] void * gnutls_malloc (size t s)

This function will allocate ’s’ bytes data, and return a pointer to memory. This function is supposed to be used by callbacks. The allocation function used is the one set by gnutls_global_set_mem_functions().

gnutls openpgp send key

[Function] void gnutls_openpgp_send_key (gnutls session t session,

gnutls openpgp key status t status) session: is a pointer to a gnutls_session_t structure. status: is one of OPENPGP KEY, or OPENPGP KEY FINGERPRINT This function will order gnutls to send the key fingerprint instead of the key in the initial handshake procedure. This should be used with care and only when there is indication or knowledge that the server can obtain the client’s key.

gnutls pem base64 decode alloc

[Function] int gnutls_pem_base64_decode_alloc (const char * header, const

gnutls datum t * b64_data, gnutls datum t * result) header: The PEM header (eg. CERTIFICATE) b64 data: contains the encoded data result: the place where decoded data lie This function will decode the given encoded data. The decoded data will be allocated, and stored into result. If the header given is non null this function will search for "—–BEGIN header" and decode only this part. Otherwise it will decode the first PEM packet found. You should use gnutls_free() to free the returned data.

gnutls pem base64 decode

[Function] int gnutls_pem_base64_decode (const char * header, const

gnutls datum t * b64_data, unsigned char * result, size t * result_size) header: A null terminated string with the PEM header (eg. CERTIFICATE) b64 data: contain the encoded data result: the place where decoded data will be copied result size: holds the size of the result This function will decode the given encoded data. If the header given is non null this function will search for "—–BEGIN header" and decode only this part. Otherwise it will decode the first PEM packet found. Returns GNUTLS E SHORT MEMORY BUFFER if the buffer given is not long enough, or 0 on success.

gnutls pem base64 encode alloc

[Function] int gnutls_pem_base64_encode_alloc (const char * msg, const

gnutls datum t * data, gnutls datum t * result) msg: is a message to be put in the encoded header data: contains the raw data result: will hold the newly allocated encoded data This function will convert the given data to printable data, using the base64 encoding. This is the encoding used in PEM messages. This function will allocate the required memory to hold the encoded data. You should use gnutls_free() to free the returned data.

gnutls pem base64 encode

[Function] int gnutls_pem_base64_encode (const char * msg, const

gnutls datum t * data, char * result, size t * result_size) msg: is a message to be put in the header data: contain the raw data result: the place where base64 data will be copied result size: holds the size of the result This function will convert the given data to printable data, using the base64 encoding. This is the encoding used in PEM messages. If the provided buffer is not long enough GNUTLS E SHORT MEMORY BUFFER is returned. The output string will be null terminated, although the size will not include the terminating null.

gnutls perror

[Function] void gnutls_perror (int error)

error: is an error returned by a gnutls function. Error is always a negative value. This function is like perror(). The only difference is that it accepts an error number returned by a gnutls function.

gnutls pk algorithm get name

[Function] const char * gnutls_pk_algorithm_get_name

(gnutls pk algorithm t algorithm) algorithm: is a pk algorithm Returns a string that contains the name of the specified public key algorithm or NULL.

gnutls prf raw

[Function] int gnutls_prf_raw (gnutls session t session, size t label_size,

const char * label, size t seed_size, const char * seed, size t outsize, char * out)

session: is a gnutls_session_t structure. label size: length of the label variable. label: label used in PRF computation, typically a short string. seed size: length of the seed variable. seed: optional extra data to seed the PRF with. outsize: size of pre-allocated output buffer to hold the output. out: pre-allocate buffer to hold the generated data. Apply the TLS Pseudo-Random-Function (PRF) using the master secret on some data. The label variable usually contain a string denoting the purpose for the generated data. The seed usually contain data such as the client and server random, perhaps together with some additional data that is added to guarantee uniqueness of the output for a particular purpose. Because the output is not guaranteed to be unique for a particular session unless seed include the client random and server random fields (the PRF would output the same data on another connection resumed from the first one), it is not recommended to use this function directly. The gnutls_prf() function seed the PRF with the client and server random fields directly, and is recommended if you want to generate pseudo random data unique for each session. Return value: Return 0 on success, or an error code.

gnutls prf

[Function] int gnutls_prf (gnutls session t session, size t label_size, const

   char * label, int server_random_first, size t extra_size, const char *
   extra, size t outsize, char * out)

session: is a gnutls_session_t structure. label size: length of the label variable. label: label used in PRF computation, typically a short string. server random first: non-0 if server random field should be first in seed extra size: length of the extra variable. extra: optional extra data to seed the PRF with. outsize: size of pre-allocated output buffer to hold the output. out: pre-allocate buffer to hold the generated data. Apply the TLS Pseudo-Random-Function (PRF) using the master secret on some data, seeded with the client and server random fields. The label variable usually contain a string denoting the purpose for the generated data. The server_random_first indicate whether the client random field or the server random field should be first in the seed. Non-0 indicate that the server random field is first, 0 that the client random field is first. The extra variable can be used to add more data to the seed, after the random variables. It can be used to tie make sure the generated output is strongly connected to some additional data (e.g., a string used in user authentication). The output is placed in *OUT, which must be pre-allocated. Return value: Return 0 on success, or an error code.

gnutls protocol get name

[Function] const char * gnutls_protocol_get_name (gnutls protocol t
   version)

version: is a (gnutls) version number Returns a string that contains the name of the specified TLS version or NULL.

gnutls protocol get version

[Function] gnutls_protocol_t gnutls_protocol_get_version

(gnutls session t session) session: is a gnutls_session_t structure. Returns the version of the currently used protocol.

gnutls protocol set priority

[Function] int gnutls_protocol_set_priority (gnutls session t session, const

int * list) session: is a gnutls_session_t structure.

list: is a 0 terminated list of gnutls protocol t elements.

Sets the priority on the protocol versions supported by gnutls. This function actually enables or disables protocols. Newer protocol versions always have highest priority.

Returns 0 on success.

gnutls psk allocate client credentials

[Function] int gnutls_psk_allocate_client_credentials

(gnutls psk client credentials t * sc)

sc: is a pointer to an gnutls_psk_server_credentials_t structure.

This structure is complex enough to manipulate directly thus this helper function is provided in order to allocate it.

Returns 0 on success.

gnutls psk allocate server credentials

[Function] int gnutls_psk_allocate_server_credentials

(gnutls psk server credentials t * sc)

sc: is a pointer to an gnutls_psk_server_credentials_t structure.

This structure is complex enough to manipulate directly thus this helper function is provided in order to allocate it.

Returns 0 on success.

gnutls psk free client credentials

[Function] void gnutls_psk_free_client_credentials

(gnutls psk client credentials t sc)

sc: is an gnutls_psk_client_credentials_t structure.

This structure is complex enough to manipulate directly thus this helper function is provided in order to free (deallocate) it.

gnutls psk free server credentials

[Function] void gnutls_psk_free_server_credentials

(gnutls psk server credentials t sc)

sc: is an gnutls_psk_server_credentials_t structure.

This structure is complex enough to manipulate directly thus this helper function is provided in order to free (deallocate) it.

gnutls psk server get username

[Function] const char * gnutls_psk_server_get_username (gnutls session t
   session)

session: is a gnutls session This function will return the username of the peer. This should only be called in case of PSK authentication and in case of a server. Returns NULL in case of an error.

gnutls psk set client credentials function

[Function] void gnutls_psk_set_client_credentials_function

(gnutls psk client credentials t cred, gnutls psk client credentials function *

   func)
cred: is a gnutls_psk_server_credentials_t structure.

func: is the callback function This function can be used to set a callback to retrieve the username and password for client PSK authentication. The callback’s function form is: int (*callback)(gnutls session t, char** username, gnutls datum* key);

The username and key must be allocated using gnutls_malloc(). username should

be ASCII strings or UTF-8 strings prepared using the "SASLprep" profile of "stringprep". The callback function will be called once per handshake. The callback function should return 0 on success. -1 indicates an error.

gnutls psk set client credentials

[Function] int gnutls_psk_set_client_credentials

(gnutls psk client credentials t res, const char * username, const gnutls datum * key, unsigned int flags)

res: is an gnutls_psk_client_credentials_t structure.

username: is the user’s zero-terminated userid key: is the user’s key This function sets the username and password, in a gnutls psk client credentials t structure. Those will be used in PSK authentication. username should be an ASCII string or UTF-8 strings prepared using the "SASLprep" profile of "stringprep". The key can be either in raw byte format or in Hex (not with the ’0x’ prefix). Returns 0 on success.

gnutls psk set params function

[Function] void gnutls_psk_set_params_function

(gnutls psk server credentials t res, gnutls params function * func) res: is a gnutls psk server credentials t structure func: is the function to be called This function will set a callback in order for the server to get the diffie hellman or RSA parameters for psk authentication. The callback should return zero on success.

gnutls psk set server credentials file

[Function] int gnutls_psk_set_server_credentials_file

(gnutls psk server credentials t res, const char * password_file)

res: is an gnutls_psk_server_credentials_t structure.

password file: is the PSK password file (passwd.psk) This function sets the password file, in a gnutls psk server credentials t structure. This password file holds usernames and keys and will be used for PSK authentication. Returns 0 on success.

gnutls psk set server credentials function

[Function] void gnutls_psk_set_server_credentials_function

(gnutls psk server credentials t cred, gnutls psk server credentials function *

   func)
cred: is a gnutls_psk_server_credentials_t structure.

func: is the callback function This function can be used to set a callback to retrieve the user’s PSK credentials. The callback’s function form is: int (*callback)(gnutls session t, const char* username, gnutls datum t* key); username contains the actual username. The key must be filled in using the gnutls_

malloc().

In case the callback returned a negative number then gnutls will assume that the username does not exist. The callback function will only be called once per handshake. The callback function should return 0 on success, while -1 indicates an error.

gnutls psk set server dh params

[Function] void gnutls_psk_set_server_dh_params

(gnutls psk server credentials t res, gnutls dh params t dh_params) res: is a gnutls psk server credentials t structure dh params: is a structure that holds diffie hellman parameters. This function will set the diffie hellman parameters for an anonymous server to use. These parameters will be used in Diffie Hellman with PSK cipher suites.

gnutls psk set server params function

[Function] void gnutls_psk_set_server_params_function

(gnutls psk server credentials t res, gnutls params function * func) res: is a gnutls certificate credentials t structure func: is the function to be called This function will set a callback in order for the server to get the diffie hellman parameters for PSK authentication. The callback should return zero on success.

gnutls record check pending

[Function] size_t gnutls_record_check_pending (gnutls session t session)

session: is a gnutls_session_t structure. This function checks if there are any data to receive in the gnutls buffers. Returns the size of that data or 0. Notice that you may also use select() to check for data in a TCP connection, instead of this function. (gnutls leaves some data in the tcp buffer in order for select to work).

gnutls record get direction

[Function] int gnutls_record_get_direction (gnutls session t session)

session: is a gnutls_session_t structure. This function provides information about the internals of the record protocol and is only useful if a prior gnutls function call (e.g. gnutls_handshake()) was interrupted for some reason, that is, if a function returned GNUTLS E INTERRUPTED or GNUTLS E AGAIN. In such a case, you might want to call select() or poll() before calling the interrupted gnutls function again. To tell you whether a file descriptor should be selected for either reading or writing, gnutls_record_get_direction() returns 0 if the interrupted function was trying to read data, and 1 if it was trying to write data.

gnutls record get max size

[Function] size_t gnutls_record_get_max_size (gnutls session t session)

session: is a gnutls_session_t structure. This function returns the maximum record packet size in this connection. The maximum record size is negotiated by the client after the first handshake message.

gnutls record recv

[Function] ssize_t gnutls_record_recv (gnutls session t session, void * data,
   size t sizeofdata)

session: is a gnutls_session_t structure. data: the buffer that the data will be read into sizeofdata: the number of requested bytes This function has the similar semantics with recv(). The only difference is that is accepts a GNUTLS session, and uses different error codes. In the special case that a server requests a renegotiation, the client may receive an error code of GNUTLS E REHANDSHAKE. This message may be simply ignored, replied with an alert containing NO RENEGOTIATION, or replied with a new handshake, depending on the client’s will. If EINTR is returned by the internal push function (the default is code{recv()}) then GNUTLS E INTERRUPTED will be returned. If GNUTLS E INTERRUPTED or GNUTLS E AGAIN is returned, you must call this function again, with the same parameters; alternatively you could provide a NULL pointer for data, and 0 for size.

cf. code{gnutls_record_get_direction()}.

A server may also receive GNUTLS E REHANDSHAKE when a client has initiated a handshake. In that case the server can only initiate a handshake or terminate the connection.

Returns the number of bytes received and zero on EOF. A negative error code is returned in case of an error. The number of bytes received might be less than

code{count}.

gnutls record send

[Function] ssize_t gnutls_record_send (gnutls session t session, const void *

   data, size t sizeofdata)

session: is a gnutls_session_t structure.

data: contains the data to send

sizeofdata: is the length of the data

This function has the similar semantics with send(). The only difference is that is accepts a GNUTLS session, and uses different error codes.

Note that if the send buffer is full, send() will block this function. See the send() documentation for full information. You can replace the default push function by us-

ing gnutls_transport_set_ptr2() with a call to send() with a MSG DONTWAIT

flag if blocking is a problem.

If the EINTR is returned by the internal push function (the default is send()} then

GNUTLS_E_INTERRUPTED will be returned. If GNUTLS_E_INTERRUPTED or GNUTLS_E_

AGAIN is returned, you must call this function again, with the same parameters; alternatively you could provide a NULL pointer for data, and 0 for size. cf. gnutls_

record_get_direction().

Returns the number of bytes sent, or a negative error code. The number of bytes sent might be less than sizeofdata. The maximum number of bytes this function can send in a single call depends on the negotiated maximum record size.

gnutls record set max size

[Function] ssize_t gnutls_record_set_max_size (gnutls session t session,

size t size) session: is a gnutls_session_t structure.

size: is the new size

This function sets the maximum record packet size in this connection. This property can only be set to clients. The server may choose not to accept the requested size.

Acceptable values are 512(=2^9), 1024(=2^10), 2048(=2^11) and 4096(=2^12). Returns 0 on success. The requested record size does get in effect immediately only while sending data. The receive part will take effect after a successful handshake.

This function uses a TLS extension called ’max record size’. Not all TLS implementations use or even understand this extension.

gnutls rehandshake

[Function] int gnutls_rehandshake (gnutls session t session)

session: is a gnutls_session_t structure. This function will renegotiate security parameters with the client. This should only be called in case of a server. This message informs the peer that we want to renegotiate parameters (perform a handshake). If this function succeeds (returns 0), you must call the gnutls_handshake() function in order to negotiate the new parameters. If the client does not wish to renegotiate parameters he will should with an alert message, thus the return code will be GNUTLS E WARNING ALERT RECEIVED and the alert will be GNUTLS A NO RENEGOTIATION. A client may also choose to ignore this message.

gnutls rsa export get modulus bits

[Function] int gnutls_rsa_export_get_modulus_bits (gnutls session t
   session)

session: is a gnutls session This function will return the bits used in the last RSA-EXPORT key exchange with the peer. Returns a negative value in case of an error.

gnutls rsa export get pubkey

[Function] int gnutls_rsa_export_get_pubkey (gnutls session t session,

gnutls datum t * exponent, gnutls datum t * modulus) session: is a gnutls session exponent: will hold the exponent. modulus: will hold the modulus. This function will return the peer’s public key exponent and modulus used in the last RSA-EXPORT authentication. The output parameters must be freed with gnutls_

free().

Returns a negative value in case of an error.

gnutls rsa params cpy

[Function] int gnutls_rsa_params_cpy (gnutls rsa params t dst,

gnutls rsa params t src) dst: Is the destination structure, which should be initialized. src: Is the source structure This function will copy the RSA parameters structure from source to destination.

gnutls rsa params deinit

[Function] void gnutls_rsa_params_deinit (gnutls rsa params t rsa_params)

rsa params: Is a structure that holds the parameters This function will deinitialize the RSA parameters structure.

gnutls rsa params export pkcs1

[Function] int gnutls_rsa_params_export_pkcs1 (gnutls rsa params t params,

gnutls x509 crt fmt t format, unsigned char * params_data, size t *

   params_data_size)

params: Holds the RSA parameters format: the format of output params. One of PEM or DER. params data: will contain a PKCS1 RSAPublicKey structure PEM or DER encoded params data size: holds the size of params data (and will be replaced by the actual size of parameters) This function will export the given RSA parameters to a PKCS1 RSAPublicKey structure. If the buffer provided is not long enough to hold the output, then GNUTLS E SHORT MEMORY BUFFER will be returned. If the structure is PEM encoded, it will have a header of "BEGIN RSA PRIVATE KEY". In case of failure a negative value will be returned, and 0 on success.

gnutls rsa params export raw

[Function] int gnutls_rsa_params_export_raw (gnutls rsa params t params,

gnutls datum t * m, gnutls datum t * e, gnutls datum t * d, gnutls datum t * p, gnutls datum t * q, gnutls datum t * u, unsigned int * bits)

params: a structure that holds the rsa parameters m: will hold the modulus e: will hold the public exponent d: will hold the private exponent p: will hold the first prime (p) q: will hold the second prime (q) u: will hold the coefficient bits: if non null will hold the prime’s number of bits This function will export the RSA parameters found in the given structure. The new parameters will be allocated using gnutls_malloc() and will be stored in the appropriate datum.

gnutls rsa params generate2

[Function] int gnutls_rsa_params_generate2 (gnutls rsa params t params,

unsigned int bits) params: The structure where the parameters will be stored bits: is the prime’s number of bits This function will generate new temporary RSA parameters for use in RSA-EXPORT ciphersuites. This function is normally slow. Note that if the parameters are to be used in export cipher suites the bits value should be 512 or less. Also note that the generation of new RSA parameters is only useful to servers. Clients use the parameters sent by the server, thus it’s no use calling this in client side.

gnutls rsa params import pkcs1

[Function] int gnutls_rsa_params_import_pkcs1 (gnutls rsa params t params,

const gnutls datum t * pkcs1_params, gnutls x509 crt fmt t format) params: A structure where the parameters will be copied to pkcs1 params: should contain a PKCS1 RSAPublicKey structure PEM or DER encoded format: the format of params. PEM or DER. This function will extract the RSAPublicKey found in a PKCS1 formatted structure. If the structure is PEM encoded, it should have a header of "BEGIN RSA PRIVATE KEY". In case of failure a negative value will be returned, and 0 on success.

gnutls rsa params import raw

[Function] int gnutls_rsa_params_import_raw (gnutls rsa params t

rsa_params, const gnutls datum t * m, const gnutls datum t * e, const gnutls datum t * d, const gnutls datum t * p, const gnutls datum t * q, const gnutls datum t * u)

rsa params: Is a structure will hold the parameters m: holds the modulus e: holds the public exponent d: holds the private exponent p: holds the first prime (p) q: holds the second prime (q) u: holds the coefficient This function will replace the parameters in the given structure. The new parameters should be stored in the appropriate gnutls datum.

gnutls rsa params init

[Function] int gnutls_rsa_params_init (gnutls rsa params t * rsa_params)

rsa params: Is a structure that will hold the parameters This function will initialize the temporary RSA parameters structure.

gnutls server name get

[Function] int gnutls_server_name_get (gnutls session t session, void * data,

size t * data_length, unsigned int * type, unsigned int indx) session: is a gnutls_session_t structure. data: will hold the data data length: will hold the data length. Must hold the maximum size of data. type: will hold the server name indicator type indx: is the index of the server name This function will allow you to get the name indication (if any), a client has sent. The name indication may be any of the enumeration gnutls server name type t. If type is GNUTLS NAME DNS, then this function is to be used by servers that support virtual hosting, and the data will be a null terminated UTF-8 string. If data has not enough size to hold the server name GNUTLS E SHORT MEMORY BUFFER is returned, and data_length will hold the required size. index is used to retrieve more than one server names (if sent by the client). The first server name has an index of 0, the second 1 and so on. If no name with the given index exists GNUTLS E REQUESTED DATA NOT AVAILABLE is returned.

gnutls server name set

[Function] int gnutls_server_name_set (gnutls session t session,

gnutls server name type t type, const void * name, size t name_length) session: is a gnutls_session_t structure. type: specifies the indicator type name: is a string that contains the server name. name length: holds the length of name This function is to be used by clients that want to inform (via a TLS extension mechanism) the server of the name they connected to. This should be used by clients that connect to servers that do virtual hosting. The value of name depends on the ind type. In case of GNUTLS NAME DNS, an ASCII or UTF-8 null terminated string, without the trailing dot, is expected. IPv4 or IPv6 addresses are not permitted.

gnutls session get client random

[Function] const void * gnutls_session_get_client_random

(gnutls session t session) session: is a gnutls_session_t structure. Return a pointer to the 32-byte client random field used in the session. The pointer must not be modified or deallocated. If a client random value has not yet been established, the output will be garbage; in particular, a NULL return value should not be expected. Return value: pointer to client random.

gnutls session get data2

[Function] int gnutls_session_get_data2 (gnutls session t session,

gnutls datum * data) session: is a gnutls_session_t structure. Returns all session parameters, in order to support resuming. The client should call this, and keep the returned session, if he wants to resume that current version later by calling gnutls_session_set_data() This function must be called after a successful handshake. The returned datum must be freed with gnutls_free(). Resuming sessions is really useful and speedups connections after a succesful one.

gnutls session get data

[Function] int gnutls_session_get_data (gnutls session t session, void *
   session_data, size t * session_data_size)

session: is a gnutls_session_t structure. session data: is a pointer to space to hold the session. session data size: is the session data’s size, or it will be set by the function. Returns all session parameters, in order to support resuming. The client should call this, and keep the returned session, if he wants to resume that current version later by calling gnutls_session_set_data() This function must be called after a successful handshake. Resuming sessions is really useful and speedups connections after a succesful one.

gnutls session get id

[Function] int gnutls_session_get_id (gnutls session t session, void *
   session_id, size t * session_id_size)

session: is a gnutls_session_t structure. session id: is a pointer to space to hold the session id. session id size: is the session id’s size, or it will be set by the function. Returns the current session id. This can be used if you want to check if the next session you tried to resume was actually resumed. This is because resumed sessions have the same sessionID with the original session. Session id is some data set by the server, that identify the current session. In TLS 1.0 and SSL 3.0 session id is always less than 32 bytes. Returns zero on success.

gnutls session get master secret

[Function] const void * gnutls_session_get_master_secret

(gnutls session t session) session: is a gnutls_session_t structure. Return a pointer to the 48-byte master secret in the session. The pointer must not be modified or deallocated. If a master secret value has not yet been established, the output will be garbage; in particular, a NULL return value should not be expected. Consider using gnutls_prf() rather than extracting the master secret and use it to derive further data. Return value: pointer to master secret.

gnutls session get ptr

[Function] void * gnutls_session_get_ptr (gnutls session t session)

session: is a gnutls_session_t structure. This function will return the user given pointer from the session structure. This is

the pointer set with gnutls_session_set_ptr().

gnutls session get server random

[Function] const void * gnutls_session_get_server_random

(gnutls session t session) session: is a gnutls_session_t structure. Return a pointer to the 32-byte server random field used in the session. The pointer must not be modified or deallocated. If a server random value has not yet been established, the output will be garbage; in particular, a NULL return value should not be expected. Return value: pointer to server random.

gnutls session is resumed

[Function] int gnutls_session_is_resumed (gnutls session t session)

session: is a gnutls_session_t structure. This function will return non zero if this session is a resumed one, or a zero if this is a new session.

gnutls session set data

[Function] int gnutls_session_set_data (gnutls session t session, const void *

   session_data, size t session_data_size)

session: is a gnutls_session_t structure. session data: is a pointer to space to hold the session. session data size: is the session’s size Sets all session parameters, in order to resume a previously established session. The session data given must be the one returned by gnutls_session_get_data(). This function should be called before gnutls_handshake(). Keep in mind that session resuming is advisory. The server may choose not to resume the session, thus a full handshake will be performed. Returns a negative value on error.

gnutls session set ptr

[Function] void gnutls_session_set_ptr (gnutls session t session, void * ptr)

session: is a gnutls_session_t structure. ptr: is the user pointer This function will set (associate) the user given pointer to the session structure. This is pointer can be accessed with gnutls_session_get_ptr().

gnutls set default export priority

[Function] int gnutls_set_default_export_priority (gnutls session t
   session)

session: is a gnutls_session_t structure.

Sets some default priority on the ciphers, key exchange methods, macs and compression methods. This is to avoid using the gnutls *_priority() functions, if these defaults are ok. This function also includes weak algorithms. The order is TLS1, SSL3 for protocols, RSA, DHE DSS, DHE RSA, RSA EXPORT for key exchange algorithms. SHA, MD5, RIPEMD160 for MAC algorithms, AES 256 CBC, AES 128 CBC, and 3DES CBC, ARCFOUR 128, ARCFOUR 40 for ciphers.

Returns 0 on success.

gnutls set default priority

[Function] int gnutls_set_default_priority (gnutls session t session)

session: is a gnutls_session_t structure.

Sets some default priority on the ciphers, key exchange methods, macs and compression methods. This is to avoid using the gnutls *_priority() functions, if these defaults are ok. You may override any of the following priorities by calling the appropriate functions.

The order is TLS1, SSL3 for protocols. RSA, DHE DSS, DHE RSA for key exchange algorithms. SHA, MD5 and RIPEMD160 for MAC algorithms. AES 128 CBC, 3DES CBC, and ARCFOUR 128 for ciphers.

Returns 0 on success.

gnutls sign algorithm get name

[Function] const char * gnutls_sign_algorithm_get_name

(gnutls sign algorithm t sign) Returns a string that contains the name of the specified sign algorithm or NULL.

gnutls srp allocate client credentials

[Function] int gnutls_srp_allocate_client_credentials

(gnutls srp client credentials t * sc)

sc: is a pointer to an gnutls_srp_server_credentials_t structure.

This structure is complex enough to manipulate directly thus this helper function is provided in order to allocate it.

Returns 0 on success.

gnutls srp allocate server credentials

[Function] int gnutls_srp_allocate_server_credentials

(gnutls srp server credentials t * sc)

sc: is a pointer to an gnutls_srp_server_credentials_t structure.

This structure is complex enough to manipulate directly thus this helper function is provided in order to allocate it.

Returns 0 on success.

gnutls srp base64 decode alloc

[Function] int gnutls_srp_base64_decode_alloc (const gnutls datum t *

b64_data, gnutls datum t * result) b64 data: contains the encoded data result: the place where decoded data lie This function will decode the given encoded data. The decoded data will be allocated, and stored into result. It will decode using the base64 algorithm found in libsrp. You should use gnutls_free() to free the returned data.

gnutls srp base64 decode

[Function] int gnutls_srp_base64_decode (const gnutls datum t * b64_data,
   char * result, size t * result_size)

b64 data: contain the encoded data result: the place where decoded data will be copied result size: holds the size of the result This function will decode the given encoded data, using the base64 encoding found in libsrp. Note that b64 data should be null terminated. Returns GNUTLS E SHORT MEMORY BUFFER if the buffer given is not long enough, or 0 on success.

gnutls srp base64 encode alloc

[Function] int gnutls_srp_base64_encode_alloc (const gnutls datum t * data,

gnutls datum t * result) data: contains the raw data result: will hold the newly allocated encoded data This function will convert the given data to printable data, using the base64 encoding. This is the encoding used in SRP password files. This function will allocate the required memory to hold the encoded data. You should use gnutls_free() to free the returned data.

gnutls srp base64 encode

[Function] int gnutls_srp_base64_encode (const gnutls datum t * data, char *

   result, size t * result_size)

data: contain the raw data result: the place where base64 data will be copied result size: holds the size of the result This function will convert the given data to printable data, using the base64 encoding, as used in the libsrp. This is the encoding used in SRP password files. If the provided buffer is not long enough GNUTLS E SHORT MEMORY BUFFER is returned.

gnutls srp free client credentials

[Function] void gnutls_srp_free_client_credentials

(gnutls srp client credentials t sc)

sc: is an gnutls_srp_client_credentials_t structure.

This structure is complex enough to manipulate directly thus this helper function is provided in order to free (deallocate) it.

gnutls srp free server credentials

[Function] void gnutls_srp_free_server_credentials

(gnutls srp server credentials t sc)

sc: is an gnutls_srp_server_credentials_t structure.

This structure is complex enough to manipulate directly thus this helper function is provided in order to free (deallocate) it.

gnutls srp server get username

[Function] const char * gnutls_srp_server_get_username (gnutls session t
   session)

session: is a gnutls session

This function will return the username of the peer. This should only be called in case of SRP authentication and in case of a server. Returns NULL in case of an error.

gnutls srp set client credentials function

[Function] void gnutls_srp_set_client_credentials_function

(gnutls srp client credentials t cred, gnutls srp client credentials function *

   func)
cred: is a gnutls_srp_server_credentials_t structure.

func: is the callback function

This function can be used to set a callback to retrieve the username and password for client SRP authentication. The callback’s function form is: int (*callback)(gnutls session t, unsigned int times, char** username, char** password);

The username and password must be allocated using gnutls_malloc(). times will

be 0 the first time called, and 1 the second. username and password should be ASCII strings or UTF-8 strings prepared using the "SASLprep" profile of "stringprep".

The callback function will be called once or twice per handshake. The first time called, is before the ciphersuite is negotiated. At that time if the callback returns a negative error code, the callback will be called again if SRP has been negotiated. This uses a special TLS-SRP idiom in order to avoid asking the user for SRP password and username if the server does not support SRP.

The callback should not return a negative error code the second time called, since the handshake procedure will be aborted.

The callback function should return 0 on success. -1 indicates an error.

gnutls srp set client credentials

[Function] int gnutls_srp_set_client_credentials

(gnutls srp client credentials t res, const char * username, const char *

   password)
res: is an gnutls_srp_client_credentials_t structure.

username: is the user’s userid password: is the user’s password This function sets the username and password, in a gnutls srp client credentials t structure. Those will be used in SRP authentication. username and password should be ASCII strings or UTF-8 strings prepared using the "SASLprep" profile of "stringprep". Returns 0 on success.

gnutls srp set server credentials file

[Function] int gnutls_srp_set_server_credentials_file

(gnutls srp server credentials t res, const char * password_file, const char

   * password_conf_file)
res: is an gnutls_srp_server_credentials_t structure.

password file: is the SRP password file (tpasswd) password conf file: is the SRP password conf file (tpasswd.conf) This function sets the password files, in a gnutls srp server credentials t structure. Those password files hold usernames and verifiers and will be used for SRP authentication. Returns 0 on success.

gnutls srp set server credentials function

[Function] void gnutls_srp_set_server_credentials_function

(gnutls srp server credentials t cred, gnutls srp server credentials function *

   func)
cred: is a gnutls_srp_server_credentials_t structure.

func: is the callback function This function can be used to set a callback to retrieve the user’s SRP credentials. The callback’s function form is: int (*callback)(gnutls session t, const char* username, gnutls datum t* salt, gnutls datum t *verifier, gnutls datum t* g, gnutls datum t* n);

username contains the actual username. The salt, verifier, generator and prime

must be filled in using the gnutls_malloc(). For convenience prime and generator may also be one of the static parameters defined in extra.h. In case the callback returned a negative number then gnutls will assume that the username does not exist. In order to prevent attackers from guessing valid usernames, if a user does not exist, g and n values should be filled in using a random user’s parameters. In that case the callback must return the special value (1).

The callback function will only be called once per handshake. The callback function should return 0 on success, while -1 indicates an error.

gnutls srp verifier

[Function] int gnutls_srp_verifier (const char * username, const char *

password, const gnutls datum t * salt, const gnutls datum t * generator, const gnutls datum t * prime, gnutls datum t * res)

username: is the user’s name password: is the user’s password salt: should be some randomly generated bytes generator: is the generator of the group prime: is the group’s prime res: where the verifier will be stored. This function will create an SRP verifier, as specified in RFC2945. The prime and generator should be one of the static parameters defined in gnutls/extra.h or may be generated using the GCRYPT functions gcry_prime_generate() and gcry_prime_ group_generator(). The verifier will be allocated with malloc and will be stored in res using binary format.

gnutls strerror

[Function] const char * gnutls_strerror (int error)

error: is an error returned by a gnutls function. Error is always a negative value. This function is similar to strerror(). Differences: it accepts an error number returned by a gnutls function; In case of an unknown error a descriptive string is sent instead of NULL.

gnutls transport get ptr2

[Function] void gnutls_transport_get_ptr2 (gnutls session t session,

gnutls transport ptr t * recv_ptr, gnutls transport ptr t * send_ptr) session: is a gnutls_session_t structure. recv ptr: will hold the value for the pull function send ptr: will hold the value for the push function Used to get the arguments of the transport functions (like PUSH and PULL). These

should have been set using gnutls_transport_set_ptr2().

gnutls transport get ptr

[Function] gnutls_transport_ptr_t gnutls_transport_get_ptr

(gnutls session t session) session: is a gnutls_session_t structure. Used to get the first argument of the transport function (like PUSH and PULL). This

must have been set using gnutls_transport_set_ptr().

gnutls transport set lowat

[Function] void gnutls_transport_set_lowat (gnutls session t session, int
   num)

session: is a gnutls_session_t structure.

num: is the low water value.

Used to set the lowat value in order for select to check if there are pending data to socket buffer. Used only if you have changed the default low water value (default is 1). Normally you will not need that function. This function is only useful if using berkeley style sockets. Otherwise it must be called and set lowat to zero.

gnutls transport set ptr2

[Function] void gnutls_transport_set_ptr2 (gnutls session t session,

gnutls transport ptr t recv_ptr, gnutls transport ptr t send_ptr) session: is a gnutls_session_t structure.

recv ptr: is the value for the pull function

send ptr: is the value for the push function

Used to set the first argument of the transport function (like PUSH and PULL). In berkeley style sockets this function will set the connection handle. With this function you can use two different pointers for receiving and sending.

gnutls transport set ptr

[Function] void gnutls_transport_set_ptr (gnutls session t session,

gnutls transport ptr t ptr) session: is a gnutls_session_t structure.

ptr: is the value.

Used to set the first argument of the transport function (like PUSH and PULL). In berkeley style sockets this function will set the connection handle.

gnutls transport set pull function

[Function] void gnutls_transport_set_pull_function (gnutls session t
   session, gnutls pull func pull_func)

session: gnutls session

pull func: a callback function similar to read()

This is the function where you set a function for gnutls to receive data. Normally, if you use berkeley style sockets, do not need to use this function since the default (recv(2)) will probably be ok.

PULL FUNC is of the form, ssize t (*gnutls pull func)(gnutls transport ptr t, void*, size t);

gnutls transport set push function

[Function] void gnutls_transport_set_push_function (gnutls session t
   session, gnutls push func push_func)

session: gnutls session push func: a callback function similar to write() This is the function where you set a push function for gnutls to use in order to send data. If you are going to use berkeley style sockets, you do not need to use this function since the default (send(2)) will probably be ok. Otherwise you should specify this function for gnutls to be able to send data. PUSH FUNC is of the form, ssize t (*gnutls push func)(gnutls transport ptr t, const void*, size t);

9.2 X.509 certificate functions

The following functions are to be used for X.509 certificate handling. Their prototypes lie

in ‘gnutls/x509.h’.

gnutls pkcs12 bag decrypt

[Function] int gnutls_pkcs12_bag_decrypt (gnutls pkcs12 bag t bag, const char

   * pass)

bag: The bag pass: The password used for encryption. This can only be ASCII. This function will decrypt the given encrypted bag and return 0 on success.

gnutls pkcs12 bag deinit

[Function] void gnutls_pkcs12_bag_deinit (gnutls pkcs12 bag t bag)

bag: The structure to be initialized This function will deinitialize a PKCS12 Bag structure.

gnutls pkcs12 bag encrypt

[Function] int gnutls_pkcs12_bag_encrypt (gnutls pkcs12 bag t bag, const char

* pass, unsigned int flags) bag: The bag pass: The password used for encryption. This can only be ASCII. flags: should be one of gnutls pkcs encrypt flags t elements bitwise or’d This function will encrypt the given bag and return 0 on success.

gnutls pkcs12 bag get count

[Function] int gnutls_pkcs12_bag_get_count (gnutls pkcs12 bag t bag)

bag: The bag This function will return the number of the elements withing the bag.

gnutls pkcs12 bag get data

[Function] int gnutls_pkcs12_bag_get_data (gnutls pkcs12 bag t bag, int

indx, gnutls datum t * data) bag: The bag indx: The element of the bag to get the data from data: where the bag’s data will be. Should be treated as constant. This function will return the bag’s data. The data is a constant that is stored into the bag. Should not be accessed after the bag is deleted. Returns 0 on success and a negative error code on error.

gnutls pkcs12 bag get friendly name

[Function] int gnutls_pkcs12_bag_get_friendly_name (gnutls pkcs12 bag t

bag, int indx, char ** name) bag: The bag indx: The bag’s element to add the id name: will hold a pointer to the name (to be treated as const) This function will return the friendly name, of the specified bag element. The key ID is usually used to distinguish the local private key and the certificate pair. Returns 0 on success, or a negative value on error.

gnutls pkcs12 bag get key id

[Function] int gnutls_pkcs12_bag_get_key_id (gnutls pkcs12 bag t bag, int

indx, gnutls datum t * id) bag: The bag indx: The bag’s element to add the id id: where the ID will be copied (to be treated as const) This function will return the key ID, of the specified bag element. The key ID is usually used to distinguish the local private key and the certificate pair. Returns 0 on success, or a negative value on error.

gnutls pkcs12 bag get type

[Function] gnutls_pkcs12_bag_type_t gnutls_pkcs12_bag_get_type

(gnutls pkcs12 bag t bag, int indx) bag: The bag indx: The element of the bag to get the type This function will return the bag’s type. One of the gnutls pkcs12 bag type t enumerations.

gnutls pkcs12 bag init

[Function] int gnutls_pkcs12_bag_init (gnutls pkcs12 bag t * bag)

bag: The structure to be initialized This function will initialize a PKCS12 bag structure. PKCS12 Bags usually contain private keys, lists of X.509 Certificates and X.509 Certificate revocation lists. Returns 0 on success.

gnutls pkcs12 bag set crl

[Function] int gnutls_pkcs12_bag_set_crl (gnutls pkcs12 bag t bag,

gnutls x509 crl t crl) bag: The bag crl: the CRL to be copied. This function will insert the given CRL into the bag. This is just a wrapper over

gnutls_pkcs12_bag_set_data().

Returns the index of the added bag on success, or a negative value on failure.

gnutls pkcs12 bag set crt

[Function] int gnutls_pkcs12_bag_set_crt (gnutls pkcs12 bag t bag,

gnutls x509 crt t crt) bag: The bag crt: the certificate to be copied. This function will insert the given certificate into the bag. This is just a wrapper over

gnutls_pkcs12_bag_set_data().

Returns the index of the added bag on success, or a negative value on failure.

gnutls pkcs12 bag set data

[Function] int gnutls_pkcs12_bag_set_data (gnutls pkcs12 bag t bag,

gnutls pkcs12 bag type t type, const gnutls datum t * data) bag: The bag type: The data’s type data: the data to be copied. This function will insert the given data of the given type into the bag. Returns the index of the added bag on success, or a negative value on error.

gnutls pkcs12 bag set friendly name

[Function] int gnutls_pkcs12_bag_set_friendly_name (gnutls pkcs12 bag t

bag, int indx, const char * name) bag: The bag indx: The bag’s element to add the id name: the name This function will add the given key friendly name, to the specified, by the index, bag element. The name will be encoded as a ’Friendly name’ bag attribute, which is usually used to set a user name to the local private key and the certificate pair. Returns 0 on success, or a negative value on error.

gnutls pkcs12 bag set key id

[Function] int gnutls_pkcs12_bag_set_key_id (gnutls pkcs12 bag t bag, int

indx, const gnutls datum t * id) bag: The bag indx: The bag’s element to add the id id: the ID This function will add the given key ID, to the specified, by the index, bag element. The key ID will be encoded as a ’Local key identifier’ bag attribute, which is usually used to distinguish the local private key and the certificate pair. Returns 0 on success, or a negative value on error.

gnutls pkcs12 deinit

[Function] void gnutls_pkcs12_deinit (gnutls pkcs12 t pkcs12)

pkcs12: The structure to be initialized This function will deinitialize a PKCS12 structure.

gnutls pkcs12 export

[Function] int gnutls_pkcs12_export (gnutls pkcs12 t pkcs12,

gnutls x509 crt fmt t format, void * output_data, size t *

   output_data_size)

pkcs12: Holds the pkcs12 structure format: the format of output params. One of PEM or DER. output data: will contain a structure PEM or DER encoded output data size: holds the size of output data (and will be replaced by the actual size of parameters) This function will export the pkcs12 structure to DER or PEM format. If the buffer provided is not long enough to hold the output, then *output data size will be updated and GNUTLS E SHORT MEMORY BUFFER will be returned. If the structure is PEM encoded, it will have a header of "BEGIN PKCS12". Return value: In case of failure a negative value will be returned, and 0 on success.

gnutls pkcs12 generate mac

[Function] int gnutls_pkcs12_generate_mac (gnutls pkcs12 t pkcs12, const

char * pass) pkcs12: should contain a gnutls pkcs12 t structure pass: The password for the MAC This function will generate a MAC for the PKCS12 structure. Returns 0 on success.

gnutls pkcs12 get bag

[Function] int gnutls_pkcs12_get_bag (gnutls pkcs12 t pkcs12, int indx,

gnutls pkcs12 bag t bag) pkcs12: should contain a gnutls pkcs12 t structure indx: contains the index of the bag to extract bag: An initialized bag, where the contents of the bag will be copied This function will return a Bag from the PKCS12 structure. Returns 0 on success. After the last Bag has been read GNUTLS E REQUESTED DATA NOT AVAILABLE will be returned.

gnutls pkcs12 import

[Function] int gnutls_pkcs12_import (gnutls pkcs12 t pkcs12, const

gnutls datum t * data, gnutls x509 crt fmt t format, unsigned int flags) pkcs12: The structure to store the parsed PKCS12. data: The DER or PEM encoded PKCS12. format: One of DER or PEM flags: an ORed sequence of gnutls privkey pkcs8 flags This function will convert the given DER or PEM encoded PKCS12 to the native gnutls pkcs12 t format. The output will be stored in ’pkcs12’. If the PKCS12 is PEM encoded it should have a header of "PKCS12". Returns 0 on success.

gnutls pkcs12 init

[Function] int gnutls_pkcs12_init (gnutls pkcs12 t * pkcs12)

pkcs12: The structure to be initialized This function will initialize a PKCS12 structure. PKCS12 structures usually contain lists of X.509 Certificates and X.509 Certificate revocation lists. Returns 0 on success.

gnutls pkcs12 set bag

[Function] int gnutls_pkcs12_set_bag (gnutls pkcs12 t pkcs12,

gnutls pkcs12 bag t bag) pkcs12: should contain a gnutls pkcs12 t structure bag: An initialized bag This function will insert a Bag into the PKCS12 structure. Returns 0 on success.

gnutls pkcs12 verify mac

[Function] int gnutls_pkcs12_verify_mac (gnutls pkcs12 t pkcs12, const char *

   pass)

pkcs12: should contain a gnutls pkcs12 t structure pass: The password for the MAC This function will verify the MAC for the PKCS12 structure. Returns 0 on success.

gnutls pkcs7 deinit

[Function] void gnutls_pkcs7_deinit (gnutls pkcs7 t pkcs7)

pkcs7: The structure to be initialized

This function will deinitialize a PKCS7 structure.

gnutls pkcs7 delete crl

[Function] int gnutls_pkcs7_delete_crl (gnutls pkcs7 t pkcs7, int indx)

indx: the index of the crl to delete

This function will delete a crl from a PKCS7 or RFC2630 crl set. Index starts from 0. Returns 0 on success.

gnutls pkcs7 delete crt

[Function] int gnutls_pkcs7_delete_crt (gnutls pkcs7 t pkcs7, int indx)

indx: the index of the certificate to delete

This function will delete a certificate from a PKCS7 or RFC2630 certificate set. Index starts from 0. Returns 0 on success.

gnutls pkcs7 export

[Function] int gnutls_pkcs7_export (gnutls pkcs7 t pkcs7,

gnutls x509 crt fmt t format, void * output_data, size t *

   output_data_size)

pkcs7: Holds the pkcs7 structure

format: the format of output params. One of PEM or DER.

output data: will contain a structure PEM or DER encoded

output data size: holds the size of output data (and will be replaced by the actual size of parameters)

This function will export the pkcs7 structure to DER or PEM format.

If the buffer provided is not long enough to hold the output, then *output data size is updated and GNUTLS E SHORT MEMORY BUFFER will be returned.

If the structure is PEM encoded, it will have a header of "BEGIN PKCS7".

Return value: In case of failure a negative value will be returned, and 0 on success.

gnutls pkcs7 get crl count

[Function] int gnutls_pkcs7_get_crl_count (gnutls pkcs7 t pkcs7)

This function will return the number of certifcates in the PKCS7 or RFC2630 crl set.

Returns a negative value on failure.

gnutls pkcs7 get crl raw

[Function] int gnutls_pkcs7_get_crl_raw (gnutls pkcs7 t pkcs7, int indx, void
   * crl, size t * crl_size)

indx: contains the index of the crl to extract crl: the contents of the crl will be copied there (may be null) crl size: should hold the size of the crl This function will return a crl of the PKCS7 or RFC2630 crl set. Returns 0 on success. If the provided buffer is not long enough, then crl_size is updated and GNUTLS E SHORT MEMORY BUFFER is returned. After the last crl has been read GNUTLS E REQUESTED DATA NOT AVAILABLE will be returned.

gnutls pkcs7 get crt count

[Function] int gnutls_pkcs7_get_crt_count (gnutls pkcs7 t pkcs7)

This function will return the number of certifcates in the PKCS7 or RFC2630 certificate set. Returns a negative value on failure.

gnutls pkcs7 get crt raw

[Function] int gnutls_pkcs7_get_crt_raw (gnutls pkcs7 t pkcs7, int indx, void
   * certificate, size t * certificate_size)

indx: contains the index of the certificate to extract certificate: the contents of the certificate will be copied there (may be null) certificate size: should hold the size of the certificate This function will return a certificate of the PKCS7 or RFC2630 certificate set. Returns 0 on success. If the provided buffer is not long enough, then certificate_size is updated and GNUTLS E SHORT MEMORY BUFFER is returned. After the last certificate has been read GNUTLS E REQUESTED DATA NOT AVAILABLE will be returned.

gnutls pkcs7 import

[Function] int gnutls_pkcs7_import (gnutls pkcs7 t pkcs7, const

gnutls datum t * data, gnutls x509 crt fmt t format) pkcs7: The structure to store the parsed PKCS7. data: The DER or PEM encoded PKCS7. format: One of DER or PEM This function will convert the given DER or PEM encoded PKCS7 to the native gnutls pkcs7 t format. The output will be stored in ’pkcs7’. If the PKCS7 is PEM encoded it should have a header of "PKCS7". Returns 0 on success.

gnutls pkcs7 init

[Function] int gnutls_pkcs7_init (gnutls pkcs7 t * pkcs7)

pkcs7: The structure to be initialized This function will initialize a PKCS7 structure. PKCS7 structures usually contain lists of X.509 Certificates and X.509 Certificate revocation lists. Returns 0 on success.

gnutls pkcs7 set crl raw

[Function] int gnutls_pkcs7_set_crl_raw (gnutls pkcs7 t pkcs7, const

gnutls datum t * crl) crl: the DER encoded crl to be added This function will add a crl to the PKCS7 or RFC2630 crl set. Returns 0 on success.

gnutls pkcs7 set crl

[Function] int gnutls_pkcs7_set_crl (gnutls pkcs7 t pkcs7, gnutls x509 crl t

   crl)

crl: the DER encoded crl to be added This function will add a parsed crl to the PKCS7 or RFC2630 crl set. Returns 0 on success.

gnutls pkcs7 set crt raw

[Function] int gnutls_pkcs7_set_crt_raw (gnutls pkcs7 t pkcs7, const

gnutls datum t * crt) crt: the DER encoded certificate to be added This function will add a certificate to the PKCS7 or RFC2630 certificate set. Returns 0 on success.

gnutls pkcs7 set crt

[Function] int gnutls_pkcs7_set_crt (gnutls pkcs7 t pkcs7, gnutls x509 crt t

   crt)

crt: the certificate to be copied. This function will add a parsed certificate to the PKCS7 or RFC2630 certificate set. This is a wrapper function over gnutls_pkcs7_set_crt_raw() . Returns 0 on success.

gnutls x509 crl check issuer

[Function] int gnutls_x509_crl_check_issuer (gnutls x509 crl t cert,

gnutls x509 crt t issuer) issuer: is the certificate of a possible issuer This function will check if the given CRL was issued by the given issuer certificate. It will return true (1) if the given CRL was issued by the given issuer, and false (0) if not. A negative value is returned in case of an error.

gnutls x509 crl deinit

[Function] void gnutls_x509_crl_deinit (gnutls x509 crl t crl)

crl: The structure to be initialized This function will deinitialize a CRL structure.

gnutls x509 crl export

[Function] int gnutls_x509_crl_export (gnutls x509 crl t crl,

gnutls x509 crt fmt t format, void * output_data, size t *

   output_data_size)

crl: Holds the revocation list format: the format of output params. One of PEM or DER. output data: will contain a private key PEM or DER encoded output data size: holds the size of output data (and will be replaced by the actual size of parameters) This function will export the revocation list to DER or PEM format. If the buffer provided is not long enough to hold the output, then GNUTLS E SHORT MEMORY BUFFER will be returned. If the structure is PEM encoded, it will have a header of "BEGIN X509 CRL". Returns 0 on success, and a negative value on failure.

gnutls x509 crl get crt count

[Function] int gnutls_x509_crl_get_crt_count (gnutls x509 crl t crl)

crl: should contain a gnutls x509 crl t structure This function will return the number of revoked certificates in the given CRL. Returns a negative value on failure.

gnutls x509 crl get crt serial

[Function] int gnutls_x509_crl_get_crt_serial (gnutls x509 crl t crl, int

indx, unsigned char * serial, size t * serial_size, time t * t) crl: should contain a gnutls x509 crl t structure indx: the index of the certificate to extract (starting from 0) serial: where the serial number will be copied serial size: initially holds the size of serial t: if non null, will hold the time this certificate was revoked This function will return the serial number of the specified, by the index, revoked certificate. Returns a negative value on failure.

gnutls x509 crl get dn oid

[Function] int gnutls_x509_crl_get_dn_oid (gnutls x509 crl t crl, int indx,
   void * oid, size t * sizeof_oid)

crl: should contain a gnutls x509 crl t structure indx: Specifies which DN OID to send. Use zero to get the first one. oid: a pointer to a structure to hold the name (may be null) sizeof oid: initially holds the size of ’oid’ This function will extract the requested OID of the name of the CRL issuer, specified by the given index. If oid is null then only the size will be filled. Returns GNUTLS E SHORT MEMORY BUFFER if the provided buffer is not long enough, and in that case the sizeof oid will be updated with the required size. On success 0 is returned.

gnutls x509 crl get issuer dn by oid

[Function] int gnutls_x509_crl_get_issuer_dn_by_oid (gnutls x509 crl t

crl, const char * oid, int indx, unsigned int raw_flag, void * buf, size t *

   sizeof_buf)

crl: should contain a gnutls x509 crl t structure oid: holds an Object Identified in null terminated string indx: In case multiple same OIDs exist in the RDN, this specifies which to send. Use zero to get the first one. raw flag: If non zero returns the raw DER data of the DN part. buf : a pointer to a structure to hold the peer’s name (may be null) sizeof buf : initially holds the size of buf This function will extract the part of the name of the CRL issuer specified by the given OID. The output will be encoded as described in RFC2253. The output string will be ASCII or UTF-8 encoded, depending on the certificate data. Some helper macros with popular OIDs can be found in gnutls/x509.h If raw flag is zero, this function will only return known OIDs as text. Other OIDs will be DER encoded, as described in RFC2253 – in hex format with a ’\#’ prefix. You can check

about known OIDs using gnutls_x509_dn_oid_known().

If buf is null then only the size will be filled. Returns GNUTLS E SHORT MEMORY BUFFER if the provided buffer is not long enough, and in that case the sizeof buf will be updated with the required size, and 0 on success.

gnutls x509 crl get issuer dn

[Function] int gnutls_x509_crl_get_issuer_dn (const gnutls x509 crl t crl,
   char * buf, size t * sizeof_buf)

crl: should contain a gnutls x509 crl t structure buf : a pointer to a structure to hold the peer’s name (may be null) sizeof buf : initially holds the size of buf This function will copy the name of the CRL issuer in the provided buffer. The name will be in the form "C=xxxx,O=yyyy,CN=zzzz" as described in RFC2253. The output string will be ASCII or UTF-8 encoded, depending on the certificate data. If buf is null then only the size will be filled. Returns GNUTLS E SHORT MEMORY BUFFER if the provided buffer is not long enough, and in that case the sizeof buf will be updated with the required size, and 0 on success.

gnutls x509 crl get next update

[Function] time_t gnutls_x509_crl_get_next_update (gnutls x509 crl t crl)

crl: should contain a gnutls x509 crl t structure This function will return the time the next CRL will be issued. This field is optional in a CRL so it might be normal to get an error instead. Returns (time t)-1 on error.

gnutls x509 crl get signature algorithm

[Function] int gnutls_x509_crl_get_signature_algorithm (gnutls x509 crl t
   crl)

crl: should contain a gnutls x509 crl t structure This function will return a value of the gnutls sign algorithm t enumeration that is the signature algorithm. Returns a negative value on error.

gnutls x509 crl get this update

[Function] time_t gnutls_x509_crl_get_this_update (gnutls x509 crl t crl)

crl: should contain a gnutls x509 crl t structure This function will return the time this CRL was issued. Returns (time t)-1 on error.

gnutls x509 crl get version

[Function] int gnutls_x509_crl_get_version (gnutls x509 crl t crl)

crl: should contain a gnutls x509 crl t structure This function will return the version of the specified CRL. Returns a negative value on error.

gnutls x509 crl import

[Function] int gnutls_x509_crl_import (gnutls x509 crl t crl, const

gnutls datum t * data, gnutls x509 crt fmt t format) crl: The structure to store the parsed CRL. data: The DER or PEM encoded CRL. format: One of DER or PEM This function will convert the given DER or PEM encoded CRL to the native gnutls x509 crl t format. The output will be stored in ’crl’. If the CRL is PEM encoded it should have a header of "X509 CRL". Returns 0 on success.

gnutls x509 crl init

[Function] int gnutls_x509_crl_init (gnutls x509 crl t * crl)

crl: The structure to be initialized This function will initialize a CRL structure. CRL stands for Certificate Revocation List. A revocation list usually contains lists of certificate serial numbers that have been revoked by an Authority. The revocation lists are always signed with the authority’s private key. Returns 0 on success.

gnutls x509 crl set crt serial

[Function] int gnutls_x509_crl_set_crt_serial (gnutls x509 crl t crl, const
   void * serial, size t serial_size, time t revocation_time)

crl: should contain a gnutls x509 crl t structure serial: The revoked certificate’s serial number serial size: Holds the size of the serial field. revocation time: The time this certificate was revoked This function will set a revoked certificate’s serial number to the CRL. Returns 0 on success, or a negative value in case of an error.

gnutls x509 crl set crt

[Function] int gnutls_x509_crl_set_crt (gnutls x509 crl t crl,

gnutls x509 crt t crt, time t revocation_time) crl: should contain a gnutls x509 crl t structure crt: should contain a gnutls x509 crt t structure with the revoked certificate revocation time: The time this certificate was revoked This function will set a revoked certificate’s serial number to the CRL. Returns 0 on success, or a negative value in case of an error.

gnutls x509 crl set next update

[Function] int gnutls_x509_crl_set_next_update (gnutls x509 crl t crl,
   time t exp_time)

crl: should contain a gnutls x509 crl t structure exp time: The actual time This function will set the time this CRL will be updated. Returns 0 on success, or a negative value in case of an error.

gnutls x509 crl set this update

[Function] int gnutls_x509_crl_set_this_update (gnutls x509 crl t crl,
   time t act_time)

crl: should contain a gnutls x509 crl t structure act time: The actual time This function will set the time this CRL was issued. Returns 0 on success, or a negative value in case of an error.

gnutls x509 crl set version

[Function] int gnutls_x509_crl_set_version (gnutls x509 crl t crl, unsigned

   int version)

crl: should contain a gnutls x509 crl t structure version: holds the version number. For CRLv1 crls must be 1. This function will set the version of the CRL. This must be one for CRL version 1, and so on. The CRLs generated by gnutls should have a version number of 2. Returns 0 on success.

gnutls x509 crl sign2

[Function] int gnutls_x509_crl_sign2 (gnutls x509 crl t crl, gnutls x509 crt t issuer, gnutls x509 privkey t issuer_key, gnutls digest algorithm t dig, unsigned int flags)

crl: should contain a gnutls x509 crl t structure issuer: is the certificate of the certificate issuer issuer key: holds the issuer’s private key dig: The message digest to use. GNUTLS DIG SHA1 is the safe choice unless you know what you’re doing. flags: must be 0 This function will sign the CRL with the issuer’s private key, and will copy the issuer’s information into the CRL. This must be the last step in a certificate CRL since all the previously set parameters are now signed. Returns 0 on success.

gnutls x509 crl sign

[Function] int gnutls_x509_crl_sign (gnutls x509 crl t crl, gnutls x509 crt t

issuer, gnutls x509 privkey t issuer_key) crl: should contain a gnutls x509 crl t structure issuer: is the certificate of the certificate issuer issuer key: holds the issuer’s private key This function is the same a gnutls_x509_crl_sign2() with no flags, and SHA1 as the hash algorithm. Returns 0 on success.

gnutls x509 crl verify

[Function] int gnutls_x509_crl_verify (gnutls x509 crl t crl, const

gnutls x509 crt t * CA_list, int CA_list_length, unsigned int flags, unsigned int * verify)

crl: is the crl to be verified CA list: is a certificate list that is considered to be trusted one CA list length: holds the number of CA certificates in CA list flags: Flags that may be used to change the verification algorithm. Use OR of the gnutls certificate verify flags enumerations. verify: will hold the crl verification output. This function will try to verify the given crl and return its status. See gnutls_x509_ crt_list_verify() for a detailed description of return values. Returns 0 on success and a negative value in case of an error.

gnutls x509 crq deinit

[Function] void gnutls_x509_crq_deinit (gnutls x509 crq t crq)

crq: The structure to be initialized This function will deinitialize a CRL structure.

gnutls x509 crq export

[Function] int gnutls_x509_crq_export (gnutls x509 crq t crq,

gnutls x509 crt fmt t format, void * output_data, size t *

   output_data_size)

crq: Holds the request format: the format of output params. One of PEM or DER. output data: will contain a certificate request PEM or DER encoded output data size: holds the size of output data (and will be replaced by the actual size of parameters) This function will export the certificate request to a PKCS10 If the buffer provided is not long enough to hold the output, then GNUTLS E SHORT MEMORY BUFFER will be returned and *output data size will be updated. If the structure is PEM encoded, it will have a header of "BEGIN NEW CERTIFI-CATE REQUEST". Return value: In case of failure a negative value will be returned, and 0 on success.

gnutls x509 crq get attribute by oid

[Function] int gnutls_x509_crq_get_attribute_by_oid (gnutls x509 crq t

crq, const char * oid, int indx, void * buf, size t * sizeof_buf) crq: should contain a gnutls x509 crq t structure oid: holds an Object Identified in null terminated string indx: In case multiple same OIDs exist in the attribute list, this specifies which to send. Use zero to get the first one. buf : a pointer to a structure to hold the attribute data (may be null) sizeof buf : initially holds the size of buf This function will return the attribute in the certificate request specified by the given Object ID. The attribute will be DER encoded. Returns 0 on success.

gnutls x509 crq get challenge password

[Function] int gnutls_x509_crq_get_challenge_password (gnutls x509 crq t
   crq, char * pass, size t * sizeof_pass)

crq: should contain a gnutls x509 crq t structure pass: will hold a null terminated password sizeof pass: Initially holds the size of pass. This function will return the challenge password in the request. Returns 0 on success.

gnutls x509 crq get dn by oid

[Function] int gnutls_x509_crq_get_dn_by_oid (gnutls x509 crq t crq, const

char * oid, int indx, unsigned int raw_flag, void * buf, size t *

   sizeof_buf)

crq: should contain a gnutls x509 crq t structure oid: holds an Object Identified in null terminated string indx: In case multiple same OIDs exist in the RDN, this specifies which to send. Use zero to get the first one. raw flag: If non zero returns the raw DER data of the DN part. buf : a pointer to a structure to hold the name (may be null) sizeof buf : initially holds the size of buf This function will extract the part of the name of the Certificate request subject, specified by the given OID. The output will be encoded as described in RFC2253. The output string will be ASCII or UTF-8 encoded, depending on the certificate data. Some helper macros with popular OIDs can be found in gnutls/x509.h If raw flag is zero, this function will only return known OIDs as text. Other OIDs will be DER encoded, as described in RFC2253 – in hex format with a ’\#’ prefix. You can check

about known OIDs using gnutls_x509_dn_oid_known().

If buf is null then only the size will be filled. Returns GNUTLS E SHORT MEMORY BUFFER if the provided buffer is not long enough, and in that case the *sizeof buf will be updated with the required size. On success 0 is returned.

gnutls x509 crq get dn oid

[Function] int gnutls_x509_crq_get_dn_oid (gnutls x509 crq t crq, int indx,
   void * oid, size t * sizeof_oid)

crq: should contain a gnutls x509 crq t structure

indx: Specifies which DN OID to send. Use zero to get the first one.

oid: a pointer to a structure to hold the name (may be null)

sizeof oid: initially holds the size of oid

This function will extract the requested OID of the name of the Certificate request subject, specified by the given index.

If oid is null then only the size will be filled.

Returns GNUTLS E SHORT MEMORY BUFFER if the provided buffer is not long enough, and in that case the *sizeof oid will be updated with the required size. On success 0 is returned.

gnutls x509 crq get dn

[Function] int gnutls_x509_crq_get_dn (gnutls x509 crq t crq, char * buf,

   size t * sizeof_buf)

crq: should contain a gnutls x509 crq t structure

buf : a pointer to a structure to hold the name (may be null)

sizeof buf : initially holds the size of buf

This function will copy the name of the Certificate request subject in the provided buffer. The name will be in the form "C=xxxx,O=yyyy,CN=zzzz" as described in RFC2253. The output string will be ASCII or UTF-8 encoded, depending on the certificate data.

If buf is null then only the size will be filled.

Returns GNUTLS E SHORT MEMORY BUFFER if the provided buffer is not long enough, and in that case the *sizeof buf will be updated with the required size. On success 0 is returned.

gnutls x509 crq get pk algorithm

[Function] int gnutls_x509_crq_get_pk_algorithm (gnutls x509 crq t crq,

unsigned int * bits) crq: should contain a gnutls x509 crq t structure

bits: if bits is non null it will hold the size of the parameters’ in bits

This function will return the public key algorithm of a PKCS \10 certificate request.

If bits is non null, it should have enough size to hold the parameters size in bits. For RSA the bits returned is the modulus. For DSA the bits returned are of the public exponent.

Returns a member of the gnutls pk algorithm t enumeration on success, or a negative value on error.

gnutls x509 crq get version

[Function] int gnutls_x509_crq_get_version (gnutls x509 crq t crq)

crq: should contain a gnutls x509 crq t structure This function will return the version of the specified Certificate request. Returns a negative value on error.

gnutls x509 crq import

[Function] int gnutls_x509_crq_import (gnutls x509 crq t crq, const

gnutls datum t * data, gnutls x509 crt fmt t format) crq: The structure to store the parsed certificate request. data: The DER or PEM encoded certificate. format: One of DER or PEM This function will convert the given DER or PEM encoded Certificate to the native gnutls x509 crq t format. The output will be stored in cert. If the Certificate is PEM encoded it should have a header of "NEW CERTIFICATE REQUEST". Returns 0 on success.

gnutls x509 crq init

[Function] int gnutls_x509_crq_init (gnutls x509 crq t * crq)

crq: The structure to be initialized This function will initialize a PKCS10 certificate request structure. Returns 0 on success.

gnutls x509 crq set attribute by oid

[Function] int gnutls_x509_crq_set_attribute_by_oid (gnutls x509 crq t

crq, const char * oid, void * buf, size t sizeof_buf) crq: should contain a gnutls x509 crq t structure oid: holds an Object Identified in null terminated string buf : a pointer to a structure that holds the attribute data sizeof buf : holds the size of buf This function will set the attribute in the certificate request specified by the given Object ID. The attribute must be be DER encoded. Returns 0 on success.

gnutls x509 crq set challenge password

[Function] int gnutls_x509_crq_set_challenge_password (gnutls x509 crq t

crq, const char * pass) crq: should contain a gnutls x509 crq t structure pass: holds a null terminated password This function will set a challenge password to be used when revoking the request. Returns 0 on success.

gnutls x509 crq set dn by oid

[Function] int gnutls_x509_crq_set_dn_by_oid (gnutls x509 crq t crq, const

char * oid, unsigned int raw_flag, const void * data, unsigned int

   sizeof_data)

crq: should contain a gnutls x509 crq t structure oid: holds an Object Identifier in a null terminated string raw flag: must be 0, or 1 if the data are DER encoded data: a pointer to the input data sizeof data: holds the size of data This function will set the part of the name of the Certificate request subject, specified by the given OID. The input string should be ASCII or UTF-8 encoded. Some helper macros with popular OIDs can be found in gnutls/x509.h With this function you can only set the known OIDs. You can test for known OIDs using gnutls_x509_dn_oid_known(). For OIDs that are not known (by gnutls) you should properly DER encode your data, and call this function with raw flag set. Returns 0 on success.

gnutls x509 crq set key

[Function] int gnutls_x509_crq_set_key (gnutls x509 crq t crq,

gnutls x509 privkey t key) crq: should contain a gnutls x509 crq t structure key: holds a private key This function will set the public parameters from the given private key to the request. Only RSA keys are currently supported. Returns 0 on success.

gnutls x509 crq set version

[Function] int gnutls_x509_crq_set_version (gnutls x509 crq t crq, unsigned

   int version)

crq: should contain a gnutls x509 crq t structure version: holds the version number. For v1 Requests must be 1. This function will set the version of the certificate request. For version 1 requests this must be one. Returns 0 on success.

gnutls x509 crq sign2

[Function] int gnutls_x509_crq_sign2 (gnutls x509 crq t crq,

gnutls x509 privkey t key, gnutls digest algorithm t dig, unsigned int flags) crq: should contain a gnutls x509 crq t structure key: holds a private key dig: The message digest to use. GNUTLS DIG SHA1 is the safe choice unless you know what you’re doing.

flags: must be 0

This function will sign the certificate request with a private key. This must be the same key as the one used in gnutls_x509_crt_set_key() since a certificate request is self signed.

This must be the last step in a certificate request generation since all the previously set parameters are now signed.

Returns 0 on success.

gnutls x509 crq sign

[Function] int gnutls_x509_crq_sign (gnutls x509 crq t crq,

gnutls x509 privkey t key) crq: should contain a gnutls x509 crq t structure

key: holds a private key

This function is the same a gnutls_x509_crq_sign2() with no flags, and SHA1 as the hash algorithm.

Returns 0 on success.

gnutls x509 crt check hostname

[Function] int gnutls_x509_crt_check_hostname (gnutls x509 crt t cert, const
   char * hostname)

cert: should contain an gnutls x509 crt t structure

hostname: A null terminated string that contains a DNS name

This function will check if the given certificate’s subject matches the given hostname. This is a basic implementation of the matching described in RFC2818 (HTTPS), which takes into account wildcards, and the subject alternative name PKIX extension.

Returns non zero on success, and zero on failure.

gnutls x509 crt check issuer

[Function] int gnutls_x509_crt_check_issuer (gnutls x509 crt t cert,

gnutls x509 crt t issuer) cert: is the certificate to be checked

issuer: is the certificate of a possible issuer

This function will check if the given certificate was issued by the given issuer. It will return true (1) if the given certificate is issued by the given issuer, and false (0) if not.

A negative value is returned in case of an error.

gnutls x509 crt check revocation

[Function] int gnutls_x509_crt_check_revocation (gnutls x509 crt t cert,

const gnutls x509 crl t * crl_list, int crl_list_length) cert: should contain a gnutls x509 crt t structure crl list: should contain a list of gnutls x509 crl t structures crl list length: the length of the crl list This function will return check if the given certificate is revoked. It is assumed that the CRLs have been verified before. Returns 0 if the certificate is NOT revoked, and 1 if it is. A negative value is returned on error.

gnutls x509 crt cpy crl dist points

[Function] int gnutls_x509_crt_cpy_crl_dist_points (gnutls x509 crt t dst,

gnutls x509 crt t src) dst: should contain a gnutls x509 crt t structure src: the certificate where the dist points will be copied from This function will copy the CRL distribution points certificate extension, from the source to the destination certificate. This may be useful to copy from a CA certificate to issued ones. Returns 0 on success.

gnutls x509 crt deinit

[Function] void gnutls_x509_crt_deinit (gnutls x509 crt t cert)

cert: The structure to be initialized This function will deinitialize a CRL structure.

gnutls x509 crt export

[Function] int gnutls_x509_crt_export (gnutls x509 crt t cert,

gnutls x509 crt fmt t format, void * output_data, size t *

   output_data_size)

cert: Holds the certificate format: the format of output params. One of PEM or DER. output data: will contain a certificate PEM or DER encoded output data size: holds the size of output data (and will be replaced by the actual size of parameters) This function will export the certificate to DER or PEM format. If the buffer provided is not long enough to hold the output, then *output data size is updated and GNUTLS E SHORT MEMORY BUFFER will be returned. If the structure is PEM encoded, it will have a header of "BEGIN CERTIFICATE". Return value: In case of failure a negative value will be returned, and 0 on success.

gnutls x509 crt get activation time

[Function] time_t gnutls_x509_crt_get_activation_time (gnutls x509 crt t
   cert)

cert: should contain a gnutls x509 crt t structure This function will return the time this Certificate was or will be activated. Returns (time t)-1 on error.

gnutls x509 crt get authority key id

[Function] int gnutls_x509_crt_get_authority_key_id (gnutls x509 crt t

cert, void * ret, size t * ret_size, unsigned int * critical) cert: should contain a gnutls x509 crt t structure critical: will be non zero if the extension is marked as critical (may be null) This function will return the X.509v3 certificate authority’s key identifier. This is obtained by the X.509 Authority Key identifier extension field (2.5.29.35). Note that this function only returns the keyIdentifier field of the extension. Returns 0 on success and a negative value in case of an error.

gnutls x509 crt get ca status

[Function] int gnutls_x509_crt_get_ca_status (gnutls x509 crt t cert,

unsigned int * critical) cert: should contain a gnutls x509 crt t structure critical: will be non zero if the extension is marked as critical This function will return certificates CA status, by reading the basicConstraints X.509 extension (2.5.29.19). If the certificate is a CA a positive value will be returned, or zero if the certificate does not have CA flag set. A negative value may be returned in case of parsing error. If the certificate does not contain the basicConstraints extension GNUTLS E REQUESTED DATA NOT AVAILABLE will be returned.

gnutls x509 crt get crl dist points

[Function] int gnutls_x509_crt_get_crl_dist_points (gnutls x509 crt t

cert, unsigned int seq, void * ret, size t * ret_size, unsigned int *

   reason_flags, unsigned int * critical)

cert: should contain a gnutls x509 crt t structure seq: specifies the sequence number of the distribution point (0 for the first one, 1 for the second etc.) ret: is the place where the distribution point will be copied to ret size: holds the size of ret. reason flags: Revocation reasons flags. critical: will be non zero if the extension is marked as critical (may be null) This function will return the CRL distribution points (2.5.29.31), contained in the given certificate. reason_flags should be an ORed sequence of GNUTLS CRL REASON UNUSED, GNUTLS CRL REASON KEY COMPROMISE, GNUTLS CRL REASON CA COMPROMISE, GNUTLS CRL REASON AFFILIATION CHANGED, GNUTLS CRL REASON SUPERSEEDED, GNUTLS CRL REASON CESSATION OF OPERATION, GNUTLS CRL REASON CERTIFICATE HOLD, GNUTLS CRL REASON PRIVILEGE WITHDRAWN, GNUTLS CRL REASON AA COMPROMISE, or zero for all possible reasons. This is specified in X509v3 Certificate Extensions. GNUTLS will return the distribution point type, or a negative error code on error. Returns GNUTLS E SHORT MEMORY BUFFER and updates &ret size if &ret size is not enough to hold the distribution point, or the type of the distribution point if everything was ok. The type is one of the enumerated gnutls x509 subject alt name t. If the certificate does not have an Alternative name with the specified sequence number then returns GNUTLS E REQUESTED DATA NOT AVAILABLE;

gnutls x509 crt get dn by oid

[Function] int gnutls_x509_crt_get_dn_by_oid (gnutls x509 crt t cert, const

char * oid, int indx, unsigned int raw_flag, void * buf, size t *

   sizeof_buf)

cert: should contain a gnutls x509 crt t structure oid: holds an Object Identified in null terminated string indx: In case multiple same OIDs exist in the RDN, this specifies which to send. Use zero to get the first one. raw flag: If non zero returns the raw DER data of the DN part. buf : a pointer where the DN part will be copied (may be null). sizeof buf : initially holds the size of buf This function will extract the part of the name of the Certificate subject specified by the given OID. The output, if the raw flag is not used, will be encoded as described in RFC2253. Thus a string that is ASCII or UTF-8 encoded, depending on the certificate data. Some helper macros with popular OIDs can be found in gnutls/x509.h If raw flag is zero, this function will only return known OIDs as text. Other OIDs will be DER encoded, as described in RFC2253 – in hex format with a ’\#’ prefix. You can check

about known OIDs using gnutls_x509_dn_oid_known().

If buf is null then only the size will be filled. Returns GNUTLS E SHORT MEMORY BUFFER if the provided buffer is not long enough, and in that case the *sizeof buf will be updated with the required size. On success 0 is returned.

gnutls x509 crt get dn oid

[Function] int gnutls_x509_crt_get_dn_oid (gnutls x509 crt t cert, int indx,
   void * oid, size t * sizeof_oid)

cert: should contain a gnutls x509 crt t structure indx: This specifies which OID to return. Use zero to get the first one. oid: a pointer to a buffer to hold the OID (may be null) sizeof oid: initially holds the size of oid This function will extract the OIDs of the name of the Certificate subject specified by the given index. If oid is null then only the size will be filled. Returns GNUTLS E SHORT MEMORY BUFFER if the provided buffer is not long enough, and in that case the *sizeof oid will be updated with the required size. On success 0 is returned.

gnutls x509 crt get dn

[Function] int gnutls_x509_crt_get_dn (gnutls x509 crt t cert, char * buf,
   size t * sizeof_buf)

cert: should contain a gnutls x509 crt t structure buf : a pointer to a structure to hold the name (may be null) sizeof buf : initially holds the size of buf This function will copy the name of the Certificate in the provided buffer. The name will be in the form "C=xxxx,O=yyyy,CN=zzzz" as described in RFC2253. The output string will be ASCII or UTF-8 encoded, depending on the certificate data. If buf is null then only the size will be filled. Returns GNUTLS E SHORT MEMORY BUFFER if the provided buffer is not long enough, and in that case the *sizeof buf will be updated with the required size. On success 0 is returned.

gnutls x509 crt get expiration time

[Function] time_t gnutls_x509_crt_get_expiration_time (gnutls x509 crt t
   cert)

cert: should contain a gnutls x509 crt t structure This function will return the time this Certificate was or will be expired. Returns (time t)-1 on error.

gnutls x509 crt get extension by oid

[Function] int gnutls_x509_crt_get_extension_by_oid (gnutls x509 crt t

cert, const char * oid, int indx, void * buf, size t * sizeof_buf, unsigned

   int * critical)

cert: should contain a gnutls x509 crt t structure oid: holds an Object Identified in null terminated string indx: In case multiple same OIDs exist in the extensions, this specifies which to send. Use zero to get the first one. buf : a pointer to a structure to hold the name (may be null) sizeof buf : initially holds the size of buf critical: will be non zero if the extension is marked as critical This function will return the extension specified by the OID in the certificate. The extensions will be returned as binary data DER encoded, in the provided buffer. A negative value may be returned in case of parsing error. If the certificate does not contain the specified extension GNUTLS E REQUESTED DATA NOT AVAILABLE will be returned.

gnutls x509 crt get extension oid

[Function] int gnutls_x509_crt_get_extension_oid (gnutls x509 crt t cert,
   int indx, void * oid, size t * sizeof_oid)

cert: should contain a gnutls x509 crt t structure indx: Specifies which extension OID to send. Use zero to get the first one. oid: a pointer to a structure to hold the OID (may be null) sizeof oid: initially holds the size of oid This function will return the requested extension OID in the certificate. The extension OID will be stored as a string in the provided buffer. A negative value may be returned in case of parsing error. If your have reached the last extension available GNUTLS E REQUESTED DATA NOT AVAILABLE will be returned.

gnutls x509 crt get fingerprint

[Function] int gnutls_x509_crt_get_fingerprint (gnutls x509 crt t cert,

gnutls digest algorithm t algo, void * buf, size t * sizeof_buf) cert: should contain a gnutls x509 crt t structure algo: is a digest algorithm buf : a pointer to a structure to hold the fingerprint (may be null) sizeof buf : initially holds the size of buf This function will calculate and copy the certificate’s fingerprint in the provided buffer. If the buffer is null then only the size will be filled. Returns GNUTLS E SHORT MEMORY BUFFER if the provided buffer is not long enough, and in that case the *sizeof buf will be updated with the required size. On success 0 is returned.

gnutls x509 crt get issuer dn by oid

[Function] int gnutls_x509_crt_get_issuer_dn_by_oid (gnutls x509 crt t

cert, const char * oid, int indx, unsigned int raw_flag, void * buf, size t *

   sizeof_buf)

cert: should contain a gnutls x509 crt t structure oid: holds an Object Identified in null terminated string indx: In case multiple same OIDs exist in the RDN, this specifies which to send. Use zero to get the first one. raw flag: If non zero returns the raw DER data of the DN part. buf : a pointer to a structure to hold the name (may be null) sizeof buf : initially holds the size of buf This function will extract the part of the name of the Certificate issuer specified by the given OID. The output, if the raw flag is not used, will be encoded as described in RFC2253. Thus a string that is ASCII or UTF-8 encoded, depending on the certificate data. Some helper macros with popular OIDs can be found in gnutls/x509.h If raw flag is zero, this function will only return known OIDs as text. Other OIDs will be DER encoded, as described in RFC2253 – in hex format with a ’\#’ prefix. You can check

about known OIDs using gnutls_x509_dn_oid_known().

If buf is null then only the size will be filled. Returns GNUTLS E SHORT MEMORY BUFFER if the provided buffer is not long enough, and in that case the *sizeof buf will be updated with the required size. On success 0 is returned.

gnutls x509 crt get issuer dn oid

[Function] int gnutls_x509_crt_get_issuer_dn_oid (gnutls x509 crt t cert,
   int indx, void * oid, size t * sizeof_oid)

cert: should contain a gnutls x509 crt t structure indx: This specifies which OID to return. Use zero to get the first one. oid: a pointer to a buffer to hold the OID (may be null) sizeof oid: initially holds the size of oid This function will extract the OIDs of the name of the Certificate issuer specified by the given index. If oid is null then only the size will be filled. Returns GNUTLS E SHORT MEMORY BUFFER if the provided buffer is not long enough, and in that case the *sizeof oid will be updated with the required size. On success 0 is returned.

gnutls x509 crt get issuer dn

[Function] int gnutls_x509_crt_get_issuer_dn (gnutls x509 crt t cert, char *
   buf, size t * sizeof_buf)

cert: should contain a gnutls x509 crt t structure buf : a pointer to a structure to hold the name (may be null) sizeof buf : initially holds the size of buf This function will copy the name of the Certificate issuer in the provided buffer. The name will be in the form "C=xxxx,O=yyyy,CN=zzzz" as described in RFC2253. The output string will be ASCII or UTF-8 encoded, depending on the certificate data. If buf is null then only the size will be filled. Returns GNUTLS E SHORT MEMORY BUFFER if the provided buffer is not long enough, and in that case the *sizeof buf will be updated with the required size. On success 0 is returned.

gnutls x509 crt get key id

[Function] int gnutls_x509_crt_get_key_id (gnutls x509 crt t crt, unsigned

   int flags, unsigned char * output_data, size t * output_data_size)

crt: Holds the certificate flags: should be 0 for now output data: will contain the key ID output data size: holds the size of output data (and will be replaced by the actual size of parameters) This function will return a unique ID the depends on the public key parameters. This ID can be used in checking whether a certificate corresponds to the given private key. If the buffer provided is not long enough to hold the output, then *output data size is updated and GNUTLS E SHORT MEMORY BUFFER will be returned. The output will normally be a SHA-1 hash output, which is 20 bytes. Return value: In case of failure a negative value will be returned, and 0 on success.

gnutls x509 crt get key purpose oid

[Function] int gnutls_x509_crt_get_key_purpose_oid (gnutls x509 crt t

cert, int indx, void * oid, size t * sizeof_oid, unsigned int * critical) cert: should contain a gnutls x509 crt t structure indx: This specifies which OID to return. Use zero to get the first one. oid: a pointer to a buffer to hold the OID (may be null) sizeof oid: initially holds the size of oid This function will extract the key purpose OIDs of the Certificate specified by the given index. These are stored in the Extended Key Usage extension (2.5.29.37) See the GNUTLS KP * definitions for human readable names. If oid is null then only the size will be filled. Returns GNUTLS E SHORT MEMORY BUFFER if the provided buffer is not long enough, and in that case the *sizeof oid will be updated with the required size. On success 0 is returned.

gnutls x509 crt get key usage

[Function] int gnutls_x509_crt_get_key_usage (gnutls x509 crt t cert,

unsigned int * key_usage, unsigned int * critical) cert: should contain a gnutls x509 crt t structure key usage: where the key usage bits will be stored critical: will be non zero if the extension is marked as critical This function will return certificate’s key usage, by reading the keyUsage X.509 extension (2.5.29.15). The key usage value will ORed values of the: GNUTLS KEY DIGITAL SIGNATURE, GNUTLS KEY NON REPUDIATION, GNUTLS KEY KEY ENCIPHERMENT, GNUTLS KEY DATA ENCIPHERMENT, GNUTLS KEY KEY AGREEMENT, GNUTLS KEY KEY CERT SIGN, GNUTLS KEY CRL SIGN, GNUTLS KEY ENCIPHER ONLY, GNUTLS KEY DECIPHER ONLY.

A negative value may be returned in case of parsing error. If the certificate does not contain the keyUsage extension GNUTLS E REQUESTED DATA NOT AVAILABLE will be returned.

gnutls x509 crt get pk algorithm

[Function] int gnutls_x509_crt_get_pk_algorithm (gnutls x509 crt t cert,

unsigned int * bits) cert: should contain a gnutls x509 crt t structure

bits: if bits is non null it will hold the size of the parameters’ in bits

This function will return the public key algorithm of an X.509 certificate.

If bits is non null, it should have enough size to hold the parameters size in bits. For RSA the bits returned is the modulus. For DSA the bits returned are of the public exponent.

Returns a member of the gnutls pk algorithm t enumeration on success, or a negative value on error.

gnutls x509 crt get pk dsa raw

[Function] int gnutls_x509_crt_get_pk_dsa_raw (gnutls x509 crt t crt,

gnutls datum t * p, gnutls datum t * q, gnutls datum t * g, gnutls datum t *

   y)

crt: Holds the certificate

p: will hold the p

q: will hold the q

g: will hold the g

y: will hold the y

This function will export the DSA private key’s parameters found in the given certificate. The new parameters will be allocated using gnutls_malloc() and will be stored in the appropriate datum.

gnutls x509 crt get pk rsa raw

[Function] int gnutls_x509_crt_get_pk_rsa_raw (gnutls x509 crt t crt,

gnutls datum t * m, gnutls datum t * e) crt: Holds the certificate

m: will hold the modulus

e: will hold the public exponent

This function will export the RSA private key’s parameters found in the given structure. The new parameters will be allocated using gnutls_malloc() and will be stored in the appropriate datum.

gnutls x509 crt get serial

[Function] int gnutls_x509_crt_get_serial (gnutls x509 crt t cert, void *
   result, size t * result_size)

cert: should contain a gnutls x509 crt t structure

result: The place where the serial number will be copied

result size: Holds the size of the result field.

This function will return the X.509 certificate’s serial number. This is obtained by the X509 Certificate serialNumber field. Serial is not always a 32 or 64bit number. Some CAs use large serial numbers, thus it may be wise to handle it as something opaque.

Returns 0 on success and a negative value in case of an error.

gnutls x509 crt get signature algorithm

[Function] int gnutls_x509_crt_get_signature_algorithm (gnutls x509 crt t
   cert)

cert: should contain a gnutls x509 crt t structure

This function will return a value of the gnutls sign algorithm t enumeration that is the signature algorithm.

Returns a negative value on error.

gnutls x509 crt get subject alt name

[Function] int gnutls_x509_crt_get_subject_alt_name (gnutls x509 crt t

cert, unsigned int seq, void * ret, size t * ret_size, unsigned int *

   critical)

cert: should contain a gnutls x509 crt t structure

seq: specifies the sequence number of the alt name (0 for the first one, 1 for the second etc.)

ret: is the place where the alternative name will be copied to

ret size: holds the size of ret.

critical: will be non zero if the extension is marked as critical (may be null)

This function will return the alternative names, contained in the given certificate.

This is specified in X509v3 Certificate Extensions. GNUTLS will return the Alternative name (2.5.29.17), or a negative error code.

Returns GNUTLS E SHORT MEMORY BUFFER if &ret size is not enough to hold the alternative name. In that case &ret size will be updated. If everything was ok the type of alternative name is returned. The type is one of the enumerated gnutls x509 subject alt name t.

If the certificate does not have an Alternative name with the specified sequence number then returns GNUTLS E REQUESTED DATA NOT AVAILABLE;

gnutls x509 crt get subject key id

[Function] int gnutls_x509_crt_get_subject_key_id (gnutls x509 crt t cert,

void * ret, size t * ret_size, unsigned int * critical) cert: should contain a gnutls x509 crt t structure critical: will be non zero if the extension is marked as critical (may be null) This function will return the X.509v3 certificate’s subject key identifier. This is obtained by the X.509 Subject Key identifier extension field (2.5.29.14). Returns 0 on success and a negative value in case of an error.

gnutls x509 crt get version

[Function] int gnutls_x509_crt_get_version (gnutls x509 crt t cert)

cert: should contain a gnutls x509 crt t structure This function will return the version of the specified Certificate. Returns a negative value on error.

gnutls x509 crt import

[Function] int gnutls_x509_crt_import (gnutls x509 crt t cert, const

gnutls datum t * data, gnutls x509 crt fmt t format) cert: The structure to store the parsed certificate. data: The DER or PEM encoded certificate. format: One of DER or PEM This function will convert the given DER or PEM encoded Certificate to the native gnutls x509 crt t format. The output will be stored in cert. If the Certificate is PEM encoded it should have a header of "X509 CERTIFICATE", or "CERTIFICATE". Returns 0 on success.

gnutls x509 crt init

[Function] int gnutls_x509_crt_init (gnutls x509 crt t * cert)

cert: The structure to be initialized This function will initialize an X.509 certificate structure. Returns 0 on success.

gnutls x509 crt list import

[Function] int gnutls_x509_crt_list_import (gnutls x509 crt t * certs,

unsigned int * cert_max, const gnutls datum t * data, gnutls x509 crt fmt t format, unsigned int flags)

certs: The structures to store the parsed certificate. Must not be initialized. cert max: Initially must hold the maximum number of certs. It will be updated with the number of certs available. data: The PEM encoded certificate. format: One of DER or PEM. flags: must be zero or an OR’d sequence of gnutls certificate import flags. This function will convert the given PEM encoded certificate list to the native gnutls x509 crt t format. The output will be stored in certs. They will be automatically initialized. If the Certificate is PEM encoded it should have a header of "X509 CERTIFICATE", or "CERTIFICATE". Returns the number of certificates read or a negative error value.

gnutls x509 crt list verify

[Function] int gnutls_x509_crt_list_verify (const gnutls x509 crt t *

   cert_list, int cert_list_length, const gnutls x509 crt t * CA_list, int

CA_list_length, const gnutls x509 crl t * CRL_list, int

   CRL_list_length, unsigned int flags, unsigned int * verify)

cert list: is the certificate list to be verified cert list length: holds the number of certificate in cert list CA list: is the CA list which will be used in verification CA list length: holds the number of CA certificate in CA list CRL list: holds a list of CRLs. CRL list length: the length of CRL list. flags: Flags that may be used to change the verification algorithm. Use OR of the gnutls certificate verify flags enumerations. verify: will hold the certificate verification output. This function will try to verify the given certificate list and return its status. Note that expiration and activation dates are not checked by this function, you should check them using the appropriate functions. If no flags are specified (0), this function will use the basicConstraints (2.5.29.19) PKIX extension. This means that only a certificate authority is allowed to sign a certificate. You must also check the peer’s name in order to check if the verified certificate belongs to the actual peer. The certificate verification output will be put in verify and will be one or more of the gnutls certificate status t enumerated elements bitwise or’d. For a more detailed verification status use gnutls_x509_crt_verify() per list element. GNUTLS CERT INVALID: the certificate chain is not valid. GNUTLS CERT REVOKED: a certificate in the chain has been revoked. Returns 0 on success and a negative value in case of an error.

gnutls x509 crt set activation time

[Function] int gnutls_x509_crt_set_activation_time (gnutls x509 crt t
   cert, time t act_time)

cert: should contain a gnutls x509 crt t structure act time: The actual time This function will set the time this Certificate was or will be activated. Returns 0 on success, or a negative value in case of an error.

gnutls x509 crt set authority key id

[Function] int gnutls_x509_crt_set_authority_key_id (gnutls x509 crt t

cert, const void * id, size t id_size) cert: should contain a gnutls x509 crt t structure id: The key ID id size: Holds the size of the serial field. This function will set the X.509 certificate’s authority key ID extension. Only the keyIdentifier field can be set with this function. Returns 0 on success, or a negative value in case of an error.

gnutls x509 crt set ca status

[Function] int gnutls_x509_crt_set_ca_status (gnutls x509 crt t crt,

unsigned int ca) crt: should contain a gnutls x509 crt t structure ca: true(1) or false(0). Depending on the Certificate authority status. This function will set the basicConstraints certificate extension. Returns 0 on success.

gnutls x509 crt set crl dist points

[Function] int gnutls_x509_crt_set_crl_dist_points (gnutls x509 crt t crt,

gnutls x509 subject alt name t type, const void * data_string, unsigned int

   reason_flags)

crt: should contain a gnutls x509 crt t structure type: is one of the gnutls x509 subject alt name t enumerations data string: The data to be set reason flags: revocation reasons This function will set the CRL distribution points certificate extension. Returns 0 on success.

gnutls x509 crt set crq

[Function] int gnutls_x509_crt_set_crq (gnutls x509 crt t crt,

gnutls x509 crq t crq) crt: should contain a gnutls x509 crt t structure crq: holds a certificate request This function will set the name and public parameters from the given certificate request to the certificate. Only RSA keys are currently supported. Returns 0 on success.

gnutls x509 crt set dn by oid

[Function] int gnutls_x509_crt_set_dn_by_oid (gnutls x509 crt t crt, const

char * oid, unsigned int raw_flag, const void * name, unsigned int

   sizeof_name)

crt: should contain a gnutls x509 crt t structure oid: holds an Object Identifier in a null terminated string raw flag: must be 0, or 1 if the data are DER encoded name: a pointer to the name sizeof name: holds the size of name This function will set the part of the name of the Certificate subject, specified by the given OID. The input string should be ASCII or UTF-8 encoded. Some helper macros with popular OIDs can be found in gnutls/x509.h With this function you can only set the known OIDs. You can test for known OIDs using gnutls_x509_dn_oid_known(). For OIDs that are not known (by gnutls) you should properly DER encode your data, and call this function with raw flag set. Returns 0 on success.

gnutls x509 crt set expiration time

[Function] int gnutls_x509_crt_set_expiration_time (gnutls x509 crt t
   cert, time t exp_time)

cert: should contain a gnutls x509 crt t structure exp time: The actual time This function will set the time this Certificate will expire. Returns 0 on success, or a negative value in case of an error.

gnutls x509 crt set extension by oid

[Function] int gnutls_x509_crt_set_extension_by_oid (gnutls x509 crt t

crt, const char * oid, const void * buf, size t sizeof_buf, unsigned int

   critical)

crt: should contain a gnutls x509 crt t structure oid: holds an Object Identified in null terminated string buf : a pointer to a DER encoded data sizeof buf : holds the size of buf critical: should be non zero if the extension is to be marked as critical This function will set an the extension, by the specified OID, in the certificate. The extension data should be binary data DER encoded. Returns 0 on success and a negative value in case of an error.

gnutls x509 crt set issuer dn by oid

[Function] int gnutls_x509_crt_set_issuer_dn_by_oid (gnutls x509 crt t

crt, const char * oid, unsigned int raw_flag, const void * name, unsigned int

   sizeof_name)

crt: should contain a gnutls x509 crt t structure oid: holds an Object Identifier in a null terminated string raw flag: must be 0, or 1 if the data are DER encoded name: a pointer to the name sizeof name: holds the size of name This function will set the part of the name of the Certificate issuer, specified by the given OID. The input string should be ASCII or UTF-8 encoded. Some helper macros with popular OIDs can be found in gnutls/x509.h With this function you can only set the known OIDs. You can test for known OIDs using gnutls_x509_dn_oid_known(). For OIDs that are not known (by gnutls) you should properly DER encode your data, and call this function with raw flag set. Normally you do not need to call this function, since the signing operation will copy the signer’s name as the issuer of the certificate. Returns 0 on success.

gnutls x509 crt set key purpose oid

[Function] int gnutls_x509_crt_set_key_purpose_oid (gnutls x509 crt t

cert, const void * oid, unsigned int critical) cert: should contain a gnutls x509 crt t structure oid: a pointer to a null terminated string that holds the OID critical: Whether this extension will be critical or not This function will set the key purpose OIDs of the Certificate. These are stored in the Extended Key Usage extension (2.5.29.37) See the GNUTLS KP * definitions for human readable names. Subsequent calls to this function will append OIDs to the OID list. On success 0 is returned.

gnutls x509 crt set key usage

[Function] int gnutls_x509_crt_set_key_usage (gnutls x509 crt t crt,

unsigned int usage) crt: should contain a gnutls x509 crt t structure usage: an ORed sequence of the GNUTLS KEY * elements. This function will set the keyUsage certificate extension. Returns 0 on success.

gnutls x509 crt set key

[Function] int gnutls_x509_crt_set_key (gnutls x509 crt t crt,

gnutls x509 privkey t key) crt: should contain a gnutls x509 crt t structure key: holds a private key This function will set the public parameters from the given private key to the certificate. Only RSA keys are currently supported. Returns 0 on success.

gnutls x509 crt set serial

[Function] int gnutls_x509_crt_set_serial (gnutls x509 crt t cert, const void

   * serial, size t serial_size)

cert: should contain a gnutls x509 crt t structure serial: The serial number serial size: Holds the size of the serial field. This function will set the X.509 certificate’s serial number. Serial is not always a 32 or 64bit number. Some CAs use large serial numbers, thus it may be wise to handle it as something opaque. Returns 0 on success, or a negative value in case of an error.

gnutls x509 crt set subject alternative name

[Function] int gnutls_x509_crt_set_subject_alternative_name

(gnutls x509 crt t crt, gnutls x509 subject alt name t type, const char *

   data_string)

crt: should contain a gnutls x509 crt t structure type: is one of the gnutls x509 subject alt name t enumerations data string: The data to be set This function will set the subject alternative name certificate extension. Returns 0 on success.

gnutls x509 crt set subject key id

[Function] int gnutls_x509_crt_set_subject_key_id (gnutls x509 crt t cert,

const void * id, size t id_size) cert: should contain a gnutls x509 crt t structure id: The key ID id size: Holds the size of the serial field. This function will set the X.509 certificate’s subject key ID extension. Returns 0 on success, or a negative value in case of an error.

gnutls x509 crt set version

[Function] int gnutls_x509_crt_set_version (gnutls x509 crt t crt, unsigned

   int version)

crt: should contain a gnutls x509 crt t structure version: holds the version number. For X.509v1 certificates must be 1. This function will set the version of the certificate. This must be one for X.509 version 1, and so on. Plain certificates without extensions must have version set to one. Returns 0 on success.

gnutls x509 crt sign2

[Function] int gnutls_x509_crt_sign2 (gnutls x509 crt t crt, gnutls x509 crt t issuer, gnutls x509 privkey t issuer_key, gnutls digest algorithm t dig, unsigned int flags)

crt: should contain a gnutls x509 crt t structure issuer: is the certificate of the certificate issuer issuer key: holds the issuer’s private key dig: The message digest to use. GNUTLS DIG SHA1 is the safe choice unless you know what you’re doing. flags: must be 0 This function will sign the certificate with the issuer’s private key, and will copy the issuer’s information into the certificate. This must be the last step in a certificate generation since all the previously set parameters are now signed. Returns 0 on success.

gnutls x509 crt sign

[Function] int gnutls_x509_crt_sign (gnutls x509 crt t crt, gnutls x509 crt t

issuer, gnutls x509 privkey t issuer_key) crt: should contain a gnutls x509 crt t structure issuer: is the certificate of the certificate issuer issuer key: holds the issuer’s private key This function is the same a gnutls_x509_crt_sign2() with no flags, and SHA1 as the hash algorithm. Returns 0 on success.

gnutls x509 crt to xml

[Function] int gnutls_x509_crt_to_xml (gnutls x509 crt t cert, gnutls datum t

   * res, int detail)

cert: should contain a gnutls x509 crt t structure res: The datum that will hold the result detail: The detail level (must be GNUTLS XML SHOW ALL or GNUTLS XML NORMAL) This function will return the XML structures of the given X.509 certificate. The XML structures are allocated internally (with malloc) and stored into res. NOTE: This function is currently not implemented. See the NEWS entry for version 1.3.5. Returns a negative error code in case of an error.

gnutls x509 crt verify data

[Function] int gnutls_x509_crt_verify_data (gnutls x509 crt t crt, unsigned

int flags, const gnutls datum t * data, const gnutls datum t * signature) crt: Holds the certificate flags: should be 0 for now data: holds the data to be signed signature: contains the signature This function will verify the given signed data, using the parameters from the certificate. In case of a verification failure 0 is returned, and 1 on success.

gnutls x509 crt verify

[Function] int gnutls_x509_crt_verify (gnutls x509 crt t cert, const

gnutls x509 crt t * CA_list, int CA_list_length, unsigned int flags, unsigned int * verify)

cert: is the certificate to be verified CA list: is one certificate that is considered to be trusted one CA list length: holds the number of CA certificate in CA list flags: Flags that may be used to change the verification algorithm. Use OR of the gnutls certificate verify flags enumerations. verify: will hold the certificate verification output. This function will try to verify the given certificate and return its status. The verification output in this functions cannot be GNUTLS CERT NOT VALID. Returns 0 on success and a negative value in case of an error.

gnutls x509 dn oid known

[Function] int gnutls_x509_dn_oid_known (const char * oid)

oid: holds an Object Identifier in a null terminated string This function will inform about known DN OIDs. This is useful since functions like gnutls_x509_crt_set_dn_by_oid() use the information on known OIDs to properly encode their input. Object Identifiers that are not known are not encoded by these functions, and their input is stored directly into the ASN.1 structure. In that case of unknown OIDs, you have the responsibility of DER encoding your data. Returns 1 on known OIDs and 0 otherwise.

gnutls x509 privkey cpy

[Function] int gnutls_x509_privkey_cpy (gnutls x509 privkey t dst,

gnutls x509 privkey t src) dst: The destination key, which should be initialized. src: The source key This function will copy a private key from source to destination key.

gnutls x509 privkey deinit

[Function] void gnutls_x509_privkey_deinit (gnutls x509 privkey t key)

key: The structure to be initialized This function will deinitialize a private key structure.

gnutls x509 privkey export dsa raw

[Function] int gnutls_x509_privkey_export_dsa_raw (gnutls x509 privkey t

key, gnutls datum t * p, gnutls datum t * q, gnutls datum t * g, gnutls datum t * y, gnutls datum t * x)

p: will hold the p q: will hold the q g: will hold the g y: will hold the y x: will hold the x This function will export the DSA private key’s parameters found in the given structure. The new parameters will be allocated using gnutls_malloc() and will be stored in the appropriate datum.

gnutls x509 privkey export pkcs8

[Function] int gnutls_x509_privkey_export_pkcs8 (gnutls x509 privkey t key,

gnutls x509 crt fmt t format, const char * password, unsigned int flags,

   void * output_data, size t * output_data_size)

key: Holds the key format: the format of output params. One of PEM or DER. password: the password that will be used to encrypt the key. flags: an ORed sequence of gnutls pkcs encrypt flags t output data: will contain a private key PEM or DER encoded output data size: holds the size of output data (and will be replaced by the actual size of parameters) This function will export the private key to a PKCS8 structure. Currently only RSA keys can be exported since there is no documented standard for other keys. If the flags do not specify the encryption cipher, then the default 3DES (PBES2) will be used.

The password can be either ASCII or UTF-8 in the default PBES2 encryption schemas, or ASCII for the PKCS12 schemas. If the buffer provided is not long enough to hold the output, then *output data size is updated and GNUTLS E SHORT MEMORY BUFFER will be returned. If the structure is PEM encoded, it will have a header of "BEGIN ENCRYPTED PRIVATE KEY" or "BEGIN PRIVATE KEY" if encryption is not used. Return value: In case of failure a negative value will be returned, and 0 on success.

gnutls x509 privkey export rsa raw

[Function] int gnutls_x509_privkey_export_rsa_raw (gnutls x509 privkey t

key, gnutls datum t * m, gnutls datum t * e, gnutls datum t * d, gnutls datum t * p, gnutls datum t * q, gnutls datum t * u)

key: a structure that holds the rsa parameters m: will hold the modulus e: will hold the public exponent d: will hold the private exponent p: will hold the first prime (p) q: will hold the second prime (q) u: will hold the coefficient This function will export the RSA private key’s parameters found in the given structure. The new parameters will be allocated using gnutls_malloc() and will be stored in the appropriate datum.

gnutls x509 privkey export

[Function] int gnutls_x509_privkey_export (gnutls x509 privkey t key,

gnutls x509 crt fmt t format, void * output_data, size t *

   output_data_size)

key: Holds the key format: the format of output params. One of PEM or DER. output data: will contain a private key PEM or DER encoded output data size: holds the size of output data (and will be replaced by the actual size of parameters) This function will export the private key to a PKCS1 structure for RSA keys, or an integer sequence for DSA keys. The DSA keys are in the same format with the parameters used by openssl. If the buffer provided is not long enough to hold the output, then *output data size is updated and GNUTLS E SHORT MEMORY BUFFER will be returned. If the structure is PEM encoded, it will have a header of "BEGIN RSA PRIVATE KEY". Return value: In case of failure a negative value will be returned, and 0 on success.

gnutls x509 privkey fix

[Function] int gnutls_x509_privkey_fix (gnutls x509 privkey t key)

key: Holds the key This function will recalculate the secondary parameters in a key. In RSA keys, this can be the coefficient and exponent1,2. Return value: In case of failure a negative value will be returned, and 0 on success.

gnutls x509 privkey generate

[Function] int gnutls_x509_privkey_generate (gnutls x509 privkey t key,

gnutls pk algorithm t algo, unsigned int bits, unsigned int flags) key: should contain a gnutls x509 privkey t structure algo: is one of RSA or DSA. bits: the size of the modulus flags: unused for now. Must be 0. This function will generate a random private key. Note that this function must be called on an empty private key. Returns 0 on success or a negative value on error.

gnutls x509 privkey get key id

[Function] int gnutls_x509_privkey_get_key_id (gnutls x509 privkey t key,

unsigned int flags, unsigned char * output_data, size t *

   output_data_size)

key: Holds the key flags: should be 0 for now output data: will contain the key ID output data size: holds the size of output data (and will be replaced by the actual size of parameters) This function will return a unique ID the depends on the public key parameters. This ID can be used in checking whether a certificate corresponds to the given key. If the buffer provided is not long enough to hold the output, then *output data size is updated and GNUTLS E SHORT MEMORY BUFFER will be returned. The output will normally be a SHA-1 hash output, which is 20 bytes. Return value: In case of failure a negative value will be returned, and 0 on success.

gnutls x509 privkey get pk algorithm

[Function] int gnutls_x509_privkey_get_pk_algorithm (gnutls x509 privkey t
   key)

key: should contain a gnutls x509 privkey t structure This function will return the public key algorithm of a private key. Returns a member of the gnutls pk algorithm t enumeration on success, or a negative value on error.

gnutls x509 privkey import dsa raw

[Function] int gnutls_x509_privkey_import_dsa_raw (gnutls x509 privkey t

key, const gnutls datum t * p, const gnutls datum t * q, const gnutls datum t * g, const gnutls datum t * y, const gnutls datum t * x)

key: The structure to store the parsed key p: holds the p q: holds the q g: holds the g y: holds the y x: holds the x This function will convert the given DSA raw parameters to the native gnutls x509 privkey t format. The output will be stored in key.

gnutls x509 privkey import pkcs8

[Function] int gnutls_x509_privkey_import_pkcs8 (gnutls x509 privkey t key,

const gnutls datum t * data, gnutls x509 crt fmt t format, const char *

   password, unsigned int flags)

key: The structure to store the parsed key data: The DER or PEM encoded key. format: One of DER or PEM password: the password to decrypt the key (if it is encrypted). flags: 0 if encrypted or GNUTLS PKCS PLAIN if not encrypted. This function will convert the given DER or PEM encoded PKCS8 2.0 encrypted key to the native gnutls x509 privkey t format. The output will be stored in key. Currently only RSA keys can be imported, and flags can only be used to indicate an unencrypted key. The password can be either ASCII or UTF-8 in the default PBES2 encryption schemas, or ASCII for the PKCS12 schemas. If the Certificate is PEM encoded it should have a header of "ENCRYPTED PRI-VATE KEY", or "PRIVATE KEY". You only need to specify the flags if the key is DER encoded, since in that case the encryption status cannot be auto-detected. Returns 0 on success.

gnutls x509 privkey import rsa raw

[Function] int gnutls_x509_privkey_import_rsa_raw (gnutls x509 privkey t

key, const gnutls datum t * m, const gnutls datum t * e, const gnutls datum t * d, const gnutls datum t * p, const gnutls datum t * q, const gnutls datum t * u)

key: The structure to store the parsed key m: holds the modulus e: holds the public exponent d: holds the private exponent p: holds the first prime (p) q: holds the second prime (q) u: holds the coefficient This function will convert the given RSA raw parameters to the native gnutls x509 privkey t format. The output will be stored in key.

gnutls x509 privkey import

[Function] int gnutls_x509_privkey_import (gnutls x509 privkey t key, const

gnutls datum t * data, gnutls x509 crt fmt t format) key: The structure to store the parsed key data: The DER or PEM encoded certificate. format: One of DER or PEM This function will convert the given DER or PEM encoded key to the native gnutls x509 privkey t format. The output will be stored in key . If the key is PEM encoded it should have a header of "RSA PRIVATE KEY", or "DSA PRIVATE KEY". Returns 0 on success.

gnutls x509 privkey init

[Function] int gnutls_x509_privkey_init (gnutls x509 privkey t * key)

key: The structure to be initialized This function will initialize an private key structure. Returns 0 on success.

gnutls x509 privkey sign data

[Function] int gnutls_x509_privkey_sign_data (gnutls x509 privkey t key,

gnutls digest algorithm t digest, unsigned int flags, const gnutls datum t *

   data, void * signature, size t * signature_size)

key: Holds the key digest: should be MD5 or SHA1 flags: should be 0 for now data: holds the data to be signed signature: will contain the signature signature size: holds the size of signature (and will be replaced by the new size) This function will sign the given data using a signature algorithm supported by the private key. Signature algorithms are always used together with a hash functions. Different hash functions may be used for the RSA algorithm, but only SHA-1 for the DSA keys. If the buffer provided is not long enough to hold the output, then *signature size is updated and GNUTLS E SHORT MEMORY BUFFER will be returned. In case of failure a negative value will be returned, and 0 on success.

gnutls x509 privkey verify data

[Function] int gnutls_x509_privkey_verify_data (gnutls x509 privkey t key,

unsigned int flags, const gnutls datum t * data, const gnutls datum t *

   signature)

key: Holds the key flags: should be 0 for now data: holds the data to be signed signature: contains the signature This function will verify the given signed data, using the parameters in the private key. In case of a verification failure 0 is returned, and 1 on success.

gnutls x509 rdn get by oid

[Function] int gnutls_x509_rdn_get_by_oid (const gnutls datum t * idn, const

char * oid, int indx, unsigned int raw_flag, void * buf, size t *

   sizeof_buf)

idn: should contain a DER encoded RDN sequence oid: an Object Identifier indx: In case multiple same OIDs exist in the RDN indicates which to send. Use 0 for the first one. raw flag: If non zero then the raw DER data are returned. buf : a pointer to a structure to hold the peer’s name sizeof buf : holds the size of buf This function will return the name of the given Object identifier, of the RDN sequence. The name will be encoded using the rules from RFC2253. Returns GNUTLS E SHORT MEMORY BUFFER and updates *sizeof buf if the provided buffer is not long enough, and 0 on success.

gnutls x509 rdn get oid

[Function] int gnutls_x509_rdn_get_oid (const gnutls datum t * idn, int indx,

   void * buf, size t * sizeof_buf)

idn: should contain a DER encoded RDN sequence indx: Indicates which OID to return. Use 0 for the first one. This function will return the specified Object identifier, of the RDN sequence. Returns GNUTLS E SHORT MEMORY BUFFER and updates *sizeof buf if the provided buffer is not long enough, and 0 on success.

gnutls x509 rdn get

[Function] int gnutls_x509_rdn_get (const gnutls datum t * idn, char * buf,

   size t * sizeof_buf)

idn: should contain a DER encoded RDN sequence buf : a pointer to a structure to hold the peer’s name sizeof buf : holds the size of buf This function will return the name of the given RDN sequence. The name will be in the form "C=xxxx,O=yyyy,CN=zzzz" as described in RFC2253. If the provided buffer is not long enough, returns GNUTLS E SHORT MEMORY BUFFER and *sizeof buf will be updated. On success 0 is returned.

9.3 GnuTLS-extra functions

These functions are only available in the GPL version of the library called gnutls-extra. The prototypes for this library lie in ‘gnutls/extra.h’.

gnutls extra check version

[Function] const char * gnutls_extra_check_version (const char *
   req_version)

req version: the version to check Check that the version of the gnutls-extra library is at minimum the requested one and return the version string; return NULL if the condition is not satisfied. If a NULL is passed to this function, no check is done, but the version string is simply returned.

gnutls global init extra

[Function] int gnutls_global_init_extra ( void)

This function initializes the global state of gnutls-extra library to defaults. Returns zero on success. Note that gnutls_global_init() has to be called before this function. If this function is not called then the gnutls-extra library will not be usable.

9.4 OpenPGP functions

The following functions are to be used for OpenPGP certificate handling. Their prototypes lie in ‘gnutls/openpgp.h’. You need to link with ‘libgnutls-extra’ to be able to use these functions (see Section 9.3 [GnuTLS-extra functions], page 184).

gnutls certificate set openpgp key file

[Function] int gnutls_certificate_set_openpgp_key_file

(gnutls certificate credentials t res, const char * certfile, const char *

   keyfile)

res: the destination context to save the data. certfile: the file that contains the public key. keyfile: the file that contains the secret key. This funtion is used to load OpenPGP keys into the GnuTLS credentials structure. It doesn’t matter whether the keys are armored or but, but the files should only contain one key which should not be encrypted.

gnutls certificate set openpgp key mem

[Function] int gnutls_certificate_set_openpgp_key_mem

(gnutls certificate credentials t res, const gnutls datum t * cert, const gnutls datum t * key)

res: the destination context to save the data.

cert: the datum that contains the public key.

key: the datum that contains the secret key.

This funtion is used to load OpenPGP keys into the GnuTLS credential structure. It doesn’t matter whether the keys are armored or but, but the files should only contain one key which should not be encrypted.

gnutls certificate set openpgp keyring file

[Function] int gnutls_certificate_set_openpgp_keyring_file

(gnutls certificate credentials t c, const char * file) c: A certificate credentials structure

file: filename of the keyring.

The function is used to set keyrings that will be used internally by various OpenPGP functions. For example to find a key when it is needed for an operations. The keyring will also be used at the verification functions.

gnutls certificate set openpgp keyring mem

[Function] int gnutls_certificate_set_openpgp_keyring_mem

(gnutls certificate credentials t c, const opaque * data, size t dlen) c: A certificate credentials structure

data: buffer with keyring data.

dlen: length of data buffer.

The function is used to set keyrings that will be used internally by various OpenPGP functions. For example to find a key when it is needed for an operations. The keyring will also be used at the verification functions.

gnutls certificate set openpgp keyserver

[Function] int gnutls_certificate_set_openpgp_keyserver

(gnutls certificate credentials t res, const char * keyserver, int port) res: the destination context to save the data.

keyserver: is the key server address

port: is the key server port to connect to

This funtion will set a key server for use with openpgp keys. This key server will only be used if the peer sends a key fingerprint instead of a key in the handshake. Using a key server may delay the handshake process.

gnutls certificate set openpgp key

[Function] int gnutls_certificate_set_openpgp_key

(gnutls certificate credentials t res, gnutls openpgp key t key, gnutls openpgp privkey t pkey)

res: is an gnutls_certificate_credentials_t structure.

key: contains an openpgp public key pkey: is an openpgp private key This function sets a certificate/private key pair in the gnutls certificate credentials t structure. This function may be called more than once (in case multiple keys/certificates exist for the server).

gnutls certificate set openpgp trustdb

[Function] int gnutls_certificate_set_openpgp_trustdb

(gnutls certificate credentials t res, const char * trustdb) res: the destination context to save the data. trustdb: is the trustdb filename This funtion will set a GnuPG trustdb which will be used in key verification functions. Only version 3 trustdb files are supported.

gnutls openpgp key check hostname

[Function] int gnutls_openpgp_key_check_hostname (gnutls openpgp key t

key, const char * hostname) key: should contain an gnutls openpgp key t structure hostname: A null terminated string that contains a DNS name This function will check if the given key’s owner matches the given hostname. This is a basic implementation of the matching described in RFC2818 (HTTPS), which takes into account wildcards. Returns non zero on success, and zero on failure.

gnutls openpgp key deinit

[Function] void gnutls_openpgp_key_deinit (gnutls openpgp key t key)

key: The structure to be initialized This function will deinitialize a key structure.

gnutls openpgp key export

[Function] int gnutls_openpgp_key_export (gnutls openpgp key t key,

gnutls openpgp key fmt t format, void * output_data, size t *

   output_data_size)

key: Holds the key. format: One of gnutls openpgp key fmt t elements. output data: will contain the key base64 encoded or raw output data size: holds the size of output data (and will be replaced by the actual size of parameters) This function will convert the given key to RAW or Base64 format. If the buffer provided is not long enough to hold the output, then GNUTLS E SHORT MEMORY BUFFER will be returned. Returns 0 on success.

gnutls openpgp key get creation time

[Function] time_t gnutls_openpgp_key_get_creation_time

(gnutls openpgp key t key) key: the structure that contains the OpenPGP public key. Returns the timestamp when the OpenPGP key was created.

gnutls openpgp key get expiration time

[Function] time_t gnutls_openpgp_key_get_expiration_time

(gnutls openpgp key t key) key: the structure that contains the OpenPGP public key. Returns the time when the OpenPGP key expires. A value of ’0’ means that the key doesn’t expire at all.

gnutls openpgp key get fingerprint

[Function] int gnutls_openpgp_key_get_fingerprint (gnutls openpgp key t

key, void * fpr, size t * fprlen) key: the raw data that contains the OpenPGP public key. fpr: the buffer to save the fingerprint. fprlen: the integer to save the length of the fingerprint. Returns the fingerprint of the OpenPGP key. Depends on the algorithm, the fingerprint can be 16 or 20 bytes.

gnutls openpgp key get id

[Function] int gnutls_openpgp_key_get_id (gnutls openpgp key t key,

unsigned char keyid[8]) key: the structure that contains the OpenPGP public key. Returns the 64-bit keyID of the OpenPGP key.

gnutls openpgp key get key usage

[Function] int gnutls_openpgp_key_get_key_usage (gnutls openpgp key t key,

unsigned int * key_usage) key: should contain a gnutls openpgp key t structure key usage: where the key usage bits will be stored This function will return certificate’s key usage, by checking the key algorithm. The key usage value will ORed values of the: GNUTLS KEY DIGITAL SIGNATURE, GNUTLS KEY KEY ENCIPHERMENT.

A negative value may be returned in case of parsing error.

gnutls openpgp key get name

[Function] int gnutls_openpgp_key_get_name (gnutls openpgp key t key, int
   idx, char * buf, size t * sizeof_buf)

key: the structure that contains the OpenPGP public key. idx: the index of the ID to extract buf : a pointer to a structure to hold the name sizeof buf : holds the size of ’buf’ Extracts the userID from the parsed OpenPGP key. Returns 0 on success, and GNUTLS E REQUESTED DATA NOT AVAILABLE if the index of the ID does not exist.

gnutls openpgp key get pk algorithm

[Function] int gnutls_openpgp_key_get_pk_algorithm (gnutls openpgp key t

key, unsigned int * bits) key: is an OpenPGP key bits: if bits is non null it will hold the size of the parameters’ in bits This function will return the public key algorithm of an OpenPGP certificate. If bits is non null, it should have enough size to hold the parameters size in bits. For RSA the bits returned is the modulus. For DSA the bits returned are of the public exponent. Returns a member of the GNUTLS PKAlgorithm enumeration on success, or a negative value on error.

gnutls openpgp key get version

[Function] int gnutls_openpgp_key_get_version (gnutls openpgp key t key)

key: the structure that contains the OpenPGP public key. Extract the version of the OpenPGP key.

gnutls openpgp key import

[Function] int gnutls_openpgp_key_import (gnutls openpgp key t key, const

gnutls datum t * data, gnutls openpgp key fmt t format) key: The structure to store the parsed key. data: The RAW or BASE64 encoded key. format: One of gnutls openpgp key fmt t elements. This function will convert the given RAW or Base64 encoded key to the native gnutls openpgp key t format. The output will be stored in ’key’. Returns 0 on success.

gnutls openpgp key init

[Function] int gnutls_openpgp_key_init (gnutls openpgp key t * key)

key: The structure to be initialized This function will initialize an OpenPGP key structure. Returns 0 on success.

gnutls openpgp key to xml

[Function] int gnutls_openpgp_key_to_xml (gnutls openpgp key t key,

gnutls datum t * xmlkey, int ext) xmlkey: he datum struct to store the XML result. ext: extension mode (1/0), 1 means include key signatures and key data. This function will return the all OpenPGP key information encapsulated as a XML string.

gnutls openpgp key verify ring

[Function] int gnutls_openpgp_key_verify_ring (gnutls openpgp key t key,

gnutls openpgp keyring t keyring, unsigned int flags, unsigned int *

   verify)

key: the structure that holds the key. keyring: holds the keyring to check against flags: unused (should be 0) verify: will hold the certificate verification output. Verify all signatures in the key, using the given set of keys (keyring). The key verification output will be put in verify and will be one or more of the gnutls certificate status t enumerated elements bitwise or’d. GNUTLS CERT INVALID: A signature on the key is invalid. GNUTLS CERT REVOKED: The key has been revoked. Note that this function does not verify using any "web of trust". You may use GnuPG for that purpose, or any other external PGP application. Returns 0 on success.

gnutls openpgp key verify self

[Function] int gnutls_openpgp_key_verify_self (gnutls openpgp key t key,

unsigned int flags, unsigned int * verify) key: the structure that holds the key. flags: unused (should be 0) verify: will hold the key verification output. Verifies the self signature in the key. The key verification output will be put in verify and will be one or more of the gnutls certificate status t enumerated elements bitwise or’d. GNUTLS CERT INVALID: The self signature on the key is invalid. Returns 0 on success.

gnutls openpgp key verify trustdb

[Function] int gnutls_openpgp_key_verify_trustdb (gnutls openpgp key t

key, gnutls openpgp trustdb t trustdb, unsigned int flags, unsigned int *

   verify)

key: the structure that holds the key. trustdb: holds the trustdb to check against flags: unused (should be 0) verify: will hold the certificate verification output. Checks if the key is revoked or disabled, in the trustdb. The verification output will be put in verify and will be one or more of the gnutls certificate status t enumerated elements bitwise or’d. GNUTLS CERT INVALID: A signature on the key is invalid. GNUTLS CERT REVOKED: The key has been revoked. Note that this function does not verify using any "web of trust". You may use GnuPG for that purpose, or any other external PGP application. Returns 0 on success.

gnutls openpgp keyring check id

[Function] int gnutls_openpgp_keyring_check_id (gnutls openpgp keyring t

ring, const unsigned char keyid[8], unsigned int flags) ring: holds the keyring to check against flags: unused (should be 0) Check if a given key ID exists in the keyring. Returns 0 on success (if keyid exists) and a negative error code on failure.

gnutls openpgp keyring deinit

[Function] void gnutls_openpgp_keyring_deinit (gnutls openpgp keyring t
   keyring)

keyring: The structure to be initialized This function will deinitialize a CRL structure.

gnutls openpgp keyring import

[Function] int gnutls_openpgp_keyring_import (gnutls openpgp keyring t

keyring, const gnutls datum t * data, gnutls openpgp key fmt t format) keyring: The structure to store the parsed key. data: The RAW or BASE64 encoded keyring. format: One of gnutls openpgp keyring fmt elements. This function will convert the given RAW or Base64 encoded keyring to the native gnutls openpgp keyring t format. The output will be stored in ’keyring’. Returns 0 on success.

gnutls openpgp keyring init

[Function] int gnutls_openpgp_keyring_init (gnutls openpgp keyring t *

   keyring)

keyring: The structure to be initialized

This function will initialize an OpenPGP keyring structure.

Returns 0 on success.

gnutls openpgp privkey deinit

[Function] void gnutls_openpgp_privkey_deinit (gnutls openpgp privkey t
   key)

key: The structure to be initialized

This function will deinitialize a key structure.

gnutls openpgp privkey get pk algorithm

[Function] int gnutls_openpgp_privkey_get_pk_algorithm

(gnutls openpgp privkey t key, unsigned int * bits) key: is an OpenPGP key

bits: if bits is non null it will hold the size of the parameters’ in bits

This function will return the public key algorithm of an OpenPGP certificate.

If bits is non null, it should have enough size to hold the parameters size in bits. For RSA the bits returned is the modulus. For DSA the bits returned are of the public exponent.

Returns a member of the GNUTLS PKAlgorithm enumeration on success, or a negative value on error.

gnutls openpgp privkey import

[Function] int gnutls_openpgp_privkey_import (gnutls openpgp privkey t key,

const gnutls datum t * data, gnutls openpgp key fmt t format, const char * pass, unsigned int flags)

key: The structure to store the parsed key.

data: The RAW or BASE64 encoded key.

format: One of gnutls openpgp key fmt t elements.

pass: Unused for now

flags: should be zero

This function will convert the given RAW or Base64 encoded key to the native gnutls openpgp privkey t format. The output will be stored in ’key’.

Returns 0 on success.

gnutls openpgp privkey init

[Function] int gnutls_openpgp_privkey_init (gnutls openpgp privkey t * key)

key: The structure to be initialized

This function will initialize an OpenPGP key structure.

Returns 0 on success.

gnutls openpgp set recv key function

[Function] void gnutls_openpgp_set_recv_key_function (gnutls session t

session, gnutls openpgp recv key func func) session: a TLS session

func: the callback

This funtion will set a key retrieval function for OpenPGP keys. This callback is only useful in server side, and will be used if the peer sent a key fingerprint instead of a full key.

gnutls openpgp trustdb deinit

[Function] void gnutls_openpgp_trustdb_deinit (gnutls openpgp trustdb t
   trustdb)

trustdb: The structure to be initialized

This function will deinitialize a CRL structure.

gnutls openpgp trustdb import file

[Function] int gnutls_openpgp_trustdb_import_file

(gnutls openpgp trustdb t trustdb, const char * file) trustdb: The structure to store the parsed key.

file: The file that holds the trustdb.

This function will convert the given RAW or Base64 encoded trustdb to the native gnutls openpgp trustdb t format. The output will be stored in ’trustdb’.

Returns 0 on success.

gnutls openpgp trustdb init

[Function] int gnutls_openpgp_trustdb_init (gnutls openpgp trustdb t *

   trustdb)

trustdb: The structure to be initialized

This function will initialize an OpenPGP trustdb structure.

Returns 0 on success.

9.5 TLS Inner Application (TLS/IA) functions

The following functions are used for TLS Inner Application (TLS/IA). Their prototypes lie in ‘gnutls/extra.h’. You need to link with ‘libgnutls-extra’ to be able to use these functions (see Section 9.3 [GnuTLS-extra functions], page 184).

The typical control flow in an TLS/IA client (that would not require an Application Phase for resumed sessions) would be similar to the following:

int client_avp (gnuls_session_t *session, void *ptr,
                const char *last, size_t lastlen,
char **new, size_t *newlen)
{
...
}
...
int main ()
{
  gnutls_ia_client_credentials_t iacred;
...
  gnutls_init (&session, GNUTLS_CLIENT);
...
  /* Enable TLS/IA. */
  gnutls_ia_allocate_client_credentials(&iacred);
  gnutls_ia_set_client_avp_function(iacred, client_avp);
  gnutls_credentials_set (session, GNUTLS_CRD_IA, iacred);
...
  ret = gnutls_handshake (session);
  // Error handling...
...
  if (gnutls_ia_handshake_p (session))
    {
      ret = gnutls_ia_handshake (session);
      // Error handling...
...

See below for detailed descriptions of all the functions used above.

The function client_avp would have to be implemented by your application. The function is responsible for handling the AVP data. See gnutls_ia_set_client_avp_function below for more information on how that function should be implemented.

The control flow in a typical server is similar to the above, use gnutls_ia_server_

credentials_t instead of gnutls_ia_client_credentials_t, and replace the call to the

client functions with the corresponding server functions.

gnutls ia allocate client credentials

[Function] int gnutls_ia_allocate_client_credentials

(gnutls ia client credentials t * sc)

sc: is a pointer to an gnutls_ia_server_credentials_t structure.

This structure is complex enough to manipulate directly thus this helper function is provided in order to allocate it. Adding this credential to a session will enable TLS/IA, and will require an Application Phase after the TLS handshake (if the server support TLS/IA). Use gnutls_ia_ require_inner_phase() to toggle the TLS/IA mode. Returns 0 on success.

gnutls ia allocate server credentials

[Function] int gnutls_ia_allocate_server_credentials

(gnutls ia server credentials t * sc)

sc: is a pointer to an gnutls_ia_server_credentials_t structure.

This structure is complex enough to manipulate directly thus this helper function is provided in order to allocate it. Adding this credential to a session will enable TLS/IA, and will require an Application Phase after the TLS handshake (if the client support TLS/IA). Use gnutls_ia_ require_inner_phase() to toggle the TLS/IA mode. Returns 0 on success.

gnutls ia enable

[Function] void gnutls_ia_enable (gnutls session t session, int

   allow_skip_on_resume)

session: is a gnutls_session_t structure. allow skip on resume: non-zero if local party allows to skip the TLS/IA application phases for a resumed session. Specify whether we must advertise support for the TLS/IA extension during the handshake. At the client side, we always advertise TLS/IA if gnutls ia enable was called before the handshake; at the server side, we also require that the client has advertised that it wants to run TLS/IA before including the advertisement, as required by the protocol. Similarly, at the client side we always advertise that we allow TLS/IA to be skipped for resumed sessions if allow_skip_on_resume is non-zero; at the server side, we also require that the session is indeed resumable and that the client has also advertised that it allows TLS/IA to be skipped for resumed sessions. After the TLS handshake, call gnutls_ia_handshake_p() to find out whether both parties agreed to do a TLS/IA handshake, before calling gnutls_ia_handshake() or one of the lower level gnutls ia * functions. Specify whether we must advertise support for the TLS/IA extension during the handshake. At the client side, we always advertise TLS/IA if gnutls ia enable was called before the handshake; at the server side, we also require that the client has advertised that it wants to run TLS/IA before including the advertisement, as required by the protocol. Similarly, at the client side we always advertise that we allow TLS/IA to be skipped for resumed sessions if allow_skip_on_resume is non-zero; at the server side, we also require that the session is indeed resumable and that the client has also advertised that it allows TLS/IA to be skipped for resumed sessions. After the TLS handshake, call gnutls_ia_handshake_p() to find out whether both parties agreed to do a TLS/IA handshake, before calling gnutls_ia_handshake() or one of the lower level gnutls ia * functions.

gnutls ia endphase send

[Function] int gnutls_ia_endphase_send (gnutls session t session, int
   final_p)

session: is a gnutls_session_t structure. final p: Set iff this should signal the final phase. Send a TLS/IA end phase message. In the client, this should only be used to acknowledge an end phase message sent by the server. In the server, this can be called instead of gnutls_ia_send() if the server wishes to end an application phase. Return value: Return 0 on success, or an error code.

gnutls ia extract inner secret

[Function] void gnutls_ia_extract_inner_secret (gnutls session t session,
   char * buffer)

session: is a gnutls_session_t structure. buffer: pre-allocated buffer to hold 48 bytes of inner secret. Copy the 48 bytes large inner secret into the specified buffer This function is typically used after the TLS/IA handshake has concluded. The TLS/IA inner secret can be used as input to a PRF to derive session keys. Do not use the inner secret directly as a session key, because for a resumed session that does not include an application phase, the inner secret will be identical to the inner secret in the original session. It is important to include, for example, the client and server randomness when deriving a sesssion key from the inner secret.

gnutls ia free client credentials

[Function] void gnutls_ia_free_client_credentials

(gnutls ia client credentials t sc)

sc: is an gnutls_ia_client_credentials_t structure.

This structure is complex enough to manipulate directly thus this helper function is provided in order to free (deallocate) it.

gnutls ia free server credentials

[Function] void gnutls_ia_free_server_credentials

(gnutls ia server credentials t sc)

sc: is an gnutls_ia_server_credentials_t structure.

This structure is complex enough to manipulate directly thus this helper function is provided in order to free (deallocate) it.

gnutls ia generate challenge

[Function] int gnutls_ia_generate_challenge (gnutls session t session, size t
   buffer_size, char * buffer)

session: is a gnutls_session_t structure.

buffer size: size of output buffer.

buffer: pre-allocated buffer to contain buffer_size bytes of output.

Generate an application challenge that the client cannot control or predict, based on the TLS/IA inner secret.

Return value: Returns 0 on success, or an negative error code.

gnutls ia get client avp ptr

[Function] void * gnutls_ia_get_client_avp_ptr

(gnutls ia client credentials t cred)

cred: is a gnutls_ia_client_credentials_t structure.

Returns the pointer that will be provided to the TLS/IA callback function as the first argument.

gnutls ia get server avp ptr

[Function] void * gnutls_ia_get_server_avp_ptr

(gnutls ia server credentials t cred)

cred: is a gnutls_ia_client_credentials_t structure.

Returns the pointer that will be provided to the TLS/IA callback function as the first argument.

gnutls ia handshake p

[Function] int gnutls_ia_handshake_p (gnutls session t session)

session: is a gnutls_session_t structure.

Predicate to be used after gnutls_handshake() to decide whether to invoke gnutls_ ia_handshake(). Usable by both clients and servers.

Return value: non-zero if TLS/IA handshake is expected, zero otherwise.

gnutls ia handshake

[Function] int gnutls_ia_handshake (gnutls session t session)

session: is a gnutls_session_t structure.

Perform a TLS/IA handshake. This should be called after gnutls_handshake() iff

gnutls_ia_handshake_p().

Return 0 on success, or an error code.

gnutls ia permute inner secret

[Function] int gnutls_ia_permute_inner_secret (gnutls session t session,
   size t session_keys_size, const char * session_keys)

session: is a gnutls_session_t structure. session keys size: Size of generated session keys (0 if none). session keys: Generated session keys, used to permute inner secret (NULL if none). Permute the inner secret using the generated session keys. This can be called in the TLS/IA AVP callback to mix any generated session keys with the TLS/IA inner secret. Return value: Return zero on success, or a negative error code.

gnutls ia recv

[Function] ssize_t gnutls_ia_recv (gnutls session t session, char * data,
   size t sizeofdata)

session: is a gnutls_session_t structure. data: the buffer that the data will be read into, must hold >= 12 bytes. sizeofdata: the number of requested bytes, must be >= 12. Receive TLS/IA data. This function has the similar semantics with recv(). The only difference is that is accepts a GNUTLS session, and uses different error codes. If the server attempt to finish an application phase, this function will return

GNUTLS_E_WARNING_IA_IPHF_RECEIVED or GNUTLS_E_WARNING_IA_FPHF_RECEIVED.

The caller should then invoke gnutls_ia_verify_endphase(), and if it runs the client side, also send an endphase message of its own using gnutls ia endphase send. If EINTR is returned by the internal push function (the default is code{recv()}) then GNUTLS E INTERRUPTED will be returned. If GNUTLS E INTERRUPTED or GNUTLS E AGAIN is returned, you must call this function again, with the same parameters; alternatively you could provide a NULL pointer for data, and 0 for size. Returns the number of bytes received. A negative error code is returned in case of

an error. The GNUTLS_E_WARNING_IA_IPHF_RECEIVED and GNUTLS_E_WARNING_IA_

FPHF_RECEIVED errors are returned when an application phase finished message has been sent by the server.

gnutls ia send

[Function] ssize_t gnutls_ia_send (gnutls session t session, const char * data,

   size t sizeofdata)

session: is a gnutls_session_t structure. data: contains the data to send sizeofdata: is the length of the data Send TLS/IA application payload data. This function has the similar semantics with send(). The only difference is that is accepts a GNUTLS session, and uses different error codes.

The TLS/IA protocol is synchronous, so you cannot send more than one packet at a time. The client always send the first packet.

To finish an application phase in the server, use gnutls_ia_endphase_send(). The client cannot end an application phase unilaterally; rather, a client is required to respond with an endphase of its own if gnutls ia recv indicates that the server has sent one.

If the EINTR is returned by the internal push function (the default is send()} then

GNUTLS_E_INTERRUPTED will be returned. If GNUTLS_E_INTERRUPTED or GNUTLS_E_

AGAIN is returned, you must call this function again, with the same parameters; alternatively you could provide a NULL pointer for data, and 0 for size.

Returns the number of bytes sent, or a negative error code.

gnutls ia set client avp function

[Function] void gnutls_ia_set_client_avp_function

(gnutls ia client credentials t cred, gnutls ia avp func avp_func)

cred: is a gnutls_ia_client_credentials_t structure.

avp func: is the callback function

Set the TLS/IA AVP callback handler used for the session.

The AVP callback is called to process AVPs received from the server, and to get a new AVP to send to the server.

The callback’s function form is: int (*avp func) (gnutls session t session, void *ptr, const char *last, size t lastlen, char **next, size t *nextlen);

The session parameter is the gnutls_session_t structure corresponding to the current session. The ptr parameter is the application hook pointer, set through gnutls_ia_set_client_avp_ptr(). The AVP received from the server is present in last of lastlen size, which will be NULL on the first invocation. The newly allocated output AVP to send to the server should be placed in *next of *nextlen size.

The callback may invoke gnutls_ia_permute_inner_secret() to mix any generated session keys with the TLS/IA inner secret.

Return 0 (GNUTLS_IA_APPLICATION_PAYLOAD) on success, or a negative error code to abort the TLS/IA handshake.

Note that the callback must use allocate the next parameter using gnutls_malloc(), because it is released via gnutls_free() by the TLS/IA handshake function.

gnutls ia set client avp ptr

[Function] void gnutls_ia_set_client_avp_ptr (gnutls ia client credentials t

   cred, void * ptr)
cred: is a gnutls_ia_client_credentials_t structure.

ptr: is the pointer

Sets the pointer that will be provided to the TLS/IA callback function as the first argument.

gnutls ia set server avp function

[Function] void gnutls_ia_set_server_avp_function

(gnutls ia server credentials t cred, gnutls ia avp func avp_func)

cred: is a gnutls_ia_server_credentials_t structure.

Set the TLS/IA AVP callback handler used for the session. The callback’s function form is: int (*avp func) (gnutls session t session, void *ptr, const char *last, size t lastlen, char **next, size t *nextlen); The session parameter is the gnutls_session_t structure corresponding to the current session. The ptr parameter is the application hook pointer, set through gnutls_ia_set_server_avp_ptr(). The AVP received from the client is present in last of lastlen size. The newly allocated output AVP to send to the client should be placed in *next of *nextlen size. The AVP callback is called to process incoming AVPs from the client, and to get a new AVP to send to the client. It can also be used to instruct the TLS/IA handshake to do go into the Intermediate or Final phases. It return a negative error code, or an

gnutls_ia_apptype_t message type.

The callback may invoke gnutls_ia_permute_inner_secret() to mix any generated session keys with the TLS/IA inner secret. Specifically, return GNUTLS_IA_APPLICATION_PAYLOAD (0) to send another AVP to

the client, return GNUTLS_IA_INTERMEDIATE_PHASE_FINISHED (1) to indicate that an

IntermediatePhaseFinished message should be sent, and return GNUTLS_IA_FINAL_ PHASE_FINISHED (2) to indicate that an FinalPhaseFinished message should be sent. In the last two cases, the contents of the next and nextlen parameter is not used. Note that the callback must use allocate the next parameter using gnutls_malloc(), because it is released via gnutls_free() by the TLS/IA handshake function.

gnutls ia set server avp ptr

[Function] void gnutls_ia_set_server_avp_ptr (gnutls ia server credentials t

   cred, void * ptr)
cred: is a gnutls_ia_client_credentials_t structure.

ptr: is the pointer Sets the pointer that will be provided to the TLS/IA callback function as the first argument.

gnutls ia verify endphase

[Function] int gnutls_ia_verify_endphase (gnutls session t session, const
   char * checksum)

session: is a gnutls_session_t structure. checksum: 12-byte checksum data, received from gnutls_ia_recv(). Verify TLS/IA end phase checksum data. If verification fails, the GNUTLS_A_INNER_ APPLICATION_VERIFICATION alert is sent to the other sie.

This function is called when gnutls_ia_recv() return GNUTLS_E_WARNING_IA_IPHF_
RECEIVED or GNUTLS_E_WARNING_IA_FPHF_RECEIVED.

Return value: Return 0 on successful verification, or an error code. If the checksum verification of the end phase message fails, GNUTLS_E_IA_VERIFY_FAILED is returned.

9.6 Error codes and descriptions

The error codes used throughout the library are described below. The return code GNUTLS_ E_SUCCESS indicate successful operation, and is guaranteed to have the value 0, so you can use it in logical expressions.

GNUTLS_E_AGAIN:

Function was interrupted.

GNUTLS_E_ASN1_DER_ERROR:

ASN1 parser: Error in DER parsing.

GNUTLS_E_ASN1_DER_OVERFLOW:

ASN1 parser: Overflow in DER parsing.

GNUTLS_E_ASN1_ELEMENT_NOT_FOUND:

ASN1 parser: Element was not found.

GNUTLS_E_ASN1_GENERIC_ERROR:

ASN1 parser: Generic parsing error.

GNUTLS_E_ASN1_IDENTIFIER_NOT_FOUND:

ASN1 parser: Identifier was not found

GNUTLS_E_ASN1_SYNTAX_ERROR:

ASN1 parser: Syntax error.

GNUTLS_E_ASN1_TAG_ERROR:

ASN1 parser: Error in TAG.

GNUTLS_E_ASN1_TAG_IMPLICIT:

ASN1 parser: error in implicit tag

GNUTLS_E_ASN1_TYPE_ANY_ERROR:

ASN1 parser: Error in type ’ANY’.

GNUTLS_E_ASN1_VALUE_NOT_FOUND:

ASN1 parser: Value was not found.

GNUTLS_E_ASN1_VALUE_NOT_VALID:

ASN1 parser: Value is not valid.

GNUTLS_E_BASE64_DECODING_ERROR:

Base64 decoding error.

GNUTLS_E_BASE64_ENCODING_ERROR:

Base64 encoding error.

GNUTLS_E_CERTIFICATE_ERROR:

Error in the certificate.

GNUTLS_E_CERTIFICATE_KEY_MISMATCH:

The certificate and the given key do not match.

GNUTLS_E_COMPRESSION_FAILED:

Compression of the TLS record packet has failed.

GNUTLS_E_CONSTRAINT_ERROR:

Some constraint limits were reached.

GNUTLS_E_DB_ERROR:

Error in Database backend.

GNUTLS_E_DECOMPRESSION_FAILED:

Decompression of the TLS record packet has failed.

GNUTLS_E_DECRYPTION_FAILED:

Decryption has failed.

GNUTLS_E_DH_PRIME_UNACCEPTABLE:

The Diffie Hellman prime sent by the server is not acceptable (not long enough).

GNUTLS_E_ENCRYPTION_FAILED:

Encryption has failed.

GNUTLS_E_ERROR_IN_FINISHED_PACKET:

An error was encountered at the TLS Finished packet calculation.

GNUTLS_E_EXPIRED:

The requested session has expired.

GNUTLS_E_FATAL_ALERT_RECEIVED:

A TLS fatal alert has been received.

GNUTLS_E_FILE_ERROR:

Error while reading file.

GNUTLS_E_GOT_APPLICATION_DATA:

TLS Application data were received, while expecting handshake data.

GNUTLS_E_HASH_FAILED:

Hashing has failed.

GNUTLS_E_IA_VERIFY_FAILED:

Verifying TLS/IA phase checksum failed

GNUTLS_E_ILLEGAL_SRP_USERNAME:

The SRP username supplied is illegal.

GNUTLS_E_INCOMPATIBLE_GCRYPT_LIBRARY:

The gcrypt library version is too old.

GNUTLS_E_INCOMPATIBLE_LIBTASN1_LIBRARY:

The tasn1 library version is too old.

GNUTLS_E_INIT_LIBEXTRA:

The initialization of GnuTLS-extra has failed.

GNUTLS_E_INSUFFICIENT_CREDENTIALS:

Insufficient credentials for that request.

GNUTLS_E_INTERNAL_ERROR:

GnuTLS internal error.

GNUTLS_E_INTERRUPTED:

Function was interrupted.

GNUTLS_E_INVALID_PASSWORD:

The given password contains invalid characters.

GNUTLS_E_INVALID_REQUEST:

The request is invalid.

GNUTLS_E_INVALID_SESSION:

The specified session has been invalidated for some reason.

GNUTLS_E_KEY_USAGE_VIOLATION:

Key usage violation in certificate has been detected.

GNUTLS_E_LARGE_PACKET:

A large TLS record packet was received.

GNUTLS_E_LIBRARY_VERSION_MISMATCH:

The GnuTLS library version does not match the GnuTLS-extra library version.

GNUTLS_E_LZO_INIT_FAILED:

The initialization of LZO has failed.

GNUTLS_E_MAC_VERIFY_FAILED:

The Message Authentication Code verification failed.

GNUTLS_E_MEMORY_ERROR:

Internal error in memory allocation.

GNUTLS_E_MPI_PRINT_FAILED:

Could not export a large integer.

GNUTLS_E_MPI_SCAN_FAILED:

The scanning of a large integer has failed.

GNUTLS_E_NO_CERTIFICATE_FOUND:

The peer did not send any certificate.

GNUTLS_E_NO_CIPHER_SUITES:

No supported cipher suites have been found.

GNUTLS_E_NO_COMPRESSION_ALGORITHMS:

No supported compression algorithms have been found.

GNUTLS_E_NO_TEMPORARY_DH_PARAMS:

No temporary DH parameters were found.

GNUTLS_E_NO_TEMPORARY_RSA_PARAMS:

No temporary RSA parameters were found.

GNUTLS_E_OPENPGP_FINGERPRINT_UNSUPPORTED:

The OpenPGP fingerprint is not supported.

GNUTLS_E_OPENPGP_GETKEY_FAILED:

Could not get OpenPGP key.

GNUTLS_E_OPENPGP_KEYRING_ERROR:

Error loading the keyring.

GNUTLS_E_OPENPGP_TRUSTDB_VERSION_UNSUPPORTED:

The specified GnuPG TrustDB version is not supported. TrustDB v4 is supported.

GNUTLS_E_PKCS1_WRONG_PAD:

Wrong padding in PKCS1 packet.

GNUTLS_E_PK_DECRYPTION_FAILED:

Public key decryption has failed.

GNUTLS_E_PK_ENCRYPTION_FAILED:

Public key encryption has failed.

GNUTLS_E_PK_SIGN_FAILED:

Public key signing has failed.

GNUTLS_E_PK_SIG_VERIFY_FAILED:

Public key signature verification has failed.

GNUTLS_E_PULL_ERROR:

Error in the pull function.

GNUTLS_E_PUSH_ERROR:

Error in the push function.

GNUTLS_E_RANDOM_FAILED:

Failed to acquire random data.

GNUTLS_E_RECEIVED_ILLEGAL_EXTENSION:

An illegal TLS extension was received.

GNUTLS_E_RECEIVED_ILLEGAL_PARAMETER:

An illegal parameter has been received.

GNUTLS_E_RECORD_LIMIT_REACHED:

The upper limit of record packet sequence numbers has been reached. Wow!

GNUTLS_E_REHANDSHAKE:

Rehandshake was requested by the peer.

GNUTLS_E_REQUESTED_DATA_NOT_AVAILABLE:

The requested data were not available.

GNUTLS_E_SHORT_MEMORY_BUFFER:

The given memory buffer is too short to hold parameters.

GNUTLS_E_SRP_PWD_ERROR:

Error in password file.

GNUTLS_E_SRP_PWD_PARSING_ERROR:

Parsing error in password file.

GNUTLS_E_SUCCESS:

Success.

GNUTLS_E_TOO_MANY_EMPTY_PACKETS:

Too many empty record packets have been received.

GNUTLS_E_UNEXPECTED_HANDSHAKE_PACKET:

An unexpected TLS handshake packet was received.

GNUTLS_E_UNEXPECTED_PACKET:

An unexpected TLS packet was received.

GNUTLS_E_UNEXPECTED_PACKET_LENGTH:

A TLS packet with unexpected length was received.

GNUTLS_E_UNKNOWN_CIPHER_SUITE:

Could not negotiate a supported cipher suite.

GNUTLS_E_UNKNOWN_CIPHER_TYPE:

The cipher type is unsupported.

GNUTLS_E_UNKNOWN_COMPRESSION_ALGORITHM:

Could not negotiate a supported compression method.

GNUTLS_E_UNKNOWN_HASH_ALGORITHM:

The hash algorithm is unknown.

GNUTLS_E_UNKNOWN_PKCS_BAG_TYPE:

The PKCS structure’s bag type is unknown.

GNUTLS_E_UNKNOWN_PKCS_CONTENT_TYPE:

The PKCS structure’s content type is unknown.

GNUTLS_E_UNKNOWN_PK_ALGORITHM:

An unknown public key algorithm was encountered.

GNUTLS_E_UNSUPPORTED_CERTIFICATE_TYPE:

The certificate type is not supported.

GNUTLS_E_UNSUPPORTED_VERSION_PACKET:

A record packet with illegal version was received.

GNUTLS_E_UNWANTED_ALGORITHM:

An algorithm that is not enabled was negotiated.

GNUTLS_E_WARNING_ALERT_RECEIVED:

A TLS warning alert has been received.

GNUTLS_E_WARNING_IA_FPHF_RECEIVED:

Received a TLS/IA Final Phase Finished message

GNUTLS_E_WARNING_IA_IPHF_RECEIVED:

Received a TLS/IA Intermediate Phase Finished message

GNUTLS_E_X509_UNKNOWN_SAN:

Unknown Subject Alternative name in X.509 certificate.

GNUTLS_E_X509_UNSUPPORTED_ATTRIBUTE:

The certificate has unsupported attributes.

GNUTLS_E_X509_UNSUPPORTED_CRITICAL_EXTENSION:

Unsupported critical extension in X.509 certificate.

GNUTLS_E_X509_UNSUPPORTED_OID:

The OID is not supported.

Chapter 10: Certificate to XML convertion functions 206

10 Certificate to XML convertion functions

This appendix contains some example output of the XML convertion functions: • [gnutls x509 crt to xml], page 176

• [gnutls openpgp key to xml], page 189

10.1 An X.509 certificate

<?xml version="1.0" encoding="UTF-8"?>
<gnutls:x509:certificate version="1.1">
<certificate type="SEQUENCE">
<tbsCertificate type="SEQUENCE">
<version type="INTEGER" encoding="HEX">02</version>
<serialNumber type="INTEGER" encoding="HEX">01</serialNumber>
<signature type="SEQUENCE">
<algorithm type="OBJECT ID">1.2.840.113549.1.1.4</algorithm>
<parameters type="ANY">
<md5WithRSAEncryption encoding="HEX">0500</md5WithRSAEncryption>
</parameters>
</signature>
<issuer type="CHOICE">
<rdnSequence type="SEQUENCE OF">
<unnamed1 type="SET OF">
<unnamed1 type="SEQUENCE">
<type type="OBJECT ID">2.5.4.6</type>
<value type="ANY">
  <X520countryName>GR</X520countryName>
</value>
</unnamed1>
</unnamed1>
<unnamed2 type="SET OF">
<unnamed1 type="SEQUENCE">
<type type="OBJECT ID">2.5.4.8</type>
<value type="ANY">
  <X520StateOrProvinceName>Attiki</X520StateOrProvinceName>
</value>
</unnamed1>
</unnamed2>
<unnamed3 type="SET OF">
<unnamed1 type="SEQUENCE">
<type type="OBJECT ID">2.5.4.7</type>
<value type="ANY">
  <X520LocalityName>Athina</X520LocalityName>
</value>
</unnamed1>
</unnamed3>
<unnamed4 type="SET OF">
<unnamed1 type="SEQUENCE">
<type type="OBJECT ID">2.5.4.10</type>
<value type="ANY">
  <X520OrganizationName>GNUTLS</X520OrganizationName>
</value>
</unnamed1>
</unnamed4>
<unnamed5 type="SET OF">
<unnamed1 type="SEQUENCE">

<type type="OBJECT ID">2.5.4.11</type>
<value type="ANY">
  <X520OrganizationalUnitName>GNUTLS dev.</X520OrganizationalUnitName>
</value>
</unnamed1>
</unnamed5>
<unnamed6 type="SET OF">
<unnamed1 type="SEQUENCE">
<type type="OBJECT ID">2.5.4.3</type>
<value type="ANY">
  <X520CommonName>GNUTLS TEST CA</X520CommonName>
</value>
</unnamed1>
</unnamed6>
<unnamed7 type="SET OF">
<unnamed1 type="SEQUENCE">
<type type="OBJECT ID">1.2.840.113549.1.9.1</type>
<value type="ANY">
  <Pkcs9email>gnutls-dev@gnupg.org</Pkcs9email>
</value>
</unnamed1>
</unnamed7>
</rdnSequence>
</issuer>
<validity type="SEQUENCE">
<notBefore type="CHOICE">
<utcTime type="TIME">010707101845Z</utcTime>
</notBefore>
<notAfter type="CHOICE">
<utcTime type="TIME">020707101845Z</utcTime>
</notAfter>
</validity>
<subject type="CHOICE">
<rdnSequence type="SEQUENCE OF">
<unnamed1 type="SET OF">
<unnamed1 type="SEQUENCE">
<type type="OBJECT ID">2.5.4.6</type>
<value type="ANY">
  <X520countryName>GR</X520countryName>
</value>
</unnamed1>
</unnamed1>
<unnamed2 type="SET OF">
<unnamed1 type="SEQUENCE">
<type type="OBJECT ID">2.5.4.8</type>
<value type="ANY">
  <X520StateOrProvinceName>Attiki</X520StateOrProvinceName>
</value>
</unnamed1>
</unnamed2>
<unnamed3 type="SET OF">
<unnamed1 type="SEQUENCE">
<type type="OBJECT ID">2.5.4.7</type>
<value type="ANY">
  <X520LocalityName>Athina</X520LocalityName>
</value>
</unnamed1>
</unnamed3>

<unnamed4 type="SET OF">
  <unnamed1 type="SEQUENCE">
    <type type="OBJECT ID">2.5.4.10</type>
    <value type="ANY">
      <X520OrganizationName>GNUTLS</X520OrganizationName>
    </value>
  </unnamed1>
</unnamed4>
<unnamed5 type="SET OF">
  <unnamed1 type="SEQUENCE">
    <type type="OBJECT ID">2.5.4.11</type>
    <value type="ANY">
      <X520OrganizationalUnitName>GNUTLS dev.</X520OrganizationalUnitName>
    </value>
  </unnamed1>
</unnamed5>
<unnamed6 type="SET OF">
  <unnamed1 type="SEQUENCE">
    <type type="OBJECT ID">2.5.4.3</type>
    <value type="ANY">
      <X520CommonName>localhost</X520CommonName>
    </value>
  </unnamed1>
</unnamed6>
<unnamed7 type="SET OF">
  <unnamed1 type="SEQUENCE">
    <type type="OBJECT ID">1.2.840.113549.1.9.1</type>
    <value type="ANY">
      <Pkcs9email>root@localhost</Pkcs9email>
    </value>
  </unnamed1>
</unnamed7>
</rdnSequence>
</subject>
<subjectPublicKeyInfo type="SEQUENCE">
<algorithm type="SEQUENCE">
<algorithm type="OBJECT ID">1.2.840.113549.1.1.1</algorithm>
<parameters type="ANY">
  <rsaEncryption encoding="HEX">0500</rsaEncryption>
</parameters>
</algorithm>
<subjectPublicKey type="BIT STRING" encoding="HEX" length="1120">
30818902818100D00B49EBB226D951F5CC57072199DDF287683D2DA1A0E
FCC96BFF73164777C78C3991E92EDA66584E7B97BAB4BE68D595D225557
E01E7E57B5C35C04B491948C5C427AD588D8C6989764996D6D44E17B65C
CFC86F3B4842DE559B730C1DE3AEF1CE1A328AFF8A357EBA911E1F7E8FC
1598E21E4BF721748C587F50CF46157D950203010001</subjectPublicKey>
</subjectPublicKeyInfo>
<extensions type="SEQUENCE OF">
<unnamed1 type="SEQUENCE">
<extnID type="OBJECT ID">2.5.29.35</extnID>
<critical type="BOOLEAN">FALSE</critical>
<extnValue type="SEQUENCE">
  <keyIdentifier type="OCTET STRING" encoding="HEX">
  EFEE94ABC8CA577F5313DB76DC1A950093BAF3C9</keyIdentifier>
</extnValue>
</unnamed1>
<unnamed2 type="SEQUENCE">

    <extnID type="OBJECT ID">2.5.29.37</extnID>
    <critical type="BOOLEAN">FALSE</critical>
    <extnValue type="SEQUENCE OF">
      <unnamed1 type="OBJECT ID">1.3.6.1.5.5.7.3.1</unnamed1>
      <unnamed2 type="OBJECT ID">1.3.6.1.5.5.7.3.2</unnamed2>
      <unnamed3 type="OBJECT ID">1.3.6.1.4.1.311.10.3.3</unnamed3>
      <unnamed4 type="OBJECT ID">2.16.840.1.113730.4.1</unnamed4>
    </extnValue>
  </unnamed2>
  <unnamed3 type="SEQUENCE">
    <extnID type="OBJECT ID">2.5.29.19</extnID>
    <critical type="BOOLEAN">TRUE</critical>
    <extnValue type="SEQUENCE">
      <cA type="BOOLEAN">FALSE</cA>
    </extnValue>
  </unnamed3>
</extensions>
</tbsCertificate>
<signatureAlgorithm type="SEQUENCE">
<algorithm type="OBJECT ID">1.2.840.113549.1.1.4</algorithm>
<parameters type="ANY">
  <md5WithRSAEncryption encoding="HEX">0500</md5WithRSAEncryption>
</parameters>
</signatureAlgorithm>
<signature type="BIT STRING" encoding="HEX" length="1024">
B73945273AF2A395EC54BF5DC669D953885A9D811A3B92909D24792D36A44EC
27E1C463AF8738BEFD29B311CCE8C6D9661BEC30911DAABB39B8813382B32D2
E259581EBCD26C495C083984763966FF35D1DEFE432891E610C85072578DA74
23244A8F5997B41A1F44E61F4F22C94375775055A5E72F25D5E4557467A91BD
4251</signature>
</certificate>
</gnutls:x509:certificate>

10.2 An OpenPGP key

<?xml version="1.0"?>
<gnutls:openpgp:key version="1.0">
<OPENPGPKEY>
<MAINKEY>
<KEYID>BD572CDCCCC07C3</KEYID>
<FINGERPRINT>BE615E88D6CFF27225B8A2E7BD572CDCCCC07C35</FINGERPRINT>
<PKALGO>DSA</PKALGO>
<KEYLEN>1024</KEYLEN>
<CREATED>1011533164</CREATED>
<REVOKED>0</REVOKED>
<KEY ENCODING="HEX"/>
<DSA-P>0400E72E76B62EEFA9A3BD594093292418050C02D7029D6CA2066E
FC34C86038627C643EB1A652A7AF1D37CF46FC505AC1E0C699B37895B4BCB
3E53541FFDA4766D6168C2B8AAFD6AB22466D06D18034D5DAC698E6993BA5
B350FF822E1CD8702A75114E8B73A6B09CB3B93CE44DBB516C9BB5F95BB66
6188602A0A1447236C0658F</DSA-P>
<DSA-Q>00A08F5B5E78D85F792CC2072F9474645726FB4D9373</DSA-Q>
<DSA-G>03FE3578D689D6606E9118E9F9A7042B963CF23F3D8F1377A273C0
F0974DBF44B3CABCBE14DD64412555863E39A9C627662D77AC36662AE4497
92C3262D3F12E9832A7565309D67BA0AE4DF25F5EDA0937056AD5BE89F406
9EBD7EC76CE432441DF5D52FFFD06D39E5F61E36947B698A77CB62AB81E4A
4122BF9050671D9946C865E</DSA-G>

<DSA-Y>0400D061437A964DDE318818C2B24DE008E60096B60DB8A684B85A
838D119FC930311889AD57A3B927F448F84EB253C623EDA73B42FF78BCE63
A6A531D75A64CE8540513808E9F5B10CE075D3417B801164918B131D3544C
8765A8ECB9971F61A09FC73D509806106B5977D211CB0E1D04D0ED96BCE89
BAE8F73D800B052139CBF8D</DSA-Y>
</MAINKEY>
<USERID>
<NAME>OpenCDK test key (Only intended for test purposes!)</NAME>
<EMAIL>opencdk@foo-bar.org</EMAIL>
<PRIMARY>0</PRIMARY>
<REVOKED>0</REVOKED>
</USERID>
<SIGNATURE>
<VERSION>4</VERSION>
<SIGCLASS>19</SIGCLASS>
<EXPIRED>0</EXPIRED>
<PKALGO>DSA</PKALGO>
<MDALGO>SHA1</MDALGO>
<CREATED>1011533164</CREATED>
<KEYID>BD572CDCCCC07C3</KEYID>
</SIGNATURE>
<SUBKEY>
<KEYID>FCB0CF3A5261E06</KEYID>
<FINGERPRINT>297B48ACC09C0FF683CA1ED1FCB0CF3A5261E067</FINGERPRINT>
<PKALGO>ELG</PKALGO>
<KEYLEN>1024</KEYLEN>
<CREATED>1011533167</CREATED>
<REVOKED>0</REVOKED>
<KEY ENCODING="HEX"/>
<ELG-P>0400E20156526069D067D24F4D71E6D38658E08BE3BF246C1ADCE0
8DB69CD8D459C1ED335738410798755AFDB79F1797CF022E70C7960F12CA6
896D27CFD24A11CD316DDE1FBCC1EA615C5C31FEC656E467078C875FC509B
1ECB99C8B56C2D875C50E2018B5B0FA378606EB6425A2533830F55FD21D64
9015615D49A1D09E9510F5F</ELG-P>
<ELG-G>000305</ELG-G>
<ELG-Y>0400D0BDADE40432758675C87D0730C360981467BAE1BEB6CC105A
3C1F366BFDBEA12E378456513238B8AD414E52A2A9661D1DF1DB6BB5F33F6
906166107556C813224330B30932DB7C8CC8225672D7AE24AF2469750E539
B661EA6475D2E03CD8D3838DC4A8AC4AFD213536FE3E96EC9D0AEA65164B5
76E01B37A8DCA89F2B257D0</ELG-Y>
</SUBKEY>
<SIGNATURE>
<VERSION>4</VERSION>
<SIGCLASS>24</SIGCLASS>
<EXPIRED>0</EXPIRED>
<PKALGO>DSA</PKALGO>
<MDALGO>SHA1</MDALGO>
<CREATED>1011533167</CREATED>
<KEYID>BD572CDCCCC07C3</KEYID>
</SIGNATURE>
</OPENPGPKEY>
</gnutls:openpgp:key>

Chapter 11: All the supported ciphersuites in GnuTLS 211

11 All the supported ciphersuites in GnuTLS

TLS_RSA_NULL_MD5  0x00 0x01  RFC 2246
TLS_ANON_DH_3DES_EDE_CBC_SHA  0x00 0x1B  RFC 2246
TLS_ANON_DH_ARCFOUR_MD5  0x00 0x18  RFC 2246
TLS_ANON_DH_AES_128_CBC_SHA  0x00 0x34  RFC 2246
TLS_ANON_DH_AES_256_CBC_SHA  0x00 0x3A  RFC 2246
TLS_RSA_ARCFOUR_SHA  0x00 0x05  RFC 2246
TLS_RSA_ARCFOUR_MD5  0x00 0x04  RFC 2246
TLS_RSA_3DES_EDE_CBC_SHA  0x00 0x0A  RFC 2246
TLS_RSA_EXPORT_ARCFOUR_40_MD5  0x00 0x03  RFC 2246
TLS_DHE_DSS_3DES_EDE_CBC_SHA  0x00 0x13  RFC 2246
TLS_DHE_RSA_3DES_EDE_CBC_SHA  0x00 0x16  RFC 2246
TLS_RSA_AES_128_CBC_SHA  0x00 0x2F  RFC 3268
TLS_RSA_AES_256_CBC_SHA  0x00 0x35  RFC 3268
TLS_DHE_DSS_AES_256_CBC_SHA  0x00 0x38  RFC 3268
TLS_DHE_DSS_AES_128_CBC_SHA  0x00 0x32  RFC 3268
TLS_DHE_RSA_AES_256_CBC_SHA  0x00 0x39  RFC 3268
TLS_DHE_RSA_AES_128_CBC_SHA  0x00 0x33  RFC 3268
TLS_SRP_SHA_3DES_EDE_CBC_SHA  0x00 0x50  draft-ietf-tls-srp
TLS_SRP_SHA_AES_128_CBC_SHA  0x00 0x53  draft-ietf-tls-srp
TLS_SRP_SHA_AES_256_CBC_SHA  0x00 0x56  draft-ietf-tls-srp
TLS_SRP_SHA_RSA_3DES_EDE_CBC_SHA  0x00 0x51  draft-ietf-tls-srp
TLS_SRP_SHA_DSS_3DES_EDE_CBC_SHA  0x00 0x52  draft-ietf-tls-srp

TLS_SRP_SHA_RSA_AES_128_CBC_SHA  0x00 0x54  draft-ietf-tls-srp
TLS_SRP_SHA_DSS_AES_128_CBC_SHA  0x00 0x55  draft-ietf-tls-srp
TLS_SRP_SHA_RSA_AES_256_CBC_SHA  0x00 0x57  draft-ietf-tls-srp
TLS_SRP_SHA_DSS_AES_256_CBC_SHA  0x00 0x58  draft-ietf-tls-srp

TLS_DHE_DSS_ARCFOUR_SHA

0x00 0x66 draft-ietf-tls-56-bit-ciphersuites

TLS_PSK_ARCFOUR_SHA  0x00 0x8A  draft-ietf-tls-psk
TLS_PSK_3DES_EDE_CBC_SHA  0x00 0x8B  draft-ietf-tls-psk
TLS_PSK_AES_128_CBC_SHA  0x00 0x8C  draft-ietf-tls-psk
TLS_PSK_AES_256_CBC_SHA  0x00 0x8D  draft-ietf-tls-psk

12 Internal architecture of GnuTLS

This chapter is to give a brief description of the way GnuTLS works. The focus is to give an idea to potential developers and those who want to know what happens inside the black box.

12.1 The TLS protocol

The main needs for the TLS protocol to be used are shown in the image below.

This is being accomplished by the following object diagram. Note that since GnuTLS is being developed in C object are just structures with attributes. The operations listed are functions that require the first parameter to be that object.

12.2 TLS Handshake protocol

The GnuTLS handshake protocol is implemented as a state machine that waits for input or returns immediately when the non-blocking transport layer functions are used. The main idea is shown in the following figure.

Also the way the input is processed varies per ciphersuite. Several implementations of the internal handlers are available and [gnutls handshake], page 119 only multiplexes the input to the appropriate handler. For example a PSK ciphersuite has a different implementation of the process_client_key_exchange than a certificate ciphersuite.

12.3 TLS authentication methods

In GnuTLS authentication methods can be implemented quite easily. Since the required changes to add a new authentication method affect only the handshake protocol, a simple interface is used. An authentication method needs only to implement the functions as seen in the figure below.

The functions that need to be implemented are the ones responsible for interpreting the handshake protocol messages. It is common for such functions to read data from one or more credentials_t structures^{1} and write data, such as certificates, usernames etc. to

auth_info_t structures.

Simple examples of existing authentication methods can be seen in auth_psk.c for PSK ciphersuites and auth_srp.c for SRP ciphersuites. After implementing these functions the structure holding its pointers has to be registered in gnutls_algorithms.c in the _gnutls_

kx_algorithms structure.

12.4 TLS Extension handling

As with authentication methods, the TLS extensions handlers can be implemented using the following interface.

Here there are two functions, one for receiving the extension data and one for sending. These functions have to check internally whether they operate in client or server side.

A simple example of an extension handler can be seen in ext_srp.c After implementing these functions, together with the extension number they handle, they have to be registered

in gnutls_extensions.c in the _gnutls_extensions structure.

1 such as the gnutls_certificate_credentials_t structures

Chapter 12: Internal architecture of GnuTLS 217

12.5 Certificate handling

What is provided by the certificate handling functions is summarized in the following diagram.

Appendix A Copying Information

A.1 GNU Free Documentation License

Version 1.2, November 2002

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2000,2001,2002 Free Software Foundation, Inc. 51 Franklin St, Fifth Floor, Boston, MA 02110-1301, USA

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

0. PREAMBLE The purpose of this License is to make a manual, textbook, or other functional and useful document free in the sense of freedom: to assure everyone the effective freedom to copy and redistribute it, with or without modifying it, either commercially or noncommercially. Secondarily, this License preserves for the author and publisher a way to get credit for their work, while not being considered responsible for modifications made by others. This License is a kind of “copyleft”, which means that derivative works of the document must themselves be free in the same sense. It complements the GNU General Public License, which is a copyleft license designed for free software. We have designed this License in order to use it for manuals for free software, because free software needs free documentation: a free program should come with manuals providing the same freedoms that the software does. But this License is not limited to software manuals; it can be used for any textual work, regardless of subject matter or whether it is published as a printed book. We recommend this License principally for works whose purpose is instruction or reference.

1. APPLICABILITY AND DEFINITIONS This License applies to any manual or other work, in any medium, that contains a notice placed by the copyright holder saying it can be distributed under the terms of this License. Such a notice grants a world-wide, royalty-free license, unlimited in duration, to use that work under the conditions stated herein. The “Document”, below, refers to any such manual or work. Any member of the public is a licensee, and is addressed as “you”. You accept the license if you copy, modify or distribute the work in a way requiring permission under copyright law. A “Modified Version” of the Document means any work containing the Document or a portion of it, either copied verbatim, or with modifications and/or translated into another language. A “Secondary Section” is a named appendix or a front-matter section of the Document that deals exclusively with the relationship of the publishers or authors of the Document to the Document’s overall subject (or to related matters) and contains nothing that could fall directly within that overall subject. (Thus, if the Document is in part a textbook of mathematics, a Secondary Section may not explain any mathematics.) The relationship could be a matter of historical connection with the subject or with related matters, or of legal, commercial, philosophical, ethical or political position regarding them.

The “Invariant Sections” are certain Secondary Sections whose titles are designated, as being those of Invariant Sections, in the notice that says that the Document is released under this License. If a section does not fit the above definition of Secondary then it is not allowed to be designated as Invariant. The Document may contain zero Invariant Sections. If the Document does not identify any Invariant Sections then there are none.

The “Cover Texts” are certain short passages of text that are listed, as Front-Cover Texts or Back-Cover Texts, in the notice that says that the Document is released under this License. A Front-Cover Text may be at most 5 words, and a Back-Cover Text may be at most 25 words.

A “Transparent” copy of the Document means a machine-readable copy, represented in a format whose specification is available to the general public, that is suitable for revising the document straightforwardly with generic text editors or (for images composed of pixels) generic paint programs or (for drawings) some widely available drawing editor, and that is suitable for input to text formatters or for automatic translation to a variety of formats suitable for input to text formatters. A copy made in an otherwise Transparent file format whose markup, or absence of markup, has been arranged to thwart or discourage subsequent modification by readers is not Transparent. An image format is not Transparent if used for any substantial amount of text. A copy that is not “Transparent” is called “Opaque”.

Examples of suitable formats for Transparent copies include plain ascii without markup, Texinfo input format, LaTEX input format, SGML or XML using a publicly available DTD, and standard-conforming simple HTML, PostScript or PDF designed for human modification. Examples of transparent image formats include PNG, XCF and JPG. Opaque formats include proprietary formats that can be read and edited only by proprietary word processors, SGML or XML for which the DTD and/or processing tools are not generally available, and the machine-generated HTML, PostScript or PDF produced by some word processors for output purposes only.

The “Title Page” means, for a printed book, the title page itself, plus such following pages as are needed to hold, legibly, the material this License requires to appear in the title page. For works in formats which do not have any title page as such, “Title Page” means the text near the most prominent appearance of the work’s title, preceding the beginning of the body of the text.

A section “Entitled XYZ” means a named subunit of the Document whose title either is precisely XYZ or contains XYZ in parentheses following text that translates XYZ in another language. (Here XYZ stands for a specific section name mentioned below, such as “Acknowledgements”, “Dedications”, “Endorsements”, or “History”.) To “Preserve the Title” of such a section when you modify the Document means that it remains a section “Entitled XYZ” according to this definition.

The Document may include Warranty Disclaimers next to the notice which states that this License applies to the Document. These Warranty Disclaimers are considered to be included by reference in this License, but only as regards disclaiming warranties: any other implication that these Warranty Disclaimers may have is void and has no effect on the meaning of this License.

2. VERBATIM COPYING You may copy and distribute the Document in any medium, either commercially or noncommercially, provided that this License, the copyright notices, and the license notice saying this License applies to the Document are reproduced in all copies, and that you add no other conditions whatsoever to those of this License. You may not use technical measures to obstruct or control the reading or further copying of the copies you make or distribute. However, you may accept compensation in exchange for copies. If you distribute a large enough number of copies you must also follow the conditions in section 3. You may also lend copies, under the same conditions stated above, and you may publicly display copies.

3. COPYING IN QUANTITY If you publish printed copies (or copies in media that commonly have printed covers) of the Document, numbering more than 100, and the Document’s license notice requires Cover Texts, you must enclose the copies in covers that carry, clearly and legibly, all these Cover Texts: Front-Cover Texts on the front cover, and Back-Cover Texts on the back cover. Both covers must also clearly and legibly identify you as the publisher of these copies. The front cover must present the full title with all words of the title equally prominent and visible. You may add other material on the covers in addition. Copying with changes limited to the covers, as long as they preserve the title of the Document and satisfy these conditions, can be treated as verbatim copying in other respects. If the required texts for either cover are too voluminous to fit legibly, you should put the first ones listed (as many as fit reasonably) on the actual cover, and continue the rest onto adjacent pages. If you publish or distribute Opaque copies of the Document numbering more than 100, you must either include a machine-readable Transparent copy along with each Opaque copy, or state in or with each Opaque copy a computer-network location from which the general network-using public has access to download using public-standard network protocols a complete Transparent copy of the Document, free of added material. If you use the latter option, you must take reasonably prudent steps, when you begin distribution of Opaque copies in quantity, to ensure that this Transparent copy will remain thus accessible at the stated location until at least one year after the last time you distribute an Opaque copy (directly or through your agents or retailers) of that edition to the public. It is requested, but not required, that you contact the authors of the Document well before redistributing any large number of copies, to give them a chance to provide you with an updated version of the Document.

4. MODIFICATIONS You may copy and distribute a Modified Version of the Document under the conditions of sections 2 and 3 above, provided that you release the Modified Version under precisely this License, with the Modified Version filling the role of the Document, thus licensing distribution and modification of the Modified Version to whoever possesses a copy of it. In addition, you must do these things in the Modified Version: A. Use in the Title Page (and on the covers, if any) a title distinct from that of the

Document, and from those of previous versions (which should, if there were any, be listed in the History section of the Document). You may use the same title as a previous version if the original publisher of that version gives permission.

B. List on the Title Page, as authors, one or more persons or entities responsible for authorship of the modifications in the Modified Version, together with at least five of the principal authors of the Document (all of its principal authors, if it has fewer than five), unless they release you from this requirement.

C. State on the Title page the name of the publisher of the Modified Version, as the

publisher. D. Preserve all the copyright notices of the Document. E. Add an appropriate copyright notice for your modifications adjacent to the other

copyright notices. F. Include, immediately after the copyright notices, a license notice giving the public

permission to use the Modified Version under the terms of this License, in the form shown in the Addendum below.

G. Preserve in that license notice the full lists of Invariant Sections and required Cover

Texts given in the Document’s license notice. H. Include an unaltered copy of this License. I. Preserve the section Entitled “History”, Preserve its Title, and add to it an item

stating at least the title, year, new authors, and publisher of the Modified Version as given on the Title Page. If there is no section Entitled “History” in the Document, create one stating the title, year, authors, and publisher of the Document as given on its Title Page, then add an item describing the Modified Version as stated in the previous sentence.

J. Preserve the network location, if any, given in the Document for public access to a Transparent copy of the Document, and likewise the network locations given in the Document for previous versions it was based on. These may be placed in the “History” section. You may omit a network location for a work that was published at least four years before the Document itself, or if the original publisher of the version it refers to gives permission.

K. For any section Entitled “Acknowledgements” or “Dedications”, Preserve the Title of the section, and preserve in the section all the substance and tone of each of the contributor acknowledgements and/or dedications given therein.

L. Preserve all the Invariant Sections of the Document, unaltered in their text and in their titles. Section numbers or the equivalent are not considered part of the section titles.

M. Delete any section Entitled “Endorsements”. Such a section may not be included

in the Modified Version. N. Do not retitle any existing section to be Entitled “Endorsements” or to conflict in

title with any Invariant Section. O. Preserve any Warranty Disclaimers.

If the Modified Version includes new front-matter sections or appendices that qualify as Secondary Sections and contain no material copied from the Document, you may at your option designate some or all of these sections as invariant. To do this, add their titles to the list of Invariant Sections in the Modified Version’s license notice. These titles must be distinct from any other section titles. You may add a section Entitled “Endorsements”, provided it contains nothing but endorsements of your Modified Version by various parties—for example, statements of peer review or that the text has been approved by an organization as the authoritative definition of a standard. You may add a passage of up to five words as a Front-Cover Text, and a passage of up to 25 words as a Back-Cover Text, to the end of the list of Cover Texts in the Modified Version. Only one passage of Front-Cover Text and one of Back-Cover Text may be added by (or through arrangements made by) any one entity. If the Document already includes a cover text for the same cover, previously added by you or by arrangement made by the same entity you are acting on behalf of, you may not add another; but you may replace the old one, on explicit permission from the previous publisher that added the old one. The author(s) and publisher(s) of the Document do not by this License give permission to use their names for publicity for or to assert or imply endorsement of any Modified Version.

5. COMBINING DOCUMENTS You may combine the Document with other documents released under this License, under the terms defined in section 4 above for modified versions, provided that you include in the combination all of the Invariant Sections of all of the original documents, unmodified, and list them all as Invariant Sections of your combined work in its license notice, and that you preserve all their Warranty Disclaimers. The combined work need only contain one copy of this License, and multiple identical Invariant Sections may be replaced with a single copy. If there are multiple Invariant Sections with the same name but different contents, make the title of each such section unique by adding at the end of it, in parentheses, the name of the original author or publisher of that section if known, or else a unique number. Make the same adjustment to the section titles in the list of Invariant Sections in the license notice of the combined work. In the combination, you must combine any sections Entitled “History” in the various original documents, forming one section Entitled “History”; likewise combine any sections Entitled “Acknowledgements”, and any sections Entitled “Dedications”. You must delete all sections Entitled “Endorsements.”

6. COLLECTIONS OF DOCUMENTS You may make a collection consisting of the Document and other documents released under this License, and replace the individual copies of this License in the various documents with a single copy that is included in the collection, provided that you follow the rules of this License for verbatim copying of each of the documents in all other respects. You may extract a single document from such a collection, and distribute it individually under this License, provided you insert a copy of this License into the extracted document, and follow this License in all other respects regarding verbatim copying of that document.

7. AGGREGATION WITH INDEPENDENT WORKS A compilation of the Document or its derivatives with other separate and independent documents or works, in or on a volume of a storage or distribution medium, is called an “aggregate” if the copyright resulting from the compilation is not used to limit the legal rights of the compilation’s users beyond what the individual works permit. When the Document is included in an aggregate, this License does not apply to the other works in the aggregate which are not themselves derivative works of the Document. If the Cover Text requirement of section 3 is applicable to these copies of the Document, then if the Document is less than one half of the entire aggregate, the Document’s Cover Texts may be placed on covers that bracket the Document within the aggregate, or the electronic equivalent of covers if the Document is in electronic form. Otherwise they must appear on printed covers that bracket the whole aggregate.

8. TRANSLATION Translation is considered a kind of modification, so you may distribute translations of the Document under the terms of section 4. Replacing Invariant Sections with translations requires special permission from their copyright holders, but you may include translations of some or all Invariant Sections in addition to the original versions of these Invariant Sections. You may include a translation of this License, and all the license notices in the Document, and any Warranty Disclaimers, provided that you also include the original English version of this License and the original versions of those notices and disclaimers. In case of a disagreement between the translation and the original version of this License or a notice or disclaimer, the original version will prevail. If a section in the Document is Entitled “Acknowledgements”, “Dedications”, or “History”, the requirement (section 4) to Preserve its Title (section 1) will typically require changing the actual title.

9. TERMINATION You may not copy, modify, sublicense, or distribute the Document except as expressly provided for under this License. Any other attempt to copy, modify, sublicense or distribute the Document is void, and will automatically terminate your rights under this License. However, parties who have received copies, or rights, from you under this License will not have their licenses terminated so long as such parties remain in full compliance.

10. FUTURE REVISIONS OF THIS LICENSE The Free Software Foundation may publish new, revised versions of the GNU Free Documentation License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

See http://www.gnu.org/copyleft/.

Each version of the License is given a distinguishing version number. If the Document specifies that a particular numbered version of this License “or any later version” applies to it, you have the option of following the terms and conditions either of that specified version or of any later version that has been published (not as a draft) by the Free Software Foundation. If the Document does not specify a version number of this License, you may choose any version ever published (not as a draft) by the Free Software Foundation.

ADDENDUM: How to use this License for your documents

To use this License in a document you have written, include a copy of the License in the document and put the following copyright and license notices just after the title page:

        Copyright (C)  year  your name.
        Permission is granted to copy, distribute and/or modify this document
        under the terms of the GNU Free Documentation License, Version 1.2
        or any later version published by the Free Software Foundation;
        with no Invariant Sections, no Front-Cover Texts, and no Back-Cover
        Texts.  A copy of the license is included in the section entitled ``GNU
        Free Documentation License''.

If you have Invariant Sections, Front-Cover Texts and Back-Cover Texts, replace the “with...Texts.” line with this:

          with the Invariant Sections being list their titles, with
          the Front-Cover Texts being list, and with the Back-Cover Texts
          being list.

If you have Invariant Sections without Cover Texts, or some other combination of the three, merge those two alternatives to suit the situation. If your document contains nontrivial examples of program code, we recommend releasing these examples in parallel under your choice of free software license, such as the GNU General Public License, to permit their use in free software.

A.2 GNU Lesser General Public License

Version 2.1, February 1999

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1991, 1999 Free Software Foundation, Inc.

51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA

Everyone is permitted to copy and distribute verbatim copies

of this license document, but changing it is not allowed.

[This is the first released version of the Lesser GPL. It also counts as the successor of the GNU Library Public License, version 2, hence the version number 2.1.]

Preamble

The licenses for most software are designed to take away your freedom to share and change it. By contrast, the GNU General Public Licenses are intended to guarantee your freedom to share and change free software—to make sure the software is free for all its users. This license, the Lesser General Public License, applies to some specially designated software—typically libraries—of the Free Software Foundation and other authors who decide to use it. You can use it too, but we suggest you first think carefully about whether this license or the ordinary General Public License is the better strategy to use in any particular case, based on the explanations below. When we speak of free software, we are referring to freedom of use, not price. Our General Public Licenses are designed to make sure that you have the freedom to distribute copies of free software (and charge for this service if you wish); that you receive source code or can get it if you want it; that you can change the software and use pieces of it in new free programs; and that you are informed that you can do these things.

To protect your rights, we need to make restrictions that forbid distributors to deny you these rights or to ask you to surrender these rights. These restrictions translate to certain responsibilities for you if you distribute copies of the library or if you modify it. For example, if you distribute copies of the library, whether gratis or for a fee, you must give the recipients all the rights that we gave you. You must make sure that they, too, receive or can get the source code. If you link other code with the library, you must provide complete object files to the recipients, so that they can relink them with the library after making changes to the library and recompiling it. And you must show them these terms so they know their rights. We protect your rights with a two-step method: (1) we copyright the library, and (2) we offer you this license, which gives you legal permission to copy, distribute and/or modify the library. To protect each distributor, we want to make it very clear that there is no warranty for the free library. Also, if the library is modified by someone else and passed on, the recipients should know that what they have is not the original version, so that the original author’s reputation will not be affected by problems that might be introduced by others. Finally, software patents pose a constant threat to the existence of any free program. We wish to make sure that a company cannot effectively restrict the users of a free program by obtaining a restrictive license from a patent holder. Therefore, we insist that any patent license obtained for a version of the library must be consistent with the full freedom of use specified in this license. Most GNU software, including some libraries, is covered by the ordinary GNU General Public License. This license, the GNU Lesser General Public License, applies to certain designated libraries, and is quite different from the ordinary General Public License. We use this license for certain libraries in order to permit linking those libraries into non-free programs. When a program is linked with a library, whether statically or using a shared library, the combination of the two is legally speaking a combined work, a derivative of the original library. The ordinary General Public License therefore permits such linking only if the entire combination fits its criteria of freedom. The Lesser General Public License permits more lax criteria for linking other code with the library. We call this license the Lesser General Public License because it does Less to protect the user’s freedom than the ordinary General Public License. It also provides other free software developers Less of an advantage over competing non-free programs. These disadvantages are the reason we use the ordinary General Public License for many libraries. However, the Lesser license provides advantages in certain special circumstances. For example, on rare occasions, there may be a special need to encourage the widest possible use of a certain library, so that it becomes a de-facto standard. To achieve this, non-free programs must be allowed to use the library. A more frequent case is that a free library does the same job as widely used non-free libraries. In this case, there is little to gain by limiting the free library to free software only, so we use the Lesser General Public License. In other cases, permission to use a particular library in non-free programs enables a greater number of people to use a large body of free software. For example, permission to use the GNU C Library in non-free programs enables many more people to use the whole GNU operating system, as well as its variant, the GNU/Linux operating system.

Although the Lesser General Public License is Less protective of the users’ freedom, it does ensure that the user of a program that is linked with the Library has the freedom and the wherewithal to run that program using a modified version of the Library.

The precise terms and conditions for copying, distribution and modification follow. Pay close attention to the difference between a “work based on the library” and a “work that uses the library”. The former contains code derived from the library, whereas the latter must be combined with the library in order to run.

TERMS AND CONDITIONS FOR COPYING, DISTRIBUTION AND MODIFICATION

0. This License Agreement applies to any software library or other program which contains a notice placed by the copyright holder or other authorized party saying it may be distributed under the terms of this Lesser General Public License (also called “this License”). Each licensee is addressed as “you”. A “library” means a collection of software functions and/or data prepared so as to be conveniently linked with application programs (which use some of those functions and data) to form executables. The “Library”, below, refers to any such software library or work which has been distributed under these terms. A “work based on the Library” means either the Library or any derivative work under copyright law: that is to say, a work containing the Library or a portion of it, either verbatim or with modifications and/or translated straightforwardly into another language. (Hereinafter, translation is included without limitation in the term “modification”.) “Source code” for a work means the preferred form of the work for making modifications to it. For a library, complete source code means all the source code for all modules it contains, plus any associated interface definition files, plus the scripts used to control compilation and installation of the library. Activities other than copying, distribution and modification are not covered by this License; they are outside its scope. The act of running a program using the Library is not restricted, and output from such a program is covered only if its contents constitute a work based on the Library (independent of the use of the Library in a tool for writing it). Whether that is true depends on what the Library does and what the program that uses the Library does.

1. You may copy and distribute verbatim copies of the Library’s complete source code as you receive it, in any medium, provided that you conspicuously and appropriately publish on each copy an appropriate copyright notice and disclaimer of warranty; keep intact all the notices that refer to this License and to the absence of any warranty; and distribute a copy of this License along with the Library. You may charge a fee for the physical act of transferring a copy, and you may at your option offer warranty protection in exchange for a fee.

2. You may modify your copy or copies of the Library or any portion of it, thus forming a work based on the Library, and copy and distribute such modifications or work under the terms of Section 1 above, provided that you also meet all of these conditions: a. The modified work must itself be a software library. b. You must cause the files modified to carry prominent notices stating that you

changed the files and the date of any change. c. You must cause the whole of the work to be licensed at no charge to all third

parties under the terms of this License. d. If a facility in the modified Library refers to a function or a table of data to

be supplied by an application program that uses the facility, other than as an argument passed when the facility is invoked, then you must make a good faith effort to ensure that, in the event an application does not supply such function or table, the facility still operates, and performs whatever part of its purpose remains meaningful. (For example, a function in a library to compute square roots has a purpose that is entirely well-defined independent of the application. Therefore, Subsection 2d requires that any application-supplied function or table used by this function must be optional: if the application does not supply it, the square root function must still compute square roots.)

These requirements apply to the modified work as a whole. If identifiable sections of that work are not derived from the Library, and can be reasonably considered independent and separate works in themselves, then this License, and its terms, do not apply to those sections when you distribute them as separate works. But when you distribute the same sections as part of a whole which is a work based on the Library, the distribution of the whole must be on the terms of this License, whose permissions for other licensees extend to the entire whole, and thus to each and every part regardless of who wrote it. Thus, it is not the intent of this section to claim rights or contest your rights to work written entirely by you; rather, the intent is to exercise the right to control the distribution of derivative or collective works based on the Library. In addition, mere aggregation of another work not based on the Library with the Library (or with a work based on the Library) on a volume of a storage or distribution medium does not bring the other work under the scope of this License.

3. You may opt to apply the terms of the ordinary GNU General Public License instead of this License to a given copy of the Library. To do this, you must alter all the notices that refer to this License, so that they refer to the ordinary GNU General Public License, version 2, instead of to this License. (If a newer version than version 2 of the ordinary GNU General Public License has appeared, then you can specify that version instead if you wish.) Do not make any other change in these notices. Once this change is made in a given copy, it is irreversible for that copy, so the ordinary GNU General Public License applies to all subsequent copies and derivative works made from that copy. This option is useful when you wish to copy part of the code of the Library into a program that is not a library.

4. You may copy and distribute the Library (or a portion or derivative of it, under Section 2) in object code or executable form under the terms of Sections 1 and 2 above provided that you accompany it with the complete corresponding machine-readable source code, which must be distributed under the terms of Sections 1 and 2 above on a medium customarily used for software interchange.

If distribution of object code is made by offering access to copy from a designated place, then offering equivalent access to copy the source code from the same place satisfies the requirement to distribute the source code, even though third parties are not compelled to copy the source along with the object code.

5. A program that contains no derivative of any portion of the Library, but is designed to work with the Library by being compiled or linked with it, is called a “work that uses the Library”. Such a work, in isolation, is not a derivative work of the Library, and therefore falls outside the scope of this License.

However, linking a “work that uses the Library” with the Library creates an executable that is a derivative of the Library (because it contains portions of the Library), rather than a “work that uses the library”. The executable is therefore covered by this License. Section 6 states terms for distribution of such executables.

When a “work that uses the Library” uses material from a header file that is part of the Library, the object code for the work may be a derivative work of the Library even though the source code is not. Whether this is true is especially significant if the work can be linked without the Library, or if the work is itself a library. The threshold for this to be true is not precisely defined by law.

If such an object file uses only numerical parameters, data structure layouts and accessors, and small macros and small inline functions (ten lines or less in length), then the use of the object file is unrestricted, regardless of whether it is legally a derivative work. (Executables containing this object code plus portions of the Library will still fall under Section 6.)

Otherwise, if the work is a derivative of the Library, you may distribute the object code for the work under the terms of Section 6. Any executables containing that work also fall under Section 6, whether or not they are linked directly with the Library itself.

6. As an exception to the Sections above, you may also combine or link a “work that uses the Library” with the Library to produce a work containing portions of the Library, and distribute that work under terms of your choice, provided that the terms permit modification of the work for the customer’s own use and reverse engineering for debugging such modifications.

You must give prominent notice with each copy of the work that the Library is used in it and that the Library and its use are covered by this License. You must supply a copy of this License. If the work during execution displays copyright notices, you must include the copyright notice for the Library among them, as well as a reference directing the user to the copy of this License. Also, you must do one of these things:

a. Accompany the work with the complete corresponding machine-readable source code for the Library including whatever changes were used in the work (which must be distributed under Sections 1 and 2 above); and, if the work is an executable linked with the Library, with the complete machine-readable “work that uses the Library”, as object code and/or source code, so that the user can modify the Library and then relink to produce a modified executable containing the modified Library. (It is understood that the user who changes the contents of definitions files in the Library will not necessarily be able to recompile the application to use the modified definitions.) b. Use a suitable shared library mechanism for linking with the Library. A suitable

mechanism is one that (1) uses at run time a copy of the library already present on the user’s computer system, rather than copying library functions into the executable, and (2) will operate properly with a modified version of the library, if the user installs one, as long as the modified version is interface-compatible with the version that the work was made with.

c. Accompany the work with a written offer, valid for at least three years, to give the same user the materials specified in Subsection 6a, above, for a charge no more than the cost of performing this distribution.

d. If distribution of the work is made by offering access to copy from a designated place, offer equivalent access to copy the above specified materials from the same place.

e. Verify that the user has already received a copy of these materials or that you have already sent this user a copy.

For an executable, the required form of the “work that uses the Library” must include any data and utility programs needed for reproducing the executable from it. However, as a special exception, the materials to be distributed need not include anything that is normally distributed (in either source or binary form) with the major components (compiler, kernel, and so on) of the operating system on which the executable runs, unless that component itself accompanies the executable. It may happen that this requirement contradicts the license restrictions of other proprietary libraries that do not normally accompany the operating system. Such a contradiction means you cannot use both them and the Library together in an executable that you distribute.

7. You may place library facilities that are a work based on the Library side-by-side in a single library together with other library facilities not covered by this License, and distribute such a combined library, provided that the separate distribution of the work based on the Library and of the other library facilities is otherwise permitted, and provided that you do these two things: a. Accompany the combined library with a copy of the same work based on the

Library, uncombined with any other library facilities. This must be distributed under the terms of the Sections above.

b. Give prominent notice with the combined library of the fact that part of it is a work based on the Library, and explaining where to find the accompanying uncombined form of the same work.

8. You may not copy, modify, sublicense, link with, or distribute the Library except as expressly provided under this License. Any attempt otherwise to copy, modify, sublicense, link with, or distribute the Library is void, and will automatically terminate your rights under this License. However, parties who have received copies, or rights, from you under this License will not have their licenses terminated so long as such parties remain in full compliance.

9. You are not required to accept this License, since you have not signed it. However, nothing else grants you permission to modify or distribute the Library or its derivative works. These actions are prohibited by law if you do not accept this License. Therefore, by modifying or distributing the Library (or any work based on the Library), you indicate your acceptance of this License to do so, and all its terms and conditions for copying, distributing or modifying the Library or works based on it.

10. Each time you redistribute the Library (or any work based on the Library), the recipient automatically receives a license from the original licensor to copy, distribute, link with or modify the Library subject to these terms and conditions. You may not impose any further restrictions on the recipients’ exercise of the rights granted herein. You are not responsible for enforcing compliance by third parties with this License.

11. If, as a consequence of a court judgment or allegation of patent infringement or for any other reason (not limited to patent issues), conditions are imposed on you (whether by court order, agreement or otherwise) that contradict the conditions of this License, they do not excuse you from the conditions of this License. If you cannot distribute so as to satisfy simultaneously your obligations under this License and any other pertinent obligations, then as a consequence you may not distribute the Library at all. For example, if a patent license would not permit royalty-free redistribution of the Library by all those who receive copies directly or indirectly through you, then the only way you could satisfy both it and this License would be to refrain entirely from distribution of the Library.

If any portion of this section is held invalid or unenforceable under any particular circumstance, the balance of the section is intended to apply, and the section as a whole is intended to apply in other circumstances.

It is not the purpose of this section to induce you to infringe any patents or other property right claims or to contest validity of any such claims; this section has the sole purpose of protecting the integrity of the free software distribution system which is implemented by public license practices. Many people have made generous contributions to the wide range of software distributed through that system in reliance on consistent application of that system; it is up to the author/donor to decide if he or she is willing to distribute software through any other system and a licensee cannot impose that choice.

This section is intended to make thoroughly clear what is believed to be a consequence of the rest of this License.

12. If the distribution and/or use of the Library is restricted in certain countries either by patents or by copyrighted interfaces, the original copyright holder who places the Library under this License may add an explicit geographical distribution limitation excluding those countries, so that distribution is permitted only in or among countries not thus excluded. In such case, this License incorporates the limitation as if written in the body of this License.

13. The Free Software Foundation may publish revised and/or new versions of the Lesser General Public License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

Each version is given a distinguishing version number. If the Library specifies a version number of this License which applies to it and “any later version”, you have the option of following the terms and conditions either of that version or of any later version published by the Free Software Foundation. If the Library does not specify a license version number, you may choose any version ever published by the Free Software Foundation.

14. If you wish to incorporate parts of the Library into other free programs whose distribution conditions are incompatible with these, write to the author to ask for permission. For software which is copyrighted by the Free Software Foundation, write to the Free Software Foundation; we sometimes make exceptions for this. Our decision will be guided by the two goals of preserving the free status of all derivatives of our free software and of promoting the sharing and reuse of software generally.

NO WARRANTY

15. BECAUSE THE LIBRARY IS LICENSED FREE OF CHARGE, THERE IS NO WARRANTY FOR THE LIBRARY, TO THE EXTENT PERMITTED BY APPLI-CABLE LAW. EXCEPT WHEN OTHERWISE STATED IN WRITING THE COPY-RIGHT HOLDERS AND/OR OTHER PARTIES PROVIDE THE LIBRARY “AS IS” WITHOUT WARRANTY OF ANY KIND, EITHER EXPRESSED OR IMPLIED, INCLUDING, BUT NOT LIMITED TO, THE IMPLIED WARRANTIES OF MER-CHANTABILITY AND FITNESS FOR A PARTICULAR PURPOSE. THE ENTIRE RISK AS TO THE QUALITY AND PERFORMANCE OF THE LIBRARY IS WITH YOU. SHOULD THE LIBRARY PROVE DEFECTIVE, YOU ASSUME THE COST OF ALL NECESSARY SERVICING, REPAIR OR CORRECTION.

16. IN NO EVENT UNLESS REQUIRED BY APPLICABLE LAW OR AGREED TO IN WRITING WILL ANY COPYRIGHT HOLDER, OR ANY OTHER PARTY WHO MAY MODIFY AND/OR REDISTRIBUTE THE LIBRARY AS PERMITTED ABOVE, BE LIABLE TO YOU FOR DAMAGES, INCLUDING ANY GENERAL, SPECIAL, INCIDENTAL OR CONSEQUENTIAL DAMAGES ARISING OUT OF THE USE OR INABILITY TO USE THE LIBRARY (INCLUDING BUT NOT LIMITED TO LOSS OF DATA OR DATA BEING RENDERED INACCURATE OR LOSSES SUSTAINED BY YOU OR THIRD PARTIES OR A FAILURE OF THE LIBRARY TO OPERATE WITH ANY OTHER SOFTWARE), EVEN IF SUCH HOLDER OR OTHER PARTY HAS BEEN ADVISED OF THE POSSIBILITY OF SUCH DAMAGES.

END OF TERMS AND CONDITIONS

How to Apply These Terms to Your New Libraries

If you develop a new library, and you want it to be of the greatest possible use to the public, we recommend making it free software that everyone can redistribute and change. You can do so by permitting redistribution under these terms (or, alternatively, under the terms of the ordinary General Public License). To apply these terms, attach the following notices to the library. It is safest to attach them to the start of each source file to most effectively convey the exclusion of warranty; and each file should have at least the “copyright” line and a pointer to where the full notice is found.

      one line to give the library's name and an idea of what it does.
      Copyright (C) year  name of author
      This library is free software; you can redistribute it and/or modify it
      under the terms of the GNU Lesser General Public License as published by
      the Free Software Foundation; either version 2.1 of the License, or (at
      your option) any later version.
      This library is distributed in the hope that it will be useful, but
      WITHOUT ANY WARRANTY; without even the implied warranty of
      MERCHANTABILITY or FITNESS FOR A PARTICULAR PURPOSE.  See the GNU
      Lesser General Public License for more details.
      You should have received a copy of the GNU Lesser General Public
      License along with this library; if not, write to the Free Software
      Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301,
      USA.

Also add information on how to contact you by electronic and paper mail. You should also get your employer (if you work as a programmer) or your school, if any, to sign a “copyright disclaimer” for the library, if necessary. Here is a sample; alter the names:

      Yoyodyne, Inc., hereby disclaims all copyright interest in the library
      `Frob' (a library for tweaking knobs) written by James Random Hacker.
      signature of Ty Coon, 1 April 1990
      Ty Coon, President of Vice

That’s all there is to it!

A.3 GNU General Public License

Version 2, June 1991

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1989, 1991 Free Software Foundation, Inc.

51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

Preamble

The licenses for most software are designed to take away your freedom to share and change it. By contrast, the GNU General Public License is intended to guarantee your freedom to share and change free software—to make sure the software is free for all its users. This General Public License applies to most of the Free Software Foundation’s software and to any other program whose authors commit to using it. (Some other Free Software Foundation software is covered by the GNU Lesser General Public License instead.) You can apply it to your programs, too. When we speak of free software, we are referring to freedom, not price. Our General Public Licenses are designed to make sure that you have the freedom to distribute copies of free software (and charge for this service if you wish), that you receive source code or can get it if you want it, that you can change the software or use pieces of it in new free programs; and that you know you can do these things. To protect your rights, we need to make restrictions that forbid anyone to deny you these rights or to ask you to surrender the rights. These restrictions translate to certain responsibilities for you if you distribute copies of the software, or if you modify it. For example, if you distribute copies of such a program, whether gratis or for a fee, you must give the recipients all the rights that you have. You must make sure that they, too, receive or can get the source code. And you must show them these terms so they know their rights. We protect your rights with two steps: (1) copyright the software, and (2) offer you this license which gives you legal permission to copy, distribute and/or modify the software. Also, for each author’s protection and ours, we want to make certain that everyone understands that there is no warranty for this free software. If the software is modified by someone else and passed on, we want its recipients to know that what they have is not the original, so that any problems introduced by others will not reflect on the original authors’ reputations. Finally, any free program is threatened constantly by software patents. We wish to avoid the danger that redistributors of a free program will individually obtain patent licenses, in effect making the program proprietary. To prevent this, we have made it clear that any patent must be licensed for everyone’s free use or not licensed at all. The precise terms and conditions for copying, distribution and modification follow.

TERMS AND CONDITIONS FOR COPYING, DISTRIBUTION AND MODIFICATION

0. This License applies to any program or other work which contains a notice placed by the copyright holder saying it may be distributed under the terms of this General Public License. The “Program”, below, refers to any such program or work, and a “work based on the Program” means either the Program or any derivative work under copyright law: that is to say, a work containing the Program or a portion of it, either verbatim or with modifications and/or translated into another language. (Hereinafter, translation is included without limitation in the term “modification”.) Each licensee is addressed as “you”. Activities other than copying, distribution and modification are not covered by this License; they are outside its scope. The act of running the Program is not restricted, and the output from the Program is covered only if its contents constitute a work based on the Program (independent of having been made by running the Program). Whether that is true depends on what the Program does.

1. You may copy and distribute verbatim copies of the Program’s source code as you receive it, in any medium, provided that you conspicuously and appropriately publish on each copy an appropriate copyright notice and disclaimer of warranty; keep intact all the notices that refer to this License and to the absence of any warranty; and give any other recipients of the Program a copy of this License along with the Program. You may charge a fee for the physical act of transferring a copy, and you may at your option offer warranty protection in exchange for a fee.

2. You may modify your copy or copies of the Program or any portion of it, thus forming a work based on the Program, and copy and distribute such modifications or work under the terms of Section 1 above, provided that you also meet all of these conditions: a. You must cause the modified files to carry prominent notices stating that you

changed the files and the date of any change. b. You must cause any work that you distribute or publish, that in whole or in part

contains or is derived from the Program or any part thereof, to be licensed as a whole at no charge to all third parties under the terms of this License.

c. If the modified program normally reads commands interactively when run, you must cause it, when started running for such interactive use in the most ordinary way, to print or display an announcement including an appropriate copyright notice and a notice that there is no warranty (or else, saying that you provide a warranty) and that users may redistribute the program under these conditions, and telling the user how to view a copy of this License. (Exception: if the Program itself is interactive but does not normally print such an announcement, your work based on the Program is not required to print an announcement.)

These requirements apply to the modified work as a whole. If identifiable sections of that work are not derived from the Program, and can be reasonably considered independent and separate works in themselves, then this License, and its terms, do not apply to those sections when you distribute them as separate works. But when you distribute the same sections as part of a whole which is a work based on the Program, the distribution of the whole must be on the terms of this License, whose permissions for other licensees extend to the entire whole, and thus to each and every part regardless of who wrote it. Thus, it is not the intent of this section to claim rights or contest your rights to work written entirely by you; rather, the intent is to exercise the right to control the distribution of derivative or collective works based on the Program. In addition, mere aggregation of another work not based on the Program with the Program (or with a work based on the Program) on a volume of a storage or distribution medium does not bring the other work under the scope of this License.

3. You may copy and distribute the Program (or a work based on it, under Section 2) in object code or executable form under the terms of Sections 1 and 2 above provided that you also do one of the following: a. Accompany it with the complete corresponding machine-readable source code,

which must be distributed under the terms of Sections 1 and 2 above on a medium customarily used for software interchange; or,

b. Accompany it with a written offer, valid for at least three years, to give any third party, for a charge no more than your cost of physically performing source distribution, a complete machine-readable copy of the corresponding source code, to be distributed under the terms of Sections 1 and 2 above on a medium customarily used for software interchange; or,

c. Accompany it with the information you received as to the offer to distribute corresponding source code. (This alternative is allowed only for noncommercial distribution and only if you received the program in object code or executable form with such an offer, in accord with Subsection b above.)

The source code for a work means the preferred form of the work for making modifications to it. For an executable work, complete source code means all the source code for all modules it contains, plus any associated interface definition files, plus the scripts used to control compilation and installation of the executable. However, as a special exception, the source code distributed need not include anything that is normally distributed (in either source or binary form) with the major components (compiler, kernel, and so on) of the operating system on which the executable runs, unless that component itself accompanies the executable.

If distribution of executable or object code is made by offering access to copy from a designated place, then offering equivalent access to copy the source code from the same place counts as distribution of the source code, even though third parties are not compelled to copy the source along with the object code.

4. You may not copy, modify, sublicense, or distribute the Program except as expressly provided under this License. Any attempt otherwise to copy, modify, sublicense or distribute the Program is void, and will automatically terminate your rights under this License. However, parties who have received copies, or rights, from you under this License will not have their licenses terminated so long as such parties remain in full compliance.

5. You are not required to accept this License, since you have not signed it. However, nothing else grants you permission to modify or distribute the Program or its derivative works. These actions are prohibited by law if you do not accept this License. Therefore, by modifying or distributing the Program (or any work based on the Program), you indicate your acceptance of this License to do so, and all its terms and conditions for copying, distributing or modifying the Program or works based on it.

6. Each time you redistribute the Program (or any work based on the Program), the recipient automatically receives a license from the original licensor to copy, distribute or modify the Program subject to these terms and conditions. You may not impose any further restrictions on the recipients’ exercise of the rights granted herein. You are not responsible for enforcing compliance by third parties to this License.

7. If, as a consequence of a court judgment or allegation of patent infringement or for any other reason (not limited to patent issues), conditions are imposed on you (whether by court order, agreement or otherwise) that contradict the conditions of this License, they do not excuse you from the conditions of this License. If you cannot distribute so as to satisfy simultaneously your obligations under this License and any other pertinent obligations, then as a consequence you may not distribute the Program at all. For example, if a patent license would not permit royalty-free redistribution of the Program by all those who receive copies directly or indirectly through you, then the only way you could satisfy both it and this License would be to refrain entirely from distribution of the Program. If any portion of this section is held invalid or unenforceable under any particular circumstance, the balance of the section is intended to apply and the section as a whole is intended to apply in other circumstances. It is not the purpose of this section to induce you to infringe any patents or other property right claims or to contest validity of any such claims; this section has the sole purpose of protecting the integrity of the free software distribution system, which is implemented by public license practices. Many people have made generous contributions to the wide range of software distributed through that system in reliance on consistent application of that system; it is up to the author/donor to decide if he or she is willing to distribute software through any other system and a licensee cannot impose that choice. This section is intended to make thoroughly clear what is believed to be a consequence of the rest of this License.

8. If the distribution and/or use of the Program is restricted in certain countries either by patents or by copyrighted interfaces, the original copyright holder who places the Program under this License may add an explicit geographical distribution limitation excluding those countries, so that distribution is permitted only in or among countries not thus excluded. In such case, this License incorporates the limitation as if written in the body of this License.

9. The Free Software Foundation may publish revised and/or new versions of the General Public License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns. Each version is given a distinguishing version number. If the Program specifies a version number of this License which applies to it and “any later version”, you have the option of following the terms and conditions either of that version or of any later version published by the Free Software Foundation. If the Program does not specify a version number of this License, you may choose any version ever published by the Free Software Foundation.

10. If you wish to incorporate parts of the Program into other free programs whose distribution conditions are different, write to the author to ask for permission. For software which is copyrighted by the Free Software Foundation, write to the Free Software Foundation; we sometimes make exceptions for this. Our decision will be guided by the two goals of preserving the free status of all derivatives of our free software and of promoting the sharing and reuse of software generally.

NO WARRANTY

11. BECAUSE THE PROGRAM IS LICENSED FREE OF CHARGE, THERE IS NO WARRANTY FOR THE PROGRAM, TO THE EXTENT PERMITTED BY APPLI-CABLE LAW. EXCEPT WHEN OTHERWISE STATED IN WRITING THE COPY-RIGHT HOLDERS AND/OR OTHER PARTIES PROVIDE THE PROGRAM “AS IS” WITHOUT WARRANTY OF ANY KIND, EITHER EXPRESSED OR IMPLIED, INCLUDING, BUT NOT LIMITED TO, THE IMPLIED WARRANTIES OF MER-CHANTABILITY AND FITNESS FOR A PARTICULAR PURPOSE. THE ENTIRE RISK AS TO THE QUALITY AND PERFORMANCE OF THE PROGRAM IS WITH YOU. SHOULD THE PROGRAM PROVE DEFECTIVE, YOU ASSUME THE COST OF ALL NECESSARY SERVICING, REPAIR OR CORRECTION.

12. IN NO EVENT UNLESS REQUIRED BY APPLICABLE LAW OR AGREED TO IN WRITING WILL ANY COPYRIGHT HOLDER, OR ANY OTHER PARTY WHO MAY MODIFY AND/OR REDISTRIBUTE THE PROGRAM AS PERMITTED ABOVE, BE LIABLE TO YOU FOR DAMAGES, INCLUDING ANY GENERAL, SPECIAL, INCIDENTAL OR CONSEQUENTIAL DAMAGES ARISING OUT OF THE USE OR INABILITY TO USE THE PROGRAM (INCLUDING BUT NOT LIMITED TO LOSS OF DATA OR DATA BEING RENDERED INACCURATE OR LOSSES SUSTAINED BY YOU OR THIRD PARTIES OR A FAILURE OF THE PROGRAM TO OPERATE WITH ANY OTHER PROGRAMS), EVEN IF SUCH HOLDER OR OTHER PARTY HAS BEEN ADVISED OF THE POSSIBILITY OF SUCH DAMAGES.

END OF TERMS AND CONDITIONS Appendix: How to Apply These Terms to Your New Programs

If you develop a new program, and you want it to be of the greatest possible use to the public, the best way to achieve this is to make it free software which everyone can redistribute and change under these terms. To do so, attach the following notices to the program. It is safest to attach them to the start of each source file to most effectively convey the exclusion of warranty; and each file should have at least the “copyright” line and a pointer to where the full notice is found.

      one line to give the program's name and a brief idea of what it does.
      Copyright (C) yyyy  name of author
      This program is free software; you can redistribute it and/or modify
      it under the terms of the GNU General Public License as published by
      the Free Software Foundation; either version 2 of the License, or
      (at your option) any later version.
      This program is distributed in the hope that it will be useful,
      but WITHOUT ANY WARRANTY; without even the implied warranty of
      MERCHANTABILITY or FITNESS FOR A PARTICULAR PURPOSE.  See the
      GNU General Public License for more details.
      You should have received a copy of the GNU General Public License
      along with this program; if not, write to the Free Software
      Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA  02110-1301, USA.

Also add information on how to contact you by electronic and paper mail. If the program is interactive, make it output a short notice like this when it starts in an interactive mode:

      Gnomovision version 69, Copyright (C) year name of author
      Gnomovision comes with ABSOLUTELY NO WARRANTY; for details type `show w'.
      This is free software, and you are welcome to redistribute it
      under certain conditions; type `show c' for details.

The hypothetical commands ‘show w’ and ‘show c’ should show the appropriate parts of the General Public License. Of course, the commands you use may be called something other than ‘show w’ and ‘show c’; they could even be mouse-clicks or menu items—whatever suits your program. You should also get your employer (if you work as a programmer) or your school, if any, to sign a “copyright disclaimer” for the program, if necessary. Here is a sample; alter the names:

     Yoyodyne, Inc., hereby disclaims all copyright interest in the program
     `Gnomovision' (which makes passes at compilers) written by James Hacker.
     signature of Ty Coon, 1 April 1989
     Ty Coon, President of Vice

This General Public License does not permit incorporating your program into proprietary programs. If your program is a subroutine library, you may consider it more useful to permit linking proprietary applications with the library. If this is what you want to do, use the GNU Lesser General Public License instead of this License.

Concept Index 239

Concept Index

A

Alert protocol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Anonymous authentication . . . . . . . . . . . . . . . . . . . 14

C

Callback functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Certificate authentication. . . . . . . . . . . . . . . . . . . . . 18 Certificate requests . . . . . . . . . . . . . . . . . . . . . . . . . . 21 Certificate to XML convertion . . . . . . . . . . . . . . . 206 certtool . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 Ciphersuites . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 Client Certificate authentication . . . . . . . . . . . . . . . 9 Compression algorithms . . . . . . . . . . . . . . . . . . . . . . . 7

D

Digital signatures . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

E

Error codes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 Example programs . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

F

FDL, GNU Free Documentation License . . . . . . 218 Function reference . . . . . . . . . . . . . . . . . . . . . . . . . . . 95

G

gnutls-cli . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 gnutls-cli-debug. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 GnuTLS-extra functions. . . . . . . . . . . . . . . . . . . . . . 184 gnutls-serv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 GPL, GNU General Public License. . . . . . . . . . . 232

H

Handshake protocol . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

I

Inner Application (TLS/IA) functions. . . . . . . . . 193 Internal architecture . . . . . . . . . . . . . . . . . . . . . . . . 213

L

LGPL, GNU Lesser General Public License … 224 License, GNU GPL . . . . . . . . . . . . . . . . . . . . . . . . . 232 License, GNU LGPL . . . . . . . . . . . . . . . . . . . . . . . . 224

M

Maximum fragment length . . . . . . . . . . . . . . . . . . . 11

O

OpenPGP functions . . . . . . . . . . . . . . . . . . . . . . . . . 184 OpenPGP Keys . . . . . . . . . . . . . . . . . . . . . . . . . . . 12, 21 OpenPGP Server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 OpenSSL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87

P

PCT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 PKCS #10 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 PKCS #12 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 PSK authentication . . . . . . . . . . . . . . . . . . . . . . . . . . 15

R

Record protocol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Resuming sessions . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

S

Server name indication . . . . . . . . . . . . . . . . . . . . . . . 11 SRP authentication . . . . . . . . . . . . . . . . . . . . . . . . . . 14 srptool . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 SSL 2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Symmetric encryption algorithms . . . . . . . . . . . . . . 7

T

TLS Extensions . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 11 TLS Inner Application (TLS/IA) functions …. 193 TLS Layers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Transport protocol . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

V

Verifying certificate paths . . . . . . . . . . . . . . . . . . . . 19

X

X.509 certificates . . . . . . . . . . . . . . . . . . . . . . . . . 12, 18 X.509 Functions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142

Function and Data Index 240

Function and Data Index

gnutls_alert_get . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 gnutls_alert_get_name . . . . . . . . . . . . . . . . . . . . . 95 gnutls_alert_send . . . . . . . . . . . . . . . . . . . . . . . . . . 95

gnutls_alert_send_appropriate . . . . . . . . . . . . . 95
gnutls_anon_allocate_client_credentials . . 96
gnutls_anon_allocate_server_credentials . . 96
gnutls_anon_free_client_credentials . . . . . . 96
gnutls_anon_free_server_credentials . . . . . . 96
gnutls_anon_set_params_function . . . . . . . . . . . 97
gnutls_anon_set_server_dh_params. . . . . . . . . . 97
gnutls_anon_set_server_params_function . . . 97
gnutls_auth_client_get_type . . . . . . . . . . . . . . . 97

gnutls_auth_get_type. . . . . . . . . . . . . . . . . . . . . . . 97

gnutls_auth_server_get_type . . . . . . . . . . . . . . . 98

gnutls_bye . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98

gnutls_certificate_activation_time_peers

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98

gnutls_certificate_allocate_credentials . . 98
gnutls_certificate_client_get_request_

status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99

gnutls_certificate_client_set_retrieve_

function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99

gnutls_certificate_expiration_time_peers

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99

gnutls_certificate_free_ca_names. . . . . . . . . 100
gnutls_certificate_free_cas . . . . . . . . . . . . . . 100
gnutls_certificate_free_credentials . . . . . 100
gnutls_certificate_free_crls . . . . . . . . . . . . . 100
gnutls_certificate_free_keys . . . . . . . . . . . . . 100
gnutls_certificate_get_ours . . . . . . . . . . . . . . 101
gnutls_certificate_get_peers . . . . . . . . . . . . . 101
gnutls_certificate_send_x509_rdn_sequence

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101

gnutls_certificate_server_set_request . . . 101
gnutls_certificate_server_set_retrieve_

function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102

gnutls_certificate_set_dh_params. . . . . . . . . 102
gnutls_certificate_set_openpgp_key . . . . . . 186
gnutls_certificate_set_openpgp_key_file

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184

gnutls_certificate_set_openpgp_key_mem . . 185
gnutls_certificate_set_openpgp_keyring_file

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185

gnutls_certificate_set_openpgp_keyring_mem

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185

gnutls_certificate_set_openpgp_keyserver

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185

gnutls_certificate_set_openpgp_trustdb . . 186
gnutls_certificate_set_params_function . . 102
gnutls_certificate_set_rsa_export_params

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102

gnutls_certificate_set_verify_flags . . . . . 103
gnutls_certificate_set_verify_limits . . . . 103
gnutls_certificate_set_x509_crl. . . . . . . . . . 104
gnutls_certificate_set_x509_crl_file . . . . 103

gnutls_certificate_set_x509_crl_mem . . . . . 103
gnutls_certificate_set_x509_key. . . . . . . . . . 105
gnutls_certificate_set_x509_key_file . . . . 104
gnutls_certificate_set_x509_key_mem . . . . . 104
gnutls_certificate_set_x509_simple_pkcs12_

file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105

gnutls_certificate_set_x509_trust . . . . . . . 106
gnutls_certificate_set_x509_trust_file . . 106
gnutls_certificate_set_x509_trust_mem . . . 106
gnutls_certificate_type_get . . . . . . . . . . . . . . 107
gnutls_certificate_type_get_name. . . . . . . . . 107
gnutls_certificate_type_set_priority . . . . 107
gnutls_certificate_verify_flags . . . . . . . . . . . 20
gnutls_certificate_verify_peers. . . . . . . . . . 108
gnutls_certificate_verify_peers2. . . . . . . . . 107

gnutls_check_version. . . . . . . . . . . . . . . . . . . . . . 108 gnutls_cipher_get . . . . . . . . . . . . . . . . . . . . . . . . . 108

gnutls_cipher_get_key_size . . . . . . . . . . . . . . . 108
gnutls_cipher_get_name . . . . . . . . . . . . . . . . . . . 108
gnutls_cipher_set_priority . . . . . . . . . . . . . . . 109
gnutls_cipher_suite_get_name . . . . . . . . . . . . . 109
gnutls_compression_get . . . . . . . . . . . . . . . . . . . 109
gnutls_compression_get_name . . . . . . . . . . . . . . 109
gnutls_compression_set_priority. . . . . . . . . . 109
gnutls_credentials_clear . . . . . . . . . . . . . . . . . 110
gnutls_credentials_set . . . . . . . . . . . . . . . . . . . 110

gnutls_db_check_entry . . . . . . . . . . . . . . . . . . . . 110 gnutls_db_get_ptr . . . . . . . . . . . . . . . . . . . . . . . . . 111

gnutls_db_remove_session . . . . . . . . . . . . . . . . . 111
gnutls_db_set_cache_expiration . . . . . . . . . . . 111

gnutls_db_set_ptr . . . . . . . . . . . . . . . . . . . . . . . . . 111

gnutls_db_set_remove_function . . . . . . . . . . . . 111
gnutls_db_set_retrieve_function. . . . . . . . . . 112
gnutls_db_set_store_function . . . . . . . . . . . . . 112

gnutls_deinit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 gnutls_dh_get_group. . . . . . . . . . . . . . . . . . . . . . . 112

gnutls_dh_get_peers_public_bits. . . . . . . . . . 113
gnutls_dh_get_prime_bits . . . . . . . . . . . . . . . . . 113

gnutls_dh_get_pubkey. . . . . . . . . . . . . . . . . . . . . . 113

gnutls_dh_get_secret_bits . . . . . . . . . . . . . . . . 113

gnutls_dh_params_cpy. . . . . . . . . . . . . . . . . . . . . . 113

gnutls_dh_params_deinit . . . . . . . . . . . . . . . . . . 113
gnutls_dh_params_export_pkcs3 . . . . . . . . . . . . 114
gnutls_dh_params_export_raw . . . . . . . . . . . . . . 114
gnutls_dh_params_generate2 . . . . . . . . . . . . . . . 114
gnutls_dh_params_import_pkcs3 . . . . . . . . . . . . 115
gnutls_dh_params_import_raw . . . . . . . . . . . . . . 115

gnutls_dh_params_init . . . . . . . . . . . . . . . . . . . . 115

gnutls_dh_set_prime_bits . . . . . . . . . . . . . . . . . 115

gnutls_error_is_fatal . . . . . . . . . . . . . . . . . . . . 116 gnutls_error_to_alert . . . . . . . . . . . . . . . . . . . . 116

gnutls_extra_check_version . . . . . . . . . . . . . . . 184

gnutls_fingerprint. . . . . . . . . . . . . . . . . . . . . . . . 116 gnutls_free . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 gnutls_global_deinit. . . . . . . . . . . . . . . . . . . . . . 117 gnutls_global_init. . . . . . . . . . . . . . . . . . . . . . . . 117

gnutls_global_init_extra . . . . . . . . . . . . . . . . . 184
gnutls_global_set_log_function . . . . . . . . . . . 117
gnutls_global_set_log_level . . . . . . . . . . . . . . 117
gnutls_global_set_mem_functions. . . . . . . . . . 118

gnutls_handshake . . . . . . . . . . . . . . . . . . . . . . . . . . 119

gnutls_handshake_get_last_in . . . . . . . . . . . . . 118
gnutls_handshake_get_last_out . . . . . . . . . . . . 118
gnutls_handshake_set_max_packet_length . . 118
gnutls_handshake_set_private_extensions

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 gnutls_hex_decode . . . . . . . . . . . . . . . . . . . . . . . . . 119 gnutls_hex_encode . . . . . . . . . . . . . . . . . . . . . . . . . 120

gnutls_ia_allocate_client_credentials . . . 193
gnutls_ia_allocate_server_credentials . . . 194

gnutls_ia_enable . . . . . . . . . . . . . . . . . . . . . . . . . . 194

gnutls_ia_endphase_send . . . . . . . . . . . . . . . . . . 195
gnutls_ia_extract_inner_secret . . . . . . . . . . . 195
gnutls_ia_free_client_credentials . . . . . . . 195
gnutls_ia_free_server_credentials . . . . . . . 195
gnutls_ia_generate_challenge . . . . . . . . . . . . . 196
gnutls_ia_get_client_avp_ptr . . . . . . . . . . . . . 196
gnutls_ia_get_server_avp_ptr . . . . . . . . . . . . . 196

gnutls_ia_handshake. . . . . . . . . . . . . . . . . . . . . . . 196 gnutls_ia_handshake_p . . . . . . . . . . . . . . . . . . . . 196

gnutls_ia_permute_inner_secret . . . . . . . . . . . 197

gnutls_ia_recv . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 gnutls_ia_send . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197

gnutls_ia_set_client_avp_function . . . . . . . 198
gnutls_ia_set_client_avp_ptr . . . . . . . . . . . . . 198
gnutls_ia_set_server_avp_function . . . . . . . 199
gnutls_ia_set_server_avp_ptr . . . . . . . . . . . . . 199
gnutls_ia_verify_endphase . . . . . . . . . . . . . . . . 199

gnutls_init . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 gnutls_kx_get . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 gnutls_kx_get_name. . . . . . . . . . . . . . . . . . . . . . . . 120

gnutls_kx_set_priority . . . . . . . . . . . . . . . . . . . 120

gnutls_mac_get . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 gnutls_mac_get_name. . . . . . . . . . . . . . . . . . . . . . . 121

gnutls_mac_set_priority . . . . . . . . . . . . . . . . . . 121

gnutls_malloc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121

gnutls_openpgp_key_check_hostname . . . . . . . 186
gnutls_openpgp_key_deinit . . . . . . . . . . . . . . . . 186
gnutls_openpgp_key_export . . . . . . . . . . . . . . . . 186
gnutls_openpgp_key_get_creation_time . . . . 187
gnutls_openpgp_key_get_expiration_time . . 187
gnutls_openpgp_key_get_fingerprint . . . . . . 187
gnutls_openpgp_key_get_id . . . . . . . . . . . . . . . . 187
gnutls_openpgp_key_get_key_usage. . . . . . . . . 187
gnutls_openpgp_key_get_name . . . . . . . . . . . . . . 188
gnutls_openpgp_key_get_pk_algorithm . . . . . 188
gnutls_openpgp_key_get_version . . . . . . . . . . . 188
gnutls_openpgp_key_import . . . . . . . . . . . . . . . . 188
gnutls_openpgp_key_init . . . . . . . . . . . . . . . . . . 189
gnutls_openpgp_key_to_xml . . . . . . . . . . . . . . . . 189
gnutls_openpgp_key_verify_ring . . . . . . . . . . . 189
gnutls_openpgp_key_verify_self . . . . . . . . . . . 189
gnutls_openpgp_key_verify_trustdb . . . . . . . 190

                                                gnutls_openpgp_keyring_check_id. . . . . . . . . . 190
                                                gnutls_openpgp_keyring_deinit . . . . . . . . . . . . 190
                                                gnutls_openpgp_keyring_import . . . . . . . . . . . . 190
                                                gnutls_openpgp_keyring_init . . . . . . . . . . . . . . 191
                                                gnutls_openpgp_privkey_deinit . . . . . . . . . . . . 191
                                                gnutls_openpgp_privkey_get_pk_algorithm

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191

                                                gnutls_openpgp_privkey_import . . . . . . . . . . . . 191
                                                gnutls_openpgp_privkey_init . . . . . . . . . . . . . . 192
                                                gnutls_openpgp_send_key . . . . . . . . . . . . . . . . . . 121
                                                gnutls_openpgp_set_recv_key_function . . . . 192
                                                gnutls_openpgp_trustdb_deinit . . . . . . . . . . . . 192
                                                gnutls_openpgp_trustdb_import_file . . . . . . 192
                                                gnutls_openpgp_trustdb_init . . . . . . . . . . . . . . 192
                                                gnutls_pem_base64_decode . . . . . . . . . . . . . . . . . 122
                                                gnutls_pem_base64_decode_alloc . . . . . . . . . . . 122
                                                gnutls_pem_base64_encode . . . . . . . . . . . . . . . . . 122
                                                gnutls_pem_base64_encode_alloc . . . . . . . . . . . 122

gnutls_perror . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123

                                                gnutls_pk_algorithm_get_name . . . . . . . . . . . . . 123
                                                gnutls_pkcs12_bag_decrypt . . . . . . . . . . . . . . . . 142
                                                gnutls_pkcs12_bag_deinit . . . . . . . . . . . . . . . . . 142
                                                gnutls_pkcs12_bag_encrypt . . . . . . . . . . . . . . . . 142
                                                gnutls_pkcs12_bag_get_count . . . . . . . . . . . . . . 142
                                                gnutls_pkcs12_bag_get_data . . . . . . . . . . . . . . . 143
                                                gnutls_pkcs12_bag_get_friendly_name . . . . . 143
                                                gnutls_pkcs12_bag_get_key_id . . . . . . . . . . . . . 143
                                                gnutls_pkcs12_bag_get_type . . . . . . . . . . . . . . . 143
                                                gnutls_pkcs12_bag_init . . . . . . . . . . . . . . . . . . . 144
                                                gnutls_pkcs12_bag_set_crl . . . . . . . . . . . . . . . . 144
                                                gnutls_pkcs12_bag_set_crt . . . . . . . . . . . . . . . . 144
                                                gnutls_pkcs12_bag_set_data . . . . . . . . . . . . . . . 144
                                                gnutls_pkcs12_bag_set_friendly_name . . . . . 144
                                                gnutls_pkcs12_bag_set_key_id . . . . . . . . . . . . . 145

gnutls_pkcs12_deinit. . . . . . . . . . . . . . . . . . . . . . 145 gnutls_pkcs12_export. . . . . . . . . . . . . . . . . . . . . . 145

                                                gnutls_pkcs12_generate_mac . . . . . . . . . . . . . . . 145

gnutls_pkcs12_get_bag . . . . . . . . . . . . . . . . . . . . 146 gnutls_pkcs12_import. . . . . . . . . . . . . . . . . . . . . . 146 gnutls_pkcs12_init. . . . . . . . . . . . . . . . . . . . . . . . 146 gnutls_pkcs12_set_bag . . . . . . . . . . . . . . . . . . . . 146

                                                gnutls_pkcs12_verify_mac . . . . . . . . . . . . . . . . . 146

gnutls_pkcs7_deinit. . . . . . . . . . . . . . . . . . . . . . . 147

                                                gnutls_pkcs7_delete_crl . . . . . . . . . . . . . . . . . . 147
                                                gnutls_pkcs7_delete_crt . . . . . . . . . . . . . . . . . . 147

gnutls_pkcs7_export. . . . . . . . . . . . . . . . . . . . . . . 147

                                                gnutls_pkcs7_get_crl_count . . . . . . . . . . . . . . . 147
                                                gnutls_pkcs7_get_crl_raw . . . . . . . . . . . . . . . . . 148
                                                gnutls_pkcs7_get_crt_count . . . . . . . . . . . . . . . 148
                                                gnutls_pkcs7_get_crt_raw . . . . . . . . . . . . . . . . . 148

gnutls_pkcs7_import. . . . . . . . . . . . . . . . . . . . . . . 148 gnutls_pkcs7_init . . . . . . . . . . . . . . . . . . . . . . . . . 149 gnutls_pkcs7_set_crl. . . . . . . . . . . . . . . . . . . . . . 149

                                                gnutls_pkcs7_set_crl_raw . . . . . . . . . . . . . . . . . 149

gnutls_pkcs7_set_crt. . . . . . . . . . . . . . . . . . . . . . 149

                                                gnutls_pkcs7_set_crt_raw . . . . . . . . . . . . . . . . . 149

gnutls_prf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 gnutls_prf_raw . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123

gnutls_protocol_get_name . . . . . . . . . . . . . . . . . 124
gnutls_protocol_get_version . . . . . . . . . . . . . . 124
gnutls_protocol_set_priority . . . . . . . . . . . . . 125
gnutls_psk_allocate_client_credentials . . 125
gnutls_psk_allocate_server_credentials . . 125
gnutls_psk_free_client_credentials . . . . . . 125
gnutls_psk_free_server_credentials . . . . . . 125
gnutls_psk_server_get_username . . . . . . . . . . . 126
gnutls_psk_set_client_credentials . . . . . . . 126
gnutls_psk_set_client_credentials_function

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126

gnutls_psk_set_params_function . . . . . . . . . . . 126
gnutls_psk_set_server_credentials_file . . 127
gnutls_psk_set_server_credentials_function

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127

gnutls_psk_set_server_dh_params. . . . . . . . . . 127
gnutls_psk_set_server_params_function . . . 127
gnutls_record_check_pending . . . . . . . . . . . . . . 128
gnutls_record_get_direction . . . . . . . . . . . . . . 128
gnutls_record_get_max_size . . . . . . . . . . . . . . . 128

gnutls_record_recv. . . . . . . . . . . . . . . . . . . . . . . . 128 gnutls_record_send. . . . . . . . . . . . . . . . . . . . . . . . 129

gnutls_record_set_max_size . . . . . . . . . . . . . . . 129

gnutls_rehandshake. . . . . . . . . . . . . . . . . . . . . . . . 130

gnutls_rsa_export_get_modulus_bits . . . . . . 130
gnutls_rsa_export_get_pubkey . . . . . . . . . . . . . 130

gnutls_rsa_params_cpy . . . . . . . . . . . . . . . . . . . . 130

gnutls_rsa_params_deinit . . . . . . . . . . . . . . . . . 130
gnutls_rsa_params_export_pkcs1 . . . . . . . . . . . 131
gnutls_rsa_params_export_raw . . . . . . . . . . . . . 131
gnutls_rsa_params_generate2 . . . . . . . . . . . . . . 131
gnutls_rsa_params_import_pkcs1 . . . . . . . . . . . 132
gnutls_rsa_params_import_raw . . . . . . . . . . . . . 132
gnutls_rsa_params_init . . . . . . . . . . . . . . . . . . . 132
gnutls_server_name_get . . . . . . . . . . . . . . . . . . . 132
gnutls_server_name_set . . . . . . . . . . . . . . . . . . . 133
gnutls_session_get_client_random. . . . . . . . . 133
gnutls_session_get_data . . . . . . . . . . . . . . . . . . 134
gnutls_session_get_data2 . . . . . . . . . . . . . . . . . 133

gnutls_session_get_id . . . . . . . . . . . . . . . . . . . . 134

gnutls_session_get_master_secret. . . . . . . . . 134
gnutls_session_get_ptr . . . . . . . . . . . . . . . . . . . 134
gnutls_session_get_server_random. . . . . . . . . 135
gnutls_session_is_resumed . . . . . . . . . . . . . . . . 135
gnutls_session_set_data . . . . . . . . . . . . . . . . . . 135
gnutls_session_set_ptr . . . . . . . . . . . . . . . . . . . 135
gnutls_set_default_export_priority . . . . . . 135
gnutls_set_default_priority . . . . . . . . . . . . . . 136
gnutls_sign_algorithm_get_name . . . . . . . . . . . 136
gnutls_srp_allocate_client_credentials . . 136
gnutls_srp_allocate_server_credentials . . 136
gnutls_srp_base64_decode . . . . . . . . . . . . . . . . . 137
gnutls_srp_base64_decode_alloc . . . . . . . . . . . 137
gnutls_srp_base64_encode . . . . . . . . . . . . . . . . . 137
gnutls_srp_base64_encode_alloc . . . . . . . . . . . 137
gnutls_srp_free_client_credentials . . . . . . 138
gnutls_srp_free_server_credentials . . . . . . 138
gnutls_srp_server_get_username . . . . . . . . . . . 138

                                                gnutls_srp_set_client_credentials . . . . . . . 139
                                                gnutls_srp_set_client_credentials_function

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138

                                                gnutls_srp_set_server_credentials_file . . 139
                                                gnutls_srp_set_server_credentials_function

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 gnutls_srp_verifier. . . . . . . . . . . . . . . . . . . . . . . 140 gnutls_strerror . . . . . . . . . . . . . . . . . . . . . . . . . . . 140

                                                gnutls_transport_get_ptr . . . . . . . . . . . . . . . . . 140
                                                gnutls_transport_get_ptr2 . . . . . . . . . . . . . . . . 140
                                                gnutls_transport_set_lowat . . . . . . . . . . . . . . . 141
                                                gnutls_transport_set_ptr . . . . . . . . . . . . . . . . . 141
                                                gnutls_transport_set_ptr2 . . . . . . . . . . . . . . . . 141
                                                gnutls_transport_set_pull_function . . . . . . 141
                                                gnutls_transport_set_push_function . . . . . . 142
                                                gnutls_x509_crl_check_issuer . . . . . . . . . . . . . 149
                                                gnutls_x509_crl_deinit . . . . . . . . . . . . . . . . . . . 150
                                                gnutls_x509_crl_export . . . . . . . . . . . . . . . . . . . 150
                                                gnutls_x509_crl_get_crt_count . . . . . . . . . . . . 150
                                                gnutls_x509_crl_get_crt_serial . . . . . . . . . . . 150
                                                gnutls_x509_crl_get_dn_oid . . . . . . . . . . . . . . . 151
                                                gnutls_x509_crl_get_issuer_dn . . . . . . . . . . . . 151
                                                gnutls_x509_crl_get_issuer_dn_by_oid . . . . 151
                                                gnutls_x509_crl_get_next_update. . . . . . . . . . 152
                                                gnutls_x509_crl_get_signature_algorithm

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152

                                                gnutls_x509_crl_get_this_update. . . . . . . . . . 152
                                                gnutls_x509_crl_get_version . . . . . . . . . . . . . . 152
                                                gnutls_x509_crl_import . . . . . . . . . . . . . . . . . . . 152

gnutls_x509_crl_init. . . . . . . . . . . . . . . . . . . . . . 153

                                                gnutls_x509_crl_set_crt . . . . . . . . . . . . . . . . . . 153
                                                gnutls_x509_crl_set_crt_serial . . . . . . . . . . . 153
                                                gnutls_x509_crl_set_next_update. . . . . . . . . . 153
                                                gnutls_x509_crl_set_this_update. . . . . . . . . . 154
                                                gnutls_x509_crl_set_version . . . . . . . . . . . . . . 154

gnutls_x509_crl_sign. . . . . . . . . . . . . . . . . . . . . . 154 gnutls_x509_crl_sign2 . . . . . . . . . . . . . . . . . . . . 154

                                                gnutls_x509_crl_verify . . . . . . . . . . . . . . . . . . . 155
                                                gnutls_x509_crq_deinit . . . . . . . . . . . . . . . . . . . 155
                                                gnutls_x509_crq_export . . . . . . . . . . . . . . . . . . . 155
                                                gnutls_x509_crq_get_attribute_by_oid . . . . 155
                                                gnutls_x509_crq_get_challenge_password . . 156
                                                gnutls_x509_crq_get_dn . . . . . . . . . . . . . . . . . . . 157
                                                gnutls_x509_crq_get_dn_by_oid . . . . . . . . . . . . 156
                                                gnutls_x509_crq_get_dn_oid . . . . . . . . . . . . . . . 157
                                                gnutls_x509_crq_get_pk_algorithm. . . . . . . . . 157
                                                gnutls_x509_crq_get_version . . . . . . . . . . . . . . 158
                                                gnutls_x509_crq_import . . . . . . . . . . . . . . . . . . . 158

gnutls_x509_crq_init. . . . . . . . . . . . . . . . . . . . . . 158

                                                gnutls_x509_crq_set_attribute_by_oid . . . . 158
                                                gnutls_x509_crq_set_challenge_password . . 158
                                                gnutls_x509_crq_set_dn_by_oid . . . . . . . . . . . . 159
                                                gnutls_x509_crq_set_key . . . . . . . . . . . . . . . . . . 159
                                                gnutls_x509_crq_set_version . . . . . . . . . . . . . . 159

gnutls_x509_crq_sign. . . . . . . . . . . . . . . . . . . . . . 160 gnutls_x509_crq_sign2 . . . . . . . . . . . . . . . . . . . . 159

                                                gnutls_x509_crt_check_hostname . . . . . . . . . . . 160
                                                gnutls_x509_crt_check_issuer . . . . . . . . . . . . . 160

gnutls_x509_crt_check_revocation. . . . . . . . . 161
gnutls_x509_crt_cpy_crl_dist_points . . . . . 161
gnutls_x509_crt_deinit . . . . . . . . . . . . . . . . . . . 161
gnutls_x509_crt_export . . . . . . . . . . . . . . . . . . . 161
gnutls_x509_crt_get_activation_time . . . . . 162
gnutls_x509_crt_get_authority_key_id . . . . 162
gnutls_x509_crt_get_ca_status . . . . . . . . . . . . 162
gnutls_x509_crt_get_crl_dist_points . . . . . 162
gnutls_x509_crt_get_dn . . . . . . . . . . . . . . . . . . . 164
gnutls_x509_crt_get_dn_by_oid . . . . . . . . . . . . 163
gnutls_x509_crt_get_dn_oid . . . . . . . . . . . . . . . 163
gnutls_x509_crt_get_expiration_time . . . . . 164
gnutls_x509_crt_get_extension_by_oid . . . . 164
gnutls_x509_crt_get_extension_oid . . . . . . . 165
gnutls_x509_crt_get_fingerprint. . . . . . . . . . 165
gnutls_x509_crt_get_issuer_dn . . . . . . . . . . . . 166
gnutls_x509_crt_get_issuer_dn_by_oid . . . . 165
gnutls_x509_crt_get_issuer_dn_oid . . . . . . . 166
gnutls_x509_crt_get_key_id . . . . . . . . . . . . . . . 167
gnutls_x509_crt_get_key_purpose_oid . . . . . 167
gnutls_x509_crt_get_key_usage . . . . . . . . . . . . 167
gnutls_x509_crt_get_pk_algorithm. . . . . . . . . 168
gnutls_x509_crt_get_pk_dsa_raw . . . . . . . . . . . 168
gnutls_x509_crt_get_pk_rsa_raw . . . . . . . . . . . 168
gnutls_x509_crt_get_serial . . . . . . . . . . . . . . . 169
gnutls_x509_crt_get_signature_algorithm

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169

gnutls_x509_crt_get_subject_alt_name . . . . 169
gnutls_x509_crt_get_subject_key_id . . . . . . 170
gnutls_x509_crt_get_version . . . . . . . . . . . . . . 170
gnutls_x509_crt_import . . . . . . . . . . . . . . . . . . . 170

gnutls_x509_crt_init. . . . . . . . . . . . . . . . . . . . . . 170

gnutls_x509_crt_list_import . . . . . . . . . . . . . . 170
gnutls_x509_crt_list_verify . . . . . . . . . . . . . . 171
gnutls_x509_crt_set_activation_time . . . . . 172
gnutls_x509_crt_set_authority_key_id . . . . 172
gnutls_x509_crt_set_ca_status . . . . . . . . . . . . 172
gnutls_x509_crt_set_crl_dist_points . . . . . 172
gnutls_x509_crt_set_crq . . . . . . . . . . . . . . . . . . 173

                                                gnutls_x509_crt_set_dn_by_oid . . . . . . . . . . . . 173
                                                gnutls_x509_crt_set_expiration_time . . . . . 173
                                                gnutls_x509_crt_set_extension_by_oid . . . . 173
                                                gnutls_x509_crt_set_issuer_dn_by_oid . . . . 174
                                                gnutls_x509_crt_set_key . . . . . . . . . . . . . . . . . . 175
                                                gnutls_x509_crt_set_key_purpose_oid . . . . . 174
                                                gnutls_x509_crt_set_key_usage . . . . . . . . . . . . 174
                                                gnutls_x509_crt_set_serial . . . . . . . . . . . . . . . 175
                                                gnutls_x509_crt_set_subject_alternative_

name. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175

                                                gnutls_x509_crt_set_subject_key_id . . . . . . 175
                                                gnutls_x509_crt_set_version . . . . . . . . . . . . . . 176

gnutls_x509_crt_sign. . . . . . . . . . . . . . . . . . . . . . 176 gnutls_x509_crt_sign2 . . . . . . . . . . . . . . . . . . . . 176

                                                gnutls_x509_crt_to_xml . . . . . . . . . . . . . . . . . . . 176
                                                gnutls_x509_crt_verify . . . . . . . . . . . . . . . . . . . 177
                                                gnutls_x509_crt_verify_data . . . . . . . . . . . . . . 177
                                                gnutls_x509_dn_oid_known . . . . . . . . . . . . . . . . . 177
                                                gnutls_x509_privkey_cpy . . . . . . . . . . . . . . . . . . 178
                                                gnutls_x509_privkey_deinit . . . . . . . . . . . . . . . 178
                                                gnutls_x509_privkey_export . . . . . . . . . . . . . . . 179
                                                gnutls_x509_privkey_export_dsa_raw . . . . . . 178
                                                gnutls_x509_privkey_export_pkcs8. . . . . . . . . 178
                                                gnutls_x509_privkey_export_rsa_raw . . . . . . 179
                                                gnutls_x509_privkey_fix . . . . . . . . . . . . . . . . . . 180
                                                gnutls_x509_privkey_generate . . . . . . . . . . . . . 180
                                                gnutls_x509_privkey_get_key_id . . . . . . . . . . . 180
                                                gnutls_x509_privkey_get_pk_algorithm . . . . 180
                                                gnutls_x509_privkey_import . . . . . . . . . . . . . . . 182
                                                gnutls_x509_privkey_import_dsa_raw . . . . . . 181
                                                gnutls_x509_privkey_import_pkcs8. . . . . . . . . 181
                                                gnutls_x509_privkey_import_rsa_raw . . . . . . 181
                                                gnutls_x509_privkey_init . . . . . . . . . . . . . . . . . 182
                                                gnutls_x509_privkey_sign_data . . . . . . . . . . . . 182
                                                gnutls_x509_privkey_verify_data. . . . . . . . . . 183

gnutls_x509_rdn_get. . . . . . . . . . . . . . . . . . . . . . . 183

                                                gnutls_x509_rdn_get_by_oid . . . . . . . . . . . . . . . 183
                                                gnutls_x509_rdn_get_oid . . . . . . . . . . . . . . . . . . 183

Bibliography

[CBCATT] Bodo Moeller, "Security of CBC Ciphersuites in SSL/TLS: Problems and Coun-

termeasures", 2002, Available from http://www.openssl.org/~bodo/tls-cbc.txt.

[GPGH] Mike Ashley, "The GNU Privacy Handbook", 2002, Available from

http://www.gnupg.org/gph/en/manual.pdf.

[GUTPKI] Peter Gutmann, "Everything you never wanted to know about PKI but were

forced to find out", Available from http://www.cs.auckland.ac.nz/~pgut001/.

[RFC2246] Tim Dierks and Christopher Allen, "The TLS Protocol Version 1.0", January

1999, Available from http://kaizi.viagenie.qc.ca/ietf/rfc/rfc2246.txt.

[RFC4346] Tim Dierks and Eric Rescorla, "The TLS Protocol Version 1.1", Match 2006,

Available from http://kaizi.viagenie.qc.ca/ietf/rfc/rfc4346.txt.

[RFC2440] Jon Callas, Lutz Donnerhacke, Hal Finney and Rodney Thayer, "OpenPGP Message Format", November 1998, Available from

http://kaizi.viagenie.qc.ca/ietf/rfc/rfc2440.txt.

[RFC4211] J. Schaad, "Internet X.509 Public Key Infrastructure Certificate Request Message Format (CRMF)", September 2005, Available from

http://kaizi.viagenie.qc.ca/ietf/rfc/rfc4211.txt.

[RFC2817] Rohit Khare and Scott Lawrence, "Upgrading to TLS Within HTTP/1.1", May

2000, Available from http://kaizi.viagenie.qc.ca/ietf/rfc/rfc2817.txt

[RFC2818] Eric Rescola, "HTTP Over TLS", May 2000, Available from

http://kaizi.viagenie.qc.ca/ietf/rfc/rfc2818.txt.

[RFC2945] Tom Wu, "The SRP Authentication and Key Exchange System", September

2000, Available from http://kaizi.viagenie.qc.ca/ietf/rfc/rfc2945.txt.

[RFC2986] Magnus Nystrom and Burt Kaliski, "PKCS 10 v1.7: Certification Request Syntax Specification", November 2000, Available from

http://kaizi.viagenie.qc.ca/ietf/rfc/rfc2986.txt.

[RFC3280] Russell Housley, Tim Polk, Warwick Ford and David Solo, "Internet X.509 Public Key Infrastructure Certificate and Certificate Revocation List (CRL) Profile", April 2002, Available from

http://kaizi.viagenie.qc.ca/ietf/rfc/rfc3280.txt.

[RFC3749] Scott Hollenbeck, "Transport Layer Security Protocol Compression Methods",

May 2004, Available from http://kaizi.viagenie.qc.ca/ietf/rfc/rfc3749.txt.

[PKCS12] RSA Laboratories, "PKCS 12 v1.0: Personal Information Exchange Syntax", June 1999, Available from http://www.rsa.com.

[RESCOLA] Eric Rescola, "SSL and TLS: Designing and Building Secure Systems", 2001

[SSL3] Alan Freier, Philip Karlton and Paul Kocher, "The SSL Protocol Version 3.0",

November 1996, Available from http://wp.netscape.com/eng/ssl3/draft302.txt.

[STEVENS] Richard Stevens, "UNIX Network Programming, Volume 1", Prentice Hall PTR, January 1998

[TLSEXT] Simon Blake-Wilson, Magnus Nystrom, David Hopwood, Jan Mikkelsen and Tim Wright, "Transport Layer Security (TLS) Extensions", June 2003, Avail-

able from http://kaizi.viagenie.qc.ca/ietf/rfc/rfc3546.txt.

[TLSPGP] Nikos Mavrogiannopoulos, "Using OpenPGP keys for TLS authentication", April 2004, Internet draft, work in progress. Available from

http://www.normos.org/ietf/draft/draft-ietf-tls-openpgp-keys-05.txt.

[TLSSRP] David Taylor, Trevor Perrin, Tom Wu and Nikos Mavrogiannopoulos, "Using SRP for TLS Authentication", August 2005, Internet draft, work in progress.

Available from http://www.normos.org/ietf/draft/draft-ietf-tls-srp-08.txt.

[TLSPSK] Pasi Eronen and Hannes Tschofenig, "Pre-shared key Ciphersuites for TLS",

December 2005, Available from http://kaizi.viagenie.qc.ca/ietf/rfc/rfc4279.txt.

[TOMSRP] Tom Wu, "The Stanford SRP Authentication Project", Available at

http://srp.stanford.edu/.

[WEGER] Arjen Lenstra and Xiaoyun Wang and Benne de Weger, "Colliding X.509 Certificates", Cryptology ePrint Archive, Report 2005/067, Available at

http://eprint.iacr.org/.